Jefferson Lab HALL A Logbook for November 08 - Full Listing

  • Main INDEX

    Sunday

  • 251549 11/30/08 23:50 bradshaw Polarization - Swing Shift
  • 251548 11/30/08 23:45 Chiranjib LHRS VDC efficiency
  • 251547 11/30/08 23:41 xqian BigBite Wire Chamber Gas Start Bubbling
  • 251546 11/30/08 23:39 a-onl Start_Run_4077,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production run, Happex Run 30488
  • 251545 11/30/08 23:38 a-onl End_of_Run_4076
  • 251544 11/30/08 23:04 xqian another Hitting efficiency plot
  • 251543 11/30/08 22:39 deng For run 4075, the Happex run is 30488
  • 251542 11/30/08 22:37 a-onl Start_Run_4076,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production run, Happex Run 30488
  • 251541 11/30/08 22:35 a-onl End_of_Run_4075
  • 251540 11/30/08 22:21 deng check list for lumis
  • 251539 11/30/08 21:33 deng happex restart at 21:30pm
  • 251538 11/30/08 21:32 a-onl Start_Run_4075,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production run with test DAQ for RICH: CRAMS_MAX_CHAN=120 back to nominal, Happex Run 30487
  • 251537 11/30/08 21:29 a-onl End_of_Run_4074
  • 251536 11/30/08 21:09 xqian Hitting efficiency for RUN 4072
  • 251535 11/30/08 21:06 xqian Data Check for the BigBite Wire Chamber Gas Problem
  • 251534 11/30/08 20:59 deng check list for lumis, run= 30487
  • 251533 11/30/08 20:43 M. Iodice Run 4073 RICH DAQ TEST hits threshold 180
  • 251532 11/30/08 20:42 Kalyan Dead time is 12% for run 4073
  • 251531 11/30/08 20:31 a-onl Start_Run_4074,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production run with test DAQ for RICH: CRAMS_MAX_CHAN=120 back to nominal, Happex Run 30487
  • 251530 11/30/08 20:27 a-onl End_of_Run_4073
  • 251529 11/30/08 20:27 adaq Target Field and Oven Temps snapshot
  • 251528 11/30/08 20:27 adaq Target Cell Temps snapshot
  • 251527 11/30/08 20:26 adaq Radiation and Ozone snapshot
  • 251526 11/30/08 20:26 adaq Hall A Tools snapshot
  • 251525 11/30/08 20:26 adaq BigBox Magnet snapshot
  • 251524 11/30/08 20:26 adaq Beam Position and Current snapshot
  • 251523 11/30/08 20:26 adaq Beam Energy in A and C snapshot
  • 251522 11/30/08 20:05 adaq Scaler rates (BigBite) snapshot
  • 251521 11/30/08 19:52 Kalyan Replaced argon/ethane gas bottles for BB wirechambers
  • 251520 11/30/08 19:23 a-onl Start_Run_4073,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production run with test DAQ for RICH: CRAMS_MAX_CHAN=180, Happex Run 30487
  • 251519 11/30/08 19:21 a-onl End_of_Run_4072
  • 251518 11/30/08 18:53 adaq Hall A general tools screen shots
  • 251517 11/30/08 18:43 H. Gao Spot++ run 4072
  • 251516 11/30/08 18:20 a-onl Start_Run_4072,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production run, Happex Run 30487
  • 251515 11/30/08 18:19 a-onl End_of_Run_4071
  • 251514 11/30/08 18:04 bradshaw Absorption Spectrum
  • 251513 11/30/08 18:02 deng check list for Lumis
  • 251512 11/30/08 17:17 deng Happex restart at 17:20 pm
  • 251511 11/30/08 17:15 a-onl Start_Run_4071,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production run, Happex Run 30487
  • 251510 11/30/08 17:13 a-onl End_of_Run_4070
  • 251509 11/30/08 16:40 adev BigBite WireChamber Chamber Turn on/off!
  • 251508 11/30/08 16:30 adev BigBite WireChamber Chamber Turn on/off!
  • 251507 11/30/08 16:18 M. Iodice DAQ rich num hits threshold scan
  • 251506 11/30/08 16:16 Kalyan BB Shower HV tripped during run 4069
  • 251505 11/30/08 16:14 a-onl Start_Run_4070,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production back to the standard config CRAMS_MAX_CHAN = 120 - Happex Run 30486
  • 251504 11/30/08 16:12 a-onl End_of_Run_4069
  • 251503 11/30/08 16:05 gaskell Day Shift Summary
  • 251502 11/30/08 16:05 adaq Scaler rates (BigBite) snapshot
  • 251501 11/30/08 16:04 adaq Beam Time Accounting
  • 251500 11/30/08 16:04 a-onl Start_Run_4069,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with new config CRAMS_MAX_CHAN = 180 - Happex Run 30486
  • 251499 11/30/08 16:02 a-onl End_of_Run_4068
  • 251498 11/30/08 15:55 a-onl Start_Run_4068,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with new config CRAMS_MAX_CHAN = 240 - Happex Run 30486
  • 251497 11/30/08 15:48 a-onl End_of_Run_4067
  • 251496 11/30/08 15:44 M. Iodice RICH HV 2150 for runs 4064 and 4065
  • 251495 11/30/08 15:34 a-onl Start_Run_4067,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with RICH at 2200 V - Happex Run 30486
  • 251494 11/30/08 15:29 a-onl End_of_Run_4066
  • 251493 11/30/08 15:26 a-onl Start_Run_4066,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Pedestal run for RICH at 2200 V - Happex Run 30486
  • 251492 11/30/08 15:20 a-onl End_of_Run_4065
  • 251491 11/30/08 14:57 xqian BigBite Wire Chamber hitting efficiency
  • 251490 11/30/08 14:49 gaskell spot++
  • 251489 11/30/08 14:33 gaskell restored links to adaql10
  • 251488 11/30/08 14:24 xqian BigBite Wire Chamber Standard Check
  • 251487 11/30/08 14:18 a-onl Start_Run_4065,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production - Happex Run 30486
  • 251486 11/30/08 14:17 a-onl End_of_Run_4064
  • 251485 11/30/08 13:59 Ole Hansen Replaced adaql10 CPU fan
  • 251484 11/30/08 13:20 M. Iodice Run 4064 - Happex is now back in DAQ
  • 251483 11/30/08 13:19 yawei ROC 23 works, happex run#30486 starts
  • 251482 11/30/08 13:18 a-onl Start_Run_4064,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production - Happex Run 30486
  • 251481 11/30/08 13:16 a-onl End_of_Run_4063
  • 251480 11/30/08 13:07 gaskell rebooted ROC 23 (parity DAQ)
  • 251479 11/30/08 12:57 M. Iodice Rich HV change from run 4063
  • 251478 11/30/08 12:55 M. Iodice Happex DAQ hanging
  • 251477 11/30/08 12:53 gaskell re-directed online analysis to adaqfs/scratch
  • 251476 11/30/08 12:41 gaskell ROC 23 down - no HAPPEX run
  • 251475 11/30/08 12:26 adaq Target Field and Oven Temps snapshot
  • 251474 11/30/08 12:26 adaq Target Cell Temps snapshot
  • 251473 11/30/08 12:26 adaq Radiation and Ozone snapshot
  • 251472 11/30/08 12:26 adaq Hall A Tools snapshot
  • 251471 11/30/08 12:26 adaq BigBox Magnet snapshot
  • 251470 11/30/08 12:26 adaq Beam Position and Current snapshot
  • 251469 11/30/08 12:26 adaq Beam Energy in A and C snapshot
  • 251468 11/30/08 12:14 a-onl Start_Run_4063,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with RICH at 2150 V - NO Happex run since ROC 23 does not work
  • 251467 11/30/08 12:05 adaq Scaler rates (BigBite) snapshot
  • 251466 11/30/08 11:59 a-onl End_of_Run_4062
  • 251465 11/30/08 11:58 gaskell disk space
  • 251464 11/30/08 11:57 a-onl Start_Run_4062,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Pedestal Run for RICH at HV=2150V - Happex Run 30485
  • 251463 11/30/08 11:55 Ole Hansen adaql10 - help is on the way
  • 251462 11/30/08 11:48 a-onl End_of_Run_4061
  • 251461 11/30/08 11:43 gaskell adaql10 stopped again
  • 251460 11/30/08 11:04 M. Iodice new happex run for runs 4059 and 4060
  • 251459 11/30/08 10:48 a-onl Start_Run_4061,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Happex Run 30485
  • 251458 11/30/08 10:47 a-onl End_of_Run_4060
  • 251456 11/30/08 10:16 Yi Qiang Updated Summary for Nov 29, Owl - Swing
  • 251455 11/30/08 10:06 yawei left arm VDC efficiency, run#4059
  • 251454 11/30/08 09:56 yawei preshower and shower energy check, run#4058
  • 251453 11/30/08 09:48 a-onl Start_Run_4060,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Happex Run 30484
  • 251452 11/30/08 09:47 a-onl End_of_Run_4059
  • 251451 11/30/08 09:33 Yi Qiang Experiment Summary for Nov 29, Owl - Swing
  • 251450 11/30/08 09:21 yawei Laser Absorption Spectrum
  • 251449 11/30/08 09:04 yawei spot++ for run #4059
  • 251448 11/30/08 08:41 a-onl Start_Run_4059,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Happex Run 30484
  • 251447 11/30/08 08:39 a-onl End_of_Run_4058
  • 251446 11/30/08 08:04 adaq Scaler rates (BigBite) snapshot
  • 251445 11/30/08 08:01 adaq Beam Time Accounting
  • 251444 11/30/08 07:58 R. Michaels owl summary
  • 251443 11/30/08 07:38 a-onl Start_Run_4058,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Happex Run 30484
  • 251442 11/30/08 07:37 a-onl End_of_Run_4057
  • 251441 11/30/08 07:07 R. Michaels tools
  • 251440 11/30/08 06:48 Yi Zhang LHRS Chamber Efficiency
  • 251439 11/30/08 06:40 a-onl Start_Run_4057,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Happex Run 30484
  • 251438 11/30/08 06:39 a-onl End_of_Run_4056
  • 251437 11/30/08 06:38 R. Michaels auto BANU problem/repair
  • 251436 11/30/08 05:47 A. Camsonne Threshold change on number of hits for RICH
  • 251435 11/30/08 05:43 a-onl Start_Run_4056,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Happex Run 30484
  • 251434 11/30/08 05:42 a-onl End_of_Run_4055
  • 251433 11/30/08 05:11 R. Michaels/ Absorption Spectrum
  • 251432 11/30/08 04:49 R. Michaels new HAPPEX run 30484
  • 251431 11/30/08 04:46 a-onl Start_Run_4055,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Happex Run 30484
  • 251430 11/30/08 04:45 a-onl End_of_Run_4054
  • 251429 11/30/08 04:41 R. Michaels DT problem back to 8% (mystery solved !!)
  • 251428 11/30/08 04:30 adev BigBite WireChamber Measured Voltage change!
  • 251427 11/30/08 04:26 adaq Target Field and Oven Temps snapshot
  • 251426 11/30/08 04:26 adaq Target Cell Temps snapshot
  • 251425 11/30/08 04:26 adaq Radiation and Ozone snapshot
  • 251424 11/30/08 04:26 adaq Hall A Tools snapshot
  • 251423 11/30/08 04:25 adaq BigBox Magnet snapshot
  • 251422 11/30/08 04:25 adaq Beam Position and Current snapshot
  • 251421 11/30/08 04:25 adaq Beam Energy in A and C snapshot
  • 251420 11/30/08 04:25 R. Michaels restarted hvgui on adaql3
  • 251419 11/30/08 04:20 adev BigBite WireChamber Measured Voltage change!
  • 251418 11/30/08 04:11 Annand 3He Target
  • 251417 11/30/08 04:09 adev BigBite WireChamber Measured Voltage change!
  • 251416 11/30/08 04:06 adaq target: Spin Flip Report: 3He Target status
  • 251415 11/30/08 04:04 adaq Scaler rates (BigBite) snapshot
  • 251414 11/30/08 04:01 R. Michaels fix start/end GUI pop-up
  • 251413 11/30/08 04:00 a-onl Start_Run_4054,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Happex Run 30483
  • 251412 11/30/08 03:56 R. Michaels DT = 25%
  • 251411 11/30/08 03:52 a-onl Start_Run_4053,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Happex Run 30483
  • 251410 11/30/08 03:50 a-onl End_of_Run_4052
  • 251409 11/30/08 03:50 R. Michaels spot++
  • 251408 11/30/08 03:49 a-onl Start_Run_4052,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Happex Run 30483
  • 251407 11/30/08 03:43 R. Michaels junk runs, power-cycled TS11
  • 251406 11/30/08 03:33 a-onl End_of_Run_4051
  • 251405 11/30/08 03:32 a-onl Start_Run_4051,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Happex Run 30483
  • 251404 11/30/08 03:25 a-onl End_of_Run_4049
  • 251403 11/30/08 03:24 adev SCALER = -1
  • 251402 11/30/08 03:18 R. Michaels high deadtime, scaler=-1
  • 251401 11/30/08 03:17 adev SCALER = -1
  • 251400 11/30/08 03:17 a-onl Start_Run_4049,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Happex Run 30483
  • 251399 11/30/08 03:16 a-onl End_of_Run_4048
  • 251398 11/30/08 03:13 a-onl Start_Run_4048,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Happex Run 30483
  • 251397 11/30/08 03:10 adev SCALER = -1
  • 251396 11/30/08 03:03 adev SCALER = -1
  • 251395 11/30/08 03:01 a-onl End_of_Run_4047
  • 251394 11/30/08 02:58 a-onl Start_Run_4047,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Happex Run 30483
  • 251393 11/30/08 02:50 a-onl End_of_Run_4046
  • 251392 11/30/08 02:45 a-onl Start_Run_4046,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Happex Run 30483
  • 251391 11/30/08 02:44 R. Michaels BB HV trip
  • 251390 11/30/08 02:38 a-onl End_of_Run_4045
  • 251389 11/30/08 02:33 a-onl Start_Run_4045,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Happex Run 30483
  • 251388 11/30/08 02:32 a-onl End_of_Run_4044
  • 251387 11/30/08 02:25 R. Michaels adaql10 died again. move to another (new) UPS.
  • 251386 11/30/08 01:39 a-onl Start_Run_4044,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Happex Run 30483
  • 251385 11/30/08 01:34 a-onl End_of_Run_4043
  • 251384 11/30/08 01:26 ycwang L-arm VDC efficiency at Run #4042
  • 251383 11/30/08 01:23 ycwang RICH Detector number of photons
  • 251382 11/30/08 01:00 Bob/Youcai hit rejection on RICH
  • 251381 11/30/08 00:37 a-onl Start_Run_4043,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Happex Run 30483
  • 251380 11/30/08 00:36 a-onl End_of_Run_4042
  • 251379 11/30/08 00:35 R. Michaels adaql10 shut itself off at 23:30
  • 251378 11/30/08 00:04 adaq Scaler rates (BigBite) snapshot
  • 251377 11/30/08 00:03 Gao Swing shift summary!
  • 251376 11/30/08 00:02 adaq Beam Time Accounting
  • Saturday

  • 251375 11/29/08 23:39 a-onl Start_Run_4042,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Happex Run 30482
  • 251374 11/29/08 23:37 a-onl End_of_Run_4041
  • 251373 11/29/08 22:41 a-onl Start_Run_4041,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Happex Run 30482
  • 251372 11/29/08 22:40 a-onl End_of_Run_4040
  • 251371 11/29/08 21:52 ycwang L-arm VDC efficiency
  • 251370 11/29/08 21:44 a-onl Start_Run_4040,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Happex Run 30482
  • 251369 11/29/08 21:44 a-onl End_of_Run_4039
  • 251368 11/29/08 20:49 Jin Golden RootFile -> 4029; Reduce replay number
  • 251367 11/29/08 20:47 a-onl Start_Run_4039,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Happex Run 30482
  • 251366 11/29/08 20:46 a-onl End_of_Run_4038
  • 251365 11/29/08 20:25 adaq Target Field and Oven Temps snapshot
  • 251364 11/29/08 20:25 adaq Target Cell Temps snapshot
  • 251363 11/29/08 20:25 adaq Radiation and Ozone snapshot
  • 251362 11/29/08 20:25 adaq Hall A Tools snapshot
  • 251361 11/29/08 20:25 adaq BigBox Magnet snapshot
  • 251360 11/29/08 20:25 adaq Beam Position and Current snapshot
  • 251359 11/29/08 20:25 adaq Beam Energy in A and C snapshot
  • 251358 11/29/08 20:04 adaq Scaler rates (BigBite) snapshot
  • 251357 11/29/08 19:50 a-onl Start_Run_4038,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Happex Run 30481
  • 251356 11/29/08 19:49 a-onl End_of_Run_4037
  • 251355 11/29/08 19:34 Jin Web Page for Target & Beam Real Time Status
  • 251354 11/29/08 19:13 M.Shabestari Target ---> TO report: Swing Shift
  • 251353 11/29/08 19:08 adaq target: Spin Flip Report: Target ---> Spin Flip Report: TO report - Swing shift
  • 251352 11/29/08 19:01 M.Shabestari Laser absorption spectrum
  • 251351 11/29/08 19:00 H. Gao Hall Tool screen
  • 251350 11/29/08 18:53 a-onl Start_Run_4037,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Happex Run 30481
  • 251349 11/29/08 18:52 a-onl End_of_Run_4036
  • 251348 11/29/08 18:47 H. Gao Spot++ on Run 4036
  • 251347 11/29/08 17:51 M.Shabestari Target Stripcharts
  • 251346 11/29/08 17:47 a-onl Start_Run_4036,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Happex Run 30481
  • 251345 11/29/08 17:46 a-onl End_of_Run_4035
  • 251344 11/29/08 17:39 adaq BigBite Wire Chamber Hitting efficiency
  • 251343 11/29/08 17:33 Jin VDC Efficiency, run # 4033
  • 251342 11/29/08 17:24 S. Sirca Replay crashes with X11 connection error
  • 251341 11/29/08 16:48 a-onl Start_Run_4035,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Happex Run 30481
  • 251340 11/29/08 16:48 M.Shabestari BigBite HV tripped
  • 251339 11/29/08 16:42 a-onl Start_Run_4034,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=JUNK JUNK JUNK
  • 251338 11/29/08 16:38 A. Camsonne Start and End of run windows
  • 251337 11/29/08 16:12 adaq Beam Time Accounting
  • 251336 11/29/08 16:08 Richard Lind Day Shift Summary
  • 251335 11/29/08 16:04 adaq Scaler rates (BigBite) snapshot
  • 251334 11/29/08 15:52 M. Iodice coda start and end run pop up window
  • 251333 11/29/08 15:51 a-onl Start_Run_4033,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=HAPPEX run 30479
  • 251332 11/29/08 15:47 xqian BigBite Wire Chamber Standard Check
  • 251331 11/29/08 15:47 a-onl End_of_Run_4032
  • 251330 11/29/08 15:41 yawei L-arm VDC efficiency, run #4029
  • 251329 11/29/08 15:37 M. Iodice runs 4030 and 4031 junk - coda problems
  • 251328 11/29/08 14:47 a-onl Start_Run_4032,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=HAPPEX run 30479
  • 251327 11/29/08 14:41 a-onl Start_Run_4031,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=HAPPEX run 30479
  • 251326 11/29/08 14:38 a-onl Start_Run_4030,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=HAPPEX run 30479
  • 251325 11/29/08 14:37 a-onl End_of_Run_4029
  • 251324 11/29/08 14:02 M. Iodice Hall A General Tool Screen
  • 251323 11/29/08 13:57 M. Iodice spot++ on run 4029
  • 251322 11/29/08 13:46 Lindgren Beam unstable
  • 251321 11/29/08 13:43 Katich Wiki - Instructions for Lumi HV Change (for experts)
  • 251320 11/29/08 13:25 a-onl Start_Run_4029,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=HAPPEX run 30479
  • 251319 11/29/08 13:24 a-onl End_of_Run_4028
  • 251318 11/29/08 12:39 Moffit Re: HAPPEX DAQ bcm4... missing asymmetry plots.
  • 251317 11/29/08 12:34 Katich Lumi Shift worker checklist updated
  • 251316 11/29/08 12:25 adaq Target Field and Oven Temps snapshot
  • 251315 11/29/08 12:25 adaq Target Cell Temps snapshot
  • 251314 11/29/08 12:25 adaq Radiation and Ozone snapshot
  • 251313 11/29/08 12:25 adaq Hall A Tools snapshot
  • 251312 11/29/08 12:25 adaq BigBox Magnet snapshot
  • 251311 11/29/08 12:25 adaq Beam Position and Current snapshot
  • 251310 11/29/08 12:24 adaq Beam Energy in A and C snapshot
  • 251309 11/29/08 12:22 a-onl Start_Run_4028,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=HAPPEX run 30479
  • 251308 11/29/08 12:20 a-onl End_of_Run_4027
  • 251307 11/29/08 12:05 Moffit HAPPEX DAQ: Feedback to use "bcm4" x10
  • 251306 11/29/08 12:04 adaq Scaler rates (BigBite) snapshot
  • 251305 11/29/08 11:21 Yi Qiang The ABU and BANU are not correct
  • 251304 11/29/08 11:21 a-onl Start_Run_4027,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=HAPPEX run 30478
  • 251303 11/29/08 11:19 a-onl End_of_Run_4026
  • 251302 11/29/08 11:18 Yi Qiang Experiment Summary for Nov 28, Owl - Swing
  • 251301 11/29/08 10:51 Lindgren Beam position monitor
  • 251300 11/29/08 10:48 Lindgren Raster size
  • 251299 11/29/08 10:45 Lindgren Hall A Tools and Beam current monitor
  • 251298 11/29/08 10:18 a-onl Start_Run_4026,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=HAPPEX run 30478
  • 251297 11/29/08 10:17 a-onl End_of_Run_4025
  • 251296 11/29/08 09:30 adaq VDC efficiency screenshot
  • 251295 11/29/08 09:17 a-onl Start_Run_4025,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=HAPPEX run 30478
  • 251294 11/29/08 09:12 a-onl End_of_Run_4024
  • 251293 11/29/08 08:31 yawei preshower and shower energy, run#4022
  • 251292 11/29/08 08:25 Lindgren Day shift beginning
  • 251291 11/29/08 08:15 a-onl Start_Run_4024,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=53,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=HAPPEX run 30478
  • 251290 11/29/08 08:14 a-onl End_of_Run_4023
  • 251289 11/29/08 08:06 adaq target: Spin Flip Report: spin flip system view
  • 251288 11/29/08 08:04 adaq Beam Time Accounting
  • 251287 11/29/08 08:04 adaq Scaler rates (BigBite) snapshot
  • 251286 11/29/08 08:03 Yi Zhang Owl Shift Summary
  • 251285 11/29/08 08:03 yawei laser absorption spectrum
  • 251284 11/29/08 07:59 Yi Zhang HallA main tools
  • 251283 11/29/08 07:57 adaq LHRS VDC efficiency
  • 251282 11/29/08 07:41 Yi Zhang spot++ @ run4023
  • 251281 11/29/08 07:32 a-onl Start_Run_4023,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=53,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=HAPPEX run 30477
  • 251280 11/29/08 07:31 a-onl End_of_Run_4022
  • 251279 11/29/08 06:26 a-onl Start_Run_4022,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=53,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=HAPPEX run 30477
  • 251278 11/29/08 06:25 a-onl End_of_Run_4021
  • 251277 11/29/08 05:22 a-onl Start_Run_4021,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=53,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=HAPPEX run 30477
  • 251276 11/29/08 05:21 a-onl End_of_Run_4020
  • 251275 11/29/08 04:24 adaq Target Field and Oven Temps snapshot
  • 251274 11/29/08 04:24 adaq Target Cell Temps snapshot
  • 251273 11/29/08 04:24 adaq Radiation and Ozone snapshot
  • 251272 11/29/08 04:24 adaq Hall A Tools snapshot
  • 251271 11/29/08 04:24 adaq BigBox Magnet snapshot
  • 251270 11/29/08 04:24 adaq Beam Position and Current snapshot
  • 251269 11/29/08 04:24 adaq Beam Energy in A and C snapshot
  • 251268 11/29/08 04:19 a-onl Start_Run_4020,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=53,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=HAPPEX run 30477
  • 251267 11/29/08 04:17 a-onl End_of_Run_4019
  • 251266 11/29/08 04:04 adaq Scaler rates (BigBite) snapshot
  • 251265 11/29/08 03:18 a-onl Start_Run_4019,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=53,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=HAPPEX run 30476
  • 251264 11/29/08 03:17 a-onl End_of_Run_4018
  • 251263 11/29/08 02:19 a-onl Start_Run_4018,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=53,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=HAPPEX run 30476
  • 251262 11/29/08 02:18 a-onl End_of_Run_4017
  • 251261 11/29/08 01:21 Katich Lumi HVs set during run 4016
  • 251260 11/29/08 01:18 a-onl Start_Run_4017,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=53,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=HAPPEX run 30476
  • 251259 11/29/08 01:16 a-onl End_of_Run_4016
  • 251258 11/29/08 00:43 Katich Lumis 1,2,5 saturated
  • 251257 11/29/08 00:14 a-onl Start_Run_4016,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=53,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Back to production
  • 251256 11/29/08 00:11 a-onl End_of_Run_4015
  • 251255 11/29/08 00:06 Chiranjib Paged Joe
  • 251254 11/29/08 00:04 adaq Scaler rates (BigBite) snapshot
  • 251253 11/29/08 00:03 adaq Beam Time Accounting
  • 251252 11/29/08 00:00 Gao swing shift summary
  • Friday

  • 251251 11/28/08 23:58 Jin Lumi 1 & 2 abnormal
  • 251250 11/28/08 23:45 M.Shabestari Laser absorption spectrum
  • 251249 11/28/08 23:43 M.Shabestari Target stripcharts
  • 251248 11/28/08 23:27 M.Shabestari Target ---> 3He
  • 251247 11/28/08 23:18 Jin Spin flip is started for production
  • 251246 11/28/08 23:12 a-onl Start_Run_4015,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=53,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Back to production
  • 251245 11/28/08 23:12 adaq target: Spin Flip Report: Spin Flip for Production Started. Pol = ~66%
  • 251244 11/28/08 22:59 Jin Target -> He3
  • 251243 11/28/08 22:57 a-onl End_of_Run_4014
  • 251242 11/28/08 22:11 Chiranjib adaql10 rebooted
  • 251241 11/28/08 22:06 M.Shabestari LASER aborption spectrum
  • 251240 11/28/08 22:05 Jin adaql10 is down
  • 251239 11/28/08 22:02 Chiranjib DISK problem /broken links
  • 251238 11/28/08 21:50 a-onl Start_Run_4014,Run_type=Calibration,target_type=Reference Cell: H2,targetpol_poldir=Transverse-in-plane,targetpol_pol=53,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Reference cell H2 135psig
  • 251237 11/28/08 21:49 a-onl End_of_Run_4013
  • 251236 11/28/08 20:31 a-onl Start_Run_4013,Run_type=Calibration,target_type=Reference Cell: H2,targetpol_poldir=Transverse-in-plane,targetpol_pol=53,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Reference cell H2 135psig
  • 251235 11/28/08 20:29 a-onl End_of_Run_4012
  • 251234 11/28/08 20:24 adaq Target Field and Oven Temps snapshot
  • 251233 11/28/08 20:24 adaq Target Cell Temps snapshot
  • 251232 11/28/08 20:24 adaq Radiation and Ozone snapshot
  • 251231 11/28/08 20:24 adaq Hall A Tools snapshot
  • 251230 11/28/08 20:24 adaq BigBox Magnet snapshot
  • 251229 11/28/08 20:24 adaq Beam Position and Current snapshot
  • 251228 11/28/08 20:24 adaq Beam Energy in A and C snapshot
  • 251227 11/28/08 20:12 M.Shabestari Target Stripcharts
  • 251226 11/28/08 20:04 adaq Scaler rates (BigBite) snapshot
  • 251225 11/28/08 19:29 a-onl Start_Run_4012,Run_type=Calibration,target_type=Reference Cell: H2,targetpol_poldir=Transverse-in-plane,targetpol_pol=53,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Reference cell H2 135psig
  • 251224 11/28/08 18:48 a-onl End_of_Run_4011
  • 251223 11/28/08 18:41 a-onl Start_Run_4011,Run_type=Calibration,target_type=Reference Cell: H2,targetpol_poldir=Transverse-in-plane,targetpol_pol=53,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Reference cell H2 135psig
  • 251222 11/28/08 18:41 M.Shabestari Target ---> H2
  • 251221 11/28/08 18:34 a-onl End_of_Run_4010
  • 251220 11/28/08 18:22 a-onl Start_Run_4010,Run_type=Calibration,target_type=Reference Cell: N2,targetpol_poldir=Transverse-in-plane,targetpol_pol=53,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Refernce cell N2 20psig
  • 251219 11/28/08 18:21 a-onl End_of_Run_4009
  • 251218 11/28/08 18:00 xqian Hitting efficiency for BigBite Wire Chamber
  • 251217 11/28/08 17:29 S. Sirca Spot++ check
  • 251216 11/28/08 17:28 S. Sirca General Tools screenshot
  • 251215 11/28/08 17:07 a-onl Start_Run_4009,Run_type=Calibration,target_type=Reference Cell: N2,targetpol_poldir=Transverse-in-plane,targetpol_pol=53,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Refernce cell N2 20psig
  • 251214 11/28/08 17:06 a-onl End_of_Run_4008
  • 251213 11/28/08 17:05 xqian BigBite Wire Chamber Standard Check
  • 251212 11/28/08 17:05 M.Shabestari Target: N2 reference
  • 251211 11/28/08 17:05 a-onl Start_Run_4008,Run_type=Calibration,target_type=Reference Cell: N2,targetpol_poldir=Transverse-in-plane,targetpol_pol=53,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Refernce cell N2 20psig
  • 251210 11/28/08 17:02 a-onl End_of_Run_4007
  • 251209 11/28/08 16:53 a-onl Start_Run_4007,Run_type=Calibration,target_type=Reference Cell: Empty,targetpol_poldir=Transverse-in-plane,targetpol_pol=53,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Refernce cell empty
  • 251208 11/28/08 16:53 M. Shabestar Target: Empty reference
  • 251207 11/28/08 16:52 a-onl End_of_Run_4006
  • 251206 11/28/08 16:50 a-onl Start_Run_4006,Run_type=Calibration,target_type=Reference Cell: Empty,targetpol_poldir=Transverse-in-plane,targetpol_pol=53,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Refernce cell empty
  • 251205 11/28/08 16:48 a-onl End_of_Run_4005
  • 251204 11/28/08 16:32 xqian BiGBite Wire Chamber Hitting efficiency
  • 251203 11/28/08 16:22 Lindgren spot++
  • 251202 11/28/08 16:08 adaq Beam Time Accounting
  • 251201 11/28/08 16:06 Lindgren day shift summary:
  • 251200 11/28/08 16:04 adaq Scaler rates (BigBite) snapshot
  • 251199 11/28/08 16:03 Kalyan Summary of reference cell runs with 5 pass upto run #3949
  • 251198 11/28/08 15:50 Lindgren Beam polarization
  • 251197 11/28/08 15:48 Lindgren Production running on reference cell
  • 251196 11/28/08 15:48 a-onl Start_Run_4005,Run_type=Calibration,target_type=Reference Cell: 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=53,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Refernce cell He-3 1.08 PSI
  • 251195 11/28/08 15:47 Lindgren Rate as Raster varies
  • 251194 11/28/08 15:46 Yi Qiang EPR Calibration for Transverse Pumping
  • 251193 11/28/08 15:43 E.Chudakov Moller results
  • 251192 11/28/08 15:40 Yi Qiang Fake spin signal enabled for reference cell runs
  • 251191 11/28/08 15:36 bradshaw Move to Reference Cell
  • 251190 11/28/08 15:29 bradshaw Moving to BeO
  • 251189 11/28/08 15:02 Lindgren Polarization measurement
  • 251188 11/28/08 14:57 bradshaw Moving Target to He3 Position
  • 251187 11/28/08 14:45 bradshaw Move to BeO
  • 251186 11/28/08 14:40 Lindgren Moller complete
  • 251185 11/28/08 13:12 E.Chudakov Moller tune
  • 251184 11/28/08 12:59 Moffit HAPPEX DAQ: investigation of dx3 signal to happex7 scaler.
  • 251183 11/28/08 12:25 Lindgren Hall A Tools
  • 251182 11/28/08 12:23 adaq Target Field and Oven Temps snapshot
  • 251181 11/28/08 12:23 adaq Target Cell Temps snapshot
  • 251180 11/28/08 12:23 adaq Radiation and Ozone snapshot
  • 251179 11/28/08 12:23 adaq Hall A Tools snapshot
  • 251178 11/28/08 12:23 adaq BigBox Magnet snapshot
  • 251177 11/28/08 12:23 adaq Beam Position and Current snapshot
  • 251176 11/28/08 12:23 adaq Beam Energy in A and C snapshot
  • 251175 11/28/08 12:10 adaq target: Spin Flip Report: Auto spin flip are ready for Transverse Pumping
  • 251174 11/28/08 12:04 adaq Scaler rates (BigBite) snapshot
  • 251173 11/28/08 12:00 Chiranjib TRANSVERSE PUMPING STARTED
  • 251172 11/28/08 11:18 A. Camsonne Compton rates stable
  • 251171 11/28/08 11:12 Kalyan BB wirechamber HV turned off for Moller measurement
  • 251170 11/28/08 11:09 Lindgren Moller Measurement
  • 251169 11/28/08 11:06 adaq target: Spin Flip Report: Report for the End of the Vertical Pumping
  • 251168 11/28/08 11:01 Yi/Chiranjib NMR Cross Calibration
  • 251167 11/28/08 10:59 E.Chudakov Start Moller tune
  • 251166 11/28/08 10:47 Lindgren Target spin change
  • 251165 11/28/08 10:45 E.Chudakov Energy lock for low currents
  • 251164 11/28/08 10:45 M. Iodice Rich Performances
  • 251163 11/28/08 10:34 yawei L-arm VDC efficiency, run #4003
  • 251162 11/28/08 10:31 Chiranjib/Yi EPR calibration
  • 251161 11/28/08 10:31 yawei preshower and shower energy, run #4003
  • 251160 11/28/08 10:30 Yi Qiang Target polarization change starts
  • 251159 11/28/08 10:30 yawei preshower and shower energy, run #4003
  • 251158 11/28/08 10:21 Yi Zhang TDC & ADC of BB Cherenkov after lowered the threshold.
  • 251157 11/28/08 09:47 bradshaw Moved Target for Moeller Test
  • 251156 11/28/08 09:45 a-onl End_of_Run_4004
  • 251155 11/28/08 09:16 a-onl Start_Run_4004,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65.2,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run with new He3 cell at 12 uA, happex run#30473
  • 251154 11/28/08 09:13 a-onl End_of_Run_4003
  • 251153 11/28/08 08:54 E.Chudakov N2 cylinder changed
  • 251152 11/28/08 08:30 Lindgen Day shift beginning
  • 251151 11/28/08 08:23 a-onl Start_Run_4003,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65.2,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run with new He3 cell at 12 uA, happex run#30473
  • 251150 11/28/08 08:16 a-onl End_of_Run_4002
  • 251149 11/28/08 08:12 a-onl Start_Run_4002,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65.2,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Pedestal run for RICH at 2200 V, happex run#30473
  • 251148 11/28/08 08:07 a-onl End_of_Run_4001
  • 251147 11/28/08 08:05 adaq Beam Time Accounting
  • 251146 11/28/08 08:04 Yi Zhang Owl Shift
  • 251145 11/28/08 08:04 adaq Scaler rates (BigBite) snapshot
  • 251144 11/28/08 07:53 Yi Zhang spot++ @ run#4001
  • 251143 11/28/08 07:43 a-onl Start_Run_4001,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65.2,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run with new He3 cell with 12 uA, happex run#30472
  • 251142 11/28/08 07:37 a-onl End_of_Run_4000
  • 251141 11/28/08 07:35 M. Iodice High Dead Time during run 4000
  • 251140 11/28/08 07:35 Yi Zhang Hall A main tools
  • 251139 11/28/08 06:06 M. Iodice Rich problem during HV scan
  • 251138 11/28/08 05:28 a-onl Start_Run_4000,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65.2,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run with new He3 cell with 12 uA and RICH HV =2200 V, happex run#30472
  • 251137 11/28/08 05:24 a-onl End_of_Run_3999
  • 251136 11/28/08 05:21 a-onl Start_Run_3999,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65.2,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Pedestal run for RICH at HV =2200 V, happex run#30472
  • 251135 11/28/08 05:17 a-onl End_of_Run_3998
  • 251134 11/28/08 04:29 Nelyubin Absorption Spectrum 11/28/08 4:25a.m.
  • 251133 11/28/08 04:24 adaq target: Spin Flip Report: SpinFlipSystemStatus
  • 251132 11/28/08 04:23 adaq Target Field and Oven Temps snapshot
  • 251131 11/28/08 04:23 adaq Target Cell Temps snapshot
  • 251130 11/28/08 04:23 adaq Radiation and Ozone snapshot
  • 251129 11/28/08 04:23 adaq Hall A Tools snapshot
  • 251128 11/28/08 04:23 adaq BigBox Magnet snapshot
  • 251127 11/28/08 04:23 adaq Beam Position and Current snapshot
  • 251126 11/28/08 04:23 adaq Beam Energy in A and C snapshot
  • 251125 11/28/08 04:09 a-onl Start_Run_3998,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65.2,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with He3 new cell at 12 uA and RICH at HV =2150 V, happex run#30472
  • 251124 11/28/08 04:04 adaq Scaler rates (BigBite) snapshot
  • 251123 11/28/08 04:03 a-onl End_of_Run_3997
  • 251122 11/28/08 04:01 a-onl Start_Run_3997,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65.2,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Pedestal run for RICH at HV =2150 V, happex run#30472
  • 251121 11/28/08 03:55 a-onl End_of_Run_3996
  • 251120 11/28/08 02:54 a-onl Start_Run_3996,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65.2,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with new cell at 12 uA and RICH at HV =2300 V, happex run#30471
  • 251119 11/28/08 02:49 a-onl End_of_Run_3995
  • 251118 11/28/08 02:46 a-onl Start_Run_3995,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65.2,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Pedestal Run for RICH at HV =2300 V, happex run#30471
  • 251117 11/28/08 02:41 a-onl End_of_Run_3994
  • 251116 11/28/08 02:39 a-onl Start_Run_3994,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65.2,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Pedestal Run for RICH at HV =2300 V, happex run#30471
  • 251115 11/28/08 02:33 a-onl End_of_Run_3993
  • 251114 11/28/08 02:10 M.Iodice Rich T/H delay scan completed
  • 251113 11/28/08 01:53 ycwang Run # 3989 for RICH pedestal run with HV=2250V, Run 3990-3992 are junk.
  • 251112 11/28/08 01:32 a-onl Start_Run_3993,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65.2,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with new cell 12 uA and RICH HV =2250, happex run#30471
  • 251111 11/28/08 01:27 a-onl End_of_Run_3992
  • 251110 11/28/08 01:27 a-onl Start_Run_3992,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65.2,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with new cell 12 uA and RICH HV =2250, happex run#30471
  • 251109 11/28/08 01:23 a-onl Start_Run_3991,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65.2,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with new cell 12 uA and RICH HV =2250, happex run#30471
  • 251108 11/28/08 01:21 a-onl Start_Run_3990,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65.2,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with new cell 12 uA and RICH HV =2250, happex run#30471
  • 251107 11/28/08 01:13 a-onl End_of_Run_3989
  • 251106 11/28/08 01:12 a-onl Start_Run_3989,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65.2,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH pedestal run for HV =2250, happex run#30471
  • 251105 11/28/08 01:06 a-onl End_of_Run_3988
  • 251104 11/28/08 01:05 a-onl Start_Run_3988,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65.2,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production 12 uA with new cell, happex run#30471
  • 251103 11/28/08 01:04 a-onl End_of_Run_3987
  • 251102 11/28/08 00:23 E.Chudakov Experiment summary for Nov 27(owl) - 27(swing)
  • 251101 11/28/08 00:04 adaq Scaler rates (BigBite) snapshot
  • 251100 11/28/08 00:01 S. Sirca Swing shift SUMMARY
  • 251099 11/28/08 00:00 a-onl Start_Run_3987,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65.2,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production 12 uA with new cell, happex run#30471
  • 251098 11/28/08 00:00 adaq Beam Time Accounting
  • 251097 11/28/08 00:00 a-onl End_of_Run_3986
  • Thursday

  • 251096 11/27/08 23:07 adaq Target (3He) states
  • 251095 11/27/08 22:57 a-onl Start_Run_3986,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65.2,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production 12 uA with new cell, new happex run#30471
  • 251094 11/27/08 22:55 a-onl End_of_Run_3985
  • 251093 11/27/08 22:26 ycwang L-arm VDC efficiency
  • 251092 11/27/08 21:54 a-onl Start_Run_3985,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65.2,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production 12 uA with new cell
  • 251091 11/27/08 21:53 a-onl End_of_Run_3984
  • 251090 11/27/08 21:24 S. Sirca General Tools screen
  • 251089 11/27/08 21:23 S. Sirca Spot++ check
  • 251088 11/27/08 21:21 ycwang new pedestal for RICH after time delay to -152 ns
  • 251087 11/27/08 21:09 ycwang pedestal run # 3983 for RICH
  • 251086 11/27/08 20:56 a-onl Start_Run_3984,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65.2,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production 12 uA with new cell and RICH time delay to -152 ns
  • 251085 11/27/08 20:52 a-onl End_of_Run_3983
  • 251084 11/27/08 20:52 a-onl Start_Run_3983,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65.2,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=pedestal run for rich after time delay change
  • 251083 11/27/08 20:48 a-onl End_of_Run_3982
  • 251082 11/27/08 20:33 adaq Target (3He) states
  • 251081 11/27/08 20:23 adaq Target Field and Oven Temps snapshot
  • 251080 11/27/08 20:23 adaq Target Cell Temps snapshot
  • 251079 11/27/08 20:22 adaq Radiation and Ozone snapshot
  • 251078 11/27/08 20:22 adaq Hall A Tools snapshot
  • 251077 11/27/08 20:22 adaq BigBox Magnet snapshot
  • 251076 11/27/08 20:22 adaq Beam Position and Current snapshot
  • 251075 11/27/08 20:22 adaq Beam Energy in A and C snapshot
  • 251074 11/27/08 20:03 adaq Scaler rates (BigBite) snapshot
  • 251073 11/27/08 19:44 a-onl Start_Run_3982,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65.2,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with new cell Maureen at 12uA, happex run#30470
  • 251072 11/27/08 19:43 a-onl End_of_Run_3981
  • 251071 11/27/08 18:43 a-onl Start_Run_3981,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65.2,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with new cell Maureen at 12uA, happex run#30470
  • 251070 11/27/08 18:43 a-onl End_of_Run_3980
  • 251069 11/27/08 17:48 yawei deadtime turn to mormal now
  • 251068 11/27/08 17:42 xqian Checked the BigBite drift time for each card
  • 251067 11/27/08 17:42 a-onl Start_Run_3980,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65.2,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with new cell Maureen at 12uA, new happex run#30470
  • 251066 11/27/08 17:41 brads BB Cerenkov beam-right PMTs turned off after run 3979
  • 251065 11/27/08 17:39 xqian Wire Map looks good
  • 251064 11/27/08 17:37 a-onl End_of_Run_3979
  • 251063 11/27/08 17:32 Kalyan Summary of reference cell runs taken after installing the new target cell
  • 251062 11/27/08 17:14 brads deadtime calculation may be thrown by the non-zero T2, T7 rates
  • 251061 11/27/08 16:52 E.Chudakov N2 cylinder replaced
  • 251060 11/27/08 16:50 yawei high deadtime
  • 251059 11/27/08 16:38 a-onl Start_Run_3979,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65.2,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with new cell Maureen at 12uA
  • 251058 11/27/08 16:37 a-onl End_of_Run_3978
  • 251057 11/27/08 16:29 brads BB Cerenkov witness PMT rates with borated plastic shielding
  • 251056 11/27/08 16:13 brads BB Cerenkov beam-right PMTs turned on during run 3978
  • 251055 11/27/08 16:03 adaq Scaler rates (BigBite) snapshot
  • 251054 11/27/08 16:01 Jin Huang Shift Summary, Nov 27 Day
  • 251053 11/27/08 16:01 adaq Beam Time Accounting
  • 251052 11/27/08 15:48 adaq target: Spin Flip Report: TO Report - Day Shift
  • 251051 11/27/08 15:26 a-onl Start_Run_3978,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.3,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with new cell Maureen at 12u
  • 251050 11/27/08 15:25 a-onl End_of_Run_3977
  • 251049 11/27/08 15:03 Kalyan VDC efficiency
  • 251048 11/27/08 14:56 Jin Hall A General Tools Screen
  • 251047 11/27/08 14:55 Jin Spot++ for Nov 27 Day shift
  • 251046 11/27/08 14:21 a-onl Start_Run_3977,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.6,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with new cell Maureen at 12u
  • 251045 11/27/08 14:19 a-onl End_of_Run_3976
  • 251044 11/27/08 13:20 Jin BB HV Error & Cleared
  • 251043 11/27/08 13:18 a-onl Start_Run_3976,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.6,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with new cell Maureen at 12u; new rich timing offset
  • 251042 11/27/08 13:17 a-onl End_of_Run_3975
  • 251041 11/27/08 12:22 adaq Target Field and Oven Temps snapshot
  • 251040 11/27/08 12:22 adaq Target Cell Temps snapshot
  • 251039 11/27/08 12:22 adaq Radiation and Ozone snapshot
  • 251038 11/27/08 12:22 adaq Hall A Tools snapshot
  • 251037 11/27/08 12:22 adaq BigBox Magnet snapshot
  • 251036 11/27/08 12:22 adaq Beam Position and Current snapshot
  • 251035 11/27/08 12:22 adaq Beam Energy in A and C snapshot
  • 251034 11/27/08 12:12 a-onl Start_Run_3975,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with new cell Maureen at 12u; new rich timing offset
  • 251033 11/27/08 12:11 a-onl End_of_Run_3974
  • 251032 11/27/08 12:03 adaq Scaler rates (BigBite) snapshot
  • 251031 11/27/08 11:12 a-onl Start_Run_3974,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with new cell Maureen at 12u; new rich timing offset
  • 251030 11/27/08 11:11 a-onl End_of_Run_3973
  • 251029 11/27/08 10:48 yawei preshower and shower energy check, run #3972
  • 251028 11/27/08 10:24 Jin Golden Root File -> Run # 3969
  • 251027 11/27/08 10:03 A. Camsonne Track and Hold delayed by 48 ns
  • 251026 11/27/08 10:03 a-onl Start_Run_3973,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with new cell Maureen at 12u; new rich timing offset
  • 251025 11/27/08 10:01 a-onl End_of_Run_3972
  • 251024 11/27/08 09:32 a-onl Start_Run_3972,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=0,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with new cell Maureen at 12u
  • 251023 11/27/08 09:31 a-onl End_of_Run_3971
  • 251022 11/27/08 08:20 a-onl Start_Run_3971,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=0,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with new cell Maureen at 12u
  • 251021 11/27/08 08:18 a-onl End_of_Run_3970
  • 251020 11/27/08 08:03 adaq Scaler rates (BigBite) snapshot
  • 251019 11/27/08 08:00 adaq Beam Time Accounting
  • 251018 11/27/08 07:56 Chiranjib OWL SHIFT SUMMARY
  • 251017 11/27/08 07:11 a-onl Start_Run_3970,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=0,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with new cell Maureen at 12u
  • 251016 11/27/08 07:09 a-onl End_of_Run_3969
  • 251015 11/27/08 06:49 M. Iodice Rich T/H delay scan
  • 251014 11/27/08 06:12 A. Camsonne T/H current setting
  • 251013 11/27/08 06:08 a-onl Start_Run_3969,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=0,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with new cell Maureen at 12u
  • 251012 11/27/08 06:05 a-onl End_of_Run_3968
  • 251011 11/27/08 04:54 a-onl Start_Run_3968,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=0,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with new cell Maureen at 12u
  • 251010 11/27/08 04:53 a-onl End_of_Run_3967
  • 251009 11/27/08 04:33 Nelyubin Absorption Spectrum
  • 251008 11/27/08 04:28 adaq target: Spin Flip Report: SpinFlipSystemControl_112708_425am
  • 251007 11/27/08 04:22 adaq Target Field and Oven Temps snapshot
  • 251006 11/27/08 04:22 adaq Target Cell Temps snapshot
  • 251005 11/27/08 04:22 adaq Radiation and Ozone snapshot
  • 251004 11/27/08 04:22 adaq Hall A Tools snapshot
  • 251003 11/27/08 04:22 adaq BigBox Magnet snapshot
  • 251002 11/27/08 04:21 adaq Beam Position and Current snapshot
  • 251001 11/27/08 04:21 adaq Beam Energy in A and C snapshot
  • 251000 11/27/08 04:03 adaq Scaler rates (BigBite) snapshot
  • 250999 11/27/08 03:59 M. Iodice Hall A General Tools Screen
  • 250998 11/27/08 03:53 adaq empty channel on BB preshower
  • 250997 11/27/08 03:52 a-onl Start_Run_3967,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=0,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with new cell Maureen at 12u
  • 250996 11/27/08 03:51 a-onl End_of_Run_3966
  • 250995 11/27/08 02:55 a-onl Start_Run_3966,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=0,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with new cell Maureen at 12u
  • 250994 11/27/08 02:54 a-onl End_of_Run_3965
  • 250993 11/27/08 02:22 Chiranjib spot++
  • 250992 11/27/08 02:00 E.Chudakov Moller on Nov 17 - possible beam energy change
  • 250991 11/27/08 01:56 a-onl Start_Run_3965,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=0,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with new cell Maureen at 12u
  • 250990 11/27/08 01:51 a-onl End_of_Run_3964
  • 250989 11/27/08 01:47 M.Iodice Rich optimization T/H delay
  • 250988 11/27/08 01:13 xqian Hitting efficiency is high
  • 250987 11/27/08 01:10 M.Iodice End of Run 3963
  • 250986 11/27/08 00:59 a-onl End_of_Run_3964
  • 250985 11/27/08 00:51 a-onl Start_Run_3964,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=0,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with new cell Maureen at 12u
  • 250984 11/27/08 00:31 E.Chudakov Experiment summary for Nov 26(owl) - 26(swing)
  • 250983 11/27/08 00:03 adaq Scaler rates (BigBite) snapshot
  • 250982 11/27/08 00:02 adaq Beam Time Accounting
  • Wednesday

  • 250981 11/26/08 23:54 J,C, Peng Swing shift summary
  • 250980 11/26/08 23:43 a-onl Start_Run_3963,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=0,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with new cell Maureen at 12u
  • 250979 11/26/08 23:40 a-onl End_of_Run_3962
  • 250978 11/26/08 23:06 Nelyubin Target System
  • 250977 11/26/08 23:04 adaq target: Spin Flip Report: SpinFlipSystemControl
  • 250976 11/26/08 22:48 adaq target: Spin Flip Report: Spin Flip System Status Report
  • 250975 11/26/08 22:43 adaq Hall A tool screen
  • 250974 11/26/08 22:37 a-onl Start_Run_3962,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=0,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with new cell Maureen at 12u
  • 250973 11/26/08 22:35 a-onl End_of_Run_3961
  • 250972 11/26/08 21:48 Peng Raster spot
  • 250971 11/26/08 21:39 Chiranjib DAQ reboot.
  • 250970 11/26/08 21:38 a-onl Start_Run_3961,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=0,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with new cell Maureen at 12u
  • 250969 11/26/08 20:40 L. Guo coda
  • 250968 11/26/08 20:33 a-onl Start_Run_3960,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=0,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with new cell Maureen at 12u
  • 250967 11/26/08 20:32 Chiranjib END of RUN 3958 problem
  • 250966 11/26/08 20:32 a-onl End_of_Run_3959
  • 250965 11/26/08 20:21 adaq Target Field and Oven Temps snapshot
  • 250964 11/26/08 20:21 adaq Target Cell Temps snapshot
  • 250963 11/26/08 20:21 adaq Radiation and Ozone snapshot
  • 250962 11/26/08 20:21 adaq Hall A Tools snapshot
  • 250961 11/26/08 20:21 adaq BigBox Magnet snapshot
  • 250960 11/26/08 20:21 adaq Beam Position and Current snapshot
  • 250959 11/26/08 20:21 adaq Beam Energy in A and C snapshot
  • 250958 11/26/08 20:17 a-onl Start_Run_3959,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=0,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with new cell Maureen at 12u
  • 250957 11/26/08 20:15 a-onl End_of_Run_3958
  • 250956 11/26/08 20:10 ycwang RUN # 3958 already started at 19:16
  • 250955 11/26/08 20:03 adaq Scaler rates (BigBite) snapshot
  • 250954 11/26/08 19:11 a-onl End_of_Run_3957
  • 250953 11/26/08 18:39 Kalyan db_DL.dat database updated with BCM and trigger information for every event
  • 250952 11/26/08 18:39 Kalyan db_DL.dat database updated with BCM and trigger information for every event
  • 250951 11/26/08 18:38 Kalyan db_DL.dat database updated with BCM and trigger information for every event
  • 250950 11/26/08 18:18 Kalyan Target polarization chart
  • 250949 11/26/08 18:07 a-onl Start_Run_3957,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=0,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with new cell Maureen at 12u
  • 250948 11/26/08 18:05 a-onl End_of_Run_3956
  • 250947 11/26/08 17:25 Moffit Re: dx3 in HAPPEX DAQ
  • 250946 11/26/08 17:23 ycwang RUN # 3951 3952 3953 is not good data for RICH. From Run #3955 the data for RICH is good.
  • 250945 11/26/08 17:10 Kalyan BCM d3 is back again in HAPPEX DAQ
  • 250944 11/26/08 17:04 a-onl Start_Run_3956,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=0,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with new cell Maureen at 12u
  • 250943 11/26/08 17:04 a-onl End_of_Run_3955
  • 250942 11/26/08 16:51 Jin Happex D3 Counting Dropping
  • 250941 11/26/08 16:27 Peng Raster profile
  • 250940 11/26/08 16:05 adaq Beam Time Accounting
  • 250939 11/26/08 16:05 a-onl Start_Run_3955,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=0,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with new cell Maureen at 12u
  • 250938 11/26/08 16:03 adaq Scaler rates (BigBite) snapshot
  • 250937 11/26/08 16:03 a-onl End_of_Run_3954
  • 250936 11/26/08 16:00 a-onl Start_Run_3954,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=0,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with new cell Maureen at 12uA , TH for RICH delayed by 120 ns
  • 250935 11/26/08 15:59 a-onl End_of_Run_3953
  • 250933 11/26/08 15:54 gaskell Day shift summary
  • 250932 11/26/08 15:52 adaq target: Spin Flip Report: TO - Day Shift
  • 250931 11/26/08 15:27 a-onl Start_Run_3953,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=0,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with new cell Maureen at 12uA
  • 250930 11/26/08 15:26 a-onl End_of_Run_3952
  • 250929 11/26/08 15:06 gaskell Spot++ (checklist)
  • 250927 11/26/08 14:35 Jin Spot++ for 1sr Production run
  • 250926 11/26/08 14:32 Jin Spot++ for 2uA; raster 5x6.6
  • 250925 11/26/08 14:31 a-onl Start_Run_3952,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=0,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with new cell Maureen at 12uA
  • 250924 11/26/08 14:31 gaskell Compton rates vs. slit position
  • 250923 11/26/08 14:30 a-onl End_of_Run_3951
  • 250922 11/26/08 14:29 a-onl Start_Run_3951,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=0,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with new cell Maureen at 12uA
  • 250921 11/26/08 14:26 Jin Trigger Screen Shot; 12uA; 3x4 raster
  • 250920 11/26/08 14:26 adaq target: Spin Flip Report: Real Spin flip started @ Pol.=~63%
  • 250919 11/26/08 14:22 gaskell raster size scan
  • 250918 11/26/08 14:20 a-onl End_of_Run_3950
  • 250917 11/26/08 14:20 Jin Trigger Screenshot; 5x6.6 raster
  • 250916 11/26/08 14:20 a-onl Start_Run_3950,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=0,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=spot++; 2uA, 5x6.6 raster
  • 250915 11/26/08 14:16 Jin Trigger Screen follow up
  • 250914 11/26/08 14:15 Jin Trigger Screenshot, 4x5 raster
  • 250913 11/26/08 14:14 Jin Trigger, 3x4
  • 250912 11/26/08 14:11 Jin Trigger 2uA, 2x2
  • 250911 11/26/08 14:07 gaskell Finished with reference cell - moving to polarized 3He
  • 250910 11/26/08 14:05 Kalyan RICH histograms doesn't look good
  • 250909 11/26/08 14:05 a-onl End_of_Run_3949
  • 250908 11/26/08 14:02 S. Sirca Change in RICH spectra
  • 250907 11/26/08 13:25 E.Chudakov Meeting today at 16:00. Current status.
  • 250906 11/26/08 13:06 a-onl Start_Run_3949,Run_type=Calibration,target_type=Reference Cell: 3He,targetpol_poldir=Vertical,targetpol_pol=0,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=12uA, raster On 3x4 MCC units. Reference cell He3 110psig - RICH T/H 0ns, HAPPEX run 30463
  • 250905 11/26/08 12:54 bradshaw Move to Reference Cell
  • 250904 11/26/08 12:50 gaskell Compton ion chamber calibration
  • 250903 11/26/08 12:21 adaq Target Field and Oven Temps snapshot
  • 250902 11/26/08 12:21 adaq Target Cell Temps snapshot
  • 250901 11/26/08 12:21 adaq Radiation and Ozone snapshot
  • 250900 11/26/08 12:21 adaq Hall A Tools snapshot
  • 250899 11/26/08 12:21 adaq BigBox Magnet snapshot
  • 250898 11/26/08 12:21 adaq Beam Position and Current snapshot
  • 250897 11/26/08 12:21 adaq Beam Energy in A and C snapshot
  • 250896 11/26/08 12:03 adaq Scaler rates (BigBite) snapshot
  • 250895 11/26/08 12:03 Jin HRS angle in database corrected
  • 250894 11/26/08 11:58 bradshaw Move to BeO
  • 250893 11/26/08 11:50 gaskell HRS ready - ask for beam
  • 250892 11/26/08 11:50 a-onl End_of_Run_3948
  • 250891 11/26/08 11:36 gaskell Back to beam permit
  • 250890 11/26/08 11:35 Kalyan Added T1, T3, T5 and T6 scaler counts in every event
  • 250889 11/26/08 11:33 Jin Fake spin signal back ON
  • 250888 11/26/08 11:28 Yi Qiang Heater is back
  • 250887 11/26/08 11:21 ewing/fansle cryo trip
  • 250886 11/26/08 11:05 Kalyan duplicate copy of T3 in L-HRS scaler
  • 250885 11/26/08 11:05 Jin He3 Heater is Down
  • 250884 11/26/08 10:49 Yi Qiang AFP Loss Test for Vertical Pumping
  • 250883 11/26/08 10:46 Yi Qiang EPR Calibration for Maureen of Vertical Pumping
  • 250882 11/26/08 10:43 E.Chudakov Faulty modules in the beam HV LeCroy
  • 250881 11/26/08 10:40 Yi Qiang New Frequency Sweep Settings for Maureen Vertical pumping
  • 250880 11/26/08 09:54 V. Sulkosky EPICS for Energy and BPMs
  • 250879 11/26/08 09:44 A. Camsonne more RICH T/H
  • 250878 11/26/08 09:41 Yi Qiang Fake spin signal stopped for calibration
  • 250877 11/26/08 09:39 A. Camsonne Track and hold delay
  • 250876 11/26/08 09:35 brads Energy and radiation EPICS data missing too
  • 250875 11/26/08 09:34 a-onl Start_Run_3948,Run_type=Cosmics,target_type=Reference Cell: 3He,targetpol_poldir=Vertical,targetpol_pol=0,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=12uA, raster On 3x4 MCC units. Reference cell He3 110psig - RICH T/H 0ns
  • 250874 11/26/08 09:33 a-onl End_of_Run_3947
  • 250873 11/26/08 09:32 a-onl Start_Run_3947,Run_type=Cosmics,target_type=Reference Cell: 3He,targetpol_poldir=Vertical,targetpol_pol=0,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=12uA, raster On 3x4 MCC units. Reference cell He3 110psig - RICH T/H 0ns
  • 250872 11/26/08 09:28 brads BPM readbacks missing since 11pm last night?
  • 250871 11/26/08 09:26 gaskell ESR down again - going to COntrolled Access
  • 250870 11/26/08 09:00 gaskell IOC vs ESR takes down HRS
  • 250869 11/26/08 08:40 a-onl End_of_Run_3946
  • 250868 11/26/08 08:22 A. Camsonne Owl shift summary
  • 250867 11/26/08 08:17 adaq Beam Time Accounting
  • 250866 11/26/08 08:05 A. Camsonne Beamline HV tripped
  • 250865 11/26/08 08:03 adaq Scaler rates (BigBite) snapshot
  • 250864 11/26/08 07:58 A. Camsonne Spot ++
  • 250863 11/26/08 07:55 M.Iodice Rates with 12 uA on 110 psig He3
  • 250862 11/26/08 07:51 A. Camsonne HV fine
  • 250861 11/26/08 07:45 M.Iodice Rich Track/Hold scan results
  • 250860 11/26/08 07:42 a-onl Start_Run_3946,Run_type=Calibration,target_type=Reference Cell: 3He,targetpol_poldir=Vertical,targetpol_pol=0,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=12uA, raster On 3x4 MCC units. Reference cell He3 110psig - RICH T/H 0ns
  • 250859 11/26/08 07:41 a-onl End_of_Run_3945
  • 250858 11/26/08 07:39 A. Camsonne HV alarm
  • 250857 11/26/08 06:50 A. Camsonne End of run failed , restarted CODA and HAPPEX DAQ
  • 250856 11/26/08 06:49 a-onl Start_Run_3945,Run_type=Calibration,target_type=Reference Cell: 3He,targetpol_poldir=Vertical,targetpol_pol=0,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=12uA, raster On 3x4 MCC units. Reference cell He3 110psig - RICH T/H 0ns
  • 250855 11/26/08 06:40 a-onl Start_Run_3944,Run_type=Calibration,target_type=Reference Cell: 3He,targetpol_poldir=Vertical,targetpol_pol=0,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=12uA, raster On 3x4 MCC units. Reference cell He3 110psig - RICH T/H 0ns
  • 250854 11/26/08 06:34 yawei Adjust BB chamber in 12uA when refrence cell filled into He3 to 110psig
  • 250853 11/26/08 06:32 yawei fill He3 into reference cell to 110psig
  • 250852 11/26/08 06:32 a-onl Start_Run_3943,Run_type=Calibration,target_type=Reference Cell: 3He,targetpol_poldir=Vertical,targetpol_pol=0,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=12uA, raster On 3x4 MCC units. Reference cell He3 110psig - RICH T/H 0ns
  • 250851 11/26/08 06:27 a-onl End_of_Run_3942
  • 250850 11/26/08 05:50 a-onl Start_Run_3942,Run_type=Calibration,target_type=Reference Cell: H2,targetpol_poldir=Vertical,targetpol_pol=0,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=12uA, raster On 3x4 MCC units. Reference cell H2 110psig - RICH T/H 0ns
  • 250849 11/26/08 05:49 A. Camsonne RICH T/H to nominal value
  • 250848 11/26/08 05:47 a-onl End_of_Run_3941
  • 250847 11/26/08 05:23 Chiranjib Laser absorption spectrum
  • 250846 11/26/08 04:42 A. Camsonne RICH T/H to -64 ns from nominal value
  • 250845 11/26/08 04:39 a-onl Start_Run_3941,Run_type=Calibration,target_type=Reference Cell: H2,targetpol_poldir=Vertical,targetpol_pol=0,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=12uA, raster On 3x4 MCC units. Reference cell H2 110psig - RICH T/H -64ns
  • 250844 11/26/08 04:37 a-onl End_of_Run_3940
  • 250843 11/26/08 04:21 adaq Target Field and Oven Temps snapshot
  • 250842 11/26/08 04:20 adaq Target Cell Temps snapshot
  • 250841 11/26/08 04:20 adaq Radiation and Ozone snapshot
  • 250840 11/26/08 04:20 adaq Hall A Tools snapshot
  • 250839 11/26/08 04:20 adaq BigBox Magnet snapshot
  • 250838 11/26/08 04:20 adaq Beam Position and Current snapshot
  • 250837 11/26/08 04:20 adaq Beam Energy in A and C snapshot
  • 250836 11/26/08 04:19 A. Camsonne Beam had trouble but now fine
  • 250835 11/26/08 04:03 adaq Scaler rates (BigBite) snapshot
  • 250834 11/26/08 03:43 A. Camsonne Compton Edet run
  • 250833 11/26/08 03:42 Chiranjib rates with 12 uA on 110 psig H2
  • 250832 11/26/08 03:39 yawei Adjust BB chamber in 12uA when reference cell filled into H2 to 110psig
  • 250831 11/26/08 03:39 a-onl Start_Run_3940,Run_type=Calibration,target_type=Reference Cell: H2,targetpol_poldir=Vertical,targetpol_pol=0,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=12uA, raster On 3x4 MCC units. Reference cell H2 110psig - RICH T/H +64ns
  • 250830 11/26/08 03:34 yawei fill H2 into refrence cell to 110psig
  • 250829 11/26/08 03:30 A. Camsonne RICH T/H moved to +64 ns from nominal
  • 250828 11/26/08 03:30 a-onl End_of_Run_3939
  • 250827 11/26/08 03:08 a-onl Start_Run_3939,Run_type=Calibration,target_type=Reference Cell: N2,targetpol_poldir=Vertical,targetpol_pol=0,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=12uA, raster On 3x4 MCC units. Reference cell N2 26psig - RICH T/H +120ns
  • 250826 11/26/08 03:07 a-onl End_of_Run_3938
  • 250825 11/26/08 03:06 a-onl Start_Run_3938,Run_type=Calibration,target_type=Reference Cell: N2,targetpol_poldir=Vertical,targetpol_pol=0,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=12uA, raster On 3x4 MCC units. Reference cell N2 26psig - RICH T/H +120ns
  • 250824 11/26/08 03:01 a-onl Start_Run_3937,Run_type=Calibration,target_type=Reference Cell: N2,targetpol_poldir=Vertical,targetpol_pol=0,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=12uA, raster On 3x4 MCC units. Reference cell N2 26psig - RICH T/H +120ns
  • 250823 11/26/08 02:58 a-onl Start_Run_3936,Run_type=Calibration,target_type=Reference Cell: N2,targetpol_poldir=Vertical,targetpol_pol=0,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=12uA, raster On 3x4 MCC units. Reference cell N2 26psig - RICH T/H +120ns
  • 250822 11/26/08 02:57 A. Camsonne ROC9 rebooted by itself during run 3935
  • 250821 11/26/08 02:55 A. Camsonne RICH T/H moved +120 ns
  • 250820 11/26/08 02:55 a-onl Start_Run_3935,Run_type=Calibration,target_type=Reference Cell: N2,targetpol_poldir=Vertical,targetpol_pol=0,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=12uA, raster On 3x4 MCC units. Reference cell N2 26psig - RICH T/H +120ns
  • 250819 11/26/08 02:54 a-onl End_of_Run_3934
  • 250818 11/26/08 02:53 Yi Zhang RE: Set angle of HRS in Hall A tools
  • 250817 11/26/08 02:41 A. Camsonne Set angle of HRS in Hall A tools
  • 250816 11/26/08 02:32 A. Camsonne RICH T/H removed 120 ns
  • 250815 11/26/08 02:31 A. Camsonne Run 3933 ROC9 crashed
  • 250814 11/26/08 02:31 a-onl Start_Run_3934,Run_type=Calibration,target_type=Reference Cell: N2,targetpol_poldir=Vertical,targetpol_pol=0,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=12uA, raster On 3x4 MCC units. Reference cell N2 26psig - RICH T/H -120ns
  • 250813 11/26/08 02:26 a-onl Start_Run_3933,Run_type=Calibration,target_type=Reference Cell: N2,targetpol_poldir=Vertical,targetpol_pol=0,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=12uA, raster On 3x4 MCC units. Reference cell N2 26psig
  • 250812 11/26/08 02:21 a-onl End_of_Run_3932
  • 250811 11/26/08 02:00 A. Camsonne Final beam position
  • 250810 11/26/08 01:47 a-onl Start_Run_3932,Run_type=Calibration,target_type=Reference Cell: N2,targetpol_poldir=Vertical,targetpol_pol=0,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=12uA, raster On 3x4 MCC units. Reference cell N2 26psig
  • 250809 11/26/08 01:45 ywang Adjust BB chamber HV in 12 uA when reference cell filled into N2 to 26psig
  • 250808 11/26/08 01:44 a-onl End_of_Run_3931
  • 250807 11/26/08 01:41 a-onl Start_Run_3931,Run_type=Calibration,target_type=Reference Cell: N2,targetpol_poldir=Vertical,targetpol_pol=0,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=12uA, raster On 3x4 MCC units. Reference cell N2 26psig
  • 250806 11/26/08 01:38 yawei fill N2 to 26 psig into reference cell
  • 250805 11/26/08 01:37 a-onl End_of_Run_3930
  • 250804 11/26/08 01:35 a-onl Start_Run_3930,Run_type=Calibration,target_type=Reference Cell: Empty,targetpol_poldir=Vertical,targetpol_pol=0,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=2uA, raster On 3x4 MCC units. Reference filling N2
  • 250803 11/26/08 01:33 A. Camsonne Empty reference cell 5 uA rates
  • 250802 11/26/08 01:32 a-onl End_of_Run_3929
  • 250801 11/26/08 01:30 A. Camsonne Spot current setting run 3929
  • 250800 11/26/08 01:20 a-onl Start_Run_3929,Run_type=Calibration,target_type=Reference Cell: Empty,targetpol_poldir=Vertical,targetpol_pol=0,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=2uA, raster On 3x4 MCC units. Reference Empty target run
  • 250799 11/26/08 01:19 a-onl End_of_Run_3928
  • 250798 11/26/08 01:17 ycwang RICH HV turned on
  • 250797 11/26/08 01:14 a-onl Start_Run_3928,Run_type=Calibration,target_type=Reference Cell: Empty,targetpol_poldir=Vertical,targetpol_pol=0,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=2uA, raster On 3x4 MCC units. Reference Empty target run
  • 250796 11/26/08 01:12 A. Camsonne Spot Run 3927
  • 250795 11/26/08 01:09 E.Chudakov iocha22 restored
  • 250794 11/26/08 01:09 a-onl End_of_Run_3927
  • 250793 11/26/08 01:08 A. Camsonne Rates with Raster 5x6.5 MCC units 2 uA
  • 250792 11/26/08 01:08 a-onl Start_Run_3927,Run_type=Test,target_type=Reference Cell: Empty,targetpol_poldir=Vertical,targetpol_pol=0,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=2uA, raster On 5x6.5 MCC units. Reference Empty target run
  • 250791 11/26/08 01:06 a-onl End_of_Run_3926
  • 250790 11/26/08 01:05 A. Camsonne Spot Run 3926
  • 250789 11/26/08 01:01 A. Camsonne Rates with raster 4x5 mm ( MCC units ) 2uA
  • 250788 11/26/08 01:00 a-onl Start_Run_3926,Run_type=Test,target_type=Reference Cell: Empty,targetpol_poldir=Vertical,targetpol_pol=0,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=2uA, raster On 4x5 MCC units. Reference Empty target run
  • 250787 11/26/08 01:00 E.Chudakov Experiment summary for Nov 24(swing) - 25(swing)
  • 250786 11/26/08 00:58 a-onl End_of_Run_3925
  • 250785 11/26/08 00:56 A. Camsonne New beam position
  • 250784 11/26/08 00:56 A. Camsonne Spot ++ Run 3925
  • 250783 11/26/08 00:53 A. Camsonne Rates new position 2 uA raster 3x4 MCC units
  • 250782 11/26/08 00:52 a-onl Start_Run_3925,Run_type=Test,target_type=Reference Cell: Empty,targetpol_poldir=Vertical,targetpol_pol=0,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=2uA, raster On. Reference Empty target run
  • 250781 11/26/08 00:48 A. Camsonne Moved to Empty reference cell
  • 250780 11/26/08 00:46 a-onl End_of_Run_3924
  • 250779 11/26/08 00:45 A. Camsonne Spot run 3924
  • 250778 11/26/08 00:43 a-onl Start_Run_3924,Run_type=Test,target_type=Multi Carbon Optics/BeO,targetpol_poldir=Vertical,targetpol_pol=0,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=1uA, No raster. BeO target run
  • 250777 11/26/08 00:28 Jin Target -> BeO
  • 250776 11/26/08 00:13 S. Sirca IOC problems
  • 250775 11/26/08 00:11 Jin Snap shot of BeO with 1uA beam
  • 250774 11/26/08 00:08 adaq Beam Time Accounting
  • 250773 11/26/08 00:06 S. Sirca Swing shift SUMMARY
  • 250772 11/26/08 00:03 adaq Scaler rates (BigBite) snapshot
  • Tuesday

  • 250771 11/25/08 23:39 Jin Target -> Ref Cell
  • 250770 11/25/08 23:15 Peng Move to BeO target
  • 250769 11/25/08 23:09 a-onl End_of_Run_3923
  • 250768 11/25/08 22:58 Re: Issues remains & Full Scalar consistency check added to online plots
  • 250767 11/25/08 22:56 a-onl Start_Run_3923,Run_type=Calibration,target_type=Reference Cell: Empty,targetpol_poldir=Vertical,targetpol_pol=0,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=5uA, 3x4 raster. Empty Ref cell run
  • 250766 11/25/08 22:53 a-onl End_of_Run_3922
  • 250765 11/25/08 22:51 a-onl Start_Run_3922,Run_type=Calibration,target_type=Reference Cell: Empty,targetpol_poldir=Vertical,targetpol_pol=0,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=5uA, 3x4 raster. Empty Ref cell run
  • 250764 11/25/08 22:50 Peng Empty reference target run
  • 250763 11/25/08 22:49 a-onl End_of_Run_3921
  • 250762 11/25/08 22:48 a-onl Start_Run_3921,Run_type=Test,target_type=Reference Cell: N2,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=spot ++, 5x7 raster, 1atm N2, tune beam
  • 250761 11/25/08 22:45 S. Sirca Spot++ for raster 5x6.6
  • 250760 11/25/08 22:42 Srish Compton Spectrum
  • 250759 11/25/08 22:42 S. Sirca Rates study with increasing raster
  • 250758 11/25/08 22:37 a-onl End_of_Run_3920
  • 250757 11/25/08 22:35 a-onl Start_Run_3920,Run_type=Test,target_type=Reference Cell: N2,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=spot ++, 5x7 raster, 1atm N2, tune beam
  • 250755 11/25/08 22:33 a-onl End_of_Run_3919
  • 250754 11/25/08 22:33 brads Firefox profiles regenerated again, also How to Fix a Broken Firefox
  • 250753 11/25/08 22:28 a-onl Start_Run_3919,Run_type=Test,target_type=Reference Cell: N2,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=spot ++, 4x5 raster, 1atm N2, tune beam
  • 250752 11/25/08 22:26 a-onl End_of_Run_3918
  • 250751 11/25/08 22:23 a-onl Start_Run_3918,Run_type=Test,target_type=Reference Cell: N2,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=spot ++, 3x4 raster, 1atm N2, tune beam
  • 250750 11/25/08 22:22 adaq Move target positions
  • 250749 11/25/08 22:20 a-onl End_of_Run_3917
  • 250748 11/25/08 22:18 a-onl Start_Run_3917,Run_type=Test,target_type=Reference Cell: N2,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=spot ++, 2x2 raster, 1atm N2, tune beam
  • 250747 11/25/08 22:17 a-onl End_of_Run_3916
  • 250746 11/25/08 22:11 S. Sirca Beam position on BeO
  • 250745 11/25/08 22:09 S. Sirca Harp scans after Compton tune
  • 250744 11/25/08 22:06 Sirish Compton Studies
  • 250743 11/25/08 22:03 a-onl Start_Run_3916,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=It works !!!!!!!!
  • 250742 11/25/08 22:02 a-onl End_of_Run_3915
  • 250741 11/25/08 21:52 S. Sirca MCC sends CW
  • 250740 11/25/08 21:43 adaq Target Field and Oven Temps snapshot
  • 250739 11/25/08 21:43 adaq Target Cell Temps snapshot
  • 250738 11/25/08 21:43 adaq Radiation and Ozone snapshot
  • 250737 11/25/08 21:43 adaq Hall A Tools snapshot
  • 250736 11/25/08 21:43 adaq BigBox Magnet snapshot
  • 250735 11/25/08 21:43 adaq Beam Position and Current snapshot
  • 250734 11/25/08 21:42 adaq Beam Energy in A and C snapshot
  • 250733 11/25/08 21:26 a-onl Start_Run_3915,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=It works !!!!!!!!
  • 250732 11/25/08 21:24 a-onl End_of_Run_3914
  • 250731 11/25/08 21:23 a-onl Start_Run_3914,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=during compton tuning test
  • 250730 11/25/08 20:03 adaq Scaler rates (BigBite) snapshot
  • 250729 11/25/08 19:33 ycwang RICH HV turned off before Beam tuning up
  • 250728 11/25/08 19:32 ycwang new rich pedestal run #3903
  • 250727 11/25/08 19:30 a-onl Start_Run_3912,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=during compton tuning
  • 250726 11/25/08 19:30 Jin Fake spin signal is ON; Target is under pumping
  • 250725 11/25/08 19:30 a-onl End_of_Run_3911
  • 250724 11/25/08 19:21 A. Camsonne Compton Electron detector current
  • 250723 11/25/08 19:13 Sirish Ref Chicane Tune
  • 250722 11/25/08 19:00 a-onl Start_Run_3911,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=during compton tuning
  • 250721 11/25/08 18:58 a-onl End_of_Run_3910
  • 250720 11/25/08 18:37 jpchen Target position survey
  • 250719 11/25/08 18:35 jpchen Reference cell
  • 250718 11/25/08 17:01 a-onl Start_Run_3910,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=
  • 250717 11/25/08 16:57 a-onl End_of_Run_3909
  • 250716 11/25/08 16:48 A. Camsonne Borrowed logic module from cabinet
  • 250715 11/25/08 16:07 brads Added shielding to BB Cerenkov
  • 250714 11/25/08 16:03 adaq Scaler rates (BigBite) snapshot
  • 250713 11/25/08 16:01 adaq Beam Time Accounting
  • 250712 11/25/08 16:01 Lingyan Zhu day shift Summary
  • 250711 11/25/08 15:25 Lingyan Zhu controlled access
  • 250710 11/25/08 14:57 a-onl Start_Run_3909,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=
  • 250709 11/25/08 14:48 a-onl End_of_Run_3908
  • 250708 11/25/08 14:47 a-onl Start_Run_3908,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=T1 1kHZ scaler check
  • 250707 11/25/08 14:44 a-onl End_of_Run_3907
  • 250706 11/25/08 14:43 a-onl Start_Run_3907,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=T1 1kHZ scaler check
  • 250705 11/25/08 14:39 a-onl End_of_Run_3906
  • 250704 11/25/08 14:38 a-onl Start_Run_3906,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=T1 1kHZ scaler check
  • 250703 11/25/08 14:30 a-onl End_of_Run_3905
  • 250702 11/25/08 13:59 a-onl Start_Run_3905,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic run
  • 250701 11/25/08 13:55 a-onl End_of_Run_3904
  • 250700 11/25/08 12:33 Kalyan Re: Issues remains & Full Scalar consistency check added to online plots
  • 250699 11/25/08 12:30 a-onl Start_Run_3904,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic run
  • 250698 11/25/08 12:28 a-onl End_of_Run_3903
  • 250697 11/25/08 12:26 a-onl Start_Run_3903,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Pedestal run for rich
  • 250696 11/25/08 12:24 a-onl End_of_Run_3902
  • 250695 11/25/08 12:23 a-onl Start_Run_3902,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Pedestal run for rich
  • 250694 11/25/08 12:21 V. Sulkosky Run 3901 Pedestal Run
  • 250693 11/25/08 12:15 a-onl End_of_Run_3901
  • 250692 11/25/08 12:13 a-onl Start_Run_3901,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Pedestal run T8
  • 250691 11/25/08 12:11 a-onl End_of_Run_3900
  • 250690 11/25/08 12:09 a-onl Start_Run_3900,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=
  • 250689 11/25/08 12:06 a-onl End_of_Run_3899
  • 250688 11/25/08 12:06 a-onl Start_Run_3899,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=SCALER CHECK bb copy in lhrs scaler
  • 250687 11/25/08 12:03 adaq Scaler rates (BigBite) snapshot
  • 250686 11/25/08 11:42 brads BB Cerenkov PMTs signal 15 and 16 swapped at ADC (fixed)
  • 250685 11/25/08 11:36 brads BB Cerenkov TDC discrim. set to -100mV
  • 250684 11/25/08 11:23 a-onl End_of_Run_3898
  • 250683 11/25/08 11:21 a-onl Start_Run_3898,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=SCALER CHECK bb copy in lhrs scaler
  • 250682 11/25/08 11:16 a-onl End_of_Run_3897
  • 250681 11/25/08 11:14 a-onl Start_Run_3897,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=SCALER CHECK bb copy in lhrs scaler
  • 250680 11/25/08 11:08 a-onl End_of_Run_3896
  • 250679 11/25/08 11:07 a-onl Start_Run_3896,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=SCALER CHECK
  • 250678 11/25/08 10:57 a-onl End_of_Run_3895
  • 250677 11/25/08 10:56 a-onl Start_Run_3895,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=SCALER CHECK
  • 250676 11/25/08 10:38 a-onl End_of_Run_3894
  • 250675 11/25/08 10:38 a-onl Start_Run_3894,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic check.. TDC replaced on ROC10 slot 2 and 3
  • 250674 11/25/08 10:35 R. Michaels adaqfs filling at alarming rate
  • 250673 11/25/08 10:33 a-onl End_of_Run_3893
  • 250672 11/25/08 10:14 a-onl Start_Run_3893,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic check.. TDC replaced on ROC10 slot 2 and 3
  • 250671 11/25/08 10:13 adev Start_Run_20180,Run_type=Test,target_type=Empty,beam_e0=2.322,beam_raster=off,beam_halfwaveplate=out,rightarm_p=0.564,rightarm_theta=105,rightarm_colli=out,comment_text=test
  • 250670 11/25/08 10:12 Chiranjib EPR RF amplifier
  • 250669 11/25/08 10:09 Kalyan Replaced 1877 TDCs on ROC10
  • 250668 11/25/08 10:09 adev Start_Run_20179,Run_type=Test,target_type=Empty,beam_e0=2.322,beam_raster=off,beam_halfwaveplate=out,rightarm_p=0.564,rightarm_theta=105,rightarm_colli=out,comment_text=test
  • 250667 11/25/08 10:06 adev End_of_Run_20179
  • 250666 11/25/08 10:02 adev End_of_Run_20178
  • 250665 11/25/08 10:01 a-onl End_of_Run_3892
  • 250664 11/25/08 10:00 a-onl Start_Run_3892,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic check.. TDC replaced on ROC10 slot 2 and 3
  • 250663 11/25/08 10:00 a-onl End_of_Run_3891
  • 250662 11/25/08 09:59 adev Start_Run_20178,Run_type=Test,target_type=Empty,beam_e0=2.322,beam_raster=off,beam_halfwaveplate=out,rightarm_p=0.564,rightarm_theta=105,rightarm_colli=out,comment_text=test
  • 250661 11/25/08 09:40 a-onl Start_Run_3891,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic check.. TDC replaced on ROC10 slot 2 and 3
  • 250660 11/25/08 09:38 Yi Qiang Lasers are Locked for cell installation
  • 250659 11/25/08 09:35 Yi Qiang NMR Cross Calibration
  • 250658 11/25/08 09:08 a-onl Start_Run_3890,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. happaex run 30460 test
  • 250657 11/25/08 09:07 a-onl End_of_Run_3889
  • 250656 11/25/08 09:02 xqian Hitting Efficiency
  • 250655 11/25/08 09:01 Chiranjib EPR Calibration
  • 250654 11/25/08 08:56 Lingyan Zhu restricted access
  • 250653 11/25/08 08:40 adaq target: Spin Flip Report: Target report after beam stops and AFP loss test. Byebye, Astral ~~~~ you did a very nice job!!
  • 250652 11/25/08 08:39 a-onl Start_Run_3889,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. happaex run 30460 test
  • 250651 11/25/08 08:37 Yi Qiang AFP Loss Test for Vertical Pumping
  • 250650 11/25/08 08:30 a-onl End_of_Run_3888
  • 250649 11/25/08 08:06 Chiranjib OWL SHIFT SUMMARY
  • 250648 11/25/08 08:06 a-onl Start_Run_3888,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. happaex run 30460 test
  • 250646 11/25/08 08:02 adaq Beam Time Accounting
  • 250645 11/25/08 08:00 a-onl End_of_Run_3887
  • 250644 11/25/08 07:50 a-onl Start_Run_3887,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. happaex run 30460 test
  • 250643 11/25/08 07:48 a-onl End_of_Run_3886
  • 250642 11/25/08 07:18 a-onl Start_Run_3886,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. happaex run 30460 test
  • 250641 11/25/08 07:07 a-onl End_of_Run_3885
  • 250640 11/25/08 06:08 a-onl Start_Run_3885,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. happaex run 30459
  • 250639 11/25/08 06:07 a-onl End_of_Run_3884
  • 250638 11/25/08 05:25 bradshaw Laser Absorption Spect- Parallel B and S
  • 250637 11/25/08 05:16 bradshaw 400 Spin Flips
  • 250636 11/25/08 05:16 adaq target: Spin Flip Report: TO - 400 Spin Flip Report
  • 250635 11/25/08 05:05 a-onl Start_Run_3884,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. happaex run 30459
  • 250634 11/25/08 05:04 a-onl End_of_Run_3883
  • 250633 11/25/08 04:03 a-onl Start_Run_3883,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. happaex run 30459
  • 250632 11/25/08 04:02 adaq Scaler rates (BigBite) snapshot
  • 250631 11/25/08 04:01 a-onl End_of_Run_3882
  • 250630 11/25/08 03:06 a-onl Start_Run_3882,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. happaex run 30459
  • 250629 11/25/08 03:05 a-onl End_of_Run_3881
  • 250628 11/25/08 02:34 Chiranjib SPOT ++
  • 250627 11/25/08 02:06 a-onl Start_Run_3881,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. happaex run 30458
  • 250626 11/25/08 02:06 a-onl End_of_Run_3880
  • 250625 11/25/08 01:14 Chiranjib COMET LASER EPICS
  • 250624 11/25/08 01:06 a-onl Start_Run_3880,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. happaex run 30458
  • 250623 11/25/08 01:06 a-onl End_of_Run_3879
  • 250622 11/25/08 00:16 Chiranjib Hall A tools / Beam current.
  • 250621 11/25/08 00:11 Chiranjib HAPPEX DAQ restarted
  • 250620 11/25/08 00:09 a-onl Start_Run_3879,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. happaex run 30458
  • 250619 11/25/08 00:08 a-onl End_of_Run_3878
  • 250618 11/25/08 00:03 S. Sirca Swing shift SUMMARY
  • 250617 11/25/08 00:02 adaq Scaler rates (BigBite) snapshot
  • 250616 11/25/08 00:02 adaq Beam Time Accounting
  • Monday

  • 250615 11/24/08 23:57 E.Chudakov Experiment summary for Nov 23(swing) - 24(day)
  • 250614 11/24/08 23:08 Chiranjib VDC TO calibration for the second period of Production
  • 250613 11/24/08 23:06 a-onl Start_Run_3878,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. happaex run 30458
  • 250612 11/24/08 23:05 a-onl End_of_Run_3877
  • 250611 11/24/08 22:52 adaq target state
  • 250610 11/24/08 22:07 a-onl Start_Run_3877,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. happaex run 30458
  • 250609 11/24/08 22:06 a-onl End_of_Run_3876
  • 250608 11/24/08 21:58 S. Sirca Spot size too small
  • 250607 11/24/08 21:53 S. Sirca Spot++ snapshot (w/o overlapping windows)
  • 250606 11/24/08 21:48 S. Sirca Spot++ snapshot
  • 250605 11/24/08 21:00 a-onl Start_Run_3876,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. happaex run 30458
  • 250604 11/24/08 20:58 a-onl End_of_Run_3875
  • 250603 11/24/08 20:20 adaq Target Field and Oven Temps snapshot
  • 250602 11/24/08 20:20 adaq Target Cell Temps snapshot
  • 250601 11/24/08 20:20 adaq Radiation and Ozone snapshot
  • 250600 11/24/08 20:20 adaq Hall A Tools snapshot
  • 250599 11/24/08 20:20 adaq BigBox Magnet snapshot
  • 250598 11/24/08 20:20 adaq Beam Position and Current snapshot
  • 250597 11/24/08 20:20 adaq Beam Energy in A and C snapshot
  • 250596 11/24/08 20:02 adaq Scaler rates (BigBite) snapshot
  • 250595 11/24/08 19:54 a-onl Start_Run_3875,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. happaex run 30458
  • 250594 11/24/08 19:51 a-onl End_of_Run_3874
  • 250593 11/24/08 18:56 a-onl Start_Run_3874,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. happaex run 30457
  • 250592 11/24/08 18:55 a-onl End_of_Run_3873
  • 250591 11/24/08 18:47 Katich Another quick Lumi update
  • 250590 11/24/08 18:19 adaq target state
  • 250589 11/24/08 18:13 adaq target state
  • 250588 11/24/08 18:00 S. Sirca Logging of runs 3871, 3872
  • 250587 11/24/08 17:56 S. Sirca hacsbc2 still in bad shape
  • 250586 11/24/08 17:52 a-onl Start_Run_3873,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. happaex run 30457
  • 250585 11/24/08 17:26 a-onl End_of_Run_3872
  • 250584 11/24/08 17:17 a-onl Start_Run_3872,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA.
  • 250583 11/24/08 16:57 D. Parno Wire chamber -- bubbling again
  • 250582 11/24/08 16:40 S. Sirca hacsbc2 behaves erratically
  • 250581 11/24/08 16:36 S. Sirca Beam energy in A and C snapshot
  • 250580 11/24/08 16:33 S. Sirca BigBite scalers snapshot
  • 250579 11/24/08 16:30 S. Sirca Hall A Tools screen
  • 250578 11/24/08 16:27 S. Sirca Spot++
  • 250577 11/24/08 16:11 a-onl Start_Run_3871,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA.
  • 250576 11/24/08 16:02 adaq Scaler rates (BigBite) snapshot
  • 250575 11/24/08 16:02 adaq Beam Time Accounting
  • 250574 11/24/08 15:57 gaskell Day Shift Summary
  • 250573 11/24/08 15:51 gaskell Beam off: magnet problem
  • 250572 11/24/08 15:49 a-onl End_of_Run_3870
  • 250571 11/24/08 15:30 brads Dump ion chamber readbacks reappeared
  • 250570 11/24/08 15:11 brads BB Cerenkov ADC vs MWDC postion, vice versa
  • 250569 11/24/08 14:53 a-onl Start_Run_3870,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA.
  • 250568 11/24/08 14:52 a-onl End_of_Run_3869
  • 250567 11/24/08 14:37 Katich Some preliminary lumi anaylsis
  • 250566 11/24/08 14:29 gaskell HALLA:p flatlined - restart BEM server
  • 250565 11/24/08 14:16 V. Sulkosky Hall A Energy Readback Flatlined
  • 250564 11/24/08 14:15 V. Sulkosky Hall A Energy Readback Flatlined
  • 250563 11/24/08 13:46 a-onl Start_Run_3869,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA.
  • 250562 11/24/08 13:45 a-onl End_of_Run_3868
  • 250561 11/24/08 12:48 a-onl Start_Run_3868,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA.
  • 250560 11/24/08 12:47 a-onl End_of_Run_3867
  • 250559 11/24/08 12:33 D. Parno Left-arm VDC efficiency
  • 250558 11/24/08 12:30 Jin Huang Issues remains & Full Scalar consistency check added to online plots
  • 250557 11/24/08 12:20 adaq Target Field and Oven Temps snapshot
  • 250556 11/24/08 12:20 adaq Target Cell Temps snapshot
  • 250555 11/24/08 12:20 adaq Radiation and Ozone snapshot
  • 250554 11/24/08 12:19 adaq Hall A Tools snapshot
  • 250553 11/24/08 12:19 adaq BigBox Magnet snapshot
  • 250552 11/24/08 12:19 adaq Beam Position and Current snapshot
  • 250551 11/24/08 12:19 adaq Beam Energy in A and C snapshot
  • 250550 11/24/08 12:06 xqian charge 2.76C
  • 250549 11/24/08 12:02 adaq Scaler rates (BigBite) snapshot
  • 250548 11/24/08 11:56 gaskell Hall A tools
  • 250547 11/24/08 11:54 gaskell Spot++
  • 250546 11/24/08 11:47 a-onl Start_Run_3867,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA.
  • 250545 11/24/08 11:47 gaskell sprinkler system test
  • 250544 11/24/08 11:46 lguo wire chamber
  • 250543 11/24/08 11:33 a-onl End_of_Run_3866
  • 250542 11/24/08 11:19 a-onl Start_Run_3866,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA.
  • 250541 11/24/08 11:17 gaskell beam was gone for RF problem
  • 250540 11/24/08 10:56 a-onl End_of_Run_3865
  • 250539 11/24/08 10:28 D. Parno Shift in BPM x
  • 250538 11/24/08 10:02 a-onl Start_Run_3865,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA.
  • 250537 11/24/08 10:00 a-onl End_of_Run_3864
  • 250536 11/24/08 09:31 yawei preshower and shower energy check, run#3864
  • 250535 11/24/08 09:01 V. Sulkosky Re: Increase in Accumulated Charge
  • 250534 11/24/08 08:49 a-onl Start_Run_3864,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA.
  • 250533 11/24/08 08:48 a-onl End_of_Run_3863
  • 250532 11/24/08 08:36 Moffit Bad BCM dx3 signal into RightHRS/happex7
  • 250531 11/24/08 08:09 Markowitz Owl shift summary
  • 250530 11/24/08 08:09 adaq Beam Time Accounting
  • 250529 11/24/08 08:02 adaq Scaler rates (BigBite) snapshot
  • 250528 11/24/08 07:59 bradshaw Laser Absorption Spect- Anti-Parallel B and S
  • 250527 11/24/08 07:52 adaq target: Spin Flip Report: TO Report - Owl Shift
  • 250526 11/24/08 07:46 a-onl Start_Run_3863,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA.
  • 250525 11/24/08 07:45 a-onl End_of_Run_3862
  • 250524 11/24/08 06:33 a-onl Start_Run_3862,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA.
  • 250523 11/24/08 06:32 a-onl End_of_Run_3861
  • 250522 11/24/08 06:24 adaq bcm & tools screenshots
  • 250521 11/24/08 05:31 Jin Huang Coda Freeze @ 5:06AM
  • 250520 11/24/08 05:24 a-onl Start_Run_3861,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA.
  • 250519 11/24/08 04:46 a-onl Start_Run_3860,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA.
  • 250518 11/24/08 04:46 a-onl End_of_Run_3859
  • 250517 11/24/08 04:19 adaq Target Field and Oven Temps snapshot
  • 250516 11/24/08 04:19 adaq Target Cell Temps snapshot
  • 250515 11/24/08 04:19 adaq Radiation and Ozone snapshot
  • 250514 11/24/08 04:19 adaq Hall A Tools snapshot
  • 250513 11/24/08 04:19 adaq BigBox Magnet snapshot
  • 250512 11/24/08 04:19 adaq Beam Position and Current snapshot
  • 250511 11/24/08 04:19 adaq Beam Energy in A and C snapshot
  • 250510 11/24/08 04:02 adaq Scaler rates (BigBite) snapshot
  • 250509 11/24/08 03:47 a-onl Start_Run_3859,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA.
  • 250508 11/24/08 03:46 a-onl End_of_Run_3858
  • 250507 11/24/08 03:41 Markowitz bcm & tools screenshots
  • 250506 11/24/08 02:51 a-onl Start_Run_3858,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA.
  • 250505 11/24/08 02:49 a-onl End_of_Run_3857
  • 250504 11/24/08 01:48 a-onl Start_Run_3857,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30452
  • 250503 11/24/08 01:47 a-onl End_of_Run_3856
  • 250502 11/24/08 00:36 a-onl Start_Run_3856,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30452
  • 250501 11/24/08 00:35 a-onl End_of_Run_3855
  • 250500 11/24/08 00:04 JEn-Chieh Pe shift summary
  • 250499 11/24/08 00:03 adaq Beam Time Accounting
  • 250498 11/24/08 00:02 adaq Scaler rates (BigBite) snapshot
  • Sunday

  • 250497 11/23/08 23:09 a-onl Start_Run_3855,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30452
  • 250496 11/23/08 23:07 a-onl End_of_Run_3854
  • 250495 11/23/08 23:02 Jin Huang RICH is integrated into Main Replay
  • 250494 11/23/08 21:58 Peng Beam is back
  • 250493 11/23/08 21:54 a-onl Start_Run_3854,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30451
  • 250492 11/23/08 21:37 Jin MCC called for 10min beam down
  • 250491 11/23/08 21:30 a-onl End_of_Run_3853
  • 250490 11/23/08 21:13 Peng Dump IOC
  • 250489 11/23/08 21:08 Peng Dump IOC
  • 250488 11/23/08 20:54 brads Dump ion chamber readbacks are zero. IOC dead?
  • 250487 11/23/08 20:33 a-onl Start_Run_3853,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30451
  • 250486 11/23/08 20:32 a-onl End_of_Run_3852
  • 250485 11/23/08 20:19 adaq Target Field and Oven Temps snapshot
  • 250484 11/23/08 20:19 adaq Target Cell Temps snapshot
  • 250483 11/23/08 20:19 adaq Radiation and Ozone snapshot
  • 250482 11/23/08 20:19 adaq Hall A Tools snapshot
  • 250481 11/23/08 20:18 adaq BigBox Magnet snapshot
  • 250480 11/23/08 20:18 adaq Beam Position and Current snapshot
  • 250479 11/23/08 20:18 adaq Beam Energy in A and C snapshot
  • 250478 11/23/08 20:02 adaq Scaler rates (BigBite) snapshot
  • 250477 11/23/08 19:34 adaq target: Spin Flip Report: 300 Successful Flips Celebration!
  • 250476 11/23/08 19:27 a-onl Start_Run_3852,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30451
  • 250475 11/23/08 19:26 a-onl End_of_Run_3851
  • 250474 11/23/08 19:01 ycwang Mauro Iodice finds the kaon's ring for the rich at run 3799
  • 250473 11/23/08 18:21 a-onl Start_Run_3851,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30451
  • 250472 11/23/08 18:20 a-onl End_of_Run_3850
  • 250471 11/23/08 17:14 a-onl Start_Run_3850,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30450
  • 250470 11/23/08 17:13 a-onl End_of_Run_3849
  • 250469 11/23/08 16:55 Peng Hall A General Tool
  • 250468 11/23/08 16:21 E.Chudakov Experiment summary for Nov 22(swing) - 23(day)
  • 250467 11/23/08 16:08 Bosted Day shift summary
  • 250466 11/23/08 16:07 adaq Beam Time Accounting
  • 250465 11/23/08 16:02 adaq Scaler rates (BigBite) snapshot
  • 250464 11/23/08 15:57 xqian Two cards in BigBite Wire Chamber need to be replaced.
  • 250463 11/23/08 15:53 a-onl Start_Run_3849,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30450
  • 250462 11/23/08 15:51 a-onl End_of_Run_3848
  • 250461 11/23/08 15:36 V. Sulkosky Re: Ignore Trigger Download
  • 250460 11/23/08 15:26 atrig Trigger-Download
  • 250459 11/23/08 15:26 adaq target: Spin Flip Report: Target log day shift
  • 250458 11/23/08 15:25 dalton Target stripchart screenshot
  • 250457 11/23/08 14:52 D. Parno HAPPEX crashed. New HAPPEX run: 30450
  • 250456 11/23/08 14:50 a-onl Start_Run_3848,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65.2,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30450
  • 250455 11/23/08 14:48 a-onl End_of_Run_3847
  • 250454 11/23/08 14:42 a-onl Start_Run_3847,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65.2,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. No HAPPEX run
  • 250453 11/23/08 14:41 a-onl End_of_Run_3846
  • 250452 11/23/08 14:40 a-onl Start_Run_3846,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65.2,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30449
  • 250451 11/23/08 14:39 a-onl End_of_Run_3845
  • 250450 11/23/08 13:37 D. Parno 3845 is production
  • 250448 11/23/08 13:35 Kalyan Run 3844 is production
  • 250447 11/23/08 13:32 a-onl Start_Run_3845,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=66.6,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30449
  • 250446 11/23/08 13:31 a-onl End_of_Run_3844
  • 250445 11/23/08 13:25 yawei preshower and shower energy check, run# 3841
  • 250444 11/23/08 12:24 Kalyan Run 3843 : pedestals for BigBite
  • 250443 11/23/08 12:23 D. Parno New HAPPEX run: 30449
  • 250442 11/23/08 12:23 a-onl Start_Run_3844,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=66.8,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30449 ( pedestal suppression back on for BB)
  • 250441 11/23/08 12:19 a-onl End_of_Run_3843
  • 250440 11/23/08 12:18 adaq Target Field and Oven Temps snapshot
  • 250439 11/23/08 12:18 adaq Target Cell Temps snapshot
  • 250438 11/23/08 12:18 adaq Radiation and Ozone snapshot
  • 250437 11/23/08 12:18 adaq Hall A Tools snapshot
  • 250436 11/23/08 12:18 adaq BigBox Magnet snapshot
  • 250435 11/23/08 12:18 adaq Beam Position and Current snapshot
  • 250434 11/23/08 12:18 adaq Beam Energy in A and C snapshot
  • 250433 11/23/08 12:10 a-onl Start_Run_3843,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=67.1,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30448 ( pedestal supression off for BB)
  • 250432 11/23/08 12:08 a-onl End_of_Run_3842
  • 250431 11/23/08 12:07 a-onl Start_Run_3842,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=67.1,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30448 ( pedestal supression off for BB)
  • 250430 11/23/08 12:03 a-onl End_of_Run_3841
  • 250429 11/23/08 12:02 adaq Scaler rates (BigBite) snapshot
  • 250428 11/23/08 11:52 D. Parno L-arm VDC efficiency
  • 250427 11/23/08 11:04 a-onl Start_Run_3841,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=67.1,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30448
  • 250426 11/23/08 10:55 a-onl End_of_Run_3840
  • 250425 11/23/08 09:56 a-onl Start_Run_3840,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=67.5,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30448
  • 250424 11/23/08 09:55 a-onl End_of_Run_3839
  • 250423 11/23/08 09:45 D. Parno New HAPPEX run: 30448
  • 250422 11/23/08 09:28 Jin run 3837 Replay is OK
  • 250421 11/23/08 09:24 Bosted Tools screen snapshot
  • 250420 11/23/08 09:17 Bosted Beam spot size
  • 250419 11/23/08 09:09 Bosted CODA reboot
  • 250418 11/23/08 08:38 a-onl Start_Run_3839,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=67.5,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30448
  • 250417 11/23/08 08:37 a-onl End_of_Run_3838
  • 250416 11/23/08 08:37 a-onl Start_Run_3838,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=67.5,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30447
  • 250415 11/23/08 08:31 a-onl End_of_Run_3837
  • 250414 11/23/08 08:13 a-onl Start_Run_3837,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=67.5,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30447
  • 250413 11/23/08 08:09 Markowitz Owl shift summary
  • 250412 11/23/08 08:07 adaq Beam Time Accounting
  • 250411 11/23/08 08:07 a-onl End_of_Run_3836
  • 250410 11/23/08 08:02 adaq Scaler rates (BigBite) snapshot
  • 250409 11/23/08 07:58 Yi Zhang Snapshot of target temperature
  • 250408 11/23/08 07:55 a-onl Start_Run_3836,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65.5,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30447
  • 250407 11/23/08 07:54 a-onl End_of_Run_3835
  • 250406 11/23/08 07:43 Yi Zhang laser spectrum
  • 250405 11/23/08 06:46 a-onl Start_Run_3835,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65.5,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30447
  • 250404 11/23/08 06:32 a-onl Start_Run_3834,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65.5,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30445
  • 250403 11/23/08 06:31 a-onl End_of_Run_3833
  • 250402 11/23/08 05:22 a-onl Start_Run_3833,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65.5,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30445
  • 250401 11/23/08 05:22 a-onl End_of_Run_3832
  • 250400 11/23/08 04:46 adaq target: Spin Flip Report: owl shift
  • 250399 11/23/08 04:18 adaq Target Field and Oven Temps snapshot
  • 250398 11/23/08 04:18 adaq Target Cell Temps snapshot
  • 250397 11/23/08 04:18 adaq Radiation and Ozone snapshot
  • 250396 11/23/08 04:18 adaq Hall A Tools snapshot
  • 250395 11/23/08 04:18 adaq BigBox Magnet snapshot
  • 250394 11/23/08 04:18 adaq Beam Position and Current snapshot
  • 250393 11/23/08 04:17 adaq Beam Energy in A and C snapshot
  • 250392 11/23/08 04:17 a-onl Start_Run_3832,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65.5,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30445
  • 250391 11/23/08 04:15 a-onl End_of_Run_3831
  • 250390 11/23/08 04:06 Markowitz poor beam quality (bcm & tools screenshots)
  • 250389 11/23/08 04:02 adaq Scaler rates (BigBite) snapshot
  • 250388 11/23/08 03:04 a-onl Start_Run_3831,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65.5,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30445
  • 250387 11/23/08 03:03 a-onl End_of_Run_3830
  • 250386 11/23/08 01:53 a-onl Start_Run_3830,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65.5,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30445
  • 250385 11/23/08 01:52 a-onl End_of_Run_3829
  • 250384 11/23/08 01:05 Markowitz spot++ screenshot
  • 250383 11/23/08 00:47 Markowitz bcm screenshot
  • 250382 11/23/08 00:43 a-onl Start_Run_3829,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65.5,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30445
  • 250381 11/23/08 00:20 E.Chudakov Experiment summary for Nov 21(swing) - 22(day)
  • 250380 11/23/08 00:02 adaq Beam Time Accounting
  • 250379 11/23/08 00:01 adaq Scaler rates (BigBite) snapshot
  • Saturday

  • 250378 11/22/08 23:58 Jen-Chieh Pe shift summary
  • 250377 11/22/08 23:38 a-onl End_of_Run_3828
  • 250376 11/22/08 23:33 Chiranjib Polarization / spin state status
  • 250375 11/22/08 23:31 a-onl Start_Run_3828,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65.5,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30445
  • 250374 11/22/08 23:30 a-onl End_of_Run_3827
  • 250373 11/22/08 22:29 a-onl Start_Run_3827,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65.5,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30445
  • 250372 11/22/08 22:28 a-onl End_of_Run_3826
  • 250371 11/22/08 21:31 a-onl Start_Run_3826,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65.5,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30445
  • 250370 11/22/08 21:29 a-onl End_of_Run_3825
  • 250369 11/22/08 20:32 a-onl Start_Run_3825,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65.5,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30444
  • 250368 11/22/08 20:32 a-onl End_of_Run_3824
  • 250367 11/22/08 20:17 adaq Target Field and Oven Temps snapshot
  • 250366 11/22/08 20:17 adaq Target Cell Temps snapshot
  • 250365 11/22/08 20:17 adaq Radiation and Ozone snapshot
  • 250364 11/22/08 20:17 adaq Hall A Tools snapshot
  • 250363 11/22/08 20:17 adaq BigBox Magnet snapshot
  • 250362 11/22/08 20:17 adaq Beam Position and Current snapshot
  • 250361 11/22/08 20:17 adaq Beam Energy in A and C snapshot
  • 250360 11/22/08 20:01 adaq Scaler rates (BigBite) snapshot
  • 250359 11/22/08 19:31 a-onl Start_Run_3824,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65.5,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30444
  • 250358 11/22/08 19:30 a-onl End_of_Run_3823
  • 250357 11/22/08 19:24 adaq target: Spin Flip Report: spin flip system control panel
  • 250356 11/22/08 19:15 yawei spin flip system control panel
  • 250355 11/22/08 19:06 yawei screen shot laser absorption spectrum
  • 250354 11/22/08 18:29 a-onl Start_Run_3823,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65.5,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30444
  • 250353 11/22/08 18:28 a-onl End_of_Run_3822
  • 250352 11/22/08 18:09 Kalyan fast clock (103.7kHz) in every event
  • 250351 11/22/08 17:46 yawei preshower and shower energy check, run#3821
  • 250350 11/22/08 17:29 a-onl Start_Run_3822,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65.5,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30444
  • 250349 11/22/08 17:27 a-onl End_of_Run_3821
  • 250348 11/22/08 16:44 Yawei Hall A general tools
  • 250347 11/22/08 16:25 a-onl Start_Run_3821,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65.5,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30443
  • 250346 11/22/08 16:24 a-onl End_of_Run_3820
  • 250345 11/22/08 16:01 adaq Scaler rates (BigBite) snapshot
  • 250344 11/22/08 16:01 adaq Beam Time Accounting
  • 250343 11/22/08 15:55 Simonetti Day shift summary
  • 250342 11/22/08 15:49 dalton Laser spectrum
  • 250341 11/22/08 15:40 adaq target: Spin Flip Report: Day shift log
  • 250340 11/22/08 15:20 D. Parno Run 3819 ended
  • 250339 11/22/08 15:18 a-onl Start_Run_3820,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65.5,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30443
  • 250338 11/22/08 15:02 Kalyan BCM u3 and d3 scaler info for every event
  • 250337 11/22/08 14:44 a-onl Start_Run_3819,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=66.8,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30443
  • 250336 11/22/08 14:43 a-onl End_of_Run_3818
  • 250335 11/22/08 13:40 D. Parno New HAPPEX run: 30443
  • 250334 11/22/08 13:40 a-onl Start_Run_3818,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=67.1,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30443
  • 250333 11/22/08 13:38 a-onl End_of_Run_3817
  • 250332 11/22/08 13:30 A. Camsonne Compton Electron detector curent
  • 250331 11/22/08 13:01 Simonetti Beam tuning ok
  • 250330 11/22/08 12:41 Simonetti Beam position tuning
  • 250329 11/22/08 12:36 a-onl Start_Run_3817,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.5,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30442
  • 250328 11/22/08 12:30 Simonetti Vacuum problem solved
  • 250327 11/22/08 12:21 Ewing Left spect. vacuum
  • 250326 11/22/08 12:18 dalton Optics target in
  • 250325 11/22/08 12:17 adaq Target Field and Oven Temps snapshot
  • 250324 11/22/08 12:17 adaq Target Cell Temps snapshot
  • 250323 11/22/08 12:17 adaq Radiation and Ozone snapshot
  • 250322 11/22/08 12:17 adaq Hall A Tools snapshot
  • 250321 11/22/08 12:17 adaq BigBox Magnet snapshot
  • 250320 11/22/08 12:17 adaq Beam Position and Current snapshot
  • 250319 11/22/08 12:17 adaq Beam Energy in A and C snapshot
  • 250318 11/22/08 12:01 adaq Scaler rates (BigBite) snapshot
  • 250317 11/22/08 11:49 Simonetti Controlled access request to check LHRS vacuum problem
  • 250316 11/22/08 11:45 a-onl End_of_Run_3816
  • 250315 11/22/08 11:18 a-onl Start_Run_3816,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=66.6,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30442
  • 250314 11/22/08 11:17 a-onl End_of_Run_3815
  • 250313 11/22/08 11:10 xqian BigBite Gas Cerenkov Study II ----- ADC
  • 250312 11/22/08 10:52 xqian Hitting efficiency
  • 250311 11/22/08 10:42 Simonetti Tech on call, Todd Ewing, will come at 12:00 to check the vacuum problem
  • 250310 11/22/08 10:27 xqian BigBite Gas Cerenkov Study I ---- TDC
  • 250309 11/22/08 10:24 Simonetti Call page number of "Tech on call" about vacuum
  • 250308 11/22/08 10:17 a-onl Start_Run_3815,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.1,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30442
  • 250307 11/22/08 10:16 a-onl End_of_Run_3814
  • 250306 11/22/08 10:09 Simonetti LHRS vacuum pressure
  • 250305 11/22/08 09:29 dalton Target screenshot
  • 250304 11/22/08 09:23 Jin 1% inconsistency between T3 scalar counts in bigbite and LHRS
  • 250303 11/22/08 09:18 D. Parno Beam Position
  • 250302 11/22/08 09:14 Simonetti LHRS vacuum pressure
  • 250301 11/22/08 09:13 D. Parno New HAPPEX run: 30442
  • 250300 11/22/08 09:12 a-onl Start_Run_3814,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=67.2,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30442
  • 250299 11/22/08 09:10 a-onl End_of_Run_3813
  • 250298 11/22/08 08:58 D. Parno Left-arm VDC efficiency
  • 250297 11/22/08 08:12 a-onl Start_Run_3813,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.6,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30441
  • 250296 11/22/08 08:10 a-onl End_of_Run_3812
  • 250295 11/22/08 08:01 adaq Scaler rates (BigBite) snapshot
  • 250294 11/22/08 08:01 Markowitz Owl shift summary
  • 250293 11/22/08 08:00 adaq Beam Time Accounting
  • 250292 11/22/08 07:48 Vilardi Laser Absorption Spectrum
  • 250291 11/22/08 07:48 adaq target: Spin Flip Report: Halor Report owl shift 11/22/2008
  • 250290 11/22/08 07:13 a-onl Start_Run_3812,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=66.6,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA.
  • 250289 11/22/08 07:12 a-onl End_of_Run_3811
  • 250288 11/22/08 06:50 Jin LHRS vacuum followup
  • 250287 11/22/08 06:39 Markowitz LHRS vacuum slowly worsens
  • 250286 11/22/08 06:14 a-onl Start_Run_3811,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=66.6,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA.
  • 250285 11/22/08 06:13 a-onl End_of_Run_3810
  • 250284 11/22/08 05:25 Jin re:re: HAPPEX run#
  • 250283 11/22/08 05:19 a-onl Start_Run_3810,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=66.6,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA.
  • 250282 11/22/08 05:16 a-onl End_of_Run_3809
  • 250281 11/22/08 05:12 Chiranjib HAPPEX DAQ restarted.
  • 250280 11/22/08 05:08 a-onl Start_Run_3809,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=66.6,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30440
  • 250279 11/22/08 05:07 a-onl End_of_Run_3808
  • 250278 11/22/08 05:03 a-onl Start_Run_3808,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=66.6,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30440
  • 250277 11/22/08 05:01 a-onl End_of_Run_3807
  • 250276 11/22/08 04:16 adaq Target Field and Oven Temps snapshot
  • 250275 11/22/08 04:16 adaq Target Cell Temps snapshot
  • 250274 11/22/08 04:16 adaq Radiation and Ozone snapshot
  • 250273 11/22/08 04:16 adaq Hall A Tools snapshot
  • 250272 11/22/08 04:16 adaq BigBox Magnet snapshot
  • 250271 11/22/08 04:16 adaq Beam Position and Current snapshot
  • 250270 11/22/08 04:16 adaq Beam Energy in A and C snapshot
  • 250269 11/22/08 04:07 R. Michaels re HAPPEX run#
  • 250268 11/22/08 04:01 adaq Scaler rates (BigBite) snapshot
  • 250267 11/22/08 04:00 a-onl Start_Run_3807,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=66.6,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30440
  • 250266 11/22/08 03:59 a-onl End_of_Run_3806
  • 250265 11/22/08 03:03 a-onl Start_Run_3806,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=66.6,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30440
  • 250264 11/22/08 03:02 a-onl End_of_Run_3805
  • 250263 11/22/08 02:31 Markowitz LHRS vacuum followup
  • 250262 11/22/08 02:08 Jin CODA not working @ run 3804
  • 250261 11/22/08 02:04 a-onl Start_Run_3805,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=66.6,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30440
  • 250260 11/22/08 02:00 a-onl Start_Run_3804,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=66.6,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30440, recover after beam down
  • 250259 11/22/08 01:59 a-onl End_of_Run_3803
  • 250258 11/22/08 01:42 Markowitz RICH screenshots
  • 250257 11/22/08 01:41 Markowitz hall A tools &bb screen
  • 250256 11/22/08 01:39 Markowitz bcm screenshot
  • 250255 11/22/08 01:02 a-onl Start_Run_3803,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=66.6,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30440, recover after beam down
  • 250254 11/22/08 00:59 a-onl End_of_Run_3802
  • 250253 11/22/08 00:05 Simonetti Beam problem solved
  • 250252 11/22/08 00:02 Simonetti Swing shift summary
  • 250251 11/22/08 00:02 adaq Beam Time Accounting
  • 250250 11/22/08 00:01 a-onl Start_Run_3802,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=66.6,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30439, recover after beam down
  • 250249 11/22/08 00:01 adaq Scaler rates (BigBite) snapshot
  • Friday

  • 250248 11/21/08 23:48 Simonetti LHSR vacuum pressure
  • 250247 11/21/08 23:45 Simonetti Beam problem
  • 250246 11/21/08 23:39 a-onl End_of_Run_3801
  • 250245 11/21/08 23:09 a-onl Start_Run_3801,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.1,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30439
  • 250244 11/21/08 23:08 a-onl End_of_Run_3800
  • 250243 11/21/08 23:08 adaq target: Spin Flip Report: Swing shift log
  • 250242 11/21/08 22:16 a-onl Start_Run_3800,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=64.1,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA. HAPPEX run 30439
  • 250241 11/21/08 22:14 a-onl End_of_Run_3799
  • 250240 11/21/08 21:40 A. Camsonne Happex run number in regular DAQ start of run
  • 250239 11/21/08 21:17 D. Parno New HAPPEX run: 30439
  • 250238 11/21/08 21:15 a-onl Start_Run_3799,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=67.4,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA
  • 250237 11/21/08 21:14 a-onl End_of_Run_3798
  • 250236 11/21/08 21:14 a-onl Start_Run_3798,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=67.4,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA
  • 250235 11/21/08 21:12 a-onl End_of_Run_3797
  • 250234 11/21/08 20:28 Simonetti Beam problem
  • 250233 11/21/08 20:16 adaq Target Field and Oven Temps snapshot
  • 250232 11/21/08 20:16 adaq Target Cell Temps snapshot
  • 250231 11/21/08 20:16 adaq Radiation and Ozone snapshot
  • 250230 11/21/08 20:16 adaq Hall A Tools snapshot
  • 250229 11/21/08 20:16 adaq BigBox Magnet snapshot
  • 250228 11/21/08 20:16 adaq Beam Position and Current snapshot
  • 250227 11/21/08 20:16 adaq Beam Energy in A and C snapshot
  • 250226 11/21/08 20:15 a-onl Start_Run_3797,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65.2,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA
  • 250225 11/21/08 20:12 a-onl End_of_Run_3796
  • 250224 11/21/08 20:01 adaq Scaler rates (BigBite) snapshot
  • 250223 11/21/08 20:01 a-onl Start_Run_3796,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=66.6,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA
  • 250222 11/21/08 19:57 a-onl End_of_Run_3795
  • 250221 11/21/08 19:47 ycwang Pions' Ring for RICH at runnumber 3782'
  • 250220 11/21/08 19:00 a-onl Start_Run_3795,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65.4,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA
  • 250219 11/21/08 18:58 a-onl End_of_Run_3794
  • 250218 11/21/08 18:55 V. Sulkosky T0 Determined for Run 3789
  • 250217 11/21/08 18:48 V. Sulkosky VDC T0 Calibration
  • 250216 11/21/08 18:46 adaq target screenshot
  • 250215 11/21/08 18:23 E.Chudakov Experiment summary for Nov 20(swing) - 21(day)
  • 250214 11/21/08 18:12 Jin Golden Run -> 3787
  • 250213 11/21/08 17:59 a-onl Start_Run_3794,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=66.9,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA
  • 250212 11/21/08 17:58 a-onl End_of_Run_3793
  • 250211 11/21/08 17:12 D. Parno Left-arm VDC efficiency
  • 250210 11/21/08 17:07 D. Parno New HAPPEX run: 30438
  • 250209 11/21/08 17:05 a-onl Start_Run_3793,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=66.8,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA
  • 250208 11/21/08 17:02 a-onl End_of_Run_3792
  • 250207 11/21/08 16:05 F. GaribaldI Day shift summary
  • 250206 11/21/08 16:04 adaq Beam Time Accounting
  • 250205 11/21/08 16:01 adaq Scaler rates (BigBite) snapshot
  • 250204 11/21/08 16:01 a-onl Start_Run_3792,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=67.6,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA
  • 250203 11/21/08 16:00 a-onl End_of_Run_3791
  • 250202 11/21/08 15:00 brads BB Cerenkov rates with front added front shielding
  • 250201 11/21/08 14:58 a-onl Start_Run_3791,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=67.6,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA
  • 250200 11/21/08 14:57 a-onl End_of_Run_3790
  • 250199 11/21/08 14:28 Holmstrom spot++
  • 250198 11/21/08 13:47 xqian What I plotted is a preshower ADC
  • 250197 11/21/08 13:37 xqian Another plot
  • 250196 11/21/08 13:21 Yi Qiang A Typo .... :S
  • 250195 11/21/08 13:17 a-onl Start_Run_3790,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=67.6,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA
  • 250194 11/21/08 13:15 Yi Qiang Some thoughts on the gas Cherenkov
  • 250193 11/21/08 13:15 ellie HAPPEX Restarted, Run 30437
  • 250192 11/21/08 13:14 a-onl End_of_Run_3789
  • 250191 11/21/08 12:32 Yi Zhang analysis on BB Cherenkov signal Fig3
  • 250190 11/21/08 12:30 Yi Zhang analysis on BB Cherenkov signal
  • 250189 11/21/08 12:16 adaq Target Field and Oven Temps snapshot
  • 250188 11/21/08 12:16 adaq Target Cell Temps snapshot
  • 250187 11/21/08 12:15 adaq Radiation and Ozone snapshot
  • 250186 11/21/08 12:15 adaq Hall A Tools snapshot
  • 250185 11/21/08 12:15 adaq BigBox Magnet snapshot
  • 250184 11/21/08 12:15 adaq Beam Position and Current snapshot
  • 250183 11/21/08 12:15 adaq Beam Energy in A and C snapshot
  • 250182 11/21/08 12:09 a-onl Start_Run_3789,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=66.8,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA
  • 250181 11/21/08 12:08 a-onl End_of_Run_3788
  • 250180 11/21/08 12:01 adaq Scaler rates (BigBite) snapshot
  • 250179 11/21/08 11:08 xqian The HV trip, and run starts before the chamber completely recover
  • 250178 11/21/08 11:05 a-onl Start_Run_3788,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=66.8,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA
  • 250177 11/21/08 11:03 a-onl End_of_Run_3787
  • 250176 11/21/08 11:02 adaq target: Spin Flip Report: Stepper Motor #1423 warning
  • 250175 11/21/08 11:01 V. Sulkosky Re: U10 BCM Double Pulsing (fixed)
  • 250174 11/21/08 10:29 ycwang Good data for the RICH after turned on HV(2200V) from last night
  • 250173 11/21/08 10:11 yawei preshower and shower energy check, run# 3785
  • 250172 11/21/08 10:07 yawei preshower and shower energy check
  • 250171 11/21/08 10:04 E.Chudakov Beam is back
  • 250170 11/21/08 09:59 a-onl Start_Run_3787,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=66.8,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run at 12uA
  • 250169 11/21/08 09:38 Correction of the time
  • 250168 11/21/08 09:36 Correction
  • 250167 11/21/08 09:12 E.Chudakov The compressor is turned off
  • 250166 11/21/08 09:11 ellie HAPPEX Restarted, Run 30436
  • 250165 11/21/08 09:09 E.Chudakov 1/2 wave plate is out
  • 250164 11/21/08 08:57 a-onl End_of_Run_3786
  • 250163 11/21/08 08:45 A. Camsonne RICH data
  • 250162 11/21/08 08:30 E.Chudakov Control access to address the smoke alert
  • 250161 11/21/08 08:23 a-onl Start_Run_3786,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Cosmic
  • 250160 11/21/08 08:21 a-onl End_of_Run_3785
  • 250159 11/21/08 08:16 Moffit Re: U10 Double pulsing.
  • 250158 11/21/08 08:15 xqian The Wire Chamber is not working at the start of this run
  • 250157 11/21/08 08:02 adaq Beam Time Accounting
  • 250156 11/21/08 08:01 adaq Scaler rates (BigBite) snapshot
  • 250155 11/21/08 07:46 Markowitz Owl shift summary
  • 250154 11/21/08 07:43 a-onl Start_Run_3785,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with 12uA and RICH, HAPPEX 30434
  • 250153 11/21/08 07:43 Vilardi La
  • 250152 11/21/08 07:43 a-onl End_of_Run_3784
  • 250151 11/21/08 07:42 adaq target: Spin Flip Report: Halog Report Owl shift 11/21/2008
  • 250150 11/21/08 07:32 adaq hitting efficiency for 3771
  • 250149 11/21/08 07:26 R. Michaels re: CRL for RICH
  • 250148 11/21/08 07:11 adaq tools screen
  • 250147 11/21/08 06:37 adev BigBite WireChamber Measured Voltage change!
  • 250146 11/21/08 06:36 Yi Zhang BB Chamber HV tripped
  • 250145 11/21/08 06:28 a-onl Start_Run_3784,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with 12uA and RICH, HAPPEX 30434
  • 250144 11/21/08 06:27 R. Michaels how to back out (RICH DAQ change)
  • 250143 11/21/08 06:26 adev BigBite WireChamber Measured Voltage change!
  • 250142 11/21/08 06:23 a-onl End_of_Run_3783
  • 250141 11/21/08 06:09 a-onl Start_Run_3783,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with 12uA and RICH, HAPPEX 30434
  • 250140 11/21/08 06:08 a-onl End_of_Run_3782
  • 250139 11/21/08 05:14 a-onl Start_Run_3782,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with 12uA and RICH, HAPPEX 30434
  • 250138 11/21/08 05:13 a-onl End_of_Run_3781
  • 250137 11/21/08 05:10 xqian Negative reading in scalers
  • 250136 11/21/08 04:50 Yi Zhang BB Cherenkov TDC Signal
  • 250135 11/21/08 04:15 adaq Target Field and Oven Temps snapshot
  • 250134 11/21/08 04:15 adaq Target Cell Temps snapshot
  • 250133 11/21/08 04:15 adaq Radiation and Ozone snapshot
  • 250132 11/21/08 04:15 adaq Hall A Tools snapshot
  • 250131 11/21/08 04:15 adaq BigBox Magnet snapshot
  • 250130 11/21/08 04:15 adaq Beam Position and Current snapshot
  • 250129 11/21/08 04:15 adaq Beam Energy in A and C snapshot
  • 250128 11/21/08 04:05 Chiranjib HAPPEX restarted
  • 250127 11/21/08 04:01 a-onl Start_Run_3781,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with 12uA and RICH, HAPPEX 30434
  • 250126 11/21/08 04:01 adaq Scaler rates (BigBite) snapshot
  • 250125 11/21/08 03:59 a-onl End_of_Run_3780
  • 250124 11/21/08 03:40 a-onl Start_Run_3780,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with 12uA and RICH, HAPPEX 30433
  • 250123 11/21/08 03:39 a-onl End_of_Run_3779
  • 250122 11/21/08 02:34 a-onl Start_Run_3779,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with 12uA and RICH, HAPPEX 30433
  • 250121 11/21/08 02:32 a-onl End_of_Run_3778
  • 250120 11/21/08 01:24 a-onl Start_Run_3778,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with 12uA and RICH, HAPPEX 30433
  • 250119 11/21/08 01:23 a-onl End_of_Run_3777
  • 250118 11/21/08 01:08 Chiranjib Internal Temperature test
  • 250117 11/21/08 00:16 a-onl Start_Run_3777,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with 12uA and RICH, HAPPEX 30433
  • 250116 11/21/08 00:15 a-onl End_of_Run_3776
  • 250115 11/21/08 00:04 adaq Beam Time Accounting
  • 250114 11/21/08 00:03 dhamilton Swing Shift Summary
  • 250113 11/21/08 00:01 adaq Scaler rates (BigBite) snapshot
  • Thursday

  • 250112 11/20/08 23:40 deng screen shot of Lumis check for run 30433
  • 250111 11/20/08 23:38 dhamilton RE: Hall A VESDAs....
  • 250110 11/20/08 23:25 Jin HAPPEX 30433 started
  • 250109 11/20/08 23:18 Kalyan BCM U10 seems to be double pulsing again on LHRS
  • 250108 11/20/08 23:14 a-onl Start_Run_3776,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with 12uA and RICH, HAPPEX 30433
  • 250107 11/20/08 23:13 a-onl End_of_Run_3775
  • 250106 11/20/08 22:54 dhamilton VESDA Warnings
  • 250105 11/20/08 22:43 Chiranjib The RUN number for the "dead time issue" was 3774
  • 250104 11/20/08 22:39 Chiranjib target spin flip screens / Laser absorption etc.
  • 250103 11/20/08 22:37 Chiranjib Target screens
  • 250102 11/20/08 22:36 dhamilton Beam position unstable around DAQ crash
  • 250101 11/20/08 22:34 Chiranjib DEAD TIME PROBLEM OCCURED / ROCs REBOOTED
  • 250100 11/20/08 22:26 a-onl Start_Run_3775,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with 12uA and RICH
  • 250099 11/20/08 21:59 LeRose Target: Laser Absorption Spectrum
  • 250098 11/20/08 21:57 deng screen shot of lumi check for run 3774
  • 250097 11/20/08 21:45 a-onl Start_Run_3774,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with 12uA and RICH
  • 250096 11/20/08 21:44 a-onl End_of_Run_3773
  • 250095 11/20/08 21:40 deng screen shot of run 3773
  • 250094 11/20/08 21:32 ycwang RICH HV turned on, try the data after run # 3774
  • 250093 11/20/08 21:22 ycwang HV tripped for RICH, probably that is the reason why data is not reasonable.
  • 250092 11/20/08 21:18 Kalyan re: CRL for RICH
  • 250091 11/20/08 21:13 dhamilton Hall A Tools
  • 250090 11/20/08 21:13 ycwang One ADC has too much noise and there is something not right for new data taking from RICH ?
  • 250089 11/20/08 21:03 R. Michaels re: CRL for RICH
  • 250088 11/20/08 20:44 a-onl Start_Run_3773,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with 12uA and RICH
  • 250087 11/20/08 20:43 a-onl End_of_Run_3772
  • 250086 11/20/08 20:30 deng screen shot of run 3772
  • 250085 11/20/08 20:15 adaq Target Field and Oven Temps snapshot
  • 250084 11/20/08 20:15 adaq Target Cell Temps snapshot
  • 250083 11/20/08 20:15 adaq Radiation and Ozone snapshot
  • 250082 11/20/08 20:14 adaq Hall A Tools snapshot
  • 250081 11/20/08 20:14 adaq BigBox Magnet snapshot
  • 250080 11/20/08 20:14 adaq Beam Position and Current snapshot
  • 250079 11/20/08 20:14 adaq Beam Energy in A and C snapshot
  • 250078 11/20/08 20:00 adaq Scaler rates (BigBite) snapshot
  • 250077 11/20/08 19:34 a-onl Start_Run_3772,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with 12uA and RICH
  • 250076 11/20/08 19:33 a-onl End_of_Run_3771
  • 250075 11/20/08 19:16 deng screen shot of coda and deadtime of run 3771
  • 250074 11/20/08 18:52 Kalyan End run failed for 3770
  • 250073 11/20/08 18:46 Kalyan Deadtime issue with DAQ
  • 250072 11/20/08 18:32 a-onl Start_Run_3771,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with 12uA and RICH
  • 250071 11/20/08 18:07 adev BigBite WireChamber Measured Voltage change!
  • 250070 11/20/08 17:55 adev BigBite WireChamber Measured Voltage change!
  • 250069 11/20/08 17:55 a-onl End_of_Run_3770
  • 250068 11/20/08 17:51 a-onl Start_Run_3770,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=65,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production with 12uA and RICH
  • 250067 11/20/08 17:51 ycwang A1 rates with 12 uA beam
  • 250066 11/20/08 17:50 a-onl End_of_Run_3769
  • 250065 11/20/08 17:49 a-onl Start_Run_3769,Run_type=Test,target_type=No Target,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=test T1, T3 and T5 and T8 with RICH
  • 250064 11/20/08 17:48 kalyan Trigger rate at 12uA on 3He
  • 250063 11/20/08 17:47 dhamilton Begin Production Running
  • 250062 11/20/08 17:42 E.Chudakov Experiment: summary for Nov 19 (swing) - 20 (day)
  • 250061 11/20/08 17:33 LeRose 3He target in
  • 250060 11/20/08 17:32 a-onl End_of_Run_3768
  • 250059 11/20/08 17:28 a-onl Start_Run_3768,Run_type=Test,target_type=No Target,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=test T1, T3 and T5 and T8 with RICH
  • 250058 11/20/08 17:26 a-onl End_of_Run_3767
  • 250057 11/20/08 17:16 A. Camsonne moved CRL file to BOB's version
  • 250056 11/20/08 17:05 adev BigBite WireChamber Measured Voltage change!
  • 250055 11/20/08 16:48 a-onl Start_Run_3767,Run_type=Test,target_type=No Target,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=test T1, T3 and T5 and T8 with RICH
  • 250054 11/20/08 16:47 a-onl End_of_Run_3766
  • 250053 11/20/08 16:47 a-onl Start_Run_3766,Run_type=Test,target_type=No Target,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=test T1, T3 and T5 and T8 no RICH
  • 250052 11/20/08 16:44 a-onl End_of_Run_3765
  • 250051 11/20/08 16:38 a-onl Start_Run_3765,Run_type=Test,target_type=No Target,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=test T1, T3 and T5 and T8 no RICH
  • 250050 11/20/08 16:37 a-onl End_of_Run_3764
  • 250049 11/20/08 16:37 a-onl Start_Run_3764,Run_type=Test,target_type=No Target,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=test T1, T3 and T5 and T8 no RICH
  • 250048 11/20/08 16:36 a-onl End_of_Run_3763
  • 250047 11/20/08 16:35 a-onl Start_Run_3763,Run_type=Test,target_type=No Target,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=test T1, T3 and T5 and T8 no RICH
  • 250046 11/20/08 16:34 a-onl End_of_Run_3762
  • 250045 11/20/08 16:31 a-onl Start_Run_3762,Run_type=Test,target_type=No Target,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=test T1, T3 and T5 and T8
  • 250044 11/20/08 16:30 a-onl End_of_Run_3761
  • 250043 11/20/08 16:30 a-onl Start_Run_3761,Run_type=Test,target_type=No Target,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=test T3 and T5 and T8
  • 250042 11/20/08 16:29 a-onl End_of_Run_3760
  • 250041 11/20/08 16:29 a-onl Start_Run_3760,Run_type=Test,target_type=No Target,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=test T3 and T5
  • 250040 11/20/08 16:27 a-onl End_of_Run_3759
  • 250039 11/20/08 16:25 a-onl Start_Run_3759,Run_type=Test,target_type=No Target,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=test T3 and T5
  • 250038 11/20/08 16:24 a-onl End_of_Run_3758
  • 250037 11/20/08 16:22 a-onl Start_Run_3758,Run_type=Test,target_type=No Target,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic run with T1 T3 and T5
  • 250036 11/20/08 16:22 a-onl End_of_Run_3757
  • 250035 11/20/08 16:21 LeRose Target BeO
  • 250034 11/20/08 16:20 a-onl Start_Run_3757,Run_type=Test,target_type=No Target,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic run with T1 T3 and T5
  • 250033 11/20/08 16:19 a-onl End_of_Run_3756
  • 250032 11/20/08 16:17 a-onl Start_Run_3756,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic run with T1 T3 and T5
  • 250031 11/20/08 16:11 Jack (Techs) Left Q1 power supply
  • 250030 11/20/08 16:04 adaq Beam Time Accounting
  • 250029 11/20/08 16:04 Saha Shift Summary
  • 250028 11/20/08 16:03 LeRose Left Q2/3 Cycled
  • 250027 11/20/08 16:00 adaq Scaler rates (BigBite) snapshot
  • 250026 11/20/08 15:58 Simonetti target moved to empty target
  • 250025 11/20/08 15:31 fansler/Ewin lead flow
  • 250024 11/20/08 15:26 A. Camsonne Edet event structure
  • 250023 11/20/08 15:24 ycwang new pedestal for rich at run 3753
  • 250022 11/20/08 15:19 a-onl Start_Run_3755,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic run with T1 T3 and T5
  • 250021 11/20/08 15:18 a-onl End_of_Run_3754
  • 250020 11/20/08 15:16 a-onl Start_Run_3754,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic run with T1 T3 and T5
  • 250019 11/20/08 15:12 a-onl End_of_Run_3753
  • 250018 11/20/08 15:11 a-onl Start_Run_3753,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH pedestal run
  • 250017 11/20/08 15:07 a-onl Start_Run_3752,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic ray with T1, T3 and T5
  • 250016 11/20/08 15:07 a-onl End_of_Run_3752
  • 250015 11/20/08 15:04 a-onl End_of_Run_3751
  • 250014 11/20/08 15:00 a-onl Start_Run_3751,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic ray with T1, T3 and T5
  • 250013 11/20/08 14:59 a-onl End_of_Run_3750
  • 250012 11/20/08 14:57 a-onl Start_Run_3750,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=pedestal run for RICH
  • 250011 11/20/08 14:46 Moffit Cleanup of some bcm cables in counting house. Mainly Dx3 and Dx10.
  • 250010 11/20/08 14:41 a-onl End_of_Run_3749
  • 250009 11/20/08 14:40 a-onl Start_Run_3749,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test cosmics
  • 250008 11/20/08 14:39 a-onl End_of_Run_3748
  • 250007 11/20/08 14:37 a-onl Start_Run_3748,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test cosmics
  • 250006 11/20/08 14:36 a-onl End_of_Run_3747
  • 250005 11/20/08 14:34 a-onl Start_Run_3747,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test cosmics
  • 250004 11/20/08 14:33 a-onl End_of_Run_3746
  • 250003 11/20/08 14:29 a-onl Start_Run_3746,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test cosmics
  • 250002 11/20/08 14:28 a-onl End_of_Run_3745
  • 250001 11/20/08 14:26 a-onl Start_Run_3745,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test cosmics
  • 250000 11/20/08 14:25 a-onl End_of_Run_3744
  • 249999 11/20/08 14:20 a-onl Start_Run_3744,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test cosmics
  • 249998 11/20/08 14:17 a-onl Start_Run_3743,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test cosmics
  • 249997 11/20/08 14:16 a-onl End_of_Run_3742
  • 249996 11/20/08 14:15 a-onl Start_Run_3742,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test cosmics
  • 249995 11/20/08 14:14 a-onl End_of_Run_3741
  • 249994 11/20/08 14:08 a-onl Start_Run_3741,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test cosmics
  • 249993 11/20/08 14:08 a-onl End_of_Run_3740
  • 249992 11/20/08 14:08 a-onl Start_Run_3740,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test cosmics
  • 249991 11/20/08 14:07 a-onl End_of_Run_3739
  • 249990 11/20/08 14:06 a-onl Start_Run_3739,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test cosmics
  • 249989 11/20/08 14:05 a-onl End_of_Run_3738
  • 249988 11/20/08 14:04 a-onl Start_Run_3738,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test cosmics
  • 249987 11/20/08 14:04 a-onl End_of_Run_3737
  • 249986 11/20/08 13:57 a-onl Start_Run_3737,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test cosmics
  • 249985 11/20/08 13:48 a-onl End_of_Run_3736
  • 249984 11/20/08 13:48 brads Pb-glass test PMTs signal & HV channels
  • 249983 11/20/08 13:46 a-onl Start_Run_3736,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test cosmics
  • 249982 11/20/08 13:44 a-onl End_of_Run_3735
  • 249981 11/20/08 13:44 a-onl Start_Run_3735,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test cosmics
  • 249980 11/20/08 13:43 a-onl End_of_Run_3734
  • 249979 11/20/08 13:31 adev Start_Run_20177,Run_type=Test,target_type=Loop 1 LH2 6 cm,beam_e0=2.322,beam_raster=off,beam_halfwaveplate=out,rightarm_p=0.564,rightarm_theta=105,rightarm_colli=out,comment_text=nothing.
  • 249978 11/20/08 13:31 simonetti Glitch during preparation of field
  • 249977 11/20/08 13:30 Yi Qiang NMR Frequency/Field Sweep Cross Calibration for Vertical Pumping
  • 249976 11/20/08 12:59 Bob/Ramesh R-HRS activity
  • 249975 11/20/08 12:57 adev Start_Run_20175,Run_type=Test,target_type=Loop 1 LH2 6 cm,beam_e0=2.322,beam_raster=off,beam_halfwaveplate=out,rightarm_p=0.564,rightarm_theta=105,rightarm_colli=out,comment_text=nothing.
  • 249974 11/20/08 12:56 adev End_of_Run_20175
  • 249973 11/20/08 12:50 adev End_of_Run_20174
  • 249972 11/20/08 12:42 adev Start_Run_20174,Run_type=Test,target_type=Loop 1 LH2 6 cm,beam_e0=2.322,beam_raster=off,beam_halfwaveplate=out,rightarm_p=0.564,rightarm_theta=105,rightarm_colli=out,comment_text=nothing.
  • 249971 11/20/08 12:39 E.Chudakov Access status
  • 249970 11/20/08 12:35 adev End_of_Run_20173
  • 249969 11/20/08 12:18 Michel Bross Electron detector dark current
  • 249968 11/20/08 12:15 adev Start_Run_20173,Run_type=Test,target_type=Loop 1 LH2 6 cm,beam_e0=2.322,beam_raster=off,beam_halfwaveplate=out,rightarm_p=0.564,rightarm_theta=105,rightarm_colli=out,comment_text=nothing.
  • 249967 11/20/08 12:14 adaq Target Field and Oven Temps snapshot
  • 249966 11/20/08 12:14 adaq Target Cell Temps snapshot
  • 249965 11/20/08 12:14 adaq Radiation and Ozone snapshot
  • 249964 11/20/08 12:14 adaq Hall A Tools snapshot
  • 249963 11/20/08 12:14 adaq BigBox Magnet snapshot
  • 249962 11/20/08 12:14 adaq Beam Position and Current snapshot
  • 249961 11/20/08 12:14 adaq Beam Energy in A and C snapshot
  • 249960 11/20/08 12:00 adaq Scaler rates (BigBite) snapshot
  • 249959 11/20/08 11:58 Yi Qiang/YI EPR Calibration for vertical+ field, anti-parallel pumping
  • 249958 11/20/08 11:56 a-onl Start_Run_3734,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test cosmics
  • 249957 11/20/08 11:10 brads BB Cerenkov witness PMT: shielding added in front
  • 249956 11/20/08 10:56 V. Sulkosky Re: U10 Double Pulsing Checked
  • 249955 11/20/08 10:45 a-onl End_of_Run_3733
  • 249954 11/20/08 10:41 Kalyan High dead time went away after I restarted coda
  • 249953 11/20/08 10:37 a-onl Start_Run_3733,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test cosmics
  • 249952 11/20/08 10:33 a-onl End_of_Run_3732
  • 249951 11/20/08 10:31 a-onl Start_Run_3732,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test cosmics
  • 249950 11/20/08 10:27 A. Camsonne High dead time with T1 even without RICH
  • 249949 11/20/08 10:26 a-onl Start_Run_3731,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test cosmics
  • 249948 11/20/08 10:25 a-onl End_of_Run_3730
  • 249947 11/20/08 10:24 a-onl Start_Run_3730,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test cosmics
  • 249946 11/20/08 10:24 a-onl End_of_Run_3729
  • 249945 11/20/08 10:22 a-onl Start_Run_3729,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test cosmics
  • 249944 11/20/08 10:21 a-onl End_of_Run_3728
  • 249943 11/20/08 10:15 A. Camsonne Serial adapter
  • 249942 11/20/08 10:10 a-onl Start_Run_3728,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=
  • 249941 11/20/08 10:04 a-onl End_of_Run_3727
  • 249940 11/20/08 10:03 a-onl Start_Run_3727,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=
  • 249939 11/20/08 10:00 a-onl End_of_Run_3726
  • 249938 11/20/08 10:00 E.Chudakov Schedule for the access
  • 249937 11/20/08 09:57 a-onl Start_Run_3726,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=
  • 249936 11/20/08 09:57 a-onl End_of_Run_3725
  • 249935 11/20/08 09:52 a-onl Start_Run_3725,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=
  • 249934 11/20/08 09:49 a-onl End_of_Run_3724
  • 249933 11/20/08 09:41 a-onl Start_Run_3724,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=
  • 249932 11/20/08 09:41 a-onl End_of_Run_3723
  • 249931 11/20/08 09:36 a-onl Start_Run_3723,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=
  • 249930 11/20/08 09:35 a-onl End_of_Run_3722
  • 249929 11/20/08 09:32 a-onl Start_Run_3722,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Beam for long time, change prescal and keep cosmics
  • 249928 11/20/08 09:31 a-onl End_of_Run_3721
  • 249927 11/20/08 09:23 E.Chudakov No unwanted dispersive mode on Nov 17 - another followup
  • 249926 11/20/08 09:09 a-onl Start_Run_3721,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Beam for long time, change prescal and keep cosmics
  • 249925 11/20/08 09:06 a-onl End_of_Run_3720
  • 249924 11/20/08 08:53 Jin Huang Polarization History for Vertical Pumping Production Run
  • 249923 11/20/08 08:39 E.Chudakov No unwanted dispersive mode on Nov 17
  • 249922 11/20/08 08:18 Jin Huang AFP Loss Test @ Vert Pumping, 70% Pol & No Correction Coil
  • 249921 11/20/08 08:15 Yi Zhang BB Cherenkov testing summary
  • 249920 11/20/08 08:13 Yi Zhang BB Cherenkov testing summary
  • 249919 11/20/08 08:03 adaq Beam Time Accounting
  • 249918 11/20/08 08:00 adaq Scaler rates (BigBite) snapshot
  • 249917 11/20/08 07:35 gilman owl shift summary
  • 249916 11/20/08 07:31 gilman ask for controlled access
  • 249915 11/20/08 07:14 a-onl Start_Run_3720,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Beam for long time, change prescal and keep cosmics
  • 249914 11/20/08 07:12 a-onl End_of_Run_3719
  • 249913 11/20/08 07:09 gilman arc shunt problem
  • 249912 11/20/08 06:59 Jin Huang Target Data is daily backed up to central Tape
  • 249911 11/20/08 06:59 Jin Huang Coincident Cut applied on physics replay
  • 249910 11/20/08 06:59 Jin Huang Protection on Golden Root File
  • 249909 11/20/08 06:59 Jin Huang Online Plots divided to Shift Worker plots and Expert Plots
  • 249908 11/20/08 06:54 a-onl Start_Run_3719,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=73,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Waiting for Beam, It's Cosmic Time
  • 249907 11/20/08 06:52 gilman magnet off
  • 249906 11/20/08 06:33 a-onl End_of_Run_3718
  • 249905 11/20/08 06:27 a-onl Start_Run_3718,Run_type=Check out,target_type=Multi Carbon Optics/BeO,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB Cherenkov test. 0.1uA
  • 249904 11/20/08 06:23 a-onl End_of_Run_3717
  • 249903 11/20/08 06:23 a-onl Start_Run_3717,Run_type=Check out,target_type=Multi Carbon Optics/BeO,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB Cherenkov test. 0.1uA
  • 249902 11/20/08 06:14 Moffit HAPPEX DAQ: Please keep runs less than 4 hours long.
  • 249901 11/20/08 06:07 a-onl End_of_Run_3716
  • 249900 11/20/08 05:55 a-onl Start_Run_3716,Run_type=Check out,target_type=Multi Carbon Optics/BeO,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB Cherenkov test. 0.5uA
  • 249899 11/20/08 05:43 a-onl Start_Run_3715,Run_type=Check out,target_type=Multi Carbon Optics/BeO,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB Cherenkov test. 0.5uA
  • 249898 11/20/08 05:33 a-onl End_of_Run_3714
  • 249897 11/20/08 05:30 gilman spot++ for shift checklist
  • 249896 11/20/08 05:26 gilman Hall A Gen Tools for shift checklist
  • 249895 11/20/08 05:21 a-onl Start_Run_3714,Run_type=Check out,target_type=Multi Carbon Optics/BeO,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB Cherenkov test. 1.0uA
  • 249894 11/20/08 05:21 a-onl End_of_Run_3713
  • 249893 11/20/08 05:06 a-onl Start_Run_3713,Run_type=Check out,target_type=Multi Carbon Optics/BeO,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB Cherenkov test. 1.0uA
  • 249892 11/20/08 05:04 a-onl End_of_Run_3712
  • 249891 11/20/08 04:59 Jin Scalar & BCM for run 3712
  • 249890 11/20/08 04:57 a-onl Start_Run_3712,Run_type=Check out,target_type=Multi Carbon Optics/BeO,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB Cherenkov test. 1.0uA
  • 249889 11/20/08 04:41 a-onl End_of_Run_3711
  • 249888 11/20/08 04:19 Jin Scalar & BCM for run 3711
  • 249887 11/20/08 04:17 a-onl Start_Run_3711,Run_type=Check out,target_type=No Target,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB Cherenkov test. 0.5uA
  • 249886 11/20/08 04:14 adaq Target Field and Oven Temps snapshot
  • 249885 11/20/08 04:14 adaq Target Cell Temps snapshot
  • 249884 11/20/08 04:14 adaq Radiation and Ozone snapshot
  • 249883 11/20/08 04:14 adaq Hall A Tools snapshot
  • 249882 11/20/08 04:13 adaq BigBox Magnet snapshot
  • 249881 11/20/08 04:13 adaq Beam Position and Current snapshot
  • 249880 11/20/08 04:13 adaq Beam Energy in A and C snapshot
  • 249879 11/20/08 04:11 a-onl End_of_Run_3710
  • 249878 11/20/08 04:10 adev BigBite WireChamber Measured Voltage change!
  • 249877 11/20/08 04:00 adaq Scaler rates (BigBite) snapshot
  • 249876 11/20/08 03:59 adev BigBite WireChamber Measured Voltage change!
  • 249875 11/20/08 03:48 gilman Moller done
  • 249874 11/20/08 03:46 E.Chudakov Moller results
  • 249873 11/20/08 03:26 a-onl Start_Run_3710,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Beam off, hall in access due to Q1 problems - test RICH -T1 T3 T5 during moller
  • 249872 11/20/08 03:26 a-onl End_of_Run_3709
  • 249871 11/20/08 03:15 a-onl Start_Run_3709,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Beam off, hall in access due to Q1 problems - test RICH -T1 T3 T5 during moller
  • 249870 11/20/08 03:14 a-onl End_of_Run_3708
  • 249869 11/20/08 03:01 a-onl Start_Run_3708,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Beam off, hall in access due to Q1 problems - test RICH -T1 T3 T5 during moller
  • 249868 11/20/08 03:00 a-onl End_of_Run_3707
  • 249867 11/20/08 02:44 a-onl Start_Run_3707,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Beam off, hall in access due to Q1 problems - test RICH -T1 T3 T5 during moller
  • 249866 11/20/08 02:43 a-onl Start_Run_3706,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Beam off, hall in access due to Q1 problems - test RICH -T1 T3 T5 during moller
  • 249865 11/20/08 02:43 a-onl End_of_Run_3706
  • 249864 11/20/08 02:24 a-onl End_of_Run_3705
  • 249863 11/20/08 02:05 a-onl Start_Run_3705,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Beam off, hall in access due to Q1 problems - test RICH -T1 T3 T5 during moller
  • 249862 11/20/08 02:04 a-onl End_of_Run_3704
  • 249861 11/20/08 01:39 a-onl Start_Run_3704,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Beam off, hall in access due to Q1 problems - test RICH -T1 T3 T5 during moller
  • 249860 11/20/08 01:38 a-onl End_of_Run_3703
  • 249859 11/20/08 01:27 A. Camsonne Trigger rates during moeller
  • 249858 11/20/08 01:26 A. Camsonne Dead time with RICH
  • 249857 11/20/08 01:21 A. Camsonne RICH readout change
  • 249856 11/20/08 01:18 a-onl Start_Run_3703,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Beam off, hall in access due to Q1 problems - test RICH -T1 T3 T5 during moller
  • 249855 11/20/08 01:18 a-onl End_of_Run_3702
  • 249854 11/20/08 01:16 a-onl Start_Run_3702,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Beam off, hall in access due to Q1 problems - test RICH -T1 T3 T5 during moller
  • 249853 11/20/08 01:15 a-onl End_of_Run_3701
  • 249852 11/20/08 01:10 a-onl Start_Run_3701,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Beam off, hall in access due to Q1 problems - test RICH -T1 T3 T5 during moller
  • 249851 11/20/08 01:09 a-onl End_of_Run_3700
  • 249850 11/20/08 01:08 a-onl Start_Run_3700,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Beam off, hall in access due to Q1 problems - test RICH -T1 T3 T5 during moller
  • 249849 11/20/08 01:07 a-onl End_of_Run_3699
  • 249848 11/20/08 01:07 a-onl Start_Run_3699,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Beam off, hall in access due to Q1 problems - test RICH -T1 T3 T5 during moller
  • 249847 11/20/08 01:06 a-onl End_of_Run_3698
  • 249846 11/20/08 01:06 a-onl Start_Run_3698,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Beam off, hall in access due to Q1 problems - test RICH -T1 T3 T5 during moller
  • 249845 11/20/08 01:05 a-onl End_of_Run_3697
  • 249844 11/20/08 00:58 E.Chudakov Moller start
  • 249843 11/20/08 00:56 a-onl Start_Run_3697,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Beam off, hall in access due to Q1 problems - test RICH -T3 only
  • 249842 11/20/08 00:46 a-onl End_of_Run_3696
  • 249841 11/20/08 00:42 A. Camsonne Compton electron detector current
  • 249840 11/20/08 00:37 A. Camsonne Compton power meter channel in odd state
  • 249839 11/20/08 00:28 a-onl Start_Run_3696,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Beam off, hall in access due to Q1 problems - test RICH -T1 only
  • 249838 11/20/08 00:27 adev BigBite WireChamber Measured Voltage change!
  • 249837 11/20/08 00:27 a-onl End_of_Run_3695
  • 249836 11/20/08 00:25 a-onl Start_Run_3695,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Beam off, hall in access due to Q1 problems - test RICH -T3
  • 249835 11/20/08 00:24 a-onl End_of_Run_3694
  • 249834 11/20/08 00:20 A. Camsonne Compton power meter channel in odd state
  • 249833 11/20/08 00:02 dhamilton Swing Shift Summary
  • 249832 11/20/08 00:02 adaq Beam Time Accounting
  • 249831 11/20/08 00:00 adaq Scaler rates (BigBite) snapshot
  • Wednesday

  • 249830 11/19/08 23:51 a-onl Start_Run_3694,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Beam off, hall in access due to Q1 problems - test RICH - T1 and T3 and T8
  • 249829 11/19/08 23:51 a-onl End_of_Run_3693
  • 249828 11/19/08 23:49 a-onl Start_Run_3693,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Beam off, hall in access due to Q1 problems - test RICH - T1 and T3
  • 249827 11/19/08 23:48 a-onl End_of_Run_3692
  • 249826 11/19/08 23:48 a-onl Start_Run_3692,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Beam off, hall in access due to Q1 problems - test RICH - T1 only
  • 249825 11/19/08 23:47 a-onl End_of_Run_3691
  • 249824 11/19/08 23:47 a-onl Start_Run_3691,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Beam off, hall in access due to Q1 problems - test RICH - T1 only
  • 249823 11/19/08 23:46 a-onl End_of_Run_3690
  • 249822 11/19/08 23:46 a-onl Start_Run_3690,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Beam off, hall in access due to Q1 problems - test RICH -T3 and T1 - limited channel
  • 249821 11/19/08 23:43 a-onl End_of_Run_3689
  • 249820 11/19/08 23:42 a-onl Start_Run_3689,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Beam off, hall in access due to Q1 problems - test RICH -T3 and T1 - limited channel
  • 249819 11/19/08 23:41 a-onl End_of_Run_3688
  • 249818 11/19/08 23:38 brads BB Cerenkov tests for tonight
  • 249817 11/19/08 23:07 a-onl Start_Run_3688,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Beam off, hall in access due to Q1 problems - test RICH -T3 and T1 - limited channel
  • 249816 11/19/08 23:02 a-onl End_of_Run_3687
  • 249815 11/19/08 23:02 a-onl Start_Run_3687,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Beam off, hall in access due to Q1 problems - test RICH -T3 and T1 - limited channel
  • 249814 11/19/08 22:56 E.Chudakov No Left Q1 for tonight
  • 249813 11/19/08 22:51 a-onl End_of_Run_3686
  • 249812 11/19/08 22:51 a-onl Start_Run_3686,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Beam off, hall in access due to Q1 problems - test RICH -T3 and T1 - limited channel
  • 249811 11/19/08 22:24 adaq target: Spin Flip Report: Shift target report
  • 249810 11/19/08 22:16 a-onl End_of_Run_3685
  • 249809 11/19/08 22:16 a-onl Start_Run_3685,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Beam off, hall in access due to Q1 problems - test RICH -T3 and T1 - limited channel
  • 249808 11/19/08 22:14 a-onl End_of_Run_3682
  • 249807 11/19/08 22:13 a-onl End_of_Run_3682
  • 249806 11/19/08 22:10 a-onl End_of_Run_3682
  • 249805 11/19/08 22:10 a-onl Start_Run_3682,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Beam off, hall in access due to Q1 problems - test RICH -T3 and T1
  • 249804 11/19/08 22:08 a-onl End_of_Run_3681
  • 249803 11/19/08 22:06 a-onl Start_Run_3681,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Beam off, hall in access due to Q1 problems - test RICH - other triggers
  • 249802 11/19/08 21:52 a-onl End_of_Run_3680
  • 249801 11/19/08 21:51 a-onl Start_Run_3680,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Beam off, hall in access due to Q1 problems - test RICH - other triggers
  • 249800 11/19/08 21:47 a-onl End_of_Run_3679
  • 249799 11/19/08 21:45 a-onl Start_Run_3679,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Beam off, hall in access due to Q1 problems - test RICH
  • 249798 11/19/08 21:26 E.Chudakov Q1 problem
  • 249797 11/19/08 20:47 a-onl Start_Run_3678,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Beam off, hall in access due to Q1 problems
  • 249796 11/19/08 20:47 a-onl End_of_Run_3677
  • 249795 11/19/08 20:35 a-onl Start_Run_3677,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Beam off, hall in access due to Q1 problems
  • 249794 11/19/08 20:33 a-onl End_of_Run_3676
  • 249793 11/19/08 20:32 E.Chudakov Experiment: summary for 11/18(day)-11/19(day)
  • 249792 11/19/08 20:30 a-onl Start_Run_3676,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Beam off, hall in access due to Q1 problems deadtime with TwoArmRICH
  • 249791 11/19/08 20:28 a-onl End_of_Run_3675
  • 249790 11/19/08 20:15 dhamilton Hall moved to controlled access
  • 249789 11/19/08 20:14 a-onl Start_Run_3675,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Beam off, hall in access due to Q1 problems
  • 249788 11/19/08 20:13 adaq Target Field and Oven Temps snapshot
  • 249787 11/19/08 20:13 adaq Target Cell Temps snapshot
  • 249786 11/19/08 20:13 adaq Radiation and Ozone snapshot
  • 249785 11/19/08 20:13 adaq Hall A Tools snapshot
  • 249784 11/19/08 20:13 adaq BigBox Magnet snapshot
  • 249783 11/19/08 20:13 adaq Beam Position and Current snapshot
  • 249782 11/19/08 20:13 adaq Beam Energy in A and C snapshot
  • 249781 11/19/08 20:00 adaq Scaler rates (BigBite) snapshot
  • 249780 11/19/08 19:46 dhamilton Q1 again - contacted Jack
  • 249779 11/19/08 19:36 dhamilton Hall moved to Beam Permit
  • 249778 11/19/08 18:55 dhamilton Hall moved to controlled access
  • 249777 11/19/08 18:48 E.Chudakov Problems with the accelerator and with Q1
  • 249776 11/19/08 18:44 dhamilton Status update
  • 249775 11/19/08 18:38 a-onl End_of_Run_3674
  • 249774 11/19/08 18:34 a-onl Start_Run_3674,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Would be production at 12 uA, but Q1 keeps quenching
  • 249773 11/19/08 18:33 a-onl End_of_Run_3673
  • 249772 11/19/08 18:32 dhamilton Q1 Status Update
  • 249771 11/19/08 18:13 a-onl Start_Run_3673,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production 12 uA
  • 249770 11/19/08 18:00 dhamilton Another Q1 Quench
  • 249769 11/19/08 17:56 xqian Coincidence events in BigBite
  • 249768 11/19/08 17:47 a-onl End_of_Run_3672
  • 249767 11/19/08 17:38 Kalyan Scintillator rate on BigBite
  • 249766 11/19/08 17:26 a-onl Start_Run_3672,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production 12 uA
  • 249765 11/19/08 17:24 dhamilton Q1 back up to 2.35 GeV/c
  • 249764 11/19/08 17:02 a-onl End_of_Run_3671
  • 249763 11/19/08 17:02 dhamilton Q1 Left Quenched again
  • 249762 11/19/08 16:48 a-onl Start_Run_3671,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production 12 uA
  • 249761 11/19/08 16:45 a-onl End_of_Run_3670
  • 249760 11/19/08 16:20 dhamilton Compton screenshot
  • 249759 11/19/08 16:19 dhamilton BPM/BCM Screenshots
  • 249758 11/19/08 16:17 dhamilton Q1 is back up
  • 249757 11/19/08 16:16 a-onl Start_Run_3670,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production 12 uA
  • 249756 11/19/08 16:07 adaq Beam Time Accounting
  • 249755 11/19/08 16:00 adaq Scaler rates (BigBite) snapshot
  • 249754 11/19/08 16:00 adaq target: Spin Flip Report: shift day 19-11
  • 249753 11/19/08 15:59 gilman day shift summary - 19 Nov 2008
  • 249752 11/19/08 15:52 brads BB detector stack's view of the target
  • 249751 11/19/08 15:49 gilman Q1 trip / quench
  • 249750 11/19/08 15:46 a-onl End_of_Run_3669
  • 249749 11/19/08 15:28 a-onl Start_Run_3669,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production 12 uA
  • 249748 11/19/08 15:25 a-onl End_of_Run_3668
  • 249747 11/19/08 15:20 Kalyan Re: U10 Double-pulsing in Left HRS
  • 249746 11/19/08 15:06 V. Sulkosky U10 Double-pulsing in Left HRS
  • 249745 11/19/08 15:01 Katich BCM U10 is absent from HAPPEX DAQ
  • 249744 11/19/08 14:39 a-onl Start_Run_3668,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production 12 uA
  • 249743 11/19/08 14:38 a-onl End_of_Run_3667
  • 249742 11/19/08 14:07 gilman spot for shift checklist
  • 249741 11/19/08 14:05 gilman Hall A Tools for shift checklist
  • 249740 11/19/08 13:45 a-onl Start_Run_3667,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production 12 uA
  • 249739 11/19/08 13:44 a-onl End_of_Run_3666
  • 249738 11/19/08 12:54 a-onl Start_Run_3666,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production 12 uA
  • 249737 11/19/08 12:54 a-onl End_of_Run_3665
  • 249736 11/19/08 12:15 Katich Lumi HV Change
  • 249735 11/19/08 12:13 adaq Target Field and Oven Temps snapshot
  • 249734 11/19/08 12:13 adaq Target Cell Temps snapshot
  • 249733 11/19/08 12:13 adaq Radiation and Ozone snapshot
  • 249732 11/19/08 12:13 adaq Hall A Tools snapshot
  • 249731 11/19/08 12:13 adaq BigBox Magnet snapshot
  • 249730 11/19/08 12:13 adaq Beam Position and Current snapshot
  • 249729 11/19/08 12:12 adaq Beam Energy in A and C snapshot
  • 249728 11/19/08 12:00 adaq Scaler rates (BigBite) snapshot
  • 249727 11/19/08 11:55 gilman beam returns ~11:43
  • 249726 11/19/08 11:44 a-onl Start_Run_3665,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production 12 uA
  • 249725 11/19/08 11:44 Moffit More Run Database (mysql) changes.
  • 249724 11/19/08 11:29 xqian Beam Charge Counting now
  • 249723 11/19/08 11:12 gilman 11 AM update
  • 249722 11/19/08 10:41 a-onl End_of_Run_3664
  • 249721 11/19/08 10:33 Ole Hansen adaql10 diagnostic
  • 249720 11/19/08 10:20 gilman beam coming back
  • 249719 11/19/08 10:15 gilman logbook end of run error
  • 249718 11/19/08 10:03 V. Sulkosky S2m Left # 4 ADC Appears Fine Now
  • 249717 11/19/08 10:03 gilman ioc issue
  • 249716 11/19/08 10:02 fansler LHRS vacuum
  • 249715 11/19/08 09:51 a-onl Start_Run_3664,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production 12 uA
  • 249714 11/19/08 09:47 V. Sulkosky Golden Root files Set to 3661
  • 249713 11/19/08 09:46 adaq target: Spin Flip Report: after a short control access
  • 249712 11/19/08 09:42 gilman controlled access done
  • 249711 11/19/08 09:30 cisbani RICH rings cumulated
  • 249710 11/19/08 09:20 a-onl End_of_Run_3663
  • 249709 11/19/08 09:10 gilman running nicely - sort off - controlled access
  • 249708 11/19/08 08:58 X. Jiang MCC is planning a RF recovery tomorow, 8:30am, four hours
  • 249707 11/19/08 08:53 a-onl Start_Run_3663,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production 12 uA
  • 249706 11/19/08 08:52 a-onl End_of_Run_3663
  • 249705 11/19/08 08:15 Moffit MYSQL Run Database... added new Table.
  • 249704 11/19/08 08:13 xqian Wire Chamber hitting efficiency per wire
  • 249703 11/19/08 08:11 Yi Zhang /adaql10/work1/e06010/ROOTfiles/ problem sloved
  • 249702 11/19/08 08:04 Simonetti Owl shift summary
  • 249701 11/19/08 08:03 Simonetti LHSR vacuum pressure update
  • 249700 11/19/08 08:02 adaq Beam Time Accounting
  • 249699 11/19/08 08:00 adaq Scaler rates (BigBite) snapshot
  • 249698 11/19/08 07:55 a-onl Start_Run_3662,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production 12 uA
  • 249697 11/19/08 07:53 Vilardi Pumping laser Absorption spectrum
  • 249696 11/19/08 07:52 a-onl End_of_Run_3661
  • 249695 11/19/08 07:50 adaq target: Spin Flip Report: Halog Report Owl shift 11/19/2008
  • 249694 11/19/08 07:49 Yi Zhang /adaql10/work1/e06010/ROOTfiles/ access problem
  • 249693 11/19/08 07:42 Moffit Lumi 7 and 8 Saturated
  • 249692 11/19/08 07:32 xqian live time without RICH is 91%
  • 249691 11/19/08 07:24 xqian BigBite Wire Chamber hitting efficiency
  • 249690 11/19/08 07:02 a-onl Start_Run_3661,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production 12 uA
  • 249689 11/19/08 07:01 a-onl End_of_Run_3660
  • 249688 11/19/08 06:22 Simonetti LHSR vacuum pressure update
  • 249687 11/19/08 06:16 a-onl Start_Run_3660,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production 12 uA
  • 249686 11/19/08 06:15 a-onl End_of_Run_3659
  • 249685 11/19/08 05:34 Simonetti LHRS vacuum pressure update
  • 249684 11/19/08 05:34 a-onl Start_Run_3659,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production 12 uA
  • 249683 11/19/08 05:32 a-onl End_of_Run_3658
  • 249682 11/19/08 05:01 Simonetti LHRS vacuum pressure read out
  • 249681 11/19/08 04:45 a-onl Start_Run_3658,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production 12 uA
  • 249680 11/19/08 04:44 a-onl End_of_Run_3657
  • 249679 11/19/08 04:12 adaq Target Field and Oven Temps snapshot
  • 249678 11/19/08 04:12 adaq Target Cell Temps snapshot
  • 249677 11/19/08 04:12 adaq Radiation and Ozone snapshot
  • 249676 11/19/08 04:12 adaq Hall A Tools snapshot
  • 249675 11/19/08 04:12 adaq BigBox Magnet snapshot
  • 249674 11/19/08 04:12 adaq Beam Position and Current snapshot
  • 249673 11/19/08 04:12 adaq Beam Energy in A and C snapshot
  • 249672 11/19/08 04:06 Chiranjib LHRS vacuum pressure read out
  • 249671 11/19/08 04:00 adaq Scaler rates (BigBite) snapshot
  • 249670 11/19/08 03:58 a-onl Start_Run_3657,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production 12 uA
  • 249669 11/19/08 03:57 a-onl End_of_Run_3656
  • 249668 11/19/08 03:17 Chiranjib We are running TWOARM FB
  • 249667 11/19/08 03:06 a-onl Start_Run_3656,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production 12 uA
  • 249666 11/19/08 03:04 a-onl End_of_Run_3655
  • 249665 11/19/08 02:15 a-onl Start_Run_3655,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production 12 uA
  • 249664 11/19/08 02:13 a-onl End_of_Run_3654
  • 249663 11/19/08 01:43 Chiranjib LHRS vacuum pressure (or something like that) read out
  • 249662 11/19/08 01:36 brads BB Cerenkov witness PMT rates with Pb surround shield
  • 249661 11/19/08 01:26 Chiranjib RUN 3653 issue
  • 249660 11/19/08 01:19 a-onl Start_Run_3654,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production 12 uA
  • 249659 11/19/08 01:15 a-onl End_of_Run_3653
  • 249658 11/19/08 00:59 Chiranjib Trigger rates/ Prescales @12 uA
  • 249657 11/19/08 00:48 Simonetti Helium alarm
  • 249656 11/19/08 00:47 a-onl Start_Run_3653,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=
  • 249655 11/19/08 00:46 a-onl End_of_Run_3652
  • 249654 11/19/08 00:46 a-onl Start_Run_3652,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=
  • 249653 11/19/08 00:43 a-onl End_of_Run_3651
  • 249652 11/19/08 00:38 a-onl Start_Run_3651,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=
  • 249651 11/19/08 00:37 Jin Trigger @ 12uA, 3x4 raster
  • 249650 11/19/08 00:35 Jin Spot++ @ 6x6 raster, go production beam
  • 249649 11/19/08 00:32 a-onl End_of_Run_3650
  • 249648 11/19/08 00:31 a-onl Start_Run_3650,Run_type=Check out,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=60,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=spot++
  • 249647 11/19/08 00:31 Simonetti raster 5x5
  • 249646 11/19/08 00:31 Jin Trigger @ 6x6 2uA
  • 249645 11/19/08 00:30 a-onl End_of_Run_3649
  • 249644 11/19/08 00:28 a-onl Start_Run_3649,Run_type=Check out,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=60,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=spot++
  • 249643 11/19/08 00:28 Jin Trigger @ 5x5 2uA
  • 249642 11/19/08 00:11 adaq Beam Time Accounting
  • 249641 11/19/08 00:00 adaq Scaler rates (BigBite) snapshot
  • Tuesday

  • 249640 11/18/08 23:57 A.Rakhman 11/18/2008 Swing Shift Summary
  • 249639 11/18/08 23:38 Chiranjib Laser parameters
  • 249638 11/18/08 23:26 Chiranjib Target screens II
  • 249637 11/18/08 23:23 Chiranjib TARGET SCREENS
  • 249636 11/18/08 23:19 A.Rakhman beam status
  • 249635 11/18/08 22:57 Jin Trigger @ 3x4, 2uA, he3
  • 249634 11/18/08 22:43 Camsonne / C RICH Activity during controlled access (Nov 18 2008)
  • 249633 11/18/08 22:21 mcanan Target moved to He-3/ started spin flip
  • 249632 11/18/08 22:15 A.Rakhman helium level alarm
  • 249631 11/18/08 22:07 mcanan target moved to optics
  • 249630 11/18/08 22:01 E.Chudakov Smoke problem in the hall
  • 249629 11/18/08 22:01 A.Rakhman hall a tools snap shots
  • 249628 11/18/08 21:47 brads BB Cerenkov ADC<->TDC mapping problem fixed
  • 249627 11/18/08 21:44 brads BB Cerenkov discriminator calibration
  • 249626 11/18/08 21:41 A.Rakhman hall a status update
  • 249625 11/18/08 21:39 brads Modified shielding around BB Cerenkov witness PMT (iteration 2)
  • 249624 11/18/08 21:37 brads Pump on left Q2 has been turned off, shift crew should watch gauge on monitor
  • 249623 11/18/08 21:32 brads The smoke-puffing pump
  • 249622 11/18/08 20:20 a-onl Start_Run_3648,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Vertical,targetpol_pol=72,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmics
  • 249621 11/18/08 20:12 adaq Target Field and Oven Temps snapshot
  • 249620 11/18/08 20:12 adaq Target Cell Temps snapshot
  • 249619 11/18/08 20:12 adaq Radiation and Ozone snapshot
  • 249618 11/18/08 20:12 adaq Hall A Tools snapshot
  • 249617 11/18/08 20:12 adaq BigBox Magnet snapshot
  • 249616 11/18/08 20:12 adaq Beam Position and Current snapshot
  • 249615 11/18/08 20:12 adaq Beam Energy in A and C snapshot
  • 249614 11/18/08 20:09 V. Sulkosky Changed S2m Left PMT # 4 ADC Delay Cable
  • 249613 11/18/08 20:00 adaq Scaler rates (BigBite) snapshot
  • 249612 11/18/08 19:41 A.Rakhman smoke detection alarm
  • 249611 11/18/08 19:36 a-onl End_of_Run_3647
  • 249610 11/18/08 19:35 a-onl Start_Run_3647,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=72,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=test for S2m Left # 4, RICH PEDESTAL true
  • 249609 11/18/08 19:33 a-onl End_of_Run_3646
  • 249608 11/18/08 19:33 a-onl Start_Run_3646,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=72,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=test for S2m Left # 4, RICH PEDESTAL
  • 249607 11/18/08 19:32 a-onl End_of_Run_3645
  • 249606 11/18/08 19:31 Jin Huang Target: Laser and RTD status at vertical pumping
  • 249605 11/18/08 19:28 A.Rakhman left arm dipole
  • 249604 11/18/08 19:23 Jack (Techs) Target IOC reboot and target motion - Make sure the beam is OFF
  • 249603 11/18/08 19:19 a-onl Start_Run_3645,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=72,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=test for S2m Left # 4, RIVH PED
  • 249602 11/18/08 19:18 a-onl End_of_Run_3644
  • 249601 11/18/08 19:17 a-onl Start_Run_3644,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=72,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=test for S2m Left # 4, RIVH PED
  • 249600 11/18/08 19:16 a-onl End_of_Run_3643
  • 249599 11/18/08 19:15 a-onl Start_Run_3643,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=72,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=test for S2m Left # 4
  • 249598 11/18/08 19:12 Yi Qiang Target Activities for field rotation from transverse to vertical
  • 249597 11/18/08 19:10 adaq target: Spin Flip Report: Vertical Pumping Auto Spin Flip Ready!
  • 249596 11/18/08 19:09 A.Rakhman Compton
  • 249595 11/18/08 19:09 a-onl Start_Run_3642,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=72,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=test for S2m Left # 4
  • 249594 11/18/08 19:08 a-onl End_of_Run_3642
  • 249593 11/18/08 19:06 Yi Qiang EPR calibration for Vertical Pumping
  • 249592 11/18/08 19:05 A.Rakhman Alarm
  • 249591 11/18/08 18:48 a-onl Start_Run_3641,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=72,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=test for S2m Left # 4
  • 249590 11/18/08 18:27 Jack (Techs) Right Q2 LHe level meter
  • 249589 11/18/08 18:22 Jack (Techs) Smoke/haze in the hall - Watch the Left HRS vacuum gauge
  • 249588 11/18/08 18:10 a-onl Start_Run_3640,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=72,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=test for S2m Left # 4
  • 249587 11/18/08 18:09 a-onl End_of_Run_3639
  • 249586 11/18/08 17:55 Yi Qiang EPR test w/ correction coil, transverse - field, - pumping
  • 249585 11/18/08 17:53 Katich BCMup3 signal restored to HAPPEX DAQ
  • 249584 11/18/08 17:51 brads BB Cerenkov PMT 5 plugged back into ADC
  • 249583 11/18/08 17:50 brads T7 (BB Cerenkov-only) unplugged from Trigger Supervisor
  • 249582 11/18/08 17:42 E.Chudakov Control access and smoke alert in the hall
  • 249581 11/18/08 17:36 Yi Qiang EPR test w/ correction coil, transverse - field, + pumping
  • 249580 11/18/08 17:30 a-onl Start_Run_3639,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=72,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=
  • 249579 11/18/08 17:27 a-onl End_of_Run_3638
  • 249578 11/18/08 17:24 a-onl Start_Run_3638,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=72,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=
  • 249577 11/18/08 17:22 Yi Qiang EPR Calibration for Transverse (anti_parallel) Pumping w/o correction coil
  • 249576 11/18/08 17:06 Yi Qiang EPR Calibration for Transverse Pumping,
  • 249575 11/18/08 17:05 A.Rakhman alarm
  • 249574 11/18/08 17:02 Yi Qiang AFP Loss Test for Transverse Pumping w/o correction coil
  • 249573 11/18/08 16:43 adev BigBite WireChamber Measured Voltage change!
  • 249572 11/18/08 16:34 adaq target: Spin Flip Report: Target Report Before Rotation to Vertical Field
  • 249571 11/18/08 16:31 adev BigBite WireChamber Measured Voltage change!
  • 249570 11/18/08 16:23 A.Rakhman hall a status
  • 249569 11/18/08 16:03 brads BB Cerenkov witness PMT rates with front shielding
  • 249568 11/18/08 16:02 a-onl End_of_Run_3637
  • 249567 11/18/08 16:00 adaq Beam Time Accounting
  • 249566 11/18/08 15:59 Saha Shift Summary
  • 249565 11/18/08 15:47 Ole Hansen Shorter runs to avoid file splitting
  • 249564 11/18/08 15:22 adaq Scaler rates (BigBite) snapshot
  • 249563 11/18/08 15:04 a-onl Start_Run_3637,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=72,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production
  • 249562 11/18/08 15:03 a-onl End_of_Run_3636
  • 249561 11/18/08 14:24 Lingyan Zhu target: polarization graph
  • 249560 11/18/08 14:24 Yi Qiang Spin Flip Error
  • 249559 11/18/08 14:08 adaq target: Spin Flip Report: Auto spin flip continues
  • 249558 11/18/08 14:06 a-onl Start_Run_3636,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=72,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production
  • 249557 11/18/08 13:52 a-onl End_of_Run_3635
  • 249556 11/18/08 13:47 a-onl Start_Run_3635,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=72,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production
  • 249555 11/18/08 13:45 a-onl End_of_Run_3634
  • 249554 11/18/08 13:24 a-onl Start_Run_3634,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=72,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production
  • 249553 11/18/08 13:22 a-onl End_of_Run_3633
  • 249552 11/18/08 12:31 D. Parno Left-arm VDC efficiency
  • 249551 11/18/08 12:17 a-onl Start_Run_3633,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=72,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production
  • 249550 11/18/08 12:15 a-onl End_of_Run_3632
  • 249549 11/18/08 12:14 A. Camsonne Switched to Rich Holmes analyzer for Compton
  • 249548 11/18/08 12:12 adaq Target Field and Oven Temps snapshot
  • 249547 11/18/08 12:11 adaq Target Cell Temps snapshot
  • 249546 11/18/08 12:11 adaq Radiation and Ozone snapshot
  • 249545 11/18/08 12:11 adaq Hall A Tools snapshot
  • 249544 11/18/08 12:11 adaq BigBox Magnet snapshot
  • 249543 11/18/08 12:11 adaq Beam Position and Current snapshot
  • 249542 11/18/08 12:11 adaq Beam Energy in A and C snapshot
  • 249541 11/18/08 12:07 xqian live time now is about 81.5%, We just keep it now
  • 249540 11/18/08 12:01 Moffit HAPPEX DAQ - Moved x10 BCM to adc15_1
  • 249539 11/18/08 11:37 xqian The BigBite Chamber HVs are not adjusted for this run
  • 249538 11/18/08 11:31 xqian Mapping problem for BB Gas Cerenkov
  • 249537 11/18/08 11:27 xqian RUN 3631 has problem with the chamber HV.
  • 249536 11/18/08 11:25 V. Sulkosky Re: Missing S2m ADC
  • 249535 11/18/08 11:24 adev BigBite WireChamber Measured Voltage change!
  • 249534 11/18/08 11:24 cisbani Run 3629 is a RICH pedestal RUN, with beam off
  • 249533 11/18/08 11:14 a-onl Start_Run_3632,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=72,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production
  • 249532 11/18/08 11:13 adev BigBite WireChamber Measured Voltage change!
  • 249531 11/18/08 11:12 a-onl End_of_Run_3631
  • 249530 11/18/08 11:10 Lingyan Zhu laster spectrum
  • 249529 11/18/08 10:58 Jin Trigger @ 12uA, he3, 3x4 raster
  • 249528 11/18/08 10:56 a-onl Start_Run_3631,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=72,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production
  • 249527 11/18/08 10:51 Jin Huang Spot++ @ 12uA, 3x4, he3 production
  • 249526 11/18/08 10:48 a-onl End_of_Run_3630
  • 249525 11/18/08 10:48 V. Sulkosky VDC T0's Versus Time (resubmission)
  • 249524 11/18/08 10:44 Ole Hansen Start Production
  • 249523 11/18/08 10:43 a-onl Start_Run_3630,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=72,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production
  • 249522 11/18/08 10:36 a-onl End_of_Run_3629
  • 249521 11/18/08 10:36 adaq target: Spin Flip Report: Target Flip ON, Ready for production run, 72% Pol.
  • 249520 11/18/08 10:33 a-onl Start_Run_3629,Run_type=Check out,target_type=Multi Carbon Optics/BeO,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=off,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB Cherenkov test. 0.2uA
  • 249519 11/18/08 10:32 Lingyan Zhu spot++ for raster 6*6
  • 249518 11/18/08 10:30 a-onl End_of_Run_3628
  • 249517 11/18/08 10:29 Jin Trigger @ raster 6x6
  • 249516 11/18/08 10:28 Lingyan Zhu spot++ for raster 5*5
  • 249515 11/18/08 10:27 a-onl Start_Run_3628,Run_type=Check out,target_type=Multi Carbon Optics/BeO,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=off,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB Cherenkov test. 0.2uA
  • 249514 11/18/08 10:26 Ole Hansen Start/stop run scripts not appearing
  • 249513 11/18/08 10:26 a-onl End_of_Run_3627
  • 249512 11/18/08 10:25 a-onl Start_Run_3627,Run_type=Check out,target_type=Multi Carbon Optics/BeO,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=off,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB Cherenkov test. 0.2uA
  • 249511 11/18/08 10:24 Lingyan Zhu spot++ for run 3626
  • 249510 11/18/08 10:23 Ole Hansen CODA restart (again)
  • 249509 11/18/08 10:22 Jin Trigger 2 Raster 5x5
  • 249508 11/18/08 10:21 a-onl End_of_Run_3626
  • 249507 11/18/08 10:20 a-onl Start_Run_3626,Run_type=Check out,target_type=Multi Carbon Optics/BeO,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=off,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB Cherenkov test. 0.2uA
  • 249506 11/18/08 10:19 Jin Trgger @ Raster = 4x4
  • 249505 11/18/08 10:17 Jin Trigger ScreenShot 3x3raster
  • 249504 11/18/08 10:13 Jin Trigger 2uA 2x2
  • 249503 11/18/08 10:13 Lingyan Zhu Target Camera Snapshot
  • 249502 11/18/08 10:10 Jin Target -> He3
  • 249501 11/18/08 10:01 Yi Zhang BB Cherenkov TDC peak
  • 249500 11/18/08 09:39 D. Parno L S2m #4 went missing, came back
  • 249499 11/18/08 08:34 a-onl End_of_Run_3625
  • 249498 11/18/08 08:34 a-onl Start_Run_3625,Run_type=Check out,target_type=Multi Carbon Optics/BeO,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=off,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB Cherenkov test. 0.2uA
  • 249497 11/18/08 08:33 a-onl End_of_Run_3624
  • 249496 11/18/08 08:28 a-onl Start_Run_3624,Run_type=Check out,target_type=Multi Carbon Optics/BeO,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=off,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB Cherenkov test. 0.2uA
  • 249495 11/18/08 08:27 Ole Hansen Run 3623 is junk
  • 249494 11/18/08 08:17 a-onl End_of_Run_3623
  • 249493 11/18/08 08:11 adaq Beam Time Accounting
  • 249492 11/18/08 08:06 Simonetti Owl shift summary
  • 249491 11/18/08 08:05 Simonetti Beam tuning to BeO target
  • 249490 11/18/08 08:04 a-onl Start_Run_3623,Run_type=Check out,target_type=Multi Carbon Optics/BeO,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=off,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB Cherenkov test. 0.6uA
  • 249489 11/18/08 08:00 a-onl End_of_Run_3622
  • 249488 11/18/08 07:53 a-onl Start_Run_3622,Run_type=Check out,target_type=Multi Carbon Optics/BeO,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=off,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB Cherenkov test. 0.2uA
  • 249487 11/18/08 07:52 a-onl End_of_Run_3621
  • 249486 11/18/08 07:51 ycwang Target motion problem has been solved
  • 249485 11/18/08 07:51 a-onl Start_Run_3621,Run_type=Check out,target_type=Multi Carbon Optics/BeO,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=off,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB Cherenkov test. 0.2uA
  • 249484 11/18/08 07:48 Simonetti Target motion problem fixed
  • 249483 11/18/08 07:37 adaq Laser Absorption Spectrum
  • 249482 11/18/08 07:36 Moffit Re: HAPPEX DAQ - fixed... please restart
  • 249481 11/18/08 07:33 Simonetti Pam works to solve target motion problem
  • 249480 11/18/08 07:31 Simonetti Beam problem
  • 249479 11/18/08 07:26 Moffit Re: HAPPEX DAQ - TS1
  • 249478 11/18/08 07:24 adaq target: Spin Flip Report: target stay ar BeO
  • 249477 11/18/08 07:20 a-onl End_of_Run_3620
  • 249476 11/18/08 06:36 Jin Trigger & BCM for run 3620
  • 249475 11/18/08 06:33 a-onl Start_Run_3620,Run_type=Check out,target_type=Multi Carbon Optics/BeO,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=off,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB Cherenkov test. 0.1uA
  • 249474 11/18/08 06:21 a-onl End_of_Run_3619
  • 249473 11/18/08 06:11 Huaibo HAPPEX DAQ problem.
  • 249472 11/18/08 06:08 Huaibo HAPPEX DAQ problem.
  • 249471 11/18/08 05:59 a-onl Start_Run_3619,Run_type=Check out,target_type=Multi Carbon Optics/BeO,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=off,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB Cherenkov test. 0.2uA
  • 249470 11/18/08 05:56 a-onl End_of_Run_3618
  • 249469 11/18/08 05:51 Jin BCM & Trigger ScreenShot for run#3618
  • 249468 11/18/08 05:45 a-onl Start_Run_3618,Run_type=Check out,target_type=Multi Carbon Optics/BeO,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=off,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB Cherenkov test. 0.2uA
  • 249467 11/18/08 05:37 a-onl End_of_Run_3617
  • 249466 11/18/08 05:28 Jin BCM & Trigger Rate for Run3617
  • 249465 11/18/08 05:24 Huaibo triggers for beam current=0.5uA.
  • 249464 11/18/08 05:21 Simonetti Beam tuning, new position for BPMA and BPMB
  • 249463 11/18/08 05:18 a-onl Start_Run_3617,Run_type=Check out,target_type=Multi Carbon Optics/BeO,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=off,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB Cherenkov test. 0.5uA
  • 249462 11/18/08 05:16 Jin Beam Tuning for 0.5uA
  • 249461 11/18/08 04:28 a-onl End_of_Run_3616
  • 249460 11/18/08 04:26 Jin Fake Spin Signal ON
  • 249459 11/18/08 04:11 adaq Target Field and Oven Temps snapshot
  • 249458 11/18/08 04:11 adaq Target Cell Temps snapshot
  • 249457 11/18/08 04:11 adaq Radiation and Ozone snapshot
  • 249456 11/18/08 04:11 adaq Hall A Tools snapshot
  • 249455 11/18/08 04:11 adaq BigBox Magnet snapshot
  • 249454 11/18/08 04:11 adaq Beam Position and Current snapshot
  • 249453 11/18/08 04:11 adaq Beam Energy in A and C snapshot
  • 249452 11/18/08 04:01 jpchen target motion problem
  • 249451 11/18/08 03:52 Fansler/Sega Target motion controls
  • 249450 11/18/08 02:16 Jin Huang hareboot24 is now accessable in counting house front room
  • 249449 11/18/08 00:12 dhamilton Swing shift summary
  • 249448 11/18/08 00:11 adaq Beam Time Accounting
  • 249447 11/18/08 00:08 a-onl Start_Run_3616,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Cosmics
  • Monday

  • 249446 11/17/08 23:37 Jin Spin Flip Exit at Spin=-1 State
  • 249445 11/17/08 23:04 a-onl Start_Run_3615,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Cosmics
  • 249444 11/17/08 23:00 Yi Qiang Upexpected target ladder movement
  • 249443 11/17/08 22:56 dhamilton Move Hall to controlled access
  • 249442 11/17/08 22:47 dhamilton Problem with target motion
  • 249441 11/17/08 22:16 adev BigBite WireChamber Measured Voltage change!
  • 249440 11/17/08 22:07 dhamilton BB HV on
  • 249439 11/17/08 22:05 dhamilton Nominal beam position and harp scan complete
  • 249438 11/17/08 21:45 Jin Huang The Scalar Counting Difference Climbing is due to missing on LHRS scalar
  • 249437 11/17/08 21:42 dhamilton Tuning through Compton finished
  • 249436 11/17/08 21:19 Sirish Compton Tune
  • 249435 11/17/08 20:11 adaq Target Field and Oven Temps snapshot
  • 249434 11/17/08 20:11 adaq Target Cell Temps snapshot
  • 249433 11/17/08 20:11 adaq Radiation and Ozone snapshot
  • 249432 11/17/08 20:11 adaq Hall A Tools snapshot
  • 249431 11/17/08 20:10 adaq BigBox Magnet snapshot
  • 249430 11/17/08 20:10 adaq Beam Position and Current snapshot
  • 249429 11/17/08 20:10 adaq Beam Energy in A and C snapshot
  • 249428 11/17/08 19:53 adaq Beam Time Accounting
  • 249427 11/17/08 19:52 Saha Shift Summary
  • 249426 11/17/08 19:47 E.Chudakov Moller results
  • 249425 11/17/08 19:42 dalton Target: optics target in
  • 249424 11/17/08 17:53 V. Sulkosky VDC T0 Comparison Run 3611
  • 249423 11/17/08 17:36 xqian BB Wire Chamber
  • 249422 11/17/08 17:31 xqian explaination
  • 249421 11/17/08 17:10 dhamilton Dipole back up to 2.35 GeV/c
  • 249420 11/17/08 15:50 gjin laser spectrum of empty target
  • 249419 11/17/08 15:43 A. Camsonne Helium level low
  • 249418 11/17/08 15:38 xqian power of retiming on BigBite
  • 249417 11/17/08 15:24 adev BigBite WireChamber Measured Current change!
  • 249416 11/17/08 15:22 a-onl End_of_Run_3614
  • 249415 11/17/08 15:20 V. Sulkosky Golden Root File Now 3611
  • 249414 11/17/08 15:14 a-onl Start_Run_3614,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH test
  • 249413 11/17/08 15:12 adev BigBite WireChamber Measured Voltage change!
  • 249412 11/17/08 15:01 adev BigBite WireChamber Measured Voltage change!
  • 249411 11/17/08 15:01 a-onl End_of_Run_3613
  • 249410 11/17/08 14:58 a-onl Start_Run_3613,Run_type=Calibration,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH Pedestal
  • 249409 11/17/08 14:58 brads BB Cerenkov PMT 5 plugged into online scope
  • 249408 11/17/08 14:57 V. Sulkosky S2m ADC # 4 Plot
  • 249407 11/17/08 14:57 xqian explaination
  • 249406 11/17/08 14:56 adaq target: Spin Flip Report: target status after stopping auto flip
  • 249405 11/17/08 14:53 a-onl End_of_Run_3612
  • 249404 11/17/08 14:51 a-onl Start_Run_3612,Run_type=Calibration,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH Pedestal
  • 249403 11/17/08 14:50 A. Camsonne RICH phase tweaked
  • 249402 11/17/08 14:46 brads Added plate in front of BB Cerenkov witness PMT
  • 249401 11/17/08 14:39 brads Cerenkov rate data taken on Sunday
  • 249400 11/17/08 14:24 a-onl End_of_Run_3611
  • 249399 11/17/08 14:22 A. Camsonne Experiment status and plan, and meeting reminder
  • 249398 11/17/08 14:18 Saha Spot++ for Run3611
  • 249397 11/17/08 14:10 a-onl Start_Run_3611,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=
  • 249396 11/17/08 14:08 a-onl End_of_Run_3610
  • 249395 11/17/08 14:07 a-onl Start_Run_3610,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test; BB DC HV down
  • 249394 11/17/08 14:06 brads Results from the paddles set-up behind the BB Shower
  • 249393 11/17/08 14:03 a-onl End_of_Run_3609
  • 249392 11/17/08 13:53 a-onl Start_Run_3609,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production
  • 249391 11/17/08 13:50 a-onl End_of_Run_3608
  • 249390 11/17/08 13:45 a-onl Start_Run_3608,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production
  • 249389 11/17/08 13:42 a-onl End_of_Run_3607
  • 249388 11/17/08 13:40 a-onl Start_Run_3607,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Raster check
  • 249387 11/17/08 13:36 a-onl End_of_Run_3606
  • 249386 11/17/08 13:36 a-onl Start_Run_3606,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Raster check
  • 249385 11/17/08 13:31 a-onl End_of_Run_3605
  • 249384 11/17/08 13:31 a-onl Start_Run_3605,Run_type=Test,target_type=Multi Carbon Optics/BeO,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Raster check
  • 249383 11/17/08 13:29 V. Sulkosky Checked Re-timing Signal with L1A
  • 249382 11/17/08 13:28 brads Photos of added shielding on BB beamline side
  • 249381 11/17/08 13:15 V. Sulkosky Moved S2m left PMT #4 Signal to a New ADC Channel
  • 249380 11/17/08 12:54 V. Sulkosky Run 3604 Pedestal run
  • 249379 11/17/08 12:51 a-onl End_of_Run_3604
  • 249378 11/17/08 12:50 a-onl Start_Run_3604,Run_type=Test,target_type=No Target,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=pedestal run with T8
  • 249377 11/17/08 12:48 a-onl End_of_Run_3603
  • 249376 11/17/08 12:47 Yi Qiang Comparison of spin up time with different pumping direction
  • 249375 11/17/08 12:37 a-onl Start_Run_3603,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test; no beam
  • 249374 11/17/08 12:26 a-onl Start_Run_3602,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test with T3 for RICH
  • 249373 11/17/08 12:25 adaq Fixed the bad scintillator TDC chan
  • 249372 11/17/08 12:22 a-onl Start_Run_3601,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test with T3 for RICH
  • 249371 11/17/08 12:21 a-onl End_of_Run_3600
  • 249370 11/17/08 12:18 a-onl Start_Run_3600,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test with T3 for RICH
  • 249369 11/17/08 12:17 a-onl End_of_Run_3599
  • 249368 11/17/08 12:10 adaq Target Field and Oven Temps snapshot
  • 249367 11/17/08 12:10 adaq Target Cell Temps snapshot
  • 249366 11/17/08 12:10 adaq Radiation and Ozone snapshot
  • 249365 11/17/08 12:10 adaq Hall A Tools snapshot
  • 249364 11/17/08 12:10 adaq BigBox Magnet snapshot
  • 249363 11/17/08 12:10 adaq Beam Position and Current snapshot
  • 249362 11/17/08 12:10 adaq Beam Energy in A and C snapshot
  • 249361 11/17/08 12:05 a-onl Start_Run_3599,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test with T3 for RICH
  • 249360 11/17/08 12:02 a-onl End_of_Run_3598
  • 249359 11/17/08 11:58 a-onl Start_Run_3598,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test with T1 and T3
  • 249358 11/17/08 11:58 a-onl End_of_Run_3597
  • 249357 11/17/08 11:50 a-onl Start_Run_3597,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test with T1 and T3
  • 249356 11/17/08 11:46 a-onl End_of_Run_3596
  • 249355 11/17/08 11:39 Fansler RQ2/LQ1
  • 249354 11/17/08 11:38 a-onl Start_Run_3596,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test with T1
  • 249353 11/17/08 10:53 Yi Zhang BB Cherenkov test ADC plots
  • 249352 11/17/08 10:52 adaq target: Spin Flip Report: target status before auto spin flip starts
  • 249351 11/17/08 10:35 A. Camsonne Energy measurement complete
  • 249350 11/17/08 10:32 Saha Arc Energy Measurement
  • 249349 11/17/08 10:30 a-onl Start_Run_3595,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test
  • 249348 11/17/08 10:20 a-onl End_of_Run_3594
  • 249347 11/17/08 10:10 a-onl Start_Run_3594,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production; Happex run#30416
  • 249346 11/17/08 09:58 a-onl End_of_Run_3593
  • 249345 11/17/08 09:43 A. Camsonne Access for fixing cryo for spectrometer
  • 249344 11/17/08 09:34 V. Sulkosky Left HRS Scaler Detector Rates Zero
  • 249343 11/17/08 09:27 a-onl Start_Run_3593,Run_type=Test,target_type=No Target,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=test; no beam
  • 249342 11/17/08 09:08 a-onl Start_Run_3592,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=test; no beam
  • 249341 11/17/08 09:05 adev SCALER = -1
  • 249340 11/17/08 09:04 a-onl End_of_Run_3591
  • 249339 11/17/08 09:01 a-onl Start_Run_3591,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=test; no beam
  • 249338 11/17/08 08:58 adev SCALER = -1
  • 249337 11/17/08 08:51 a-onl End_of_Run_3589
  • 249336 11/17/08 08:48 adaq target: Spin Flip Report: test
  • 249335 11/17/08 08:46 a-onl Start_Run_3589,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=test; no beam
  • 249334 11/17/08 08:23 a-onl End_of_Run_3588
  • 249333 11/17/08 08:02 adaq Beam Time Accounting
  • 249332 11/17/08 08:01 Cisbani Owl Shift Summary
  • 249331 11/17/08 08:01 Simonetti Laser spectrum and spin flip screen
  • 249330 11/17/08 07:54 adaq target: Spin Flip Report: End Shift, Spin flip report
  • 249329 11/17/08 07:43 cisbani BB Cherenkov tests
  • 249328 11/17/08 07:30 a-onl Start_Run_3588,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Big Bite Cherenkov test, 6 microAmp
  • 249327 11/17/08 07:27 a-onl End_of_Run_3587
  • 249326 11/17/08 07:09 Vilardi Missing channel, Left Arm S2 back again
  • 249325 11/17/08 07:09 Vilardi Missing channel, Left Arm S2 back again
  • 249324 11/17/08 06:43 a-onl Start_Run_3587,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Big Bite Cherenkov test, 1 microAmp
  • 249323 11/17/08 06:36 a-onl End_of_Run_3586
  • 249322 11/17/08 05:58 a-onl Start_Run_3586,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=12uA Production.
  • 249321 11/17/08 05:58 a-onl End_of_Run_3585
  • 249320 11/17/08 05:44 vilardi I. Missing channel
  • 249319 11/17/08 05:17 cisbani ROC5, 8, 9, 10 and 15 not responding
  • 249318 11/17/08 05:15 a-onl Start_Run_3585,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=12uA Production.
  • 249317 11/17/08 05:09 a-onl Start_Run_3584,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=12uA Production.
  • 249316 11/17/08 04:56 a-onl Start_Run_3583,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=12uA Production.
  • 249315 11/17/08 04:42 adaq target: Spin Flip Report: Pol. Stablized to ~66%
  • 249314 11/17/08 04:36 Jin Correction : Pol. History of Session # 241
  • 249313 11/17/08 04:25 Jin Huang Target Polarization History during Production
  • 249312 11/17/08 04:10 adaq Target Field and Oven Temps snapshot
  • 249311 11/17/08 04:10 adaq Target Cell Temps snapshot
  • 249310 11/17/08 04:10 adaq Radiation and Ozone snapshot
  • 249309 11/17/08 04:10 adaq Hall A Tools snapshot
  • 249308 11/17/08 04:10 adaq BigBox Magnet snapshot
  • 249307 11/17/08 04:10 adaq Beam Position and Current snapshot
  • 249306 11/17/08 04:10 adaq Beam Energy in A and C snapshot
  • 249305 11/17/08 04:02 a-onl Start_Run_3582,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=12uA Production.
  • 249304 11/17/08 04:01 a-onl End_of_Run_3581
  • 249303 11/17/08 03:42 adaq Hall A Tools
  • 249302 11/17/08 03:25 a-onl Start_Run_3581,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=12uA Production.
  • 249301 11/17/08 03:24 a-onl End_of_Run_3580
  • 249300 11/17/08 02:46 a-onl Start_Run_3580,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=12uA Production.
  • 249299 11/17/08 02:44 a-onl End_of_Run_3579
  • 249298 11/17/08 02:01 a-onl Start_Run_3579,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=12uA Production.
  • 249297 11/17/08 02:00 a-onl End_of_Run_3578
  • 249296 11/17/08 01:19 a-onl Start_Run_3578,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=12uA Production.
  • 249295 11/17/08 01:18 a-onl End_of_Run_3577
  • 249294 11/17/08 00:42 a-onl Start_Run_3577,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=12uA Production.
  • 249293 11/17/08 00:41 a-onl End_of_Run_3576
  • 249292 11/17/08 00:02 a-onl Start_Run_3576,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=12uA Production.
  • 249291 11/17/08 00:01 adaq Beam Time Accounting
  • 249290 11/17/08 00:01 D. Dutta Shift Summary
  • 249289 11/17/08 00:01 a-onl End_of_Run_3575
  • 249288 11/17/08 00:00 narayan Target Polarization
  • 249287 11/17/08 00:00 adaq target: Spin Flip Report: Target General Update: the NMR looks bad
  • Sunday

  • 249286 11/16/08 23:32 xqian Background study at 12 uA
  • 249285 11/16/08 23:22 a-onl Start_Run_3575,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=12uA Production.
  • 249284 11/16/08 23:21 a-onl End_of_Run_3574
  • 249283 11/16/08 23:15 Yi Qiang Really Masing !!!
  • 249282 11/16/08 23:04 A. Camsonne RICH data for ROC15
  • 249281 11/16/08 23:01 Kalyan re: Dead time issue again,
  • 249280 11/16/08 22:58 Jin Huang Update on scalar synchronization check
  • 249279 11/16/08 22:46 Chiranjib Effect of correction coil
  • 249278 11/16/08 22:45 a-onl Start_Run_3574,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=12uA Production. During target tests.
  • 249277 11/16/08 22:44 a-onl End_of_Run_3573
  • 249276 11/16/08 22:43 D. Parno L-HRS S2m ADC 4 is back
  • 249275 11/16/08 22:28 Chiranjib CORRECTION COIL TURNED OFF
  • 249274 11/16/08 22:18 xqian Dead time issue again,
  • 249273 11/16/08 22:12 D. Parno Data from S2m Left ADC 4
  • 249272 11/16/08 22:08 narayan target test performed
  • 249271 11/16/08 22:06 adaq target: Spin Flip Report: Correction Coil Off before Flip#28
  • 249270 11/16/08 22:05 a-onl Start_Run_3573,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=12uA Production. During target tests.
  • 249269 11/16/08 22:03 a-onl End_of_Run_3572
  • 249268 11/16/08 21:29 adaq Q1 Helium flow alarm
  • 249267 11/16/08 21:25 a-onl Start_Run_3572,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=12uA Production
  • 249266 11/16/08 21:23 a-onl End_of_Run_3571
  • 249265 11/16/08 20:57 adev BigBite WireChamber Measured Voltage change!
  • 249264 11/16/08 20:45 a-onl Start_Run_3571,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=12uA Production
  • 249263 11/16/08 20:44 adaq BCM/BPM strip scan
  • 249262 11/16/08 20:44 a-onl End_of_Run_3570
  • 249261 11/16/08 20:42 adaq Hall A tools scan
  • 249260 11/16/08 20:09 adaq Target Field and Oven Temps snapshot
  • 249259 11/16/08 20:09 adaq Target Cell Temps snapshot
  • 249258 11/16/08 20:09 adaq Radiation and Ozone snapshot
  • 249257 11/16/08 20:09 adaq Hall A Tools snapshot
  • 249256 11/16/08 20:09 adaq BigBox Magnet snapshot
  • 249255 11/16/08 20:09 adaq Beam Position and Current snapshot
  • 249254 11/16/08 20:09 adaq Beam Energy in A and C snapshot
  • 249253 11/16/08 20:08 a-onl Start_Run_3570,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=12uA Production
  • 249252 11/16/08 20:07 a-onl End_of_Run_3569
  • 249251 11/16/08 19:42 a-onl End_of_Run_3569
  • 249250 11/16/08 19:31 a-onl Start_Run_3569,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=12uA Production
  • 249249 11/16/08 19:30 adaq target: Spin Flip Report: Regular Update: We initiated taking data a little, Regular update
  • 249248 11/16/08 18:55 a-onl Start_Run_3568,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=12uA Production
  • 249247 11/16/08 18:18 Kalyan BB Shower vs preshower for T1 and T6
  • 249246 11/16/08 17:20 a-onl End_of_Run_3567
  • 249245 11/16/08 17:03 Kalyan Run 3566 is production
  • 249244 11/16/08 16:51 Jin Golden Run -> 3565
  • 249243 11/16/08 16:45 a-onl Start_Run_3567,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=8uA Production, stable with MWDC HV adjusted
  • 249242 11/16/08 16:43 a-onl End_of_Run_3566
  • 249241 11/16/08 16:22 Chiranjib Polarization just before the production @12uA
  • 249240 11/16/08 16:20 Chiranjib Target moved to He-3/ started spin flip
  • 249239 11/16/08 16:18 Chiranjib BB wire ch dark current
  • 249238 11/16/08 16:18 Chiranjib BB wire chamber dark currents
  • 249237 11/16/08 16:16 Kalyan Trigger rates with 12uA current on 3He
  • 249236 11/16/08 16:15 a-onl Start_Run_3566,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=8uA Production, stable with MWDC HV adjusted
  • 249235 11/16/08 16:03 dhamilton Day Shift Summary
  • 249234 11/16/08 16:02 adaq Beam Time Accounting
  • 249233 11/16/08 15:52 Zhihong Ye Laser Spetrum Screen for Nov 16th Day shift
  • 249232 11/16/08 15:20 Zhihong Ye Target moved to Optics Target Position
  • 249231 11/16/08 14:41 Yi Qiang Target Strip Charts
  • 249230 11/16/08 14:32 adev BigBite WireChamber Measured Voltage change!
  • 249229 11/16/08 14:24 a-onl End_of_Run_3565
  • 249228 11/16/08 14:16 Kalyan BB scintillator right chan 10 TDC is missing
  • 249227 11/16/08 14:04 Kalyan Run 3565 is production and not test
  • 249226 11/16/08 13:54 Yi Qiang Target Spin Flip Code Restarted
  • 249225 11/16/08 13:40 Kalyan Test for LHRS retiming
  • 249224 11/16/08 13:34 D. Parno Paddle Timing
  • 249223 11/16/08 13:34 adaq target: Spin Flip Report: Target Auto Spin Flip is back on line
  • 249222 11/16/08 13:32 a-onl Start_Run_3565,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=8uA Production, stable with MWDC HV adjusted
  • 249221 11/16/08 13:26 a-onl End_of_Run_3564
  • 249220 11/16/08 13:24 a-onl Start_Run_3564,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=8uA Production, stable with MWDC HV adjusted
  • 249219 11/16/08 13:14 Kalyan Plot of HRS target z - BB target z for good e-pi coincidences
  • 249218 11/16/08 13:10 a-onl End_of_Run_3563
  • 249217 11/16/08 12:51 D. Parno Timing Shifts -- Individual Paddles
  • 249216 11/16/08 12:50 Kalyan Coincidence time for run 3562
  • 249215 11/16/08 12:38 Kalyan S2m Left ADC 4 is missing
  • 249214 11/16/08 12:37 a-onl Start_Run_3563,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=8uA Production, stable with MWDC HV adjusted
  • 249213 11/16/08 12:36 a-onl End_of_Run_3562
  • 249212 11/16/08 12:13 D. Parno L-HRS TDC still shifted after trigger download
  • 249211 11/16/08 12:09 adaq Target Field and Oven Temps snapshot
  • 249210 11/16/08 12:09 adaq Target Cell Temps snapshot
  • 249209 11/16/08 12:09 adaq Radiation and Ozone snapshot
  • 249208 11/16/08 12:09 adaq Hall A Tools snapshot
  • 249207 11/16/08 12:09 adaq BigBox Magnet snapshot
  • 249206 11/16/08 12:09 adaq Beam Position and Current snapshot
  • 249205 11/16/08 12:09 adaq Beam Energy in A and C snapshot
  • 249204 11/16/08 11:57 a-onl Start_Run_3562,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=8uA Production, stable with MWDC HV adjusted, after trigger redownload
  • 249203 11/16/08 11:56 A. Camsonne Redownload of the HRS trigger
  • 249202 11/16/08 11:54 atrig Trigger-Download
  • 249201 11/16/08 11:54 a-onl End_of_Run_3561
  • 249200 11/16/08 11:39 a-onl Start_Run_3561,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=8uA Production, stable with MWDC HV adjusted
  • 249199 11/16/08 11:31 a-onl End_of_Run_3560
  • 249198 11/16/08 11:08 xqian Here is the plot
  • 249197 11/16/08 11:03 Kalyan T3 delay was set to zero on 11th nov (log 248234 )
  • 249196 11/16/08 10:42 a-onl Start_Run_3560,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=8uA Production, stable with MWDC HV adjusted
  • 249195 11/16/08 10:40 a-onl End_of_Run_3559
  • 249194 11/16/08 10:24 D. Parno Shift in L-HRS TDC Spectrum
  • 249193 11/16/08 10:22 xqian RUN 3559 and the run before that all habe timing problems in left HRS!!!
  • 249192 11/16/08 10:21 xqian LEFT HRS out of window again!!!
  • 249191 11/16/08 10:12 a-onl Start_Run_3559,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=8uA Production, stable with MWDC HV adjusted
  • 249190 11/16/08 10:09 a-onl Start_Run_3558,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=8uA Production, stable with MWDC HV adjusted
  • 249189 11/16/08 10:04 a-onl End_of_Run_3557
  • 249188 11/16/08 09:55 xqian live time again for RUN 3556
  • 249187 11/16/08 09:52 a-onl Start_Run_3557,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=8uA Production, stable with MWDC HV adjusted
  • 249186 11/16/08 09:52 xqian Update the drift time t0 for the new run , I created a new database of 20081113
  • 249185 11/16/08 09:51 a-onl End_of_Run_3556
  • 249184 11/16/08 09:44 xqian Update in background rates
  • 249183 11/16/08 09:21 dhamilton BPM/BCM strip charts
  • 249182 11/16/08 09:20 dhamilton Hall A Tools
  • 249181 11/16/08 09:08 a-onl Start_Run_3556,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=8uA Production, stable with MWDC HV adjusted
  • 249180 11/16/08 09:05 a-onl End_of_Run_3555
  • 249179 11/16/08 08:28 xqian RUN 3553,3554,3555 need special treatment
  • 249178 11/16/08 08:20 a-onl Start_Run_3555,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=waiting 8uA
  • 249177 11/16/08 08:19 a-onl End_of_Run_3554
  • 249176 11/16/08 08:16 a-onl Start_Run_3554,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=8uA
  • 249175 11/16/08 08:14 a-onl End_of_Run_3553
  • 249174 11/16/08 08:13 a-onl Start_Run_3553,Run_type=Calibration,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=rich pedestal 3-6 uA (approx., expecting 8.) junk
  • 249173 11/16/08 08:06 a-onl End_of_Run_3552
  • 249172 11/16/08 08:06 Simonetti Laser spectrum and Spin flip system control
  • 249171 11/16/08 08:04 adaq Beam Time Accounting
  • 249170 11/16/08 08:04 cisbani Owl Shift Summary
  • 249169 11/16/08 08:04 a-onl Start_Run_3552,Run_type=Calibration,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=rich pedestal 3-6 uA (approx., expecting 8.)
  • 249168 11/16/08 08:02 a-onl End_of_Run_3551
  • 249167 11/16/08 08:01 a-onl Start_Run_3551,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=rich pedestal 6 uA (approx., expecting 8.)
  • 249166 11/16/08 07:56 xqian live time for RUN 3550
  • 249165 11/16/08 07:54 adaq target: Spin Flip Report: Laser spectrum
  • 249164 11/16/08 07:48 Simonetti Target activity
  • 249163 11/16/08 07:42 a-onl End_of_Run_3550
  • 249162 11/16/08 07:22 xqian Effect of our new collimator
  • 249161 11/16/08 07:17 Jin Huang Scalar Fast Clock Synchronization Check
  • 249160 11/16/08 07:09 xqian Prescale study
  • 249159 11/16/08 07:04 Jin Huang New Replay, New Online Plots
  • 249158 11/16/08 07:03 a-onl Start_Run_3550,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=production run, 4 micro A
  • 249157 11/16/08 06:53 cisbani On RICH deadtime during cosmics in Nov/15 swing shift
  • 249156 11/16/08 06:50 a-onl End_of_Run_3549
  • 249155 11/16/08 06:43 a-onl Start_Run_3549,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH in , test P1 and P6 prescale
  • 249154 11/16/08 06:43 a-onl End_of_Run_3548
  • 249153 11/16/08 06:41 a-onl Start_Run_3548,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH in , test P1 and P6 prescale
  • 249152 11/16/08 06:39 a-onl End_of_Run_3547
  • 249151 11/16/08 06:38 a-onl Start_Run_3547,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=No RICH, the rates is about 700 Hz, 10% dead time
  • 249150 11/16/08 06:36 a-onl End_of_Run_3546
  • 249149 11/16/08 06:35 a-onl Start_Run_3546,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=prescale the T1 and T6 away!!! only 4-5% dead time
  • 249148 11/16/08 06:32 a-onl End_of_Run_3545
  • 249147 11/16/08 06:23 xqian Background study after more shielding and collimator
  • 249146 11/16/08 06:09 a-onl End_of_Run_3545
  • 249145 11/16/08 06:06 xqian BigBite Retming works, Great!!!
  • 249144 11/16/08 06:02 xqian Live time for 2uA 3He, see plot
  • 249143 11/16/08 05:55 a-onl Start_Run_3545,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=
  • 249142 11/16/08 05:31 cisbani spot++ 3He TPoc
  • 249141 11/16/08 05:14 Jin Huang Real Spin Flip Start. Pol.=58%
  • 249140 11/16/08 05:14 a-onl Start_Run_3544,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=
  • 249139 11/16/08 05:10 Simonetti Temperatures of He3 target during polarization
  • 249138 11/16/08 04:58 Simonetti moving target to He3 cell
  • 249137 11/16/08 04:55 a-onl End_of_Run_3543
  • 249136 11/16/08 04:44 xqian Live time in RUN3541
  • 249135 11/16/08 04:32 xqian RUN 3541,3542,3543, have a lot of F1 TDC overflow. Why?
  • 249134 11/16/08 04:24 a-onl Start_Run_3543,Run_type=Test,target_type=Reference Cell: Empty,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=
  • 249133 11/16/08 04:23 cisbani BB chamber tripped again, run 3542
  • 249132 11/16/08 04:21 a-onl End_of_Run_3542
  • 249131 11/16/08 04:12 cisbani BB Chamber HV Trip, run 3542
  • 249130 11/16/08 04:10 adaq spot++ on BeO 3x4 MCC
  • 249129 11/16/08 04:09 adaq Target Field and Oven Temps snapshot
  • 249128 11/16/08 04:09 adaq Target Cell Temps snapshot
  • 249127 11/16/08 04:08 adaq Radiation and Ozone snapshot
  • 249126 11/16/08 04:08 adaq Hall A Tools snapshot
  • 249125 11/16/08 04:08 adaq BigBox Magnet snapshot
  • 249124 11/16/08 04:08 adaq Beam Position and Current snapshot
  • 249123 11/16/08 04:08 adaq Beam Energy in A and C snapshot
  • 249122 11/16/08 03:59 Simonetti Hall A Target Cell Temperature
  • 249121 11/16/08 03:57 a-onl Start_Run_3542,Run_type=Production,target_type=Reference Cell: Empty,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=
  • 249120 11/16/08 03:56 a-onl End_of_Run_3541
  • 249119 11/16/08 03:50 xqian Some thoughts about the dead time with our current trigger setup
  • 249118 11/16/08 03:44 adaq target: Spin Flip Report: No Shift Worker Comment
  • 249117 11/16/08 03:32 Simonetti Beam tuning and run with empty reference cell
  • 249116 11/16/08 03:16 a-onl Start_Run_3541,Run_type=Production,target_type=Reference Cell: Empty,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=
  • 249115 11/16/08 03:15 a-onl End_of_Run_3540
  • 249114 11/16/08 03:12 a-onl Start_Run_3540,Run_type=Production,target_type=Reference Cell: Empty,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=
  • 249113 11/16/08 03:10 a-onl End_of_Run_3539
  • 249112 11/16/08 03:04 a-onl Start_Run_3539,Run_type=Production,target_type=Reference Cell: Empty,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=
  • 249111 11/16/08 03:03 a-onl End_of_Run_3538
  • 249110 11/16/08 03:01 a-onl Start_Run_3538,Run_type=Production,target_type=Reference Cell: Empty,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=
  • 249109 11/16/08 02:57 a-onl End_of_Run_3537
  • 249108 11/16/08 02:49 a-onl Start_Run_3537,Run_type=Test,target_type=Multi Carbon Optics/BeO,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=rich deadtime check!!
  • 249107 11/16/08 02:48 a-onl End_of_Run_3536
  • 249106 11/16/08 02:46 a-onl Start_Run_3536,Run_type=Test,target_type=Multi Carbon Optics/BeO,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=
  • 249105 11/16/08 02:41 a-onl End_of_Run_3535
  • 249104 11/16/08 02:40 a-onl Start_Run_3535,Run_type=Cosmics,target_type=No Target,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=check!!
  • 249103 11/16/08 02:32 ycwang RICH HV tripped when tuning up beam
  • 249102 11/16/08 02:27 a-onl End_of_Run_3534
  • 249101 11/16/08 02:20 a-onl Start_Run_3534,Run_type=Cosmics,target_type=No Target,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=check!!
  • 249100 11/16/08 02:19 a-onl End_of_Run_3533
  • 249099 11/16/08 02:14 a-onl Start_Run_3533,Run_type=Cosmics,target_type=No Target,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=check!!
  • 249098 11/16/08 02:04 a-onl End_of_Run_3532
  • 249097 11/16/08 02:03 a-onl Start_Run_3532,Run_type=Cosmics,target_type=No Target,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=pedestal run for rich
  • 249096 11/16/08 01:57 a-onl End_of_Run_3531
  • 249095 11/16/08 01:56 a-onl Start_Run_3531,Run_type=Cosmics,target_type=No Target,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=check!!!
  • 249094 11/16/08 01:54 a-onl End_of_Run_3530
  • 249093 11/16/08 01:54 adaq Harp scan problem solved, beam profile within 200 um
  • 249092 11/16/08 01:53 a-onl Start_Run_3530,Run_type=Cosmics,target_type=No Target,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=check!!!
  • 249091 11/16/08 01:10 A. Camsonne BCM strip chart
  • 249090 11/16/08 00:42 cisbani HallA Tools
  • 249089 11/16/08 00:24 Kalyan BB thresholds set to beam conditions
  • 249088 11/16/08 00:19 a-onl End_of_Run_3529
  • 249087 11/16/08 00:10 a-onl Start_Run_3529,Run_type=Cosmics,target_type=No Target,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=check!!!
  • 249086 11/16/08 00:10 adaq target: Spin Flip Report: test
  • 249085 11/16/08 00:07 adaq Beam Time Accounting
  • 1 11/16/08 00:07 a-onl End_of_Run_3528
  • Saturday

  • 249084 11/15/08 23:52 a-onl Start_Run_3528,Run_type=Cosmics,target_type=No Target,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=check!!!
  • 249083 11/15/08 23:49 R. Michaels notes after raster repair (TECH)
  • 249082 11/15/08 23:29 a-onl End_of_Run_3527
  • 249081 11/15/08 23:26 a-onl Start_Run_3527,Run_type=Cosmics,target_type=No Target,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=check!!!
  • 249080 11/15/08 23:25 a-onl End_of_Run_3526
  • 249079 11/15/08 23:24 a-onl Start_Run_3526,Run_type=Cosmics,target_type=No Target,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=check!!!
  • 249078 11/15/08 23:23 a-onl End_of_Run_3525
  • 249077 11/15/08 23:22 a-onl Start_Run_3525,Run_type=Cosmics,target_type=No Target,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=check!!!
  • 249076 11/15/08 23:21 a-onl End_of_Run_3524
  • 249075 11/15/08 23:21 a-onl Start_Run_3524,Run_type=Cosmics,target_type=No Target,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=check!!!
  • 249074 11/15/08 23:20 a-onl End_of_Run_3523
  • 249073 11/15/08 23:20 a-onl Start_Run_3523,Run_type=Cosmics,target_type=No Target,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=check!!!
  • 249072 11/15/08 23:19 a-onl End_of_Run_3522
  • 249071 11/15/08 23:18 a-onl Start_Run_3522,Run_type=Cosmics,target_type=No Target,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=check!!!
  • 249070 11/15/08 23:17 a-onl End_of_Run_3521
  • 249069 11/15/08 23:17 adev SCALER = -1
  • 249068 11/15/08 23:17 a-onl Start_Run_3521,Run_type=Cosmics,target_type=No Target,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=check!!!
  • 249067 11/15/08 23:14 a-onl End_of_Run_3520
  • 249066 11/15/08 23:13 a-onl Start_Run_3520,Run_type=Cosmics,target_type=No Target,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=check!!!
  • 249065 11/15/08 23:12 a-onl End_of_Run_3519
  • 249064 11/15/08 23:11 a-onl Start_Run_3519,Run_type=Cosmics,target_type=No Target,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=check!!!
  • 249063 11/15/08 23:10 a-onl End_of_Run_3518
  • 249062 11/15/08 23:08 adev SCALER = -1
  • 249061 11/15/08 23:08 a-onl Start_Run_3518,Run_type=Cosmics,target_type=No Target,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=check!!!
  • 249060 11/15/08 23:07 a-onl End_of_Run_3517
  • 249059 11/15/08 23:06 a-onl Start_Run_3517,Run_type=Cosmics,target_type=No Target,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=check!!!
  • 249058 11/15/08 23:06 a-onl End_of_Run_3516
  • 249057 11/15/08 22:59 adev SCALER = -1
  • 249056 11/15/08 22:58 a-onl Start_Run_3516,Run_type=Cosmics,target_type=No Target,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=check!!!
  • 249055 11/15/08 22:57 a-onl End_of_Run_3515
  • 249054 11/15/08 22:57 a-onl Start_Run_3515,Run_type=Cosmics,target_type=No Target,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=check!!!
  • 249053 11/15/08 22:47 a-onl End_of_Run_3514
  • 249052 11/15/08 22:47 a-onl Start_Run_3514,Run_type=Cosmics,target_type=No Target,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=check!!!
  • 249051 11/15/08 22:46 a-onl End_of_Run_3513
  • 249050 11/15/08 22:45 a-onl Start_Run_3513,Run_type=Cosmics,target_type=No Target,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=check!!!
  • 249049 11/15/08 22:43 a-onl End_of_Run_3512
  • 249048 11/15/08 22:42 a-onl Start_Run_3512,Run_type=Cosmics,target_type=No Target,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=check!!!
  • 249047 11/15/08 22:32 a-onl End_of_Run_3510
  • 249046 11/15/08 22:32 a-onl Start_Run_3510,Run_type=Cosmics,target_type=No Target,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=check!!!
  • 249045 11/15/08 22:30 a-onl End_of_Run_3509
  • 249044 11/15/08 22:30 a-onl Start_Run_3509,Run_type=Cosmics,target_type=No Target,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=check!!!
  • 249043 11/15/08 22:21 a-onl End_of_Run_3508
  • 249042 11/15/08 22:20 a-onl Start_Run_3508,Run_type=Cosmics,target_type=No Target,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=check!!!
  • 249041 11/15/08 22:18 A. Camsonne Update about raster
  • 249040 11/15/08 22:15 a-onl End_of_Run_3507
  • 249039 11/15/08 22:14 a-onl Start_Run_3507,Run_type=Cosmics,target_type=No Target,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=check!!!
  • 249038 11/15/08 22:12 a-onl End_of_Run_3506
  • 249037 11/15/08 22:11 a-onl Start_Run_3506,Run_type=Cosmics,target_type=No Target,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=check!!!
  • 249036 11/15/08 22:07 a-onl Start_Run_3505,Run_type=Cosmics,target_type=No Target,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=check!!!
  • 249035 11/15/08 22:06 a-onl End_of_Run_3504
  • 249034 11/15/08 22:04 a-onl Start_Run_3504,Run_type=Cosmics,target_type=No Target,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=check!!!
  • 249033 11/15/08 22:03 a-onl End_of_Run_3503
  • 249032 11/15/08 22:01 a-onl Start_Run_3503,Run_type=Cosmics,target_type=No Target,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=check!!!
  • 249031 11/15/08 22:01 a-onl End_of_Run_3502
  • 249030 11/15/08 21:56 a-onl Start_Run_3502,Run_type=Cosmics,target_type=No Target,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=check!!!
  • 249029 11/15/08 21:56 a-onl End_of_Run_3501
  • 249028 11/15/08 21:56 a-onl Start_Run_3501,Run_type=Cosmics,target_type=No Target,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=bb shower cosmics check
  • 249027 11/15/08 21:33 A. Camsonne Horizontal raster need to be replaced
  • 249026 11/15/08 21:04 R. Michaels re: raster repair
  • 249025 11/15/08 20:29 a-onl End_of_Run_3500
  • 249024 11/15/08 20:25 D. Dutta Raster problem
  • 249023 11/15/08 20:08 adaq Target Field and Oven Temps snapshot
  • 249022 11/15/08 20:08 adaq Target Cell Temps snapshot
  • 249021 11/15/08 20:08 adaq Radiation and Ozone snapshot
  • 249020 11/15/08 20:08 adaq Hall A Tools snapshot
  • 249019 11/15/08 20:08 adaq BigBox Magnet snapshot
  • 249018 11/15/08 20:08 adaq Beam Position and Current snapshot
  • 249017 11/15/08 20:08 adaq Beam Energy in A and C snapshot
  • 249016 11/15/08 20:01 Dutta Beam is back
  • 249015 11/15/08 19:37 Jin Fake Spin Signal ON till Production
  • 249014 11/15/08 19:06 adaq target: Spin Flip Report: Target Moved to Empty position. Auto Flip was already disabled, I fliped the target once, before moving it.
  • 249013 11/15/08 19:02 Chiranjib EPR correction
  • 249012 11/15/08 17:22 Chiranjib LHRS TDC spectra with T3 and T5
  • 249011 11/15/08 17:05 a-onl Start_Run_3500,Run_type=Cosmics,target_type=No Target,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=bb shower cosmics check
  • 249010 11/15/08 17:05 a-onl End_of_Run_3499
  • 249009 11/15/08 16:55 A. Camsonne Left Q1 tripped while ramping down right arm
  • 249008 11/15/08 16:04 adaq Beam Time Accounting
  • 249007 11/15/08 16:04 Lindgren Summary
  • 249006 11/15/08 15:03 Chiranjib EPR Calibration/ SPIN UP for TRANSVERSE + / POLARIZATION etc.
  • 249005 11/15/08 13:42 ycwang A1 should be OK for now
  • 249004 11/15/08 13:12 a-onl Start_Run_3499,Run_type=Cosmics,target_type=No Target,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=bb shower cosmics check
  • 249003 11/15/08 13:11 a-onl End_of_Run_3498
  • 249002 11/15/08 13:10 a-onl Start_Run_3498,Run_type=Cosmics,target_type=No Target,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=bb shower cosmics check
  • 249001 11/15/08 13:07 a-onl End_of_Run_3497
  • 249000 11/15/08 12:54 a-onl Start_Run_3497,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic test
  • 248999 11/15/08 12:49 a-onl End_of_Run_3496
  • 248998 11/15/08 12:48 Lindgren Hall A Tools
  • 248997 11/15/08 12:43 a-onl Start_Run_3496,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic test
  • 248996 11/15/08 12:43 a-onl End_of_Run_3495
  • 248995 11/15/08 12:39 Lindgren Hall A Entry
  • 248994 11/15/08 12:20 a-onl Start_Run_3495,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic test
  • 248993 11/15/08 12:19 D. Parno Scaler problem fixed (new Compton photon DAQ)
  • 248992 11/15/08 12:19 a-onl End_of_Run_3494
  • 248991 11/15/08 12:11 Lindgren Magnets
  • 248990 11/15/08 12:08 adaq Target Field and Oven Temps snapshot
  • 248989 11/15/08 12:08 adaq Target Cell Temps snapshot
  • 248988 11/15/08 12:08 adaq Radiation and Ozone snapshot
  • 248987 11/15/08 12:08 adaq Hall A Tools snapshot
  • 248986 11/15/08 12:08 adaq BigBox Magnet snapshot
  • 248985 11/15/08 12:08 adaq Beam Position and Current snapshot
  • 248984 11/15/08 12:07 adaq Beam Energy in A and C snapshot
  • 248983 11/15/08 12:04 a-onl Start_Run_3494,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic test
  • 248982 11/15/08 12:04 a-onl End_of_Run_3493
  • 248981 11/15/08 11:56 a-onl Start_Run_3493,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic test
  • 248980 11/15/08 11:56 a-onl End_of_Run_3492
  • 248979 11/15/08 11:38 a-onl Start_Run_3492,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic test
  • 248978 11/15/08 11:36 a-onl End_of_Run_3491
  • 248977 11/15/08 11:32 a-onl Start_Run_3491,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic test
  • 248976 11/15/08 11:29 Lindgren Hall A exit
  • 248975 11/15/08 11:14 A. Camsonne Compton IOC on hareboot4 port 3
  • 248974 11/15/08 11:01 A. Camsonne APC remote controlled plug
  • 248973 11/15/08 11:00 a-onl End_of_Run_3490
  • 248972 11/15/08 11:00 A. Camsonne Compton webcams
  • 248971 11/15/08 10:55 a-onl Start_Run_3490,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Pedestal RICH, TS=0 ns, Clock=1.8 MHz
  • 248970 11/15/08 10:54 a-onl End_of_Run_3489
  • 248969 11/15/08 10:53 adev BigBite WireChamber Measured Voltage change!
  • 248968 11/15/08 10:45 a-onl Start_Run_3489,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Pedestal RICH, TS=0 ns, Clock=1.8 MHz
  • 248967 11/15/08 10:32 a-onl End_of_Run_3488
  • 248966 11/15/08 10:31 a-onl Start_Run_3488,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic test RICH, TS=0 ns
  • 248965 11/15/08 10:29 a-onl End_of_Run_3487
  • 248964 11/15/08 10:29 a-onl Start_Run_3487,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic test RICH, TS=20 ns
  • 248963 11/15/08 10:29 xqian I think that we should be able to run 12uA
  • 248962 11/15/08 10:28 a-onl End_of_Run_3486
  • 248961 11/15/08 10:28 a-onl Start_Run_3486,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic test RICH, TS=change ns
  • 248960 11/15/08 10:26 a-onl End_of_Run_3485
  • 248959 11/15/08 10:25 a-onl Start_Run_3485,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic test RICH, TS=140-150 ns
  • 248958 11/15/08 10:25 a-onl End_of_Run_3484
  • 248957 11/15/08 10:24 a-onl Start_Run_3484,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic test RICH, TS=change ns
  • 248956 11/15/08 10:21 a-onl End_of_Run_3483
  • 248955 11/15/08 10:21 a-onl Start_Run_3483,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic test RICH, TS=50 ns
  • 248954 11/15/08 10:20 a-onl End_of_Run_3482
  • 248953 11/15/08 10:20 a-onl Start_Run_3482,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic test RICH, TS=change ns
  • 248952 11/15/08 10:17 a-onl End_of_Run_3481
  • 248951 11/15/08 10:17 a-onl Start_Run_3481,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic test RICH, TS=150 ns
  • 248950 11/15/08 10:16 a-onl End_of_Run_3480
  • 248949 11/15/08 10:15 a-onl Start_Run_3480,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic test RICH
  • 248948 11/15/08 10:11 a-onl End_of_Run_3479
  • 248947 11/15/08 10:09 a-onl Start_Run_3479,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic test RICH
  • 248946 11/15/08 10:07 Lindgren Hall A Entry
  • 248945 11/15/08 10:01 a-onl End_of_Run_3478
  • 248944 11/15/08 09:40 Lindgren Day shift begins
  • 248943 11/15/08 09:40 a-onl Start_Run_3478,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic test
  • 248942 11/15/08 09:39 a-onl End_of_Run_3477
  • 248941 11/15/08 08:54 A. Camsonne Remote controlled plug for compton electron detector
  • 248940 11/15/08 08:52 A. Camsonne experiment status
  • 248939 11/15/08 08:45 a-onl Start_Run_3477,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=A1 cosmic test after modifying pedestal
  • 248938 11/15/08 04:07 adaq Target Field and Oven Temps snapshot
  • 248937 11/15/08 04:07 adaq Target Cell Temps snapshot
  • 248936 11/15/08 04:07 adaq Radiation and Ozone snapshot
  • 248935 11/15/08 04:07 adaq Hall A Tools snapshot
  • 248934 11/15/08 04:07 adaq BigBox Magnet snapshot
  • 248933 11/15/08 04:07 adaq Beam Position and Current snapshot
  • 248932 11/15/08 04:07 adaq Beam Energy in A and C snapshot
  • 248931 11/15/08 03:25 a-onl End_of_Run_3476
  • 248930 11/15/08 03:23 a-onl Start_Run_3476,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH ped, clock = 1.800 MHz, TS=0 ns
  • 248929 11/15/08 03:20 a-onl End_of_Run_3475
  • 248928 11/15/08 03:19 a-onl Start_Run_3475,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH ped, clock = 1.800 MHz, TS=0 ns
  • 248927 11/15/08 03:12 a-onl End_of_Run_3474
  • 248926 11/15/08 03:11 a-onl Start_Run_3474,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH ped, clock = 1.800 MHz, TS=0 ns
  • 248925 11/15/08 03:01 a-onl End_of_Run_3473
  • 248924 11/15/08 03:00 a-onl Start_Run_3473,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH ped, clock = 1.800 MHz, TS=0 ns
  • 248923 11/15/08 02:58 a-onl Start_Run_3472,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH ped, clock = 1.800 MHz, TS=100 ns
  • 248922 11/15/08 02:57 a-onl End_of_Run_3471
  • 248921 11/15/08 02:56 a-onl Start_Run_3471,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH ped, clock = 1.800 MHz, TS=100 ns
  • 248920 11/15/08 02:53 a-onl End_of_Run_3470
  • 248919 11/15/08 02:51 a-onl Start_Run_3470,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH ped, clock = 1.800 MHz, TS=100 ns
  • 248918 11/15/08 02:50 a-onl End_of_Run_3469
  • 248917 11/15/08 02:48 a-onl Start_Run_3469,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH ped, clock = 1.800 MHz, TS=change ns
  • 248916 11/15/08 02:42 a-onl End_of_Run_3468
  • 248915 11/15/08 02:37 a-onl Start_Run_3468,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH ped, clock = 1.800 MHz, TS=200 ns
  • 248914 11/15/08 02:36 a-onl End_of_Run_3467
  • 248913 11/15/08 02:31 a-onl Start_Run_3467,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH ped, clock = 1.800 MHz, TS=change ns, all ADC
  • 248912 11/15/08 02:23 a-onl End_of_Run_3466
  • 248911 11/15/08 02:19 a-onl Start_Run_3466,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH ped, clock = 1.800 MHz, TS=100 ns, all ADC
  • 248910 11/15/08 02:18 a-onl End_of_Run_3465
  • 248909 11/15/08 02:18 a-onl Start_Run_3465,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH ped, clock = 1.800 MHz, TS=change ns
  • 248908 11/15/08 02:15 a-onl End_of_Run_3464
  • 248907 11/15/08 02:14 a-onl Start_Run_3464,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH ped, clock = 1.800 MHz, TS=200 ns
  • 248906 11/15/08 02:13 a-onl End_of_Run_3463
  • 248905 11/15/08 02:13 a-onl Start_Run_3463,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH ped, clock = 1.800 MHz, TS=change ns
  • 248904 11/15/08 02:12 a-onl End_of_Run_3462
  • 248903 11/15/08 02:11 a-onl Start_Run_3462,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH ped, clock = 1.800 MHz, TS=150 ns
  • 248902 11/15/08 02:10 a-onl End_of_Run_3461
  • 248901 11/15/08 02:10 a-onl Start_Run_3461,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH ped, clock = 1.800 MHz, TS=change ns
  • 248900 11/15/08 02:09 a-onl End_of_Run_3460
  • 248899 11/15/08 02:08 a-onl Start_Run_3460,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH ped, clock = 1.800 MHz, TS=100 ns
  • 248898 11/15/08 02:08 a-onl End_of_Run_3459
  • 248897 11/15/08 02:07 a-onl Start_Run_3459,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH ped, clock = 1.800 MHz, TS=change ns
  • 248896 11/15/08 02:07 a-onl End_of_Run_3458
  • 248895 11/15/08 02:06 a-onl Start_Run_3458,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH ped, clock = 1.800 MHz, TS=50 ns
  • 248894 11/15/08 02:05 a-onl End_of_Run_3457
  • 248893 11/15/08 02:05 a-onl Start_Run_3457,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH ped, clock = 1.800 MHz, TS=change ns
  • 248892 11/15/08 02:04 a-onl End_of_Run_3456
  • 248891 11/15/08 02:03 a-onl Start_Run_3456,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH ped, clock = 1.800 MHz, TS=0 ns
  • 248890 11/15/08 01:59 a-onl Start_Run_3455,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH ped, clock = 1.800 MHz, TS=change ns
  • 248889 11/15/08 01:58 a-onl End_of_Run_3454
  • 248888 11/15/08 01:57 a-onl Start_Run_3454,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH ped, clock = 1.800 MHz, TS=0 ns
  • 248887 11/15/08 01:57 a-onl End_of_Run_3453
  • 248886 11/15/08 01:54 a-onl Start_Run_3453,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH ped, clock = 2.004 MHz, TS=change ns
  • 248885 11/15/08 01:53 a-onl End_of_Run_3452
  • 248884 11/15/08 01:53 a-onl Start_Run_3452,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH ped, clock = 2.004 MHz, TS=130-140 ns
  • 248883 11/15/08 01:52 a-onl End_of_Run_3451
  • 248882 11/15/08 01:51 a-onl Start_Run_3451,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH ped, clock = 2.004 MHz, TS=change ns
  • 248881 11/15/08 01:47 a-onl End_of_Run_3450
  • 248880 11/15/08 01:44 a-onl Start_Run_3450,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH ped, clock = 2.004 MHz, TS=180-190 ns
  • 248879 11/15/08 01:44 a-onl End_of_Run_3449
  • 248878 11/15/08 01:44 a-onl Start_Run_3449,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH ped, clock = 2.004 MHz, TS=change ns
  • 248877 11/15/08 01:43 a-onl End_of_Run_3448
  • 248876 11/15/08 01:41 a-onl Start_Run_3448,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH ped, clock = 2.004 MHz, TS=160-170 ns
  • 248875 11/15/08 01:40 a-onl End_of_Run_3447
  • 248874 11/15/08 01:40 a-onl Start_Run_3447,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH ped, clock = 2.004 MHz, TS=change ns
  • 248873 11/15/08 01:38 a-onl End_of_Run_3446
  • 248872 11/15/08 01:36 a-onl Start_Run_3446,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH ped, clock = 2.004 MHz, TS=250 ns
  • 248871 11/15/08 01:35 a-onl End_of_Run_3445
  • 248870 11/15/08 01:35 a-onl Start_Run_3445,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH ped, clock = 2.004 MHz, TS=change ns
  • 248869 11/15/08 01:33 a-onl End_of_Run_3444
  • 248868 11/15/08 01:31 brads BB Cerenkov gain check results using LED (w/ snapshots)
  • 248867 11/15/08 01:31 a-onl Start_Run_3444,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH ped, clock = 2.004 MHz, TS=225 ns
  • 248866 11/15/08 01:30 a-onl End_of_Run_3443
  • 248865 11/15/08 01:30 a-onl Start_Run_3443,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH ped, clock = 2.004 MHz, TS=change ns
  • 248864 11/15/08 01:29 a-onl End_of_Run_3442
  • 248863 11/15/08 01:27 a-onl Start_Run_3442,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH ped, clock = 2.004 MHz, TS=200 ns
  • 248862 11/15/08 01:26 a-onl End_of_Run_3441
  • 248861 11/15/08 01:26 a-onl Start_Run_3441,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH ped, clock = 2.004 MHz, TS=change ns
  • 248860 11/15/08 01:25 a-onl End_of_Run_3440
  • 248859 11/15/08 01:23 a-onl Start_Run_3440,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH ped, clock = 2.004 MHz, TS=175 ns
  • 248858 11/15/08 01:22 a-onl End_of_Run_3439
  • 248857 11/15/08 01:22 a-onl Start_Run_3439,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH ped, clock = 2.004 MHz, TS=change ns
  • 248856 11/15/08 01:21 a-onl End_of_Run_3438
  • 248855 11/15/08 01:20 brads BB Cerenkov gain check results using LED
  • 248854 11/15/08 01:18 a-onl Start_Run_3438,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH ped, clock = 2.004 MHz, TS=150 ns
  • 248853 11/15/08 01:17 a-onl End_of_Run_3437
  • 248852 11/15/08 01:17 a-onl Start_Run_3437,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH ped, clock = 2.004 MHz, TS=change ns
  • 248851 11/15/08 01:16 a-onl End_of_Run_3436
  • 248850 11/15/08 01:14 a-onl Start_Run_3436,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH ped, clock = 2.004 MHz, TS=125 ns
  • 248849 11/15/08 01:13 a-onl End_of_Run_3435
  • 248848 11/15/08 01:13 a-onl Start_Run_3435,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH ped, clock = 2.004 MHz, TS=1000 ns
  • 248847 11/15/08 01:12 a-onl End_of_Run_3434
  • 248846 11/15/08 01:10 a-onl Start_Run_3434,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH ped, clock = 2.004 MHz, TS=1000 ns
  • 248845 11/15/08 01:09 a-onl End_of_Run_3433
  • 248844 11/15/08 01:08 a-onl Start_Run_3433,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH ped, clock = 2.004 MHz, TS=change ns
  • 248843 11/15/08 01:08 a-onl End_of_Run_3432
  • 248842 11/15/08 01:04 a-onl Start_Run_3432,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH ped, clock = 2.004 MHz, TS=50 ns
  • 248841 11/15/08 01:03 a-onl End_of_Run_3431
  • 248840 11/15/08 01:03 a-onl Start_Run_3431,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH ped, clock = 2.004 MHz, TS=change ns
  • 248839 11/15/08 01:02 a-onl End_of_Run_3430
  • 248838 11/15/08 01:00 a-onl Start_Run_3430,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH ped, clock = 2.004 MHz, TS=0 ns
  • 248837 11/15/08 00:59 a-onl End_of_Run_3429
  • 248836 11/15/08 00:55 a-onl Start_Run_3429,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH test, check ROC14 sequencer
  • 248835 11/15/08 00:52 a-onl End_of_Run_3428
  • 248834 11/15/08 00:50 a-onl Start_Run_3428,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH ped run, Clock=2 MHz, TS=0 ns
  • 248833 11/15/08 00:48 a-onl End_of_Run_3427
  • 248832 11/15/08 00:47 a-onl Start_Run_3427,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH ped run, Clock=2 MHz, TS=0 ns
  • 248831 11/15/08 00:46 a-onl End_of_Run_3426
  • 248830 11/15/08 00:42 a-onl Start_Run_3426,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH test
  • 248829 11/15/08 00:39 brads Cerenkov work finished, DAQ now being used by RICH guys
  • 248828 11/15/08 00:24 a-onl End_of_Run_3425
  • 248827 11/15/08 00:22 a-onl Start_Run_3425,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB cerenkov test
  • 248826 11/15/08 00:21 a-onl End_of_Run_3424
  • 248825 11/15/08 00:21 a-onl Start_Run_3424,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB cerenkov test
  • 248824 11/15/08 00:17 a-onl End_of_Run_3423
  • 248823 11/15/08 00:15 a-onl Start_Run_3423,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB cerenkov test
  • 248822 11/15/08 00:02 a-onl End_of_Run_3422
  • 248821 11/15/08 00:01 a-onl Start_Run_3422,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB cerenkov test
  • Friday

  • 248820 11/14/08 23:54 a-onl End_of_Run_3421
  • 248819 11/14/08 23:53 a-onl Start_Run_3421,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB cerenkov test
  • 248818 11/14/08 23:43 a-onl End_of_Run_3420
  • 248817 11/14/08 23:42 a-onl Start_Run_3420,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB cerenkov test
  • 248816 11/14/08 23:23 a-onl End_of_Run_3419
  • 248815 11/14/08 23:22 a-onl Start_Run_3419,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB cerenkov test
  • 248814 11/14/08 23:15 a-onl End_of_Run_3418
  • 248813 11/14/08 23:14 a-onl Start_Run_3418,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB cerenkov test
  • 248812 11/14/08 23:09 a-onl End_of_Run_3417
  • 248811 11/14/08 23:07 a-onl Start_Run_3417,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB cerenkov test
  • 248810 11/14/08 22:26 a-onl End_of_Run_3416
  • 248809 11/14/08 22:25 a-onl Start_Run_3416,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB cerenkov test
  • 248808 11/14/08 22:24 a-onl End_of_Run_3415
  • 248807 11/14/08 22:23 a-onl Start_Run_3415,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB cerenkov test
  • 248806 11/14/08 22:02 brads Run 3411 stopped for Cerenkov check out
  • 248805 11/14/08 22:01 a-onl End_of_Run_3414
  • 248804 11/14/08 21:59 a-onl Start_Run_3414,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB cerenkov test
  • 248803 11/14/08 21:37 a-onl End_of_Run_3413
  • 248802 11/14/08 21:35 a-onl Start_Run_3413,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB cerenkov test
  • 248801 11/14/08 21:35 a-onl End_of_Run_3412
  • 248800 11/14/08 21:25 a-onl Start_Run_3412,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB cerenkov test
  • 248799 11/14/08 21:16 Jin Spin Up Test @ Spin + State
  • 248798 11/14/08 20:56 brads Upper BigBite HV crate is beeping
  • 248797 11/14/08 20:10 a-onl Start_Run_3411,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB cosmic check
  • 248796 11/14/08 20:09 a-onl End_of_Run_3410
  • 248795 11/14/08 20:07 adaq Target Field and Oven Temps snapshot
  • 248794 11/14/08 20:07 adaq Target Cell Temps snapshot
  • 248793 11/14/08 20:07 adaq Radiation and Ozone snapshot
  • 248792 11/14/08 20:07 adaq Hall A Tools snapshot
  • 248791 11/14/08 20:07 adaq BigBox Magnet snapshot
  • 248790 11/14/08 20:07 adaq Beam Position and Current snapshot
  • 248789 11/14/08 20:07 adaq Beam Energy in A and C snapshot
  • 248788 11/14/08 19:46 Chiranjib/Ji QUICK optics enclosure inspection.
  • 248787 11/14/08 19:34 Yi Qiang AFP Loss Test
  • 248786 11/14/08 18:57 V. Sulkosky Re: ADC Spectra with T3 and T5
  • 248785 11/14/08 18:32 brads BigBite HV GUI now in VNC wrapper
  • 248784 11/14/08 18:26 adev BigBite WireChamber Measured Voltage change!
  • 248783 11/14/08 18:15 adev BigBite WireChamber Measured Voltage change!
  • 248782 11/14/08 18:15 D. Parno BCM output wiring change
  • 248781 11/14/08 17:53 D. Parno Compton VME crate moved to remote power control
  • 248780 11/14/08 17:44 Chiranjib Temperature Test on ASTRAL
  • 248779 11/14/08 17:22 Chiranjib/ Y EPR measurement
  • 248778 11/14/08 17:14 brads BB Cerenkov bad signal cable on PMT 20 fixed
  • 248777 11/14/08 17:00 a-onl Start_Run_3410,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB cosmic check
  • 248776 11/14/08 16:58 a-onl End_of_Run_3409
  • 248775 11/14/08 16:53 a-onl Start_Run_3409,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH pedestal
  • 248774 11/14/08 16:52 a-onl End_of_Run_3408
  • 248773 11/14/08 16:49 a-onl Start_Run_3408,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH pedestal
  • 248772 11/14/08 16:12 a-onl End_of_Run_3407
  • 248771 11/14/08 16:11 a-onl Start_Run_3407,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=A1 pedestal run
  • 248770 11/14/08 16:00 a-onl End_of_Run_3406
  • 248769 11/14/08 16:00 a-onl Start_Run_3406,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=ok ..this is the real ROC8 pedestal run
  • 248768 11/14/08 15:57 a-onl End_of_Run_3405
  • 248767 11/14/08 15:57 a-onl Start_Run_3405,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=ROC8 pedestal run
  • 248766 11/14/08 15:49 a-onl End_of_Run_3404
  • 248765 11/14/08 15:48 a-onl Start_Run_3404,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=A1 pedestal run
  • 248764 11/14/08 15:23 A. Camsonne Webcam
  • 248763 11/14/08 15:10 a-onl End_of_Run_3403
  • 248762 11/14/08 15:10 a-onl Start_Run_3403,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=A1 pedestal run
  • 248761 11/14/08 14:49 a-onl End_of_Run_3402
  • 248760 11/14/08 14:48 Kalyan Run 3402 ROC8 pedestals
  • 248759 11/14/08 14:47 a-onl Start_Run_3402,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB ROC8 pedestal run
  • 248758 11/14/08 14:45 a-onl End_of_Run_3401
  • 248757 11/14/08 14:42 a-onl Start_Run_3401,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=DAQ test
  • 248756 11/14/08 14:39 a-onl End_of_Run_3400
  • 248755 11/14/08 14:37 a-onl Start_Run_3400,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=DAQ test
  • 248754 11/14/08 14:33 a-onl End_of_Run_3399
  • 248753 11/14/08 14:17 A. Camsonne Compton electron detector interlock checked out
  • 248752 11/14/08 14:13 A. Camsonne Left Q1 in trouble
  • 248751 11/14/08 14:07 Higinbotham Right HRS Q2 and Dipole Test
  • 248750 11/14/08 13:53 a-onl Start_Run_3399,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=DAQ test
  • 248749 11/14/08 13:45 a-onl End_of_Run_3398
  • 248748 11/14/08 13:03 R. Michaels HAPPEX DAQ to "safe" mode
  • 248747 11/14/08 12:07 adaq Target Field and Oven Temps snapshot
  • 248746 11/14/08 12:07 adaq Target Cell Temps snapshot
  • 248745 11/14/08 12:06 adaq Radiation and Ozone snapshot
  • 248744 11/14/08 12:06 adaq Hall A Tools snapshot
  • 248743 11/14/08 12:06 adaq BigBox Magnet snapshot
  • 248742 11/14/08 12:06 adaq Beam Position and Current snapshot
  • 248741 11/14/08 12:06 adaq Beam Energy in A and C snapshot
  • 248740 11/14/08 10:13 V. Sulkosky Coefficients for Left Gas Cerenkov Updated
  • 248739 11/14/08 09:28 a-onl Start_Run_3398,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic check after BB retiming. BB Cherenkov HV is on
  • 248738 11/14/08 09:27 a-onl End_of_Run_3397
  • 248737 11/14/08 09:26 A. Camsonne HRS ramped down to 0 Amps
  • 248736 11/14/08 09:20 A. Camsonne proxy for accelerator website
  • 248735 11/14/08 08:57 ycwang The previous halog about A1 rates are without beam.
  • 248734 11/14/08 08:52 ycwang A1 rate with 8uA beam
  • 248733 11/14/08 08:30 A. Camsonne CHL still down
  • 248732 11/14/08 07:57 Yi Zhang B Cherenkov ADC in low HV, 5 pass beam
  • 248731 11/14/08 07:34 adev BigBite WireChamber Measured Voltage change!
  • 248730 11/14/08 07:32 adev BigBite WireChamber Measured Voltage change!
  • 248729 11/14/08 07:23 adev BigBite WireChamber Measured Current change!
  • 248728 11/14/08 07:21 a-onl Start_Run_3397,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic check after BB retiming. BB Cherenkov HV is on
  • 248727 11/14/08 07:20 a-onl End_of_Run_3396
  • 248726 11/14/08 07:13 adev BigBite WireChamber Measured Current change!
  • 248725 11/14/08 06:14 Yi Zhang BB Cherenkov ADC in low current, 5 pass beam
  • 248724 11/14/08 05:14 a-onl Start_Run_3396,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic check after BB retiming. BB Cherenkov HV is on
  • 248723 11/14/08 05:13 a-onl End_of_Run_3395
  • 248722 11/14/08 04:06 adaq Target Field and Oven Temps snapshot
  • 248721 11/14/08 04:06 adaq Target Cell Temps snapshot
  • 248720 11/14/08 04:06 adaq Radiation and Ozone snapshot
  • 248719 11/14/08 04:06 adaq Hall A Tools snapshot
  • 248718 11/14/08 04:06 adaq BigBox Magnet snapshot
  • 248717 11/14/08 04:06 adaq Beam Position and Current snapshot
  • 248716 11/14/08 04:06 adaq Beam Energy in A and C snapshot
  • 248715 11/14/08 04:02 Jin Target Activity of Nov-14 Owl
  • 248714 11/14/08 03:08 a-onl Start_Run_3395,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic check after BB retiming. BB Cherenkov HV is on
  • 248713 11/14/08 03:07 a-onl End_of_Run_3394
  • 248712 11/14/08 01:24 adev BigBite WireChamber Measured Voltage change!
  • 248711 11/14/08 01:24 a-onl Start_Run_3394,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic check after BB retiming. BB Cherenkov HV is on
  • 248710 11/14/08 01:22 a-onl End_of_Run_3393
  • 248709 11/14/08 01:14 adev BigBite WireChamber Measured Voltage change!
  • 248708 11/14/08 00:11 adev BigBite WireChamber Measured Voltage change!
  • 248707 11/14/08 00:11 brads BB Cerenkov work and todo list
  • 248706 11/14/08 00:06 dhamilton Status Update - no owl shift
  • 248705 11/14/08 00:01 Kalyan Change BB threshold to beam conditions
  • Thursday

  • 248704 11/13/08 23:49 adev BigBite GUI Stopped!
  • 248703 11/13/08 23:48 Jin Roc15 reboot, start run # 3393
  • 248702 11/13/08 23:46 adev BigBite GUI Stopped!
  • 248701 11/13/08 23:45 a-onl Start_Run_3393,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic check after BB retiming
  • 248700 11/13/08 23:20 Jin Bigbite -> 710A
  • 248699 11/13/08 22:32 a-onl Start_Run_3392,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic check after BB retiming
  • 248698 11/13/08 22:25 xqian I may interprete wrong
  • 248697 11/13/08 21:44 brads/kalyan BigBite retiming, potential gotchas with T6/T2
  • 248696 11/13/08 21:43 xqian Pion+ with the gap on the top of our acceptance
  • 248695 11/13/08 20:41 a-onl Start_Run_3391,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic check after BB retiming
  • 248694 11/13/08 20:23 a-onl End_of_Run_3390
  • 248693 11/13/08 20:15 a-onl Start_Run_3390,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB retiming check
  • 248692 11/13/08 20:06 adaq Target Field and Oven Temps snapshot
  • 248691 11/13/08 20:06 adaq Target Cell Temps snapshot
  • 248690 11/13/08 20:06 adaq Radiation and Ozone snapshot
  • 248689 11/13/08 20:06 adaq Hall A Tools snapshot
  • 248688 11/13/08 20:06 adaq BigBox Magnet snapshot
  • 248687 11/13/08 20:06 adaq Beam Position and Current snapshot
  • 248686 11/13/08 20:05 adaq Beam Energy in A and C snapshot
  • 248685 11/13/08 19:57 a-onl End_of_Run_3389
  • 248684 11/13/08 19:36 a-onl Start_Run_3389,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB retiming check
  • 248683 11/13/08 19:35 a-onl End_of_Run_3388
  • 248682 11/13/08 19:35 a-onl Start_Run_3388,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB retiming check
  • 248681 11/13/08 19:34 camsonne / c RICH activity during Nov/13/2008 access
  • 248680 11/13/08 19:01 a-onl End_of_Run_3387
  • 248679 11/13/08 19:01 a-onl Start_Run_3387,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB retiming check
  • 248678 11/13/08 18:59 a-onl Start_Run_3386,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB retiming check
  • 248677 11/13/08 18:59 a-onl End_of_Run_3386
  • 248676 11/13/08 18:49 Chiranjib/yi Fiber -comet combo
  • 248675 11/13/08 18:48 a-onl End_of_Run_3385
  • 248674 11/13/08 18:48 a-onl Start_Run_3385,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB retiming check
  • 248673 11/13/08 18:47 a-onl End_of_Run_3384
  • 248672 11/13/08 18:46 adev BigBite WireChamber Measured Voltage change!
  • 248671 11/13/08 18:46 adev BigBite WireChamber Measured Voltage change!
  • 248670 11/13/08 18:39 Yi/Jin/Chira Comet Laser Spectrum in Vertical Pumping
  • 248669 11/13/08 18:35 adev BigBite WireChamber Measured Voltage change!
  • 248668 11/13/08 18:21 a-onl Start_Run_3384,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB retiming check
  • 248667 11/13/08 18:20 a-onl End_of_Run_3383
  • 248666 11/13/08 18:15 a-onl Start_Run_3383,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB retiming check
  • 248665 11/13/08 18:15 a-onl End_of_Run_3382
  • 248664 11/13/08 17:54 a-onl Start_Run_3382,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB retiming check
  • 248663 11/13/08 17:54 a-onl End_of_Run_3381
  • 248662 11/13/08 17:46 ycwang A1 channel 03 come back
  • 248661 11/13/08 17:43 V. Sulkosky Strobe was Missing in Scalers
  • 248660 11/13/08 17:42 Yi Qiang Lasers are now UNLOCKED
  • 248659 11/13/08 17:35 V. Sulkosky Re: Changes in Left HRS Trigger Signals in DecData Class
  • 248658 11/13/08 17:28 a-onl Start_Run_3381,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB retiming check
  • 248657 11/13/08 17:17 Jack Left, not Right, Q1
  • 248656 11/13/08 17:10 LeRose Wien Angle Change
  • 248655 11/13/08 17:10 a-onl End_of_Run_3380
  • 248654 11/13/08 17:08 a-onl Start_Run_3380,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=rich pedestal, use ROC15 sequencer, clock at 1.775 MHz, ts=150 ns
  • 248653 11/13/08 17:03 Jack (Techs) Right spectrometer status
  • 248652 11/13/08 16:57 a-onl End_of_Run_3378
  • 248651 11/13/08 16:55 Jack (Techs) Right Q1 status
  • 248650 11/13/08 16:54 a-onl Start_Run_3378,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=rich pedestal, use ROC15 sequencer, clock at 1.775 MHz, ts=150 ns, clock cable panel 0 and 1 swap
  • 248649 11/13/08 16:51 a-onl End_of_Run_3377
  • 248648 11/13/08 16:50 a-onl Start_Run_3377,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=rich pedestal, use ROC15 sequencer, clock at 1.775 MHz, ts=150 ns, clock cable panel 0 and 1 swap
  • 248647 11/13/08 16:49 a-onl End_of_Run_3376
  • 248646 11/13/08 16:48 a-onl Start_Run_3376,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=rich pedestal, use ROC15 sequencer, clock at 1.75 MHz, ts=150 ns, clock cable panel 0 and 1 swap
  • 248645 11/13/08 16:47 a-onl End_of_Run_3375
  • 248644 11/13/08 16:46 a-onl Start_Run_3375,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=rich pedestal, use ROC15 sequencer, clock at 1.75 MHz, ts=150 ns, clock cable panel 0 and 1 swap
  • 248643 11/13/08 16:42 a-onl End_of_Run_3374
  • 248642 11/13/08 16:41 a-onl Start_Run_3374,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=rich pedestal, use ROC15 sequencer for clock/convert, clock at 1.75 MHz, ts=140 ns
  • 248641 11/13/08 16:39 a-onl End_of_Run_3373
  • 248640 11/13/08 16:38 a-onl Start_Run_3373,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=rich pedestal, use ROC15 sequencer for clock/convert, clock at 1.75 MHz, ts=140 ns
  • 248639 11/13/08 16:37 a-onl End_of_Run_3372
  • 248638 11/13/08 16:35 a-onl Start_Run_3372,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=rich pedestal, use ROC15 sequencer for clock/convert, clock at 1.75 MHz, ts=200 ns
  • 248637 11/13/08 16:34 a-onl End_of_Run_3371
  • 248636 11/13/08 16:34 a-onl Start_Run_3371,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=rich pedestal, use ROC15 sequencer for clock/convert, clock at 1.75 MHz, ts=200 ns
  • 248635 11/13/08 16:32 a-onl End_of_Run_3370
  • 248634 11/13/08 16:32 Techs Check list
  • 248633 11/13/08 16:31 a-onl Start_Run_3370,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=rich pedestal, use ROC15 sequencer for clock/convert, clock at 1.75 MHz, ts=150 ns
  • 248632 11/13/08 16:31 a-onl End_of_Run_3369
  • 248631 11/13/08 16:30 a-onl Start_Run_3369,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=rich pedestal, use ROC15 sequencer for clock/convert, clock at 1.75 MHz, ts=100 ns
  • 248630 11/13/08 16:28 a-onl End_of_Run_3368
  • 248629 11/13/08 16:27 a-onl Start_Run_3368,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=rich pedestal, use ROC15 sequencer for clock/convert, clock at 1.75 MHz, ts=50 ns
  • 248628 11/13/08 16:26 a-onl End_of_Run_3367
  • 248627 11/13/08 16:25 a-onl Start_Run_3367,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=rich pedestal, use ROC15 sequencer for clock/convert, clock at 1.75 MHz, ts=0 ns
  • 248626 11/13/08 16:24 a-onl End_of_Run_3366
  • 248625 11/13/08 16:22 a-onl Start_Run_3366,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=rich pedestal, use ROC15 sequencer for clock/convert, clock at 1 MHz, ts=500 ns
  • 248624 11/13/08 16:21 a-onl End_of_Run_3365
  • 248623 11/13/08 16:20 a-onl Start_Run_3365,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=rich pedestal, use ROC15 sequencer for clock/convert, clock at 1 MHz, ts=400 ns
  • 248622 11/13/08 16:19 a-onl End_of_Run_3364
  • 248621 11/13/08 16:18 a-onl Start_Run_3364,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=rich pedestal, use ROC15 sequencer for clock/convert, clock at 1 MHz, ts=300 ns
  • 248620 11/13/08 16:17 a-onl End_of_Run_3363
  • 248619 11/13/08 16:16 a-onl Start_Run_3363,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=rich pedestal, use ROC15 sequencer for clock/convert, clock at 1 MHz, ts=200 ns
  • 248618 11/13/08 16:15 a-onl End_of_Run_3362
  • 248617 11/13/08 16:14 a-onl Start_Run_3362,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=rich pedestal, use ROC15 sequencer for clock/convert, clock at 1 MHz, ts=100 ns
  • 248616 11/13/08 16:13 a-onl End_of_Run_3361
  • 248615 11/13/08 16:11 a-onl Start_Run_3361,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=rich pedestal, use ROC15 sequencer for clock/convert, clock at 1 MHz, ts=0
  • 248614 11/13/08 16:10 a-onl End_of_Run_3360
  • 248613 11/13/08 16:09 a-onl Start_Run_3360,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=rich pedestal, use ROC15 sequencer for clock/convert, clock at 1 MHz, ts moved to left
  • 248612 11/13/08 16:06 a-onl End_of_Run_3359
  • 248611 11/13/08 16:05 a-onl Start_Run_3359,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=rich pedestal, use ROC15 sequencer for clock/convert, clock at 1 MHz, ts moved to left
  • 248610 11/13/08 15:56 a-onl Start_Run_3358,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=rich pedestal, use ROC15 sequencer for clock/convert, clock at 1 MHz, ts moved to left
  • 248609 11/13/08 15:54 a-onl Start_Run_3357,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=rich pedestal, use ROC15 sequencer for clock/convert, clock at 1 MHz, ts moved to left
  • 248608 11/13/08 15:51 a-onl End_of_Run_3356
  • 248607 11/13/08 15:50 a-onl Start_Run_3356,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=rich pedestal, use ROC15 sequencer for clock/convert, clock at 1 MHz, ts moved to left
  • 248606 11/13/08 15:49 a-onl End_of_Run_3355
  • 248605 11/13/08 15:48 a-onl Start_Run_3355,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=rich pedestal, use ROC15 sequencer for clock/convert, clock at 1 MHz
  • 248604 11/13/08 15:46 a-onl End_of_Run_3354
  • 248603 11/13/08 15:46 a-onl Start_Run_3354,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=rich pedestal, use ROC15 sequencer for clock/convert, clock at 1 MHz
  • 248602 11/13/08 15:46 a-onl End_of_Run_3353
  • 248601 11/13/08 15:42 V. Sulkosky Re: RF Time Signal
  • 248600 11/13/08 15:37 a-onl Start_Run_3353,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=rich pedestal, use ROC15 sequencer for clock/convert, clock at 1 MHz
  • 248599 11/13/08 15:33 a-onl End_of_Run_3352
  • 248598 11/13/08 15:32 a-onl Start_Run_3352,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=rich pedestal, use ROC15 sequencer for clock/convert, ps1 really prescaled; clock at 1 MHz
  • 248597 11/13/08 15:31 a-onl End_of_Run_3351
  • 248596 11/13/08 15:29 a-onl Start_Run_3351,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=rich pedestal, use ROC15 sequencer for clock/convert, ps1 really prescaled; clock at 1 MHz
  • 248595 11/13/08 15:26 a-onl End_of_Run_3350
  • 248594 11/13/08 15:25 a-onl Start_Run_3350,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=rich pedestal, use ROC15 sequencer for clock/convert, ps1 really prescaled; clock at 1 MHz
  • 248593 11/13/08 15:24 a-onl End_of_Run_3349
  • 248592 11/13/08 15:16 a-onl Start_Run_3349,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=rich pedestal, use ROC15 sequencer for clock/convert, ps1 prescaled
  • 248591 11/13/08 15:16 a-onl End_of_Run_3348
  • 248590 11/13/08 15:14 a-onl Start_Run_3348,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=rich pedestal, use ROC15 sequencer for clock/convert
  • 248589 11/13/08 15:13 a-onl End_of_Run_3347
  • 248588 11/13/08 15:05 a-onl Start_Run_3347,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=rf test
  • 248587 11/13/08 15:05 a-onl End_of_Run_3346
  • 248586 11/13/08 15:02 a-onl Start_Run_3346,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH pedestal, Ts=-100 ns
  • 248585 11/13/08 15:01 a-onl End_of_Run_3345
  • 248584 11/13/08 14:57 a-onl Start_Run_3345,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH pedestal, Ts=max
  • 248583 11/13/08 14:57 a-onl End_of_Run_3344
  • 248582 11/13/08 14:54 a-onl Start_Run_3344,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=rf test
  • 248581 11/13/08 14:53 a-onl End_of_Run_3343
  • 248580 11/13/08 14:52 a-onl Start_Run_3343,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH pedestal
  • 248579 11/13/08 14:51 a-onl End_of_Run_3342
  • 248578 11/13/08 14:36 a-onl Start_Run_3342,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=rf test
  • 248577 11/13/08 14:35 a-onl End_of_Run_3341
  • 248576 11/13/08 14:34 a-onl Start_Run_3341,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=run pedestal
  • 248575 11/13/08 14:33 a-onl End_of_Run_3340
  • 248574 11/13/08 14:28 a-onl Start_Run_3340,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=rf test
  • 248573 11/13/08 14:27 a-onl End_of_Run_3339
  • 248572 11/13/08 14:24 a-onl Start_Run_3339,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Cosmics for A1
  • 248571 11/13/08 14:21 a-onl End_of_Run_3338
  • 248570 11/13/08 14:18 a-onl Start_Run_3338,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Cosmics for A1
  • 248569 11/13/08 14:18 a-onl End_of_Run_3337
  • 248568 11/13/08 14:16 a-onl Start_Run_3337,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Cosmics for A1
  • 248567 11/13/08 14:15 a-onl End_of_Run_3336
  • 248566 11/13/08 14:10 A. Camsonne Beam dowm for CHL work expected return of beam around 8PM today
  • 248565 11/13/08 14:08 a-onl Start_Run_3336,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Cosmics for A1
  • 248564 11/13/08 14:08 a-onl End_of_Run_3335
  • 248563 11/13/08 13:30 a-onl Start_Run_3335,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Cosmics for A1
  • 248562 11/13/08 13:29 a-onl End_of_Run_3334
  • 248561 11/13/08 13:21 Kalyan HV on two shower channels back to normal
  • 248560 11/13/08 12:47 a-onl Start_Run_3334,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Cosmics for A1
  • 248559 11/13/08 12:46 a-onl End_of_Run_3333
  • 248558 11/13/08 12:44 a-onl Start_Run_3333,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH Pedestal
  • 248557 11/13/08 12:43 a-onl End_of_Run_3332
  • 248556 11/13/08 12:15 a-onl Start_Run_3332,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic run for A1
  • 248555 11/13/08 12:14 a-onl End_of_Run_3331
  • 248554 11/13/08 12:05 adaq Target Field and Oven Temps snapshot
  • 248553 11/13/08 12:05 adaq Target Cell Temps snapshot
  • 248552 11/13/08 12:05 adaq Radiation and Ozone snapshot
  • 248551 11/13/08 12:05 adaq Hall A Tools snapshot
  • 248550 11/13/08 12:05 adaq BigBox Magnet snapshot
  • 248549 11/13/08 12:05 adaq Beam Position and Current snapshot
  • 248548 11/13/08 12:05 adaq Beam Energy in A and C snapshot
  • 248547 11/13/08 12:03 Yi Qiang BigBite Magnets is OFF
  • 248546 11/13/08 12:01 a-onl Start_Run_3331,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic run for A1
  • 248545 11/13/08 12:00 Yi Qiang Lasers are locked for putting in new collimator
  • 248544 11/13/08 11:58 a-onl End_of_Run_3330
  • 248543 11/13/08 11:57 Yi Qiang EPR Calibration
  • 248542 11/13/08 11:16 a-onl Start_Run_3330,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic run for A1
  • 248541 11/13/08 11:15 a-onl End_of_Run_3329
  • 248540 11/13/08 10:58 adaq target: Spin Flip Report: Target Report before EPR measurement
  • 248539 11/13/08 10:19 ycwang The base has problem for previous No.3 in A1
  • 248538 11/13/08 09:59 a-onl Start_Run_3329,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic run for A1
  • 248537 11/13/08 09:57 a-onl End_of_Run_3328
  • 248536 11/13/08 09:46 a-onl Start_Run_3328,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic run for A1
  • 248535 11/13/08 09:38 a-onl Start_Run_3326,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=pedestal run for A1
  • 248534 11/13/08 09:38 a-onl End_of_Run_3326
  • 248533 11/13/08 09:18 Kalyan BB Shower display for background
  • 248532 11/13/08 08:28 a-onl End_of_Run_3325
  • 248531 11/13/08 08:28 a-onl Start_Run_3325,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=pedestal run for A1
  • 248530 11/13/08 08:15 Z.-E. Mezian Shift Summary
  • 248529 11/13/08 08:05 a-onl Start_Run_3324,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Cosmic run
  • 248528 11/13/08 08:05 adaq Beam Time Accounting
  • 248527 11/13/08 08:00 Yi Zhang Auto spin flip off
  • 248526 11/13/08 07:52 a-onl End_of_Run_3323
  • 248525 11/13/08 07:38 a-onl Start_Run_3323,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production run at 8uA
  • 248524 11/13/08 07:37 a-onl End_of_Run_3322
  • 248523 11/13/08 06:55 a-onl Start_Run_3322,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production run at 8uA
  • 248522 11/13/08 06:54 a-onl End_of_Run_3321
  • 248521 11/13/08 06:52 a-onl Start_Run_3321,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production run at 8uA
  • 248520 11/13/08 06:49 a-onl End_of_Run_3320
  • 248519 11/13/08 06:20 Yi Zhang BB Cherenkov HV set
  • 248518 11/13/08 06:19 a-onl Start_Run_3320,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production run at 8uA
  • 248517 11/13/08 06:15 Yi Zhang trig rates while beam on (8uA)
  • 248516 11/13/08 06:08 a-onl End_of_Run_3319
  • 248515 11/13/08 06:03 a-onl Start_Run_3319,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=61,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production run at 8uA (W/. RICH detector)
  • 248514 11/13/08 05:55 a-onl End_of_Run_3318
  • 248513 11/13/08 05:52 a-onl Start_Run_3318,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=61,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production at 8uA
  • 248512 11/13/08 05:51 a-onl End_of_Run_3317
  • 248511 11/13/08 05:34 Chiranjib Target Screens II
  • 248510 11/13/08 05:32 Chiranjib Target Screens I
  • 248509 11/13/08 05:20 a-onl Start_Run_3317,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=62,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production at 8uA
  • 248508 11/13/08 05:19 a-onl End_of_Run_3316
  • 248507 11/13/08 04:54 Yi Zhang BB Cherenkov HV in channel L12.4
  • 248506 11/13/08 04:52 a-onl Start_Run_3316,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=63,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production at 8uA
  • 248505 11/13/08 04:51 a-onl End_of_Run_3315
  • 248504 11/13/08 04:45 Yi Zhang BB Cherenkov signal
  • 248503 11/13/08 04:44 xqian Update the BigBite library
  • 248502 11/13/08 04:20 Yi Zhang BB Cerenkov HV
  • 248501 11/13/08 04:17 xqian plots
  • 248500 11/13/08 04:16 a-onl Start_Run_3315,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=58,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production at 8uA
  • 248499 11/13/08 04:13 xqian BigBite Wire Chamber Background study
  • 248498 11/13/08 04:05 a-onl End_of_Run_3314
  • 248497 11/13/08 04:05 adaq Target Field and Oven Temps snapshot
  • 248496 11/13/08 04:05 adaq Target Cell Temps snapshot
  • 248495 11/13/08 04:05 adaq Radiation and Ozone snapshot
  • 248494 11/13/08 04:05 adaq Hall A Tools snapshot
  • 248493 11/13/08 04:05 adaq BigBox Magnet snapshot
  • 248492 11/13/08 04:05 adaq Beam Position and Current snapshot
  • 248491 11/13/08 04:05 adaq Beam Energy in A and C snapshot
  • 248490 11/13/08 03:33 a-onl Start_Run_3314,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=58,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production at 8uA
  • 248489 11/13/08 03:32 a-onl End_of_Run_3313
  • 248488 11/13/08 03:04 a-onl Start_Run_3313,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=58,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production at 8uA (After rebooting ROC8 & ROC9 crashed in the last run 3312)
  • 248487 11/13/08 02:38 a-onl Start_Run_3312,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=58,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production at 8uA
  • 248486 11/13/08 02:35 a-onl End_of_Run_3311
  • 248485 11/13/08 02:17 Yi Zhang BB Cerenkov HV tuning
  • 248484 11/13/08 01:24 a-onl Start_Run_3311,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=58,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production at 1uA to test BB CH. signal
  • 248483 11/13/08 01:19 a-onl End_of_Run_3310
  • 248482 11/13/08 01:10 Chiranjib Target RTDs
  • 248481 11/13/08 01:08 Chiranjib Spin flip screens and Laser spectrum
  • 248480 11/13/08 00:47 a-onl Start_Run_3310,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=58,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production at 8uA
  • 248479 11/13/08 00:46 a-onl End_of_Run_3309
  • 248478 11/13/08 00:38 Chiranjib S2 TDCs
  • 248477 11/13/08 00:26 Chiranjb S1 TDCs
  • 248476 11/13/08 00:11 a-onl Start_Run_3309,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=58,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production at 8uA
  • 248475 11/13/08 00:10 a-onl End_of_Run_3308
  • 248474 11/13/08 00:04 cisbani Swing shift summary
  • 248473 11/13/08 00:03 adaq Beam Time Accounting
  • Wednesday

  • 248472 11/12/08 23:50 Chiranjb S1 ADCs look okay
  • 248471 11/12/08 23:41 Yi Zhang golden run sets to 3306
  • 248470 11/12/08 23:22 a-onl Start_Run_3308,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=58,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production at 8uA
  • 248469 11/12/08 23:08 adaq target: Spin Flip Report: spin flip status
  • 248468 11/12/08 23:07 a-onl End_of_Run_3307
  • 248467 11/12/08 22:56 nelyubin laser_line_spectrum
  • 248466 11/12/08 22:55 a-onl Start_Run_3307,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=58,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production at 8uA
  • 248465 11/12/08 22:54 a-onl End_of_Run_3306
  • 248464 11/12/08 22:23 Chiranjib RUN 3304 is a check for LHRS
  • 248463 11/12/08 22:22 a-onl Start_Run_3306,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=58,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production at 8uA
  • 248462 11/12/08 22:21 adev BigBite WireChamber Measured Voltage change!
  • 248461 11/12/08 22:18 a-onl End_of_Run_3305
  • 248460 11/12/08 22:16 a-onl Start_Run_3305,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=58,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production at 8uA
  • 248459 11/12/08 22:14 a-onl End_of_Run_3304
  • 248458 11/12/08 21:52 cisbani HallA Tools
  • 248457 11/12/08 21:44 a-onl Start_Run_3304,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=58,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production at 8uA
  • 248456 11/12/08 21:41 a-onl End_of_Run_3303
  • 248455 11/12/08 21:38 a-onl Start_Run_3303,Run_type=Calibration,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=58,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH Pedestal at 8uA
  • 248454 11/12/08 21:34 a-onl End_of_Run_3302
  • 248453 11/12/08 21:31 a-onl Start_Run_3302,Run_type=Calibration,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=58,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH Pedestal at 2uA
  • 248452 11/12/08 21:27 a-onl End_of_Run_3301
  • 248451 11/12/08 21:26 cisbani spot raster 6.6x6.6 mcc
  • 248450 11/12/08 21:23 cisbani trigger rates, raster 6.6x6.6 mcc
  • 248449 11/12/08 21:23 a-onl Start_Run_3301,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=58,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=test run for spot++
  • 248448 11/12/08 21:21 a-onl End_of_Run_3300
  • 248447 11/12/08 21:17 a-onl Start_Run_3300,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=58,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=test run for spot++
  • 248446 11/12/08 21:16 cisbani trigger rates raster 6x6 mcc
  • 248445 11/12/08 21:15 a-onl End_of_Run_3299
  • 248444 11/12/08 21:14 cisbani spot, raster 5x5
  • 248443 11/12/08 21:12 a-onl Start_Run_3299,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=58,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=test run for spot++
  • 248442 11/12/08 21:11 cisbani trigger rates, raster 5x5 mcc
  • 248441 11/12/08 21:09 a-onl End_of_Run_3298
  • 248440 11/12/08 21:08 cisbani spot, raster 4x4 mcc
  • 248439 11/12/08 21:05 a-onl Start_Run_3298,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=58,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=test run for spot++
  • 248438 11/12/08 21:04 cisbani trigger rates, raster 4x4 mcc
  • 248437 11/12/08 21:02 a-onl End_of_Run_3297
  • 248436 11/12/08 21:01 cisbani spot, raster 3x3 mcc
  • 248435 11/12/08 20:57 a-onl Start_Run_3297,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=58,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=test run for spot++
  • 248434 11/12/08 20:56 cisbani trigger rates, raster 3x3 mcc
  • 248433 11/12/08 20:55 brads BB Cerenkov HVs turned on.
  • 248432 11/12/08 20:55 a-onl End_of_Run_3296
  • 248431 11/12/08 20:54 cisbani spot, raster 2x2 mcc
  • 248430 11/12/08 20:52 cisbani trigger rates raster 2x2 mcc
  • 248429 11/12/08 20:51 a-onl Start_Run_3296,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=58,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=test run for spot++
  • 248428 11/12/08 20:46 adev BigBite WireChamber Measured Voltage change!
  • 248427 11/12/08 20:41 E.Chudakov Moller results
  • 248426 11/12/08 20:05 adaq Target Field and Oven Temps snapshot
  • 248425 11/12/08 20:05 adaq Target Cell Temps snapshot
  • 248424 11/12/08 20:04 adaq Radiation and Ozone snapshot
  • 248423 11/12/08 20:04 adaq Hall A Tools snapshot
  • 248422 11/12/08 20:04 adaq BigBox Magnet snapshot
  • 248421 11/12/08 20:04 adaq Beam Position and Current snapshot
  • 248420 11/12/08 20:04 adaq Beam Energy in A and C snapshot
  • 248419 11/12/08 18:18 V. Sulkosky Left HRS Momentum 2.35 GeV/c
  • 248418 11/12/08 17:13 Moffit HAPPEX DAQ Working... Bring it back up.
  • 248417 11/12/08 16:31 Moffit Re: HAPPEX DAQ ROC23
  • 248416 11/12/08 16:31 V. Sulkosky Left Q1 Tripped off
  • 248415 11/12/08 16:12 a-onl End_of_Run_3295
  • 248414 11/12/08 16:04 averett Day Shift Summary
  • 248413 11/12/08 16:00 a-onl Start_Run_3295,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=58,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=test run for spot++
  • 248412 11/12/08 16:00 adaq Beam Time Accounting
  • 248411 11/12/08 15:57 bradshaw Spectra at Optics
  • 248410 11/12/08 15:55 nelyubin absorption spectrum
  • 248409 11/12/08 15:47 Sirish Compton Commissioning
  • 248408 11/12/08 15:26 A. Camsonne HAPPEX DAQ down
  • 248407 11/12/08 12:43 adev BigBite WireChamber Measured Voltage change!
  • 248406 11/12/08 12:41 Kalyan wirechamber HV trip
  • 248405 11/12/08 12:37 bradshaw Move Target to Optics
  • 248404 11/12/08 12:33 adev BigBite WireChamber Measured Voltage change!
  • 248403 11/12/08 12:04 adaq Target Field and Oven Temps snapshot
  • 248402 11/12/08 12:04 adaq Target Cell Temps snapshot
  • 248401 11/12/08 12:04 adaq Radiation and Ozone snapshot
  • 248400 11/12/08 12:04 adaq Hall A Tools snapshot
  • 248399 11/12/08 12:04 adaq BigBox Magnet snapshot
  • 248398 11/12/08 12:04 adaq Beam Position and Current snapshot
  • 248397 11/12/08 12:04 adaq Beam Energy in A and C snapshot
  • 248396 11/12/08 11:04 averett target to empty
  • 248395 11/12/08 11:04 adaq target: Spin Flip Report: Day Shift Target Report
  • 248394 11/12/08 10:55 a-onl End_of_Run_3294
  • 248393 11/12/08 10:48 Yi Qiang Target Laser IOC is back online
  • 248392 11/12/08 10:41 Kalyan coincidence time run 3291
  • 248391 11/12/08 10:37 A. Camsonne RICH configuration in
  • 248390 11/12/08 10:27 ycwang A1 rates with beam 8uA, previous one beam Trip
  • 248389 11/12/08 10:21 a-onl Start_Run_3294,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=58,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run; Two Arm Rich ;raster 3X3 mm
  • 248388 11/12/08 10:18 a-onl End_of_Run_3293
  • 248387 11/12/08 10:15 ycwang Channel 03 has less events than the neighbors
  • 248386 11/12/08 10:10 A. Camsonne HAPPEX DAQ crashing
  • 248385 11/12/08 10:09 A. Camsonne HAPPEX DAQ crashing
  • 248384 11/12/08 09:59 V. Sulkosky New Trigger Channels Working
  • 248383 11/12/08 09:58 camsonne / c RICH activity during Nov/11/2008 Tue access
  • 248382 11/12/08 09:35 a-onl Start_Run_3293,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=58,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production Run; Two Arm Rich ;raster 3X3 mm
  • 248381 11/12/08 09:30 a-onl End_of_Run_3292
  • 248380 11/12/08 09:29 a-onl Start_Run_3292,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=58,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Two Arm Rich ;raster 3X3 mm
  • 248379 11/12/08 09:02 a-onl End_of_Run_3291
  • 248378 11/12/08 08:46 ycwang A1 rate with beam 8uA
  • 248377 11/12/08 08:32 Jin Target Temperature screen shot
  • 248376 11/12/08 08:30 Jin Huang Target Activity Before Production Running
  • 248375 11/12/08 08:29 a-onl Start_Run_3291,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=58,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=production run; raster 3X3 mm
  • 248374 11/12/08 08:19 Z.E. Meziani Shift Summary
  • 248373 11/12/08 08:17 a-onl End_of_Run_3290
  • 248372 11/12/08 08:14 a-onl Start_Run_3290,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=58,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Spot++ test run with polarized 3He target with 5x6 mm raster (MCC unit)
  • 248371 11/12/08 08:14 Lamiaa Spot++ for run 3290 w/. 5x6 mm (MCC unit)
  • 248370 11/12/08 08:11 a-onl End_of_Run_3289
  • 248369 11/12/08 08:10 Lamiaa Spot++ for run 3289 w/. 4x5 mm raster (MCC unit)
  • 248368 11/12/08 08:08 a-onl Start_Run_3289,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=58,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Spot++ test run with polarized 3He target with 4x5 mm raster (MCC unit)
  • 248367 11/12/08 08:06 adaq Beam Time Accounting
  • 248366 11/12/08 08:06 adaq target: Spin Flip Report: Spin FLip Subsystem Switched On and Take Over The Target.
  • 248365 11/12/08 08:05 a-onl End_of_Run_3288
  • 248364 11/12/08 08:04 Lamiaa spot++ for run 3288 w/. 3x4 mm raster (MCC unit)
  • 248363 11/12/08 08:02 a-onl Start_Run_3288,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=58,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Spot++ test run with polarized 3He target with 3x4 mm raster (MCC unit)
  • 248362 11/12/08 07:59 a-onl End_of_Run_3287
  • 248361 11/12/08 07:58 Lamiaa Spot++ result for run 3287 w/. 2.5x3 mm raster (MCC unit)
  • 248360 11/12/08 07:53 a-onl Start_Run_3287,Run_type=Other,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=60,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Spot++ test run with polarized 3He target with 2.5x3 mm raster (MCC unit)
  • 248359 11/12/08 07:52 Yi Zhang maybe a timing problem?
  • 248358 11/12/08 07:38 a-onl End_of_Run_3286
  • 248357 11/12/08 07:06 a-onl Start_Run_3286,Run_type=Calibration,target_type=Reference Cell: H2,targetpol_poldir=Transverse-in-plane,targetpol_pol=-1,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Ref cell H2 calibration run with 3x4 mm raster MCC unit & BPMX is -1.5 mm
  • 248356 11/12/08 07:05 a-onl End_of_Run_3285
  • 248355 11/12/08 06:44 Yi Zhang Pion Rejector HV
  • 248354 11/12/08 06:37 a-onl Start_Run_3285,Run_type=Calibration,target_type=Reference Cell: H2,targetpol_poldir=Transverse-in-plane,targetpol_pol=-1,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Ref cell H2 calibration run with 3x4 mm raster MCC unit & BPMX is -1.5 mm
  • 248353 11/12/08 06:33 a-onl End_of_Run_3284
  • 248352 11/12/08 06:27 a-onl Start_Run_3284,Run_type=Calibration,target_type=Reference Cell: H2,targetpol_poldir=Transverse-in-plane,targetpol_pol=-1,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Ref cell H2 calibration run with 3x4 mm raster MCC unit & BPMX is -1.5 mm
  • 248351 11/12/08 06:23 a-onl End_of_Run_3283
  • 248350 11/12/08 06:20 a-onl Start_Run_3283,Run_type=Calibration,target_type=Reference Cell: H2,targetpol_poldir=Transverse-in-plane,targetpol_pol=-1,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Ref cell H2 calibration run with 3x4 mm raster MCC unit & BPMX is -1.5 mm
  • 248349 11/12/08 06:12 Jin Vac Ref Cell, Fill with H2 -> 100psig
  • 248348 11/12/08 06:06 a-onl End_of_Run_3282
  • 248347 11/12/08 05:38 a-onl Start_Run_3282,Run_type=Calibration,target_type=Reference Cell: N2,targetpol_poldir=Transverse-in-plane,targetpol_pol=-1,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Ref cell N2 calibration run with 3x4 mm raster MCC unit, BPMX is -1.5 mm & BB MWDC HV has been changed
  • 248346 11/12/08 05:34 a-onl End_of_Run_3281
  • 248345 11/12/08 05:18 Jin Huang Beam Tuning for Reference cell run
  • 248344 11/12/08 05:09 a-onl Start_Run_3281,Run_type=Calibration,target_type=Reference Cell: N2,targetpol_poldir=Transverse-in-plane,targetpol_pol=-1,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Ref cell N2 calibration run with 3x4 mm raster MCC unit, BPMX is -1.5 mm & BB MWDC HV has been changed
  • 248343 11/12/08 05:02 adev BigBite WireChamber Measured Voltage change!
  • 248342 11/12/08 04:56 a-onl End_of_Run_3279
  • 248341 11/12/08 04:53 a-onl Start_Run_3279,Run_type=Check out,target_type=Reference Cell: N2,targetpol_poldir=Transverse-in-plane,targetpol_pol=-1,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Spot++ test run with 3x4 mm raster MCC unit, Ref cell N2 @ 100 psig & BPMX is -1.5 mm
  • 248340 11/12/08 04:52 Yi Zhang BPM & raster size after 5 pass beam tuning.
  • 248339 11/12/08 04:50 Jin Ref Cell -> 100 psig N2
  • 248338 11/12/08 04:50 a-onl End_of_Run_3278
  • 248337 11/12/08 04:50 adaq target: Spin Flip Report: Fake Spin Signal ON, 20min/flip. True Target Pol. > 40%
  • 248336 11/12/08 04:47 a-onl Start_Run_3278,Run_type=Check out,target_type=Reference Cell: N2,targetpol_poldir=Transverse-in-plane,targetpol_pol=-1,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Spot++ test run with 3x4 mm raster MCC unit, Ref cell N2 @ 31 psig & BPMX is -1.5 mm
  • 248335 11/12/08 04:40 Jin Target -> Ref Cell
  • 248334 11/12/08 04:40 Jin Target -> BeO
  • 248333 11/12/08 04:38 a-onl End_of_Run_3277
  • 248332 11/12/08 04:35 a-onl Start_Run_3277,Run_type=Check out,target_type=Multi Carbon Optics/BeO,targetpol_poldir=Transverse-in-plane,targetpol_pol=-1,beam_e0=5.892,beam_raster=off,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Spot++ test run with 5x6 mm raster MCC unit & BPMX is -1.5 mm
  • 248331 11/12/08 04:25 a-onl End_of_Run_3276
  • 248330 11/12/08 04:25 Jin Target -> Carbon with Hole -> BeO
  • 248329 11/12/08 04:22 a-onl Start_Run_3276,Run_type=Check out,target_type=Single Carbon w/ Hole,targetpol_poldir=Transverse-in-plane,targetpol_pol=-1,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Spot++ test run with 5x6 mm raster MCC unit & BPMX is -1.4 mm
  • 248328 11/12/08 04:17 a-onl End_of_Run_3275
  • 248327 11/12/08 04:15 a-onl Start_Run_3275,Run_type=Check out,target_type=Reference Cell: Empty,targetpol_poldir=Transverse-in-plane,targetpol_pol=-1,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Spot++ test run with 5x6 mm raster MCC unit & BPMX is -1.4 mm
  • 248326 11/12/08 04:12 a-onl End_of_Run_3274
  • 248325 11/12/08 04:10 a-onl Start_Run_3274,Run_type=Check out,target_type=Reference Cell: Empty,targetpol_poldir=Transverse-in-plane,targetpol_pol=-1,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Spot++ test run with 4x5 mm raster MCC unit & BPMX is -1.4 mm
  • 248324 11/12/08 04:07 a-onl End_of_Run_3273
  • 248323 11/12/08 04:04 a-onl Start_Run_3273,Run_type=Check out,target_type=Reference Cell: Empty,targetpol_poldir=Transverse-in-plane,targetpol_pol=-1,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Spot++ test run with 4x5 mm raster MCC unit & BPMX is -1 mm
  • 248322 11/12/08 04:04 adaq Target Field and Oven Temps snapshot
  • 248321 11/12/08 04:04 adaq Target Cell Temps snapshot
  • 248320 11/12/08 04:04 adaq Radiation and Ozone snapshot
  • 248319 11/12/08 04:04 adaq Hall A Tools snapshot
  • 248318 11/12/08 04:04 adaq BigBox Magnet snapshot
  • 248317 11/12/08 04:03 adaq Beam Position and Current snapshot
  • 248316 11/12/08 04:03 adaq Beam Energy in A and C snapshot
  • 248315 11/12/08 04:01 a-onl End_of_Run_3272
  • 248314 11/12/08 04:00 a-onl Start_Run_3272,Run_type=Check out,target_type=Reference Cell: Empty,targetpol_poldir=Transverse-in-plane,targetpol_pol=-1,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Spot++ test run with 3x4 mm raster MCC unit & BPMX is -1 mm
  • 248313 11/12/08 03:57 a-onl End_of_Run_3271
  • 248312 11/12/08 03:53 Lamiaa Trigger screen shot. Empth ref, cell with raster 2.5x3 mm & BPMBX -1 mm
  • 248311 11/12/08 03:49 a-onl Start_Run_3271,Run_type=Check out,target_type=Reference Cell: Empty,targetpol_poldir=Transverse-in-plane,targetpol_pol=-1,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Spot++ test run with 2.5x3 mm raster MCC unit & BPMX is -1 mm
  • 248310 11/12/08 03:41 a-onl End_of_Run_3270
  • 248309 11/12/08 03:38 a-onl Start_Run_3270,Run_type=Check out,target_type=Reference Cell: Empty,targetpol_poldir=Transverse-in-plane,targetpol_pol=-1,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Spot++ test run with 2.5x3 mm raster MCC unit & BPMX is -0.5 mm
  • 248308 11/12/08 03:33 a-onl End_of_Run_3269
  • 248307 11/12/08 03:30 a-onl Start_Run_3269,Run_type=Check out,target_type=Reference Cell: Empty,targetpol_poldir=Transverse-in-plane,targetpol_pol=-1,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Spot++ test run with 4x5 raster MCC unit
  • 248306 11/12/08 03:29 a-onl End_of_Run_3268
  • 248305 11/12/08 03:28 adev BigBite WireChamber Measured Voltage change!
  • 248304 11/12/08 03:28 Jin Trigger ScreenShot, Empty Ref, 3x3 MCC raster, 2uA
  • 248303 11/12/08 03:27 a-onl Start_Run_3268,Run_type=Check out,target_type=Reference Cell: Empty,targetpol_poldir=Transverse-in-plane,targetpol_pol=-1,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Spot++ test run with 3x4 raster MCC unit
  • 248302 11/12/08 03:25 a-onl End_of_Run_3267
  • 248301 11/12/08 03:24 Lamiaa Spot++ for run 3267
  • 248300 11/12/08 03:21 a-onl Start_Run_3267,Run_type=Check out,target_type=Reference Cell: Empty,targetpol_poldir=Transverse-in-plane,targetpol_pol=-1,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Spot++ test run with 2.5x3 raster
  • 248299 11/12/08 03:20 Jin Trigger Screen Shot, Empty Ref Cell, 2.5x3 MCC Raster, 2uA
  • 248298 11/12/08 03:19 a-onl End_of_Run_3266
  • 248297 11/12/08 03:17 adev BigBite WireChamber Measured Voltage change!
  • 248296 11/12/08 03:12 Lamiaa Spot++ with Holy Carbon target
  • 248295 11/12/08 03:12 a-onl Start_Run_3266,Run_type=Check out,target_type=Reference Cell: Empty,targetpol_poldir=Transverse-in-plane,targetpol_pol=-1,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Spot++ test run with 2x2 raster
  • 248294 11/12/08 03:10 xqian How does retiming work, I will explain this with the left HRS data.
  • 248293 11/12/08 03:09 Jin For last screen shot, we asked for 3x3 to MCC
  • 248292 11/12/08 03:06 Jin Trigger Screen Shot @ ref cell 2x2 2uA
  • 248291 11/12/08 03:04 a-onl End_of_Run_3265
  • 248290 11/12/08 03:02 Jin Target -> Ref Cell
  • 248289 11/12/08 03:02 xqian Trigger timing problem is still there. Solved half.
  • 248288 11/12/08 02:41 a-onl Start_Run_3265,Run_type=Check out,target_type=Single Carbon w/ Hole,targetpol_poldir=Transverse-in-plane,targetpol_pol=-1,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Spot++ test run
  • 248287 11/12/08 02:36 Yi Zhang spot++ check
  • 248286 11/12/08 02:35 Jin Target -> Carbon with Hole
  • 248285 11/12/08 02:29 a-onl End_of_Run_3264
  • 248284 11/12/08 02:26 a-onl Start_Run_3264,Run_type=Optics,target_type=Multi Carbon Optics/BeO,targetpol_poldir=Transverse-in-plane,targetpol_pol=-1,beam_e0=5.892,beam_raster=off,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test run with Beo optics target
  • 248283 11/12/08 01:43 a-onl End_of_Run_3263
  • 248282 11/12/08 01:36 a-onl Start_Run_3263,Run_type=Optics,target_type=Multi Carbon Optics/BeO,targetpol_poldir=Transverse-in-plane,targetpol_pol=-1,beam_e0=5.892,beam_raster=off,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test run with Beo optics target
  • 248281 11/12/08 01:32 adev BigBite WireChamber Measured Voltage change!
  • 248280 11/12/08 01:22 adev BigBite WireChamber Measured Voltage change!
  • 248279 11/12/08 01:12 Jin Target -> BeO
  • 248278 11/12/08 00:43 a-onl End_of_Run_3262
  • 248277 11/12/08 00:40 a-onl Start_Run_3262,Run_type=Calibration,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH test
  • 248276 11/12/08 00:15 a-onl End_of_Run_3261
  • 248275 11/12/08 00:12 a-onl Start_Run_3261,Run_type=Calibration,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH Pedestal (in TwoArmRICH config, ps1 prescaled)
  • 248274 11/12/08 00:10 adaq Beam Time Accounting
  • 248273 11/12/08 00:09 cisbani Swing Shift Summary
  • 248272 11/12/08 00:08 a-onl End_of_Run_3260
  • 248271 11/12/08 00:05 a-onl Start_Run_3260,Run_type=Calibration,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH Pedestal (in RICHPedestal config, ps1 prescaled)
  • 248270 11/12/08 00:02 a-onl End_of_Run_3259
  • 248269 11/12/08 00:02 a-onl Start_Run_3259,Run_type=Calibration,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH Pedestal (in RICHPedestal config)
  • Tuesday

  • 248268 11/11/08 23:57 a-onl End_of_Run_3258
  • 248267 11/11/08 23:38 a-onl Start_Run_3258,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Cosmics run for Bigbite T0 optimization
  • 248266 11/11/08 23:35 a-onl End_of_Run_3257
  • 248265 11/11/08 23:35 a-onl Start_Run_3257,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test RICH deadtime (P3=P5=P8=1)
  • 248264 11/11/08 23:33 a-onl End_of_Run_3256
  • 248263 11/11/08 23:32 a-onl Start_Run_3256,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test RICH deadtime (P1=P3=P5=P8=1)
  • 248262 11/11/08 23:31 a-onl End_of_Run_3255
  • 248261 11/11/08 23:30 a-onl Start_Run_3255,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test RICH deadtime (P8=1)
  • 248260 11/11/08 23:30 a-onl End_of_Run_3254
  • 248259 11/11/08 23:08 V. Sulkosky Run 3253 Pedestal Run
  • 248258 11/11/08 23:05 Yi Qiang Laser IOC is offline
  • 248257 11/11/08 23:03 Jin/Chiranji Comet Laser Spectrum/Spin up start at ~23%
  • 248256 11/11/08 22:52 Yi Qiang 2 COMET Lasers are ON for transverse pumping
  • 248255 11/11/08 22:52 Yi Qiang 2 COMET Lasers are ON for transverse pumping
  • 248254 11/11/08 22:44 a-onl Start_Run_3254,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Cosmic run for Bigbite T0 Optimization
  • 248253 11/11/08 22:41 a-onl End_of_Run_3253
  • 248252 11/11/08 22:39 a-onl Start_Run_3253,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=pedestal run for left HRS
  • 248251 11/11/08 22:34 V. Sulkosky New Trigger Signals in 1877 Empty
  • 248250 11/11/08 22:31 V. Sulkosky Quick Access to fix RF Time Signals
  • 248249 11/11/08 22:23 brads Plugged BB Cerenkov PMT 5 into networked scope
  • 248248 11/11/08 22:18 a-onl End_of_Run_3252
  • 248247 11/11/08 22:15 a-onl Start_Run_3252,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Cosmics for the BigBite Wire Chamber T0 calibration
  • 248246 11/11/08 22:12 a-onl Start_Run_3251,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Cosmics for the BigBite Wire Chamber T0 calibration
  • 248245 11/11/08 22:12 a-onl End_of_Run_3251
  • 248244 11/11/08 22:08 a-onl Start_Run_3250,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Cosmics for the BigBite Wire Chamber T0 calibration
  • 248243 11/11/08 22:08 a-onl End_of_Run_3250
  • 248242 11/11/08 22:02 a-onl End_of_Run_3249
  • 248241 11/11/08 21:59 a-onl Start_Run_3249,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Cosmics for the BigBite Wire Chamber T0 calibration
  • 248240 11/11/08 21:59 a-onl End_of_Run_3249
  • 248239 11/11/08 21:53 adaq HallA Tools
  • 248238 11/11/08 21:50 brads re: RHRS VDC discriminator problem
  • 248237 11/11/08 21:07 V. Sulkosky Re: Checked RT on Left HRS
  • 248236 11/11/08 21:06 Jin Target Holding field messed up by wired Bigbite setting, Pol -> 10%, target -> empty, test pumping up with new laser
  • 248235 11/11/08 20:50 V. Sulkosky Added T1 and T3 to 1875
  • 248234 11/11/08 20:47 Kalyan DAQ activity today
  • 248233 11/11/08 20:41 V. Sulkosky Added TDC to ROC 4 in Slot 9 for Trigger Signals
  • 248232 11/11/08 20:19 a-onl Start_Run_3248,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Cosmics for the BigBite Wire Chamber T0 calibration
  • 248231 11/11/08 20:15 a-onl End_of_Run_3247
  • 248230 11/11/08 20:03 adaq Target Field and Oven Temps snapshot
  • 248229 11/11/08 20:03 adaq Target Cell Temps snapshot
  • 248228 11/11/08 20:03 adaq Radiation and Ozone snapshot
  • 248227 11/11/08 20:03 adaq Hall A Tools snapshot
  • 248226 11/11/08 20:03 adaq BigBox Magnet snapshot
  • 248225 11/11/08 20:03 adaq Beam Position and Current snapshot
  • 248224 11/11/08 20:03 adaq Beam Energy in A and C snapshot
  • 248223 11/11/08 20:01 a-onl Start_Run_3247,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmics
  • 248222 11/11/08 20:00 Jin Target -> He3 Cell
  • 248221 11/11/08 19:56 adaq target: Spin Flip Report: Software Update Done
  • 248220 11/11/08 19:39 A. Camsonne Webcam Compton
  • 248219 11/11/08 19:28 a-onl End_of_Run_3245
  • 248218 11/11/08 19:23 a-onl Start_Run_3245,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=TDC test
  • 248217 11/11/08 19:20 a-onl End_of_Run_3244
  • 248216 11/11/08 19:19 a-onl Start_Run_3244,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=TDC test
  • 248215 11/11/08 19:16 a-onl End_of_Run_3243
  • 248214 11/11/08 19:15 brads BB Cerenkov PMTs 5, 6,7 shielded from beamline
  • 248213 11/11/08 19:13 a-onl Start_Run_3243,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=TDC test
  • 248212 11/11/08 19:12 brads Two paddles set-up downstream of BB Shower detector
  • 248211 11/11/08 19:11 a-onl End_of_Run_3242
  • 248210 11/11/08 19:11 a-onl Start_Run_3242,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=TDC test
  • 248209 11/11/08 18:45 Yi Qiang Lasers are now UNLOCKED
  • 248208 11/11/08 18:40 a-onl Start_Run_3239,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=TDC test
  • 248207 11/11/08 18:36 a-onl End_of_Run_3238
  • 248206 11/11/08 18:36 a-onl Start_Run_3238,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=TDC test
  • 248205 11/11/08 18:34 a-onl End_of_Run_3237
  • 248204 11/11/08 18:34 a-onl Start_Run_3237,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=TDC test
  • 248203 11/11/08 18:33 a-onl End_of_Run_3236
  • 248202 11/11/08 18:28 ycwang The rate of A1 for different channnel
  • 248201 11/11/08 18:21 ycwang A1 channel 03 has signal after we change base
  • 248200 11/11/08 18:11 a-onl Start_Run_3236,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=bbite test
  • 248199 11/11/08 18:08 a-onl End_of_Run_3235
  • 248198 11/11/08 17:58 a-onl Start_Run_3235,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=
  • 248197 11/11/08 17:50 a-onl End_of_Run_3234
  • 248196 11/11/08 17:47 a-onl Start_Run_3234,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Left HRS and BigBite Scaler consistency check
  • 248195 11/11/08 17:47 a-onl End_of_Run_3233
  • 248194 11/11/08 17:32 a-onl Start_Run_3233,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB gates testing with T5
  • 248193 11/11/08 17:31 a-onl End_of_Run_3232
  • 248192 11/11/08 17:04 a-onl Start_Run_3232,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB gates testing
  • 248191 11/11/08 16:13 a-onl Start_Run_3231,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Cosmics check run
  • 248190 11/11/08 16:12 a-onl End_of_Run_3230
  • 248189 11/11/08 16:02 adaq Beam Time Accounting
  • 248188 11/11/08 16:01 Saha Shift Summary
  • 248187 11/11/08 15:44 Techs Checklist
  • 248186 11/11/08 15:19 a-onl Start_Run_3230,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Cosmics check run
  • 248185 11/11/08 15:18 a-onl End_of_Run_3229
  • 248184 11/11/08 15:17 a-onl Start_Run_3229,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RT check with T4
  • 248183 11/11/08 15:15 a-onl End_of_Run_3228
  • 248182 11/11/08 15:14 a-onl Start_Run_3228,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RT check with T4
  • 248181 11/11/08 15:13 a-onl End_of_Run_3227
  • 248180 11/11/08 15:11 a-onl Start_Run_3227,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RT check with T3
  • 248179 11/11/08 15:11 a-onl End_of_Run_3226
  • 248178 11/11/08 15:08 a-onl Start_Run_3226,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RT check with T5
  • 248177 11/11/08 15:07 a-onl End_of_Run_3225
  • 248176 11/11/08 15:06 a-onl Start_Run_3225,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RT check with T1
  • 248175 11/11/08 15:05 a-onl End_of_Run_3224
  • 248174 11/11/08 15:03 a-onl Start_Run_3224,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RT check with T3
  • 248173 11/11/08 15:02 a-onl End_of_Run_3223
  • 248172 11/11/08 15:00 a-onl Start_Run_3223,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RT check with T5
  • 248171 11/11/08 14:59 a-onl End_of_Run_3222
  • 248170 11/11/08 14:38 a-onl Start_Run_3222,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH cosmics
  • 248169 11/11/08 14:29 a-onl End_of_Run_3221
  • 248168 11/11/08 14:28 a-onl Start_Run_3221,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=T3 test for A1 and RICH, T1 not prescaled, L1A generates READEVT on T/H only
  • 248167 11/11/08 14:27 a-onl End_of_Run_3220
  • 248166 11/11/08 14:26 Jin Target Computer is rebooted
  • 248165 11/11/08 14:26 a-onl Start_Run_3220,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=T3 test for A1 and RICH, T1 not prescaled, L1A generates READEVT
  • 248164 11/11/08 14:25 a-onl End_of_Run_3219
  • 248163 11/11/08 14:21 a-onl Start_Run_3219,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=T3 test for A1 and RICH, T1 not prescaled
  • 248162 11/11/08 14:20 a-onl End_of_Run_3218
  • 248161 11/11/08 14:01 a-onl Start_Run_3218,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=T3 test for A1 and RICH, T1 not prescaled
  • 248160 11/11/08 13:42 a-onl End_of_Run_3216
  • 248159 11/11/08 13:41 a-onl Start_Run_3216,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=T3 test for A1 and RICH, T1 not prescaled
  • 248158 11/11/08 13:40 a-onl End_of_Run_3215
  • 248157 11/11/08 13:39 a-onl Start_Run_3215,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=T3 test for A1 and RICH
  • 248156 11/11/08 13:32 a-onl End_of_Run_3214
  • 248155 11/11/08 13:30 a-onl Start_Run_3214,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=T3 test for A1 and RICH pedestal
  • 248154 11/11/08 13:29 a-onl End_of_Run_3213
  • 248153 11/11/08 13:24 a-onl Start_Run_3213,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=T3 test for A1 and RICH, T1 not prescaled
  • 248152 11/11/08 13:23 a-onl End_of_Run_3212
  • 248151 11/11/08 13:07 ycwang Need a new base for A1
  • 248150 11/11/08 12:52 a-onl Start_Run_3212,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=T3 test for A1 and RICH, T1 not prescaled
  • 248149 11/11/08 12:50 a-onl End_of_Run_3211
  • 248148 11/11/08 12:49 ycwang After swap base 3 and base 4
  • 248147 11/11/08 12:48 a-onl Start_Run_3211,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=T3 test for A1 and RICH
  • 248146 11/11/08 12:48 a-onl End_of_Run_3210
  • 248145 11/11/08 12:35 a-onl Start_Run_3210,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=T3 test for A1 and RICH T1 not prescaled
  • 248144 11/11/08 12:34 a-onl End_of_Run_3209
  • 248143 11/11/08 12:32 a-onl Start_Run_3209,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=T3 test for A1 and RICH T1 not prescaled
  • 248142 11/11/08 12:31 a-onl End_of_Run_3208
  • 248141 11/11/08 12:30 adaq target: Spin Flip Report: ScreenShot before quit
  • 248140 11/11/08 12:03 adaq Target Field and Oven Temps snapshot
  • 248139 11/11/08 12:03 adaq Target Cell Temps snapshot
  • 248138 11/11/08 12:03 adaq Radiation and Ozone snapshot
  • 248137 11/11/08 12:03 adaq Hall A Tools snapshot
  • 248136 11/11/08 12:03 adaq BigBox Magnet snapshot
  • 248135 11/11/08 12:03 adaq Beam Position and Current snapshot
  • 248134 11/11/08 12:03 adaq Beam Energy in A and C snapshot
  • 248133 11/11/08 12:01 a-onl Start_Run_3208,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=T3 test for A1 and RICH
  • 248132 11/11/08 11:52 a-onl Start_Run_3207,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=T3 test for A1
  • 248131 11/11/08 11:51 a-onl End_of_Run_3207
  • 248130 11/11/08 11:40 a-onl Start_Run_3207,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=T3 test for A1
  • 248129 11/11/08 11:40 a-onl End_of_Run_3207
  • 248128 11/11/08 11:26 a-onl Start_Run_3207,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=T5 test with RT
  • 248127 11/11/08 11:26 a-onl End_of_Run_3207
  • 248126 11/11/08 11:14 a-onl Start_Run_3207,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=T3 test with RT
  • 248125 11/11/08 11:13 a-onl End_of_Run_3207
  • 248124 11/11/08 10:56 a-onl Start_Run_3207,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=rich test
  • 248123 11/11/08 10:51 a-onl End_of_Run_3207
  • 248122 11/11/08 10:25 xqian BigBite Shower Threshold to cosmics
  • 248121 11/11/08 10:24 a-onl Start_Run_3207,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=
  • 248120 11/11/08 10:22 a-onl End_of_Run_3206
  • 248119 11/11/08 10:21 Yi Qiang Field rotated to vertical for correction coil check
  • 248118 11/11/08 10:18 Yi Qiang EPR Correction
  • 248117 11/11/08 10:11 Yi Qiang EPR Correction
  • 248116 11/11/08 10:07 a-onl Start_Run_3206,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH test
  • 248115 11/11/08 10:02 a-onl End_of_Run_3205
  • 248114 11/11/08 10:01 a-onl Start_Run_3205,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH checks
  • 248113 11/11/08 09:41 Chiranjb/Yi EPR measurement
  • 248112 11/11/08 09:39 Yi Qiang Laser locked
  • 248111 11/11/08 09:30 a-onl Start_Run_3204,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BigBite Wire Chamber cosmics!!!
  • 248110 11/11/08 09:29 adev BigBite WireChamber Measured Voltage change!
  • 248109 11/11/08 09:21 A. Camsonne Powering BigBite Magnet down
  • 248108 11/11/08 09:01 Lingyan Zhu Laser spectrum
  • 248107 11/11/08 08:56 a-onl End_of_Run_3203
  • 248106 11/11/08 08:45 R. Michaels re: coinc. timing
  • 248105 11/11/08 08:40 xqian Just want to clarify
  • 248104 11/11/08 08:34 R. Michaels re: trigger timing issues
  • 248103 11/11/08 08:33 a-onl Start_Run_3203,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Start new run without HAPPEX waiting for ROC 26 on host happex7 to be rebooted
  • 248102 11/11/08 08:33 a-onl End_of_Run_3202
  • 248101 11/11/08 08:31 A. Camsonne Target work hazard analysis and plan
  • 248100 11/11/08 08:20 Kalyan/Xin Trying to understand coincidence timing spectra
  • 248099 11/11/08 08:18 xqian BigBite Wire Chamber T0 offset,
  • 248098 11/11/08 08:10 a-onl Start_Run_3202,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Start new run without HAPPEX waiting for ROC 26 on host happex7 to be rebooted
  • 248097 11/11/08 08:10 a-onl End_of_Run_3201
  • 248096 11/11/08 08:02 Yi Zhang field direction
  • 248095 11/11/08 08:02 Meziani Shift Summary
  • 248094 11/11/08 08:01 adaq Beam Time Accounting
  • 248093 11/11/08 07:56 Yi Zhang RTD & laser screen
  • 248092 11/11/08 07:52 Yi Zhang Laser spectrum
  • 248091 11/11/08 07:52 xqian BigBite Gas Cerenkov ADC signal
  • 248090 11/11/08 07:52 Yi Zhang Laser spectrum
  • 248089 11/11/08 07:50 adaq target: Spin Flip Report: owl shift on Nov.11
  • 248088 11/11/08 07:40 a-onl Start_Run_3201,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Start new run without HAPPEX waiting for ROC 26 on host happex7 to be rebooted
  • 248087 11/11/08 07:39 a-onl End_of_Run_3200
  • 248086 11/11/08 07:06 a-onl Start_Run_3200,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Start new run without HAPPEX waiting for ROC 26 on host happex7 to be rebooted
  • 248085 11/11/08 07:04 a-onl End_of_Run_3199
  • 248084 11/11/08 06:37 a-onl Start_Run_3199,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Start new run without HAPPEX waiting for ROC 26 on host happex7 to be rebooted
  • 248083 11/11/08 06:32 Lamiaa Update: ROC synchronization distributions for run 3197.
  • 248082 11/11/08 06:11 a-onl End_of_Run_3198
  • 248081 11/11/08 06:09 meziani Alarm handler acknowledment window not opening
  • 248080 11/11/08 06:08 a-onl Start_Run_3198,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Same HAPPEX run 30399
  • 248079 11/11/08 06:02 a-onl End_of_Run_3197
  • 248078 11/11/08 05:49 Lamiaa Empty ROC synchronization distributions for run 3196
  • 248077 11/11/08 05:35 a-onl Start_Run_3197,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Same HAPPEX run 30399
  • 248076 11/11/08 05:34 a-onl End_of_Run_3196
  • 248075 11/11/08 05:18 Moffit HAPPEX GUI: changed database
  • 248074 11/11/08 05:07 a-onl Start_Run_3196,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Same HAPPEX run 30399
  • 248073 11/11/08 05:06 a-onl End_of_Run_3195
  • 248072 11/11/08 04:24 a-onl Start_Run_3195,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Same HAPPEX run 30399
  • 248071 11/11/08 04:23 a-onl End_of_Run_3194
  • 248070 11/11/08 04:03 adaq Target Field and Oven Temps snapshot
  • 248069 11/11/08 04:03 adaq Target Cell Temps snapshot
  • 248068 11/11/08 04:02 adaq Radiation and Ozone snapshot
  • 248067 11/11/08 04:02 adaq Hall A Tools snapshot
  • 248066 11/11/08 04:02 adaq BigBox Magnet snapshot
  • 248065 11/11/08 04:02 adaq Beam Position and Current snapshot
  • 248064 11/11/08 04:02 adaq Beam Energy in A and C snapshot
  • 248063 11/11/08 03:49 a-onl Start_Run_3194,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Same HAPPEX run 30399
  • 248062 11/11/08 03:48 a-onl End_of_Run_3193
  • 248061 11/11/08 03:24 xqian Possible timing problem; some suggestions
  • 248060 11/11/08 03:14 a-onl Start_Run_3193,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Same HAPPEX run 30399
  • 248059 11/11/08 03:13 a-onl End_of_Run_3192
  • 248058 11/11/08 02:40 a-onl Start_Run_3192,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=New HAPPEX run 30399
  • 248057 11/11/08 02:33 a-onl Start_Run_3191,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=New HAPPEX run 30399
  • 248056 11/11/08 02:30 a-onl End_of_Run_3190
  • 248055 11/11/08 02:01 a-onl Start_Run_3190,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Same HAPPEX run 30398
  • 248054 11/11/08 02:00 a-onl End_of_Run_3189
  • 248053 11/11/08 01:25 a-onl Start_Run_3189,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Same HAPPEX run 30398
  • 248052 11/11/08 01:24 a-onl End_of_Run_3188
  • 248051 11/11/08 01:13 Chiranjib Vertex Z II
  • 248050 11/11/08 00:55 a-onl Start_Run_3188,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Same HAPPEX run 30398
  • 248049 11/11/08 00:53 a-onl End_of_Run_3187
  • 248048 11/11/08 00:29 Jin Huang Update of Online Analysis Scripts. Golden run -> Run # 3171. More Events to replay
  • 248047 11/11/08 00:22 a-onl Start_Run_3187,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Same HAPPEX run 30398, delay = 44 ns same as previous data taking
  • 248046 11/11/08 00:21 a-onl End_of_Run_3186
  • 248045 11/11/08 00:20 a-onl Start_Run_3186,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Same HAPPEX run 30398, delay = 44 ns
  • 248044 11/11/08 00:18 a-onl End_of_Run_3185
  • 248043 11/11/08 00:17 a-onl Start_Run_3185,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Same HAPPEX run 30398, delay = 35 ns
  • 248042 11/11/08 00:16 a-onl End_of_Run_3184
  • 248041 11/11/08 00:15 a-onl Start_Run_3184,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Same HAPPEX run 30398, delay = 0 ns
  • 248040 11/11/08 00:12 a-onl End_of_Run_3183
  • 248039 11/11/08 00:04 E. Cisbani Swing Shift summary
  • 248038 11/11/08 00:04 adaq Beam Time Accounting
  • 248037 11/11/08 00:03 Kalyan Target z plot
  • Monday

  • 248036 11/10/08 23:49 Chiranjib EPICS variable for the target ladder IN
  • 248035 11/10/08 23:42 a-onl Start_Run_3183,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Same HAPPEX run 30398
  • 248034 11/10/08 23:41 a-onl End_of_Run_3182
  • 248033 11/10/08 23:14 a-onl Start_Run_3182,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Same HAPPEX run 30398
  • 248032 11/10/08 23:13 a-onl End_of_Run_3181
  • 248031 11/10/08 22:41 a-onl Start_Run_3181,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Same HAPPEX run 30398
  • 248030 11/10/08 22:35 cisbani spot++ after long beam off
  • 248029 11/10/08 22:34 a-onl End_of_Run_3180
  • 248028 11/10/08 22:33 a-onl Start_Run_3180,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text= Same HAPPEX run 30398, spot test
  • 248027 11/10/08 22:29 Jin Target -> He3, Spin Flip recovers
  • 248026 11/10/08 22:11 a-onl End_of_Run_3179
  • 248025 11/10/08 21:51 Kalyan/Xin Trying to understand coincidence timing spectra
  • 248024 11/10/08 20:37 cisbani hallA tools
  • 248023 11/10/08 20:02 adaq Target Field and Oven Temps snapshot
  • 248022 11/10/08 20:02 adaq Target Cell Temps snapshot
  • 248021 11/10/08 20:02 adaq Radiation and Ozone snapshot
  • 248020 11/10/08 20:02 adaq Hall A Tools snapshot
  • 248019 11/10/08 20:02 adaq BigBox Magnet snapshot
  • 248018 11/10/08 20:02 adaq Beam Position and Current snapshot
  • 248017 11/10/08 20:02 adaq Beam Energy in A and C snapshot
  • 248016 11/10/08 19:47 Jin Target -> Optics
  • 248015 11/10/08 19:46 Sirish Compton Status
  • 248014 11/10/08 18:25 ycwang A1 all sum in 1 pass run # 2512 and 2 pass run #2777
  • 248013 11/10/08 18:22 ycwang A1 all sum in 1 pass run # 2512 and 2 pass run #2777
  • 248012 11/10/08 17:50 ycwang A1 rates Channel 10
  • 248011 11/10/08 17:48 ycwang A1 rates when Channel 4 HV turn off and on
  • 248010 11/10/08 17:42 a-onl Start_Run_3179,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic ray, Same HAPPEX run 30397. Rich pedestal config
  • 248009 11/10/08 17:39 a-onl End_of_Run_3178
  • 248008 11/10/08 17:39 a-onl Start_Run_3178,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic ray, Same HAPPEX run 30397. Rich pedestal config
  • 248007 11/10/08 17:34 a-onl Start_Run_3177,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic ray, Same HAPPEX run 30397. Rich pedestal config
  • 248006 11/10/08 17:33 a-onl End_of_Run_3176
  • 248005 11/10/08 17:33 a-onl Start_Run_3176,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic ray, Same HAPPEX run 30397. Rich pedestal config
  • 248004 11/10/08 17:29 a-onl End_of_Run_3175
  • 248003 11/10/08 17:28 a-onl Start_Run_3175,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text= cosmic ray, Same HAPPEX run 30397.
  • 248002 11/10/08 17:25 a-onl End_of_Run_3174
  • 248001 11/10/08 17:23 a-onl Start_Run_3174,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH pedestal run with cosmic (ps3=1) Same HAPPEX run 30397.
  • 248000 11/10/08 17:21 a-onl End_of_Run_3173
  • 247999 11/10/08 17:19 ycwang A1 rates in cosmic state
  • 247998 11/10/08 17:15 a-onl Start_Run_3173,Run_type=Test,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH pedestal run with cosmic Same HAPPEX run 30397.
  • 247997 11/10/08 17:11 Yi Qiang Auto spin flip paused for beam down
  • 247996 11/10/08 17:00 Jin MWDC hit drop highly depends on events range replayed
  • 247995 11/10/08 16:06 a-onl End_of_Run_3172
  • 247994 11/10/08 16:01 adaq Beam Time Accounting
  • 247993 11/10/08 16:00 Saha Day Shift Summary
  • 247992 11/10/08 15:55 a-onl Start_Run_3172,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Same HAPPEX run 30397.
  • 247991 11/10/08 15:54 a-onl End_of_Run_3171
  • 247990 11/10/08 15:49 narayan Screen Shot Pumping Laser Absorption Spectrum
  • 247989 11/10/08 15:27 a-onl Start_Run_3171,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Same HAPPEX run 30397.
  • 247988 11/10/08 15:24 a-onl End_of_Run_3170
  • 247987 11/10/08 15:18 Kalyan Re: LHRS aero A1 CH#3 dead
  • 247986 11/10/08 15:08 Kalyan Al left HV screen
  • 247985 11/10/08 14:47 yuan,Dimitri LHRS aero A1 CH#3 dead
  • 247984 11/10/08 14:41 a-onl Start_Run_3170,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Same HAPPEX run 30397.
  • 247983 11/10/08 14:37 a-onl End_of_Run_3169
  • 247982 11/10/08 14:09 a-onl Start_Run_3169,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Same HAPPEX run 30397.
  • 247981 11/10/08 14:02 a-onl End_of_Run_3168
  • 247980 11/10/08 13:31 yuan SPOT++ for Run 3167
  • 247979 11/10/08 13:30 a-onl Start_Run_3168,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Same HAPPEX run 30396.
  • 247978 11/10/08 13:29 a-onl End_of_Run_3167
  • 247977 11/10/08 13:16 yuan BPM and BCM charts
  • 247976 11/10/08 13:13 Saha Hall A Tools Screen
  • 247975 11/10/08 12:57 a-onl Start_Run_3167,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Same HAPPEX run 30396.
  • 247974 11/10/08 12:56 a-onl End_of_Run_3166
  • 247973 11/10/08 12:39 Kalyan shower channels
  • 247972 11/10/08 12:28 a-onl Start_Run_3166,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Same HAPPEX run 30396.
  • 247971 11/10/08 12:23 narayan target moved back
  • 247970 11/10/08 12:14 narayan target moved
  • 247969 11/10/08 12:10 narayan narayan
  • 247968 11/10/08 12:02 adaq Target Field and Oven Temps snapshot
  • 247967 11/10/08 12:02 adaq Target Cell Temps snapshot
  • 247966 11/10/08 12:02 adaq Radiation and Ozone snapshot
  • 247965 11/10/08 12:02 adaq Hall A Tools snapshot
  • 247964 11/10/08 12:02 adaq BigBox Magnet snapshot
  • 247963 11/10/08 12:02 adaq Beam Position and Current snapshot
  • 247962 11/10/08 12:01 adaq Beam Energy in A and C snapshot
  • 247961 11/10/08 11:57 a-onl End_of_Run_3165
  • 247960 11/10/08 11:42 adaq target: Spin Flip Report: General Shift Update
  • 247959 11/10/08 11:28 a-onl Start_Run_3165,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Same HAPPEX run 30396.
  • 247958 11/10/08 11:27 a-onl End_of_Run_3164
  • 247957 11/10/08 11:04 Kalyan increased HV on three shower blocks
  • 247956 11/10/08 10:58 a-onl Start_Run_3164,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Same HAPPEX run 30396. HV for Shower channel changed
  • 247955 11/10/08 10:52 a-onl End_of_Run_3163
  • 247954 11/10/08 10:20 a-onl Start_Run_3163,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Same HAPPEX run 30396
  • 247953 11/10/08 10:19 a-onl End_of_Run_3162
  • 247952 11/10/08 10:13 Kalyan BB wirechamber signals improved and back to normal again
  • 247951 11/10/08 09:49 a-onl Start_Run_3162,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Same HAPPEX run 30395
  • 247950 11/10/08 09:48 a-onl End_of_Run_3161
  • 247949 11/10/08 09:18 a-onl Start_Run_3161,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Same HAPPEX run 30395
  • 247948 11/10/08 09:17 a-onl End_of_Run_3160
  • 247947 11/10/08 09:03 A. Camsonne Checked Argon/Ethane bottle
  • 247946 11/10/08 08:34 adaq target: Spin Flip Report: for test
  • 247945 11/10/08 08:32 a-onl Start_Run_3160,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.892,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Same HAPPEX run 30395
  • 247944 11/10/08 08:30 a-onl End_of_Run_3159
  • 247943 11/10/08 08:08 adaq Beam Time Accounting
  • 247942 11/10/08 08:07 Yi Zhang Owl shift summary
  • 247941 11/10/08 07:56 a-onl Start_Run_3159,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Same HAPPEX run 30395
  • 247940 11/10/08 07:55 a-onl End_of_Run_3158
  • 247939 11/10/08 07:43 ycwang something different with Bigbite wire Chamber RUN# 3152 3157
  • 247938 11/10/08 07:30 ycwang pions and coincidence pions
  • 247937 11/10/08 07:27 a-onl Start_Run_3158,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=New HAPPEX run 30395
  • 247936 11/10/08 07:25 a-onl End_of_Run_3157
  • 247935 11/10/08 07:09 ycwang can gas Cherenkov make very clear electrons cut?
  • 247934 11/10/08 07:06 Yi Zhang ROC9 rebooted
  • 247933 11/10/08 06:54 a-onl Start_Run_3157,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Same HAPPEX run 30394
  • 247932 11/10/08 06:52 a-onl End_of_Run_3156
  • 247931 11/10/08 06:51 Yi Zhang target-> 3He
  • 247930 11/10/08 06:50 a-onl Start_Run_3156,Run_type=Check out,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=raster check
  • 247929 11/10/08 06:31 a-onl Start_Run_3155,Run_type=Check out,target_type=Multi Carbon Optics/BeO,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=raster check
  • 247928 11/10/08 06:28 a-onl Start_Run_3154,Run_type=Check out,target_type=Multi Carbon Optics/BeO,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=raster check
  • 247927 11/10/08 06:25 R. Michaels re: Qasy in BCM
  • 247921 11/10/08 05:40 xqian About the BCM
  • 247920 11/10/08 05:38 xqian It seems that there are a lot of "faked" trigger generated by the shower
  • 247919 11/10/08 05:30 Yi Zhang Move target to BeO
  • 247918 11/10/08 05:21 L. El Fassi Update : Hot Spot In BB Shower (run 3153 seems okay)
  • 247917 11/10/08 05:14 Yi Zhang stop spin flip for beam trip
  • 247916 11/10/08 05:12 a-onl End_of_Run_3153
  • 247915 11/10/08 05:00 L. El Fassi Update : Hot Spot In BB Shower (run 3152)
  • 247914 11/10/08 04:46 a-onl Start_Run_3153,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Same HAPPEX run 30394
  • 247913 11/10/08 04:44 a-onl End_of_Run_3152
  • 247912 11/10/08 04:14 a-onl Start_Run_3152,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Same HAPPEX run 30394
  • 247911 11/10/08 04:12 a-onl End_of_Run_3151
  • 247910 11/10/08 04:01 adaq Target Field and Oven Temps snapshot
  • 247909 11/10/08 04:01 adaq Target Cell Temps snapshot
  • 247908 11/10/08 04:01 adaq Radiation and Ozone snapshot
  • 247907 11/10/08 04:01 adaq Hall A Tools snapshot
  • 247906 11/10/08 04:01 adaq BigBox Magnet snapshot
  • 247905 11/10/08 04:01 adaq Beam Position and Current snapshot
  • 247904 11/10/08 04:01 adaq Beam Energy in A and C snapshot
  • 247903 11/10/08 03:40 a-onl Start_Run_3151,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Same HAPPEX run 30394
  • 247902 11/10/08 03:38 a-onl End_of_Run_3150
  • 247901 11/10/08 03:33 Jin Laser Speacturm ScreenShot
  • 247900 11/10/08 03:33 Jin Oven Temp and Laser ScreenShot
  • 247899 11/10/08 03:33 a-onl Start_Run_3150,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=New HAPPEX run 30394
  • 247898 11/10/08 03:32 adaq target: Spin Flip Report: Spin Flip Working OK
  • 247897 11/10/08 03:30 a-onl End_of_Run_3149
  • 247896 11/10/08 03:16 Moffit Re: Feedback not working... Beam Charge asymmetry still small.
  • 247895 11/10/08 03:01 a-onl Start_Run_3149,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=45,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=, HAPPEX run 30393
  • 247894 11/10/08 02:59 a-onl End_of_Run_3148
  • 247893 11/10/08 02:57 Moffit HAPPEX Feedback: now using "bcm3"
  • 247892 11/10/08 02:48 Jin Update : Hot Spot In BB Shower
  • 247891 11/10/08 02:47 Moffit HAPPEX Feedback: not working
  • 247890 11/10/08 02:31 a-onl Start_Run_3148,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=46,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=, HAPPEX run 30393
  • 247889 11/10/08 02:29 a-onl End_of_Run_3147
  • 247888 11/10/08 02:21 adaq Hot spot in BB shower
  • 247887 11/10/08 02:03 brads Target cell temperature changes a lot depending on ladder postion (w/o beam)
  • 247886 11/10/08 02:01 a-onl Start_Run_3147,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=46,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=, HAPAX run 30393
  • 247885 11/10/08 01:53 a-onl End_of_Run_3146
  • 247884 11/10/08 01:51 brads BB ADC gates walking by 40ns+
  • 247883 11/10/08 01:28 a-onl Start_Run_3146,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=46,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=, HAPAX run 30393
  • 247882 11/10/08 01:25 a-onl End_of_Run_3145
  • 247881 11/10/08 00:45 a-onl Start_Run_3145,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=46.5,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=, HAPAX run 303933
  • 247880 11/10/08 00:42 a-onl End_of_Run_3144
  • 247879 11/10/08 00:25 a-onl Start_Run_3144,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=46.5,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=, HAPAX run 303933
  • 247878 11/10/08 00:12 a-onl End_of_Run_3143
  • 247877 11/10/08 00:01 adaq Beam Time Accounting
  • Sunday

  • 247876 11/09/08 23:53 xqian Swing Day Shift Summary
  • 247875 11/09/08 23:50 adev BigBite WireChamber Measured Voltage change!
  • 247874 11/09/08 23:44 a-onl Start_Run_3143,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=46.5,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=, HAPAX run 303933
  • 247873 11/09/08 23:42 adaq target: Spin Flip Report: Flip Paused due to beam off. Actived again before production
  • 247872 11/09/08 23:40 a-onl End_of_Run_3142
  • 247871 11/09/08 23:39 a-onl Start_Run_3142,Run_type=Check out,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=46.5,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=check out BB WC out, HAPAX run 303933
  • 247870 11/09/08 23:35 a-onl End_of_Run_3141
  • 247869 11/09/08 23:34 Zhihong Ye Target moved to Polarized He3
  • 247868 11/09/08 23:30 a-onl Start_Run_3141,Run_type=Check out,target_type=Multi Carbon Optics/BeO,targetpol_poldir=Transverse-in-plane,targetpol_pol=46.5,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=check out BB WC out, HAPAX run 303933
  • 247867 11/09/08 23:29 adev BigBite WireChamber Measured Voltage change!
  • 247866 11/09/08 23:19 Zhihong Ye Target moved to Optics
  • 247865 11/09/08 23:16 brads Target cell temperature fluctuations during beam off period
  • 247864 11/09/08 23:10 brads Power cycled ROC8, ROC9, and ROC10
  • 247863 11/09/08 23:07 a-onl End_of_Run_3140
  • 247862 11/09/08 23:06 a-onl Start_Run_3140,Run_type=Check out,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=46.5,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=production by the Left VDC may not recover completely, HAPAX run 30392,T3 delay set to 44 ns instead of 60 ns previously
  • 247861 11/09/08 23:05 a-onl End_of_Run_3139
  • 247860 11/09/08 23:05 a-onl Start_Run_3139,Run_type=Check out,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=46.5,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=production by the Left VDC may not recover completely, HAPAX run 30392,T3 delay set to 44 ns instead of 60 ns previously
  • 247859 11/09/08 22:56 brads Coda rate troubleshooting
  • 247858 11/09/08 22:51 a-onl End_of_Run_3138
  • 247857 11/09/08 22:50 a-onl Start_Run_3138,Run_type=Check out,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=46.5,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=production by the Left VDC may not recover completely, HAPAX run 30392,T3 delay set to 44 ns instead of 60 ns previously
  • 247856 11/09/08 22:39 a-onl End_of_Run_3137
  • 247855 11/09/08 22:39 a-onl Start_Run_3137,Run_type=Check out,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=46.5,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=production by the Left VDC may not recover completely, HAPAX run 30392,T3 delay set to 44 ns instead of 60 ns previously
  • 247854 11/09/08 22:34 a-onl End_of_Run_3136
  • 247853 11/09/08 22:33 a-onl Start_Run_3136,Run_type=Check out,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=46.5,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=production by the Left VDC may not recover completely, HAPAX run 30392,T3 delay set to 44 ns instead of 60 ns previously
  • 247852 11/09/08 22:27 a-onl End_of_Run_3135
  • 247851 11/09/08 22:26 a-onl Start_Run_3135,Run_type=Check out,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=46.5,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=production by the Left VDC may not recover completely, HAPAX run 30392,T3 delay set to 44 ns instead of 60 ns previously
  • 247850 11/09/08 22:25 a-onl End_of_Run_3134
  • 247849 11/09/08 22:25 a-onl Start_Run_3134,Run_type=Check out,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=46.5,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=production by the Left VDC may not recover completely, HAPAX run 30392,T3 delay set to 44 ns instead of 60 ns previously
  • 247848 11/09/08 22:23 a-onl End_of_Run_3133
  • 247847 11/09/08 22:22 a-onl Start_Run_3133,Run_type=Check out,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=46.5,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=production by the Left VDC may not recover completely, HAPAX run 30392,T3 delay set to 44 ns instead of 60 ns previously
  • 247846 11/09/08 22:22 a-onl Start_Run_3132,Run_type=Check out,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=46.5,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=production by the Left VDC may not recover completely, HAPAX run 30392,T3 delay set to 44 ns instead of 60 ns previously
  • 247845 11/09/08 22:22 a-onl End_of_Run_3132
  • 247844 11/09/08 22:17 brads Coda rate slowed to a crawl
  • 247843 11/09/08 22:11 Jin High Dead Time!
  • 247842 11/09/08 22:06 a-onl Start_Run_3131,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=46.5,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=production by the Left VDC may not recover completely, HAPAX run 30392,T3 delay set to 44 ns instead of 60 ns previously
  • 247841 11/09/08 22:00 Kalyan Re; ADC gate
  • 247840 11/09/08 21:59 a-onl End_of_Run_3130
  • 247839 11/09/08 21:59 a-onl Start_Run_3130,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=46.5,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=production by the Left VDC may not recover completely, HAPAX run 30392,T3 delay set to 44 ns instead of 60 ns previously
  • 247838 11/09/08 21:58 a-onl End_of_Run_3129
  • 247837 11/09/08 21:45 K Wang halla tools
  • 247836 11/09/08 21:37 A. Camsonne ADC gate
  • 247835 11/09/08 21:29 a-onl Start_Run_3129,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=46.5,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=production by the Left VDC may not recover completely, HAPAX run 30392,T3 delay set to 44 ns instead of 60 ns previously
  • 247834 11/09/08 21:27 a-onl End_of_Run_3128
  • 247833 11/09/08 21:04 xqian BCM coefficient
  • 247832 11/09/08 21:00 xqian It seems that only S1 ADC TDC got problems
  • 247831 11/09/08 20:59 xqian ROC3 and 4 rebooted by themselves
  • 247830 11/09/08 20:59 a-onl Start_Run_3128,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=46.5,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=production by the Left VDC may not recover completely, HAPAX run 30392,T3 delay set to 44 ns instead of 60 ns previously
  • 247829 11/09/08 20:55 a-onl Start_Run_3127,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=46.5,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=production by the Left VDC may not recover completely, HAPAX run 30392,T3 delay set to 44 ns instead of 60 ns previously
  • 247828 11/09/08 20:54 a-onl End_of_Run_3126
  • 247827 11/09/08 20:48 brads Please log the new BCM coeffs
  • 247826 11/09/08 20:30 brads ADC gates should not be affected by coin. timing change(!?)
  • 247825 11/09/08 20:24 a-onl Start_Run_3126,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=46.5,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=production by the Left VDC may not recover completely, HAPAX run 30392,T3 delay set to 44 ns instead of 60 ns previously
  • 247824 11/09/08 20:23 a-onl End_of_Run_3125
  • 247823 11/09/08 20:11 brads FFB on standby, should it be?
  • 247822 11/09/08 20:01 adaq Target Field and Oven Temps snapshot
  • 247821 11/09/08 20:01 adaq Target Cell Temps snapshot
  • 247820 11/09/08 20:01 adaq Radiation and Ozone snapshot
  • 247819 11/09/08 20:01 adaq Hall A Tools snapshot
  • 247818 11/09/08 20:01 adaq BigBox Magnet snapshot
  • 247817 11/09/08 20:01 adaq Beam Position and Current snapshot
  • 247816 11/09/08 20:01 adaq Beam Energy in A and C snapshot
  • 247815 11/09/08 19:55 xqian Timing may screwed up in this run
  • 247814 11/09/08 19:51 A. Camsonne T3 delay change
  • 247813 11/09/08 19:50 a-onl Start_Run_3125,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=46.5,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=production by the Left VDC may not recover completely, HAPAX run 30392,T3 delay set to 44 ns instead of 60 ns previously
  • 247812 11/09/08 19:44 Zhihong Ye Laser Pumping Spectrum Screen-shot
  • 247811 11/09/08 19:44 a-onl End_of_Run_3124
  • 247810 11/09/08 19:41 Jin Target RTD & Laser ScreenShot @ Beam Comes
  • 247809 11/09/08 19:23 xqian BCM Number Changed
  • 247808 11/09/08 19:17 a-onl Start_Run_3124,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=46.5,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=production byt the Left VDC may not recover completely, HAPAX run 30392
  • 247807 11/09/08 19:15 a-onl End_of_Run_3123
  • 247806 11/09/08 19:13 a-onl Start_Run_3123,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=46.5,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=production byt the Left VDC may not recover completely
  • 247805 11/09/08 19:08 adev BigBite WireChamber Measured Voltage change!
  • 247804 11/09/08 19:05 a-onl End_of_Run_3122
  • 247803 11/09/08 19:04 a-onl Start_Run_3122,Run_type=Check out,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=44.5,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=check up
  • 247802 11/09/08 19:03 a-onl End_of_Run_3121
  • 247801 11/09/08 19:02 a-onl Start_Run_3121,Run_type=Check out,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=44.5,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=check up
  • 247800 11/09/08 18:57 Zhihong Ye Target moved to He3, and Spin Flip
  • 247799 11/09/08 18:48 Zhihong Ye Target moved to Optics
  • 247798 11/09/08 18:36 adev BigBite WireChamber Measured Voltage change!
  • 247797 11/09/08 18:33 a-onl End_of_Run_3120
  • 247796 11/09/08 18:32 a-onl Start_Run_3120,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=44.5,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic test! HAPPEX DAQ run#=30391.
  • 247795 11/09/08 18:31 Zhihong Ye Target Moved to Empty Target
  • 247794 11/09/08 18:19 a-onl End_of_Run_3119
  • 247793 11/09/08 18:04 A. Camsonne Experiment status update : Hall A dump issue
  • 247792 11/09/08 17:38 a-onl Start_Run_3119,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=44.5,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic test! HAPPEX DAQ run#=30391.
  • 247791 11/09/08 17:31 a-onl End_of_Run_3118
  • 247790 11/09/08 17:23 Kalyan Controlled access activity
  • 247789 11/09/08 17:18 Yi Qiang Please following the special procedures to move target from/to He3 position
  • 247788 11/09/08 16:50 A. Camsonne Requested controlled access while no beam
  • 247787 11/09/08 16:48 a-onl Start_Run_3118,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=44.5,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic test! HAPPEX DAQ run#=30391.
  • 247786 11/09/08 16:45 a-onl End_of_Run_3117
  • 247785 11/09/08 16:08 Zhihong Ye Target Moved to Optics w/o hole
  • 247784 11/09/08 16:05 a-onl Start_Run_3117,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=44.5,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic test! HAPPEX DAQ run#=30391.
  • 247783 11/09/08 16:03 a-onl End_of_Run_3116
  • 247782 11/09/08 16:02 adaq Beam Time Accounting
  • 247781 11/09/08 16:02 Yi Day Shift Summary
  • 247780 11/09/08 15:57 a-onl Start_Run_3116,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=44.5,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic test! HAPPEX DAQ run#=30391.
  • 247779 11/09/08 15:56 a-onl End_of_Run_3115
  • 247778 11/09/08 15:52 chench summary of effciency low peak position moving
  • 247777 11/09/08 15:50 a-onl Start_Run_3115,Run_type=Cosmics,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=44.5,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic test! HAPPEX DAQ run#=30391.
  • 247776 11/09/08 15:49 Chiranjib Hall A Tools
  • 247775 11/09/08 15:32 adaq target: Spin Flip Report: Day shift report, transverse pumping
  • 247774 11/09/08 15:19 Katich Hall A Tools Screen - Day shift
  • 247773 11/09/08 15:17 V. Sulkosky Pion Cuts in Left HRS
  • 247772 11/09/08 15:00 Chiranjib NO BEAM for almost more than 2 hours
  • 247771 11/09/08 14:43 V. Sulkosky Pion Rejector Layer 2
  • 247770 11/09/08 14:39 V. Sulkosky Pion Rejector Layer 1 Plots
  • 247769 11/09/08 14:31 a-onl End_of_Run_3114
  • 247768 11/09/08 14:31 a-onl Start_Run_3114,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=44.5,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=junk test! HAPPEX DAQ run#=30391.
  • 247767 11/09/08 14:14 a-onl End_of_Run_3113
  • 247766 11/09/08 14:14 a-onl Start_Run_3113,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=44.5,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=junk test! HAPPEX DAQ run#=30390.
  • 247765 11/09/08 13:08 a-onl End_of_Run_3112
  • 247764 11/09/08 13:06 chench run3112 left arm VDC efficiency
  • 247763 11/09/08 12:45 a-onl Start_Run_3112,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=44.5,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=production run! HAPPEX DAQ run#=30389.
  • 247762 11/09/08 12:44 a-onl End_of_Run_3111
  • 247761 11/09/08 12:38 A. Camsonne Expect drop in efficiency because of gas shut down
  • 247760 11/09/08 12:28 chench left arm efficiency for run3111
  • 247759 11/09/08 12:11 Jack (Techs) Gas system shutdown
  • 247758 11/09/08 12:09 a-onl Start_Run_3111,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=44.5,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=production run! HAPPEX DAQ run#=30389.
  • 247757 11/09/08 12:08 a-onl End_of_Run_3110
  • 247756 11/09/08 12:03 Kalyan Plot of reaction point at target
  • 247755 11/09/08 12:01 Yi Qiang A huge drop of left VDC effiencicy
  • 247754 11/09/08 12:01 adaq Target Field and Oven Temps, snapshot
  • 247753 11/09/08 12:01 adaq Target Cell Temps snapshot
  • 247752 11/09/08 12:00 adaq Radiation and Ozone snapshot
  • 247751 11/09/08 12:00 adaq Hall A Tools snapshot
  • 247750 11/09/08 12:00 adaq BigBox Magnet snapshot
  • 247749 11/09/08 12:00 adaq Beam Position and Current snapshot
  • 247748 11/09/08 12:00 adaq Beam Energy in A and C snapshot
  • 247747 11/09/08 11:55 chench run3110 left arm efficiency
  • 247746 11/09/08 11:35 a-onl Start_Run_3110,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=44.5,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=production run! HAPPEX DAQ run#=30389.
  • 247745 11/09/08 11:35 a-onl End_of_Run_3109
  • 247744 11/09/08 11:32 chench run3109 VDC efficiency
  • 247743 11/09/08 11:18 Katich Correction to lumi 2 HV
  • 247742 11/09/08 11:14 Katich Reset Lumi HVs
  • 247741 11/09/08 11:12 chench run3108 left arm efficiency
  • 247740 11/09/08 11:07 a-onl Start_Run_3109,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=44.5,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=production run! HAPPEX DAQ run#=30389.
  • 247739 11/09/08 11:06 a-onl End_of_Run_3108
  • 247738 11/09/08 10:39 a-onl Start_Run_3108,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=44.5,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=production run! HAPPEX DAQ run#=30389.
  • 247737 11/09/08 10:38 a-onl End_of_Run_3107
  • 247736 11/09/08 10:08 a-onl Start_Run_3107,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=44.5,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=production run! HAPPEX DAQ run#=30387.
  • 247735 11/09/08 10:06 a-onl End_of_Run_3106
  • 247734 11/09/08 10:00 a-onl Start_Run_3106,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=44.5,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=production run! HAPPEX DAQ run#=30387.
  • 247733 11/09/08 09:59 a-onl End_of_Run_3105
  • 247732 11/09/08 09:27 a-onl Start_Run_3105,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=Transverse-in-plane,targetpol_pol=44.5,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=production run! HAPPEX DAQ run#=30386.
  • 247731 11/09/08 09:26 a-onl End_of_Run_3104
  • 247730 11/09/08 09:25 A. Camsonne Gas bottle swapped
  • 247729 11/09/08 09:00 V. Sulkosky Start-of-Run tcl Modified
  • 247728 11/09/08 08:58 a-onl Start_Run_3104,Run_type=Production,target_type=Polarized 3He,targetpol_poldir=,targetpol_pol=,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=production run! HAPPEX DAQ run#=30385.
  • 247727 11/09/08 08:55 a-onl End_of_Run_3103
  • 247726 11/09/08 08:25 a-onl Start_Run_3103,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=production run! HAPPEX DAQ run#=30385.
  • 247725 11/09/08 08:23 a-onl End_of_Run_3102
  • 247724 11/09/08 08:15 cisbani Owl Shift Summary
  • 247723 11/09/08 08:14 adaq Beam Time Accounting
  • 247722 11/09/08 07:55 a-onl Start_Run_3102,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=production run! HAPPEX DAQ run#=30385.
  • 247721 11/09/08 07:53 a-onl End_of_Run_3101
  • 247720 11/09/08 07:39 yawei dead time sometimes up to 44,usually less than 11
  • 247719 11/09/08 07:23 a-onl Start_Run_3101,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=production run! HAPPEX DAQ run#=30384.
  • 247718 11/09/08 07:22 a-onl End_of_Run_3100
  • 247717 11/09/08 06:56 a-onl Start_Run_3100,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=production run! HAPPEX DAQ run#=30384.
  • 247716 11/09/08 06:55 a-onl End_of_Run_3099
  • 247715 11/09/08 06:46 cisbani Argon and CO2 gas pressure level low
  • 247714 11/09/08 06:29 a-onl Start_Run_3099,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=production run! HAPPEX DAQ run#=30384.
  • 247713 11/09/08 06:28 a-onl End_of_Run_3098
  • 247712 11/09/08 06:01 a-onl Start_Run_3098,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=production run! HAPPEX DAQ run#=30384.
  • 247711 11/09/08 06:01 a-onl End_of_Run_3097
  • 247710 11/09/08 05:43 ycwang Coincidence peak in LHRS in Run3086
  • 247709 11/09/08 05:28 a-onl Start_Run_3097,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=production run! HAPPEX DAQ run#=30384.
  • 247708 11/09/08 05:27 a-onl End_of_Run_3096
  • 247707 11/09/08 05:22 Jin Huang RTD & Laser ScreenShot
  • 247706 11/09/08 05:22 Jin Huang Laser absorption Screenshot
  • 247705 11/09/08 05:20 adaq target: Spin Flip Report: Running OK.
  • 247704 11/09/08 05:15 Jin Huang Physics Replay released, Run#3092 -> Golden Run
  • 247703 11/09/08 04:57 a-onl Start_Run_3096,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=production run! HAPPEX DAQ run#=30384.
  • 247702 11/09/08 04:56 a-onl End_of_Run_3095
  • 247701 11/09/08 04:33 cisbani spot++ run 3094
  • 247700 11/09/08 04:30 a-onl Start_Run_3095,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=production run! HAPPEX DAQ run#=30384.
  • 247699 11/09/08 04:29 a-onl End_of_Run_3094
  • 247698 11/09/08 04:12 cisbani HallA Tools
  • 247697 11/09/08 04:10 cisbani bbps1:2001 HVOFF alarm, run 3094
  • 247696 11/09/08 04:00 adaq Target Field and Oven Temps, snapshot
  • 247695 11/09/08 04:00 adaq Target Cell Temps snapshot
  • 247694 11/09/08 04:00 adaq Radiation and Ozone snapshot
  • 247693 11/09/08 04:00 adaq Hall A Tools snapshot
  • 247692 11/09/08 04:00 adaq BigBox Magnet snapshot
  • 247691 11/09/08 04:00 adaq Beam Position and Current snapshot
  • 247690 11/09/08 04:00 adaq Beam Energy in A and C snapshot
  • 247689 11/09/08 03:57 a-onl Start_Run_3094,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=production run! HAPPEX DAQ run#=30384.
  • 247688 11/09/08 03:56 a-onl End_of_Run_3093
  • 247687 11/09/08 03:28 a-onl Start_Run_3093,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=production run! HAPPEX DAQ run#=30384.
  • 247686 11/09/08 03:26 a-onl End_of_Run_3092
  • 247685 11/09/08 02:59 a-onl Start_Run_3092,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=production run! HAPPEX DAQ run#=30383.
  • 247684 11/09/08 02:56 a-onl End_of_Run_3091
  • 247683 11/09/08 02:55 a-onl Start_Run_3091,Run_type=Test,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH pedstal run! HAPPEX DAQ run#=30383.
  • 247682 11/09/08 02:48 a-onl Start_Run_3090,Run_type=Test,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH pedstal run! HAPPEX DAQ run#=30383.
  • 247681 11/09/08 02:41 a-onl End_of_Run_3089
  • 247680 11/09/08 02:14 a-onl Start_Run_3089,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production! HAPPEX DAQ run#=30383.
  • 247679 11/09/08 02:13 a-onl End_of_Run_3088
  • 247678 11/09/08 01:46 a-onl Start_Run_3088,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production! HAPPEX DAQ run#=30383.
  • 247677 11/09/08 01:46 a-onl End_of_Run_3087
  • 247676 11/09/08 01:09 a-onl Start_Run_3087,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production! HAPPEX DAQ run#=30383.
  • 247675 11/09/08 01:08 a-onl End_of_Run_3086
  • 247674 11/09/08 00:32 a-onl Start_Run_3086,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production! HAPPEX DAQ run#=30383.
  • 247673 11/09/08 00:31 a-onl End_of_Run_3085
  • 247672 11/09/08 00:02 adaq Beam Time Accounting
  • 247671 11/09/08 00:01 Kebin Wang swing shift summary Nov 8
  • 247670 11/09/08 00:00 a-onl Start_Run_3085,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production! HAPPEX DAQ run#=30383.
  • Saturday

  • 247669 11/08/08 23:58 a-onl End_of_Run_3084
  • 247668 11/08/08 23:36 adaq target: Spin Flip Report: Running Smoothly
  • 247667 11/08/08 23:34 Jin Laser Spectrum Screen Shot
  • 247666 11/08/08 23:30 a-onl Start_Run_3084,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production! HAPPEX DAQ run#=30383.
  • 247665 11/08/08 23:28 a-onl End_of_Run_3083
  • 247664 11/08/08 23:17 brads Added Target EPICs screens to auto_snap script
  • 247663 11/08/08 22:59 a-onl Start_Run_3083,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production! HAPPEX DAQ run#=30382.
  • 247662 11/08/08 22:57 a-onl End_of_Run_3082
  • 247661 11/08/08 22:31 xqian DEAD TIME on T5
  • 247660 11/08/08 22:30 a-onl Start_Run_3082,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production! HAPPEX DAQ run#=30382. T1 = 400, T6 = 200.
  • 247659 11/08/08 22:25 a-onl End_of_Run_3081
  • 247658 11/08/08 22:16 Yi Qiang EPICS strip chart screen shot
  • 247657 11/08/08 22:02 a-onl Start_Run_3081,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production! HAPPEX DAQ run#=30382.
  • 247656 11/08/08 21:58 a-onl End_of_Run_3080
  • 247655 11/08/08 21:35 a-onl Start_Run_3080,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production! HAPPEX DAQ run#=30382. set TS1=500, so PS1= 140
  • 247654 11/08/08 21:32 a-onl End_of_Run_3079
  • 247653 11/08/08 20:58 a-onl Start_Run_3079,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production! HAPPEX DAQ run#=30382
  • 247652 11/08/08 20:58 a-onl End_of_Run_3078
  • 247651 11/08/08 20:35 a-onl Start_Run_3078,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production! HAPPEX DAQ run#=30382
  • 247650 11/08/08 20:34 a-onl End_of_Run_3077
  • 247649 11/08/08 20:29 a-onl Start_Run_3077,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production! HAPPEX DAQ run#=30382
  • 247648 11/08/08 20:22 K Wang Beamline HV tripped
  • 247647 11/08/08 20:14 a-onl End_of_Run_3076
  • 247646 11/08/08 20:04 adaq Radiation and Ozone snapshot
  • 247645 11/08/08 20:04 adaq Hall A Tools snapshot
  • 247644 11/08/08 20:04 adaq BigBox Magnet snapshot
  • 247643 11/08/08 20:04 adaq Beam Position and Current snapshot
  • 247642 11/08/08 20:04 adaq Beam Energy in A and C snapshot
  • 247641 11/08/08 19:52 a-onl Start_Run_3076,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production! HAPPEX DAQ run#=30382
  • 247640 11/08/08 19:50 a-onl End_of_Run_3075
  • 247639 11/08/08 19:36 Yi Qiang Suggestions on target spin flip while moving target
  • 247638 11/08/08 19:32 a-onl Start_Run_3075,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production! HAPPEX DAQ run#=30382,
  • 247637 11/08/08 19:27 a-onl End_of_Run_3074
  • 247636 11/08/08 19:06 a-onl Start_Run_3074,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production! Stop and start HAPPEX, run# 30382.
  • 247635 11/08/08 19:04 a-onl End_of_Run_3073
  • 247634 11/08/08 18:54 Kalyan Coincidence time with 3He target
  • 247633 11/08/08 18:43 a-onl Start_Run_3073,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production!
  • 247632 11/08/08 18:43 a-onl End_of_Run_3073
  • 247631 11/08/08 18:31 Huaibo rate T3, T6
  • 247630 11/08/08 18:20 a-onl Start_Run_3072,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production!
  • 247629 11/08/08 18:20 a-onl End_of_Run_3071
  • 247628 11/08/08 17:57 a-onl Start_Run_3071,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production!
  • 247627 11/08/08 17:56 a-onl End_of_Run_3070
  • 247626 11/08/08 17:50 V. Sulkosky BPM Calibration
  • 247625 11/08/08 17:43 adaq Swiong shift tools
  • 247624 11/08/08 17:37 a-onl Start_Run_3070,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production!
  • 247623 11/08/08 17:36 a-onl End_of_Run_3069
  • 247622 11/08/08 17:05 a-onl Start_Run_3069,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production!
  • 247621 11/08/08 17:04 a-onl End_of_Run_3068
  • 247620 11/08/08 16:52 Chiranjib EPR RF amplifier OK
  • 247619 11/08/08 16:42 a-onl Start_Run_3068,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production!
  • 247618 11/08/08 16:41 a-onl End_of_Run_3067
  • 247617 11/08/08 16:31 Chiranjib RTD readout screens
  • 247616 11/08/08 16:29 Chiranjib Target spin flip screens/ laser absorption
  • 247615 11/08/08 16:25 V. Sulkosky Harp Positions Versus EPICS (11072008 Bull's eye scan)
  • 247614 11/08/08 16:21 a-onl Start_Run_3067,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production!
  • 247613 11/08/08 16:20 Chiranjib timing offset for s2m
  • 247612 11/08/08 16:19 a-onl End_of_Run_3066
  • 247611 11/08/08 16:00 adaq Beam Time Accounting
  • 247610 11/08/08 15:56 a-onl Start_Run_3066,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production!
  • 247609 11/08/08 15:55 E. Schulte Day Shift Summary, November 8, 2008
  • 247608 11/08/08 15:55 a-onl End_of_Run_3065
  • 247607 11/08/08 15:55 E. Schulte Run Numbers for Day Shift starting at 11:40
  • 247606 11/08/08 15:34 a-onl Start_Run_3065,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production!
  • 247605 11/08/08 15:33 a-onl End_of_Run_3064
  • 247604 11/08/08 15:29 deng target reference cell RTD readbacks
  • 247603 11/08/08 15:28 adaq Polarized He3 target RTD readbacks
  • 247602 11/08/08 15:18 chench replay for run 3063
  • 247601 11/08/08 15:11 a-onl Start_Run_3064,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production!
  • 247600 11/08/08 15:08 a-onl End_of_Run_3063
  • 247599 11/08/08 14:51 a-onl Start_Run_3063,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production!
  • 247598 11/08/08 14:44 a-onl End_of_Run_3062
  • 247597 11/08/08 14:22 a-onl Start_Run_3062,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production!
  • 247596 11/08/08 14:21 a-onl End_of_Run_3061
  • 247595 11/08/08 13:58 a-onl Start_Run_3061,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production!
  • 247594 11/08/08 13:57 a-onl End_of_Run_3060
  • 247593 11/08/08 13:48 adaq target: Spin Flip Report: 111
  • 247592 11/08/08 13:39 E. Schulte Tools Screen
  • 247591 11/08/08 13:34 a-onl Start_Run_3060,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production!
  • 247590 11/08/08 13:33 a-onl End_of_Run_3059
  • 247589 11/08/08 13:11 V. Sulkosky Pion Rejector Layer 2 HV
  • 247588 11/08/08 13:05 a-onl Start_Run_3059,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production!
  • 247587 11/08/08 13:04 a-onl End_of_Run_3058
  • 247586 11/08/08 12:54 Kalyan Coincidence time with 3He target
  • 247585 11/08/08 12:43 a-onl Start_Run_3058,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production!
  • 247584 11/08/08 12:42 a-onl End_of_Run_3057
  • 247583 11/08/08 12:25 Kalyan BigBite scalers not counting T7
  • 247582 11/08/08 12:14 a-onl Start_Run_3057,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production!
  • 247581 11/08/08 12:12 a-onl End_of_Run_3056
  • 247580 11/08/08 12:12 debg Laser spectrum
  • 247579 11/08/08 12:09 kalyan Trigger rates at 7uA on 3He target
  • 247578 11/08/08 12:06 E. Schulte MCC having RF problems
  • 247577 11/08/08 12:04 adaq Radiation and Ozone snapshot
  • 247576 11/08/08 12:04 adaq Hall A Tools snapshot
  • 247575 11/08/08 12:04 adaq BigBox Magnet snapshot
  • 247574 11/08/08 12:04 adaq Beam Position and Current snapshot
  • 247573 11/08/08 12:04 adaq Beam Energy in A and C snapshot
  • 247572 11/08/08 11:48 E. Schulte Beam excersion on return of beam
  • 247571 11/08/08 11:44 Kalyan T6 threshold changed to -120mV
  • 247570 11/08/08 11:42 a-onl Start_Run_3056,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production!
  • 247569 11/08/08 11:39 E. Schulte Spot++ After Extended Beam Down.
  • 247568 11/08/08 11:37 a-onl End_of_Run_3055
  • 247567 11/08/08 11:37 a-onl Start_Run_3055,Run_type=Test,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=raster test
  • 247566 11/08/08 11:29 deng Target moved to He3 cell
  • 247565 11/08/08 11:11 Kalyan changed HV on few shower blocks
  • 247564 11/08/08 11:10 a-onl End_of_Run_3054
  • 247563 11/08/08 11:10 a-onl Start_Run_3054,Run_type=Cosmics,target_type=Multi Carbon Optics/BeO,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH run with pulser ps8=1
  • 247562 11/08/08 10:46 E. Hall back in Beam Permit, but Accelerator has a Dead Magnet
  • 247561 11/08/08 10:25 deng target moved to optics
  • 247560 11/08/08 10:15 a-onl End_of_Run_3052
  • 247559 11/08/08 10:14 a-onl End_of_Run_3052
  • 247558 11/08/08 10:08 Kalyan Pedestals updated for BigBite ADC channels
  • 247557 11/08/08 10:07 a-onl Start_Run_3052,Run_type=Cosmics,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH run with pulser ps8=1
  • 247556 11/08/08 10:05 a-onl End_of_Run_3051
  • 247555 11/08/08 10:05 a-onl Start_Run_3051,Run_type=Cosmics,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH run with pulser ps8=1
  • 247554 11/08/08 10:04 a-onl End_of_Run_3050
  • 247553 11/08/08 09:57 a-onl Start_Run_3050,Run_type=Cosmics,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH run with pulser ps8=1
  • 247552 11/08/08 09:56 a-onl End_of_Run_3049
  • 247551 11/08/08 09:54 a-onl Start_Run_3049,Run_type=Cosmics,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH cosmics to check dead time with scope
  • 247550 11/08/08 09:40 Chiranjib Target spin flip screen
  • 247549 11/08/08 09:39 E. Schulte Hall in Controlled Access
  • 247548 11/08/08 09:26 adaq target: Spin Flip Report: day shift
  • 247547 11/08/08 09:17 E. Schulte MCC Needs to Work on Cooling Water Leak
  • 247546 11/08/08 09:15 a-onl End_of_Run_3048
  • 247545 11/08/08 09:14 a-onl Start_Run_3048,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production run!
  • 247544 11/08/08 09:13 a-onl End_of_Run_3047
  • 247543 11/08/08 08:56 a-onl Start_Run_3047,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production run!
  • 247542 11/08/08 08:53 a-onl End_of_Run_3046
  • 247541 11/08/08 08:34 a-onl Start_Run_3046,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production run!
  • 247540 11/08/08 08:33 a-onl End_of_Run_3045
  • 247539 11/08/08 08:09 a-onl Start_Run_3045,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Production run!
  • 247538 11/08/08 08:07 a-onl End_of_Run_3044
  • 247537 11/08/08 08:06 a-onl Start_Run_3044,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=pedestal run for RICH
  • 247536 11/08/08 08:03 adaq Beam Time Accounting
  • 247535 11/08/08 08:02 Cisbani Owl Shift Summary
  • 247534 11/08/08 07:32 samuel NMR indication of transfer polarization is 42.7%
  • 247533 11/08/08 07:01 Jin Huang Target Activity Summary @ Starting of Production
  • 247532 11/08/08 06:55 ycwang Coincidence peak in run #3030
  • 247531 11/08/08 06:54 a-onl Start_Run_3043,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=production with RICH
  • 247530 11/08/08 06:52 a-onl End_of_Run_3042
  • 247529 11/08/08 06:32 a-onl Start_Run_3042,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=
  • 247528 11/08/08 06:31 a-onl End_of_Run_3041
  • 247527 11/08/08 06:13 a-onl Start_Run_3041,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=
  • 247526 11/08/08 06:11 a-onl End_of_Run_3040
  • 247525 11/08/08 05:53 xqian BigBite Wire Chamber background study
  • 247524 11/08/08 05:16 online run3034,3035, 1700V chamber HV test,not 7000V
  • 247523 11/08/08 05:14 adev BigBite WireChamber Measured Voltage change!
  • 247522 11/08/08 05:14 a-onl Start_Run_3040,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=production with RICH
  • 247521 11/08/08 05:11 a-onl End_of_Run_3039
  • 247520 11/08/08 05:10 a-onl Start_Run_3039,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=
  • 247519 11/08/08 05:03 adev BigBite WireChamber Measured Voltage change!
  • 247518 11/08/08 05:03 a-onl End_of_Run_3038
  • 247517 11/08/08 05:01 a-onl Start_Run_3038,Run_type=Calibration,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=1550V chamber HV test
  • 247516 11/08/08 04:59 a-onl End_of_Run_3037
  • 247515 11/08/08 04:57 a-onl Start_Run_3037,Run_type=Calibration,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=1600V chamber HV test
  • 247514 11/08/08 04:54 online 1700V chamber HV test
  • 247513 11/08/08 04:54 a-onl End_of_Run_3036
  • 247512 11/08/08 04:53 a-onl Start_Run_3036,Run_type=Calibration,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=1700 V chamber HV test
  • 247511 11/08/08 04:52 adev BigBite WireChamber Measured Voltage change!
  • 247510 11/08/08 04:50 a-onl End_of_Run_3035
  • 247509 11/08/08 04:48 a-onl Start_Run_3035,Run_type=Calibration,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=7000 V chamber HV test
  • 247508 11/08/08 04:45 a-onl End_of_Run_3034
  • 247507 11/08/08 04:44 samuel Polarization reach to 40.5%
  • 247506 11/08/08 04:43 a-onl Start_Run_3034,Run_type=Calibration,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=7000 V chamber HV test
  • 247505 11/08/08 04:42 adev BigBite WireChamber Measured Voltage change!
  • 247504 11/08/08 04:40 a-onl End_of_Run_3033
  • 247503 11/08/08 04:38 a-onl Start_Run_3033,Run_type=Calibration,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=1650 V chamber HV test
  • 247502 11/08/08 04:35 a-onl End_of_Run_3032
  • 247501 11/08/08 04:33 a-onl Start_Run_3032,Run_type=Calibration,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=1600 V chamber HV test
  • 247500 11/08/08 04:32 a-onl End_of_Run_3031
  • 247499 11/08/08 04:32 a-onl Start_Run_3031,Run_type=Calibration,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=
  • 247498 11/08/08 04:31 adev BigBite WireChamber Measured Voltage change!
  • 247497 11/08/08 04:31 adaq HallA Tools
  • 247496 11/08/08 04:28 a-onl End_of_Run_3030
  • 247495 11/08/08 04:28 Cisbani spot++
  • 247494 11/08/08 04:16 a-onl Start_Run_3030,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=
  • 247493 11/08/08 04:15 a-onl End_of_Run_3029
  • 247492 11/08/08 04:04 adaq Radiation and Ozone snapshot
  • 247491 11/08/08 04:03 adaq Hall A Tools snapshot
  • 247490 11/08/08 04:03 adaq BigBox Magnet snapshot
  • 247489 11/08/08 04:03 adaq Beam Position and Current snapshot
  • 247488 11/08/08 04:03 adaq Beam Energy in A and C snapshot
  • 247487 11/08/08 03:53 a-onl Start_Run_3029,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=
  • 247486 11/08/08 03:49 a-onl End_of_Run_3028
  • 247485 11/08/08 03:48 a-onl Start_Run_3028,Run_type=Test,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH pedestal run
  • 247484 11/08/08 03:27 a-onl End_of_Run_3027
  • 247483 11/08/08 03:27 a-onl Start_Run_3027,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=JUNK
  • 247482 11/08/08 03:24 a-onl End_of_Run_3026
  • 247481 11/08/08 03:12 Jin Temp. Golden Root File
  • 247480 11/08/08 03:03 a-onl Start_Run_3026,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=
  • 247479 11/08/08 03:02 a-onl End_of_Run_3025
  • 247478 11/08/08 03:00 a-onl Start_Run_3025,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=first production run (happex run 30381)
  • 247477 11/08/08 02:46 adev BigBite WireChamber Measured Voltage change!
  • 247476 11/08/08 02:46 adaq target: Spin Flip Report: "Production run starts"; 20min Auto Spin Flip Actived
  • 247475 11/08/08 02:44 a-onl End_of_Run_3024
  • 247474 11/08/08 02:42 a-onl Start_Run_3024,Run_type=Production,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=
  • 247473 11/08/08 02:38 a-onl End_of_Run_3023
  • 247472 11/08/08 02:35 adev BigBite WireChamber Measured Voltage change!
  • 247471 11/08/08 02:35 a-onl Start_Run_3023,Run_type=Calibration,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=12uA Wire Chamber test
  • 247470 11/08/08 02:32 adaq target: Spin Flip Report: Successful Spin Flip in 12 uA Beam. Pol.=36%
  • 247469 11/08/08 02:28 a-onl End_of_Run_3022
  • 247468 11/08/08 02:27 a-onl Start_Run_3022,Run_type=Calibration,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=10uA Wire Chamber test
  • 247467 11/08/08 02:25 adev BigBite WireChamber Measured Voltage change!
  • 247466 11/08/08 02:21 a-onl End_of_Run_3021
  • 247465 11/08/08 02:19 a-onl Start_Run_3021,Run_type=Calibration,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=8uA Wire Chamber test
  • 247464 11/08/08 02:14 a-onl End_of_Run_3020
  • 247463 11/08/08 02:14 adev BigBite WireChamber Measured Voltage change!
  • 247462 11/08/08 02:13 Jin Huang Spin Flip software taking over the target
  • 247461 11/08/08 02:12 a-onl Start_Run_3020,Run_type=Calibration,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=6uA Wire Chamber test
  • 247460 11/08/08 02:09 a-onl End_of_Run_3019
  • 247459 11/08/08 02:06 a-onl Start_Run_3019,Run_type=Calibration,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=4uA Wire Chamber test
  • 247458 11/08/08 02:03 a-onl End_of_Run_3018
  • 247457 11/08/08 02:03 adev BigBite WireChamber Measured Voltage change!
  • 247456 11/08/08 02:02 a-onl Start_Run_3018,Run_type=Calibration,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=2uA Wire Chamber test
  • 247455 11/08/08 02:02 adaq target: Spin Flip Report: Spin Flip System is taking over the target
  • 247454 11/08/08 01:56 a-onl End_of_Run_3017
  • 247453 11/08/08 01:52 cisbani rate at raster 6.5x6 mcc
  • 247452 11/08/08 01:52 a-onl Start_Run_3017,Run_type=Calibration,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=6.5x6 MCC,raster on
  • 247451 11/08/08 01:48 a-onl End_of_Run_3016
  • 247450 11/08/08 01:43 xqian BB_PSh threshold is changed to 80mV from 20mV
  • 247449 11/08/08 01:40 cisbani rate at raster 6x5 mcc
  • 247448 11/08/08 01:40 a-onl Start_Run_3016,Run_type=Calibration,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=6x5 MCC,raster on
  • 247447 11/08/08 01:37 a-onl End_of_Run_3015
  • 247446 11/08/08 01:36 a-onl Start_Run_3015,Run_type=Calibration,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=4.5x4 MCC,raster on
  • 247445 11/08/08 01:35 a-onl End_of_Run_3014
  • 247444 11/08/08 01:33 cisbani Rate at Raster 4.5x4mcc
  • 247443 11/08/08 01:28 a-onl Start_Run_3014,Run_type=Calibration,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=3x3 MCC,raster on
  • 247442 11/08/08 01:25 a-onl End_of_Run_3013
  • 247441 11/08/08 01:24 adaq Rate at raster 3x3 mcc
  • 247440 11/08/08 01:22 a-onl Start_Run_3013,Run_type=Calibration,target_type=Polarized 3He,beam_e0=5.9,beam_raster=on,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=3x3 MCC,raster on
  • 247439 11/08/08 01:13 samuel move to He3 target position
  • 247438 11/08/08 00:52 samuel move target to empty position
  • 247437 11/08/08 00:52 Holmstrom / Bull-s Eyes Scan
  • 247436 11/08/08 00:51 a-onl End_of_Run_3012
  • 247435 11/08/08 00:49 a-onl Start_Run_3012,Run_type=Calibration,target_type=Multi Carbon Optics/BeO,beam_e0=5.9,beam_raster=off,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=bull's eye, raster off, x=0,y=0 ,optics target
  • 247434 11/08/08 00:43 a-onl End_of_Run_3011
  • 247433 11/08/08 00:42 a-onl Start_Run_3011,Run_type=Calibration,target_type=Multi Carbon Optics/BeO,beam_e0=5.889,beam_raster=off,leftarm_p=2.35,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=bull's eye, raster off, x=+2,y=-2 ,optics target
  • 247432 11/08/08 00:32 a-onl End_of_Run_3010
  • 247431 11/08/08 00:28 a-onl Start_Run_3010,Run_type=Calibration,target_type=Multi Carbon Optics/BeO,beam_e0=5.889,beam_raster=off,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=bull's eye, raster off, x=-2,y=-2 ,optics target
  • 247430 11/08/08 00:24 a-onl End_of_Run_3009
  • 247429 11/08/08 00:24 a-onl Start_Run_3009,Run_type=Calibration,target_type=Multi Carbon Optics/BeO,beam_e0=5.889,beam_raster=off,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=bull's eye, raster off, x=-2,y=-2 ,optics target
  • 247428 11/08/08 00:01 Holmstrom Swing Shift Summary
  • 247427 11/08/08 00:00 adaq Beam Time Accounting
  • Friday

  • 247426 11/07/08 23:50 adaq target: Spin Flip Report: New Code Test
  • 247425 11/07/08 23:35 a-onl End_of_Run_3008
  • 247424 11/07/08 23:33 a-onl Start_Run_3008,Run_type=Calibration,target_type=Multi Carbon Optics/BeO,beam_e0=5.889,beam_raster=off,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=bull's eye, raster off, x=-2,y=2 ,optics target
  • 247423 11/07/08 23:33 a-onl End_of_Run_3007
  • 247422 11/07/08 23:24 a-onl Start_Run_3007,Run_type=Calibration,target_type=Multi Carbon Optics/BeO,beam_e0=5.889,beam_raster=off,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=bull's eye, raster off, x=-2,y=2 ,optics target
  • 247421 11/07/08 23:11 a-onl End_of_Run_3006
  • 247420 11/07/08 23:09 a-onl Start_Run_3006,Run_type=Calibration,target_type=Multi Carbon Optics/BeO,beam_e0=5.889,beam_raster=off,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=bull's eye, raster off, x,y=2 ,optics target
  • 247419 11/07/08 23:07 a-onl End_of_Run_3005
  • 247418 11/07/08 22:55 a-onl Start_Run_3005,Run_type=Calibration,target_type=Multi Carbon Optics/BeO,beam_e0=5.889,beam_raster=off,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=bull's eye, raster off, x,y=2 ,optics target
  • 247417 11/07/08 22:47 Kalyan Coincidence time peak with 5 pass beam
  • 247416 11/07/08 22:16 a-onl End_of_Run_3004
  • 247415 11/07/08 22:13 a-onl Start_Run_3004,Run_type=Calibration,target_type=Multi Carbon Optics/BeO,beam_e0=5.889,beam_raster=off,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=bull's eye, raster off, nominal beam postion ,optics target
  • 247414 11/07/08 21:59 yez Target Moved to Optics w/o hole
  • 247413 11/07/08 21:55 A. Camsonne Reference cell well aligned good for setting ion chamber trip point
  • 247412 11/07/08 21:50 a-onl End_of_Run_3003
  • 247411 11/07/08 21:50 A. Camsonne Rates reference cell N2 50 psi 2uA
  • 247410 11/07/08 21:49 a-onl Start_Run_3003,Run_type=Check out,target_type=Reference Cell: N2,beam_e0=5.889,beam_raster=on,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=2uA raster 5x5
  • 247409 11/07/08 21:47 A. Camsonne Rates N2 50 psi 2 uA raster 4x4mm
  • 247408 11/07/08 21:46 a-onl End_of_Run_3002
  • 247407 11/07/08 21:43 a-onl Start_Run_3002,Run_type=Check out,target_type=Reference Cell: N2,beam_e0=5.889,beam_raster=on,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=2uA raster 4x4
  • 247406 11/07/08 21:42 a-onl End_of_Run_3001
  • 247405 11/07/08 21:41 Yi Qiang Correction Tests
  • 247404 11/07/08 21:40 Yi Qiang Correction Tests
  • 247403 11/07/08 21:39 A. Camsonne Rates reference cell raster 3x3mm
  • 247402 11/07/08 21:39 yez Target Moved to Refference Cell
  • 247401 11/07/08 21:39 a-onl Start_Run_3001,Run_type=Check out,target_type=Reference Cell: N2,beam_e0=5.889,beam_raster=on,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=2uA raster 3x3
  • 247400 11/07/08 21:34 yez Target Moved to Refference Cell
  • 247399 11/07/08 21:30 Yi Qiang Filled Reference Cell to 25 psig with N2
  • 247398 11/07/08 21:19 a-onl End_of_Run_3000
  • 247397 11/07/08 21:18 a-onl Start_Run_3000,Run_type=Check out,target_type=Multi Carbon Optics/BeO,beam_e0=5.889,beam_raster=on,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=5uA spot check
  • 247396 11/07/08 21:17 a-onl End_of_Run_2999
  • 247395 11/07/08 21:03 a-onl Start_Run_2999,Run_type=Check out,target_type=Multi Carbon Optics/BeO,beam_e0=5.889,beam_raster=on,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=9 uA to check pion rejector and t3 dealyed by 10 more ..command t3_delay 60 140
  • 247394 11/07/08 21:02 brads (Tech) RHRS VDC discriminator problem?
  • 247393 11/07/08 21:00 a-onl End_of_Run_2998
  • 247392 11/07/08 20:28 a-onl Start_Run_2998,Run_type=Check out,target_type=Multi Carbon Optics/BeO,beam_e0=5.889,beam_raster=on,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=9 uA to check pion rejector
  • 247391 11/07/08 20:25 a-onl End_of_Run_2997
  • 247390 11/07/08 20:03 adaq Radiation and Ozone snapshot
  • 247389 11/07/08 20:03 adaq Hall A Tools snapshot
  • 247388 11/07/08 20:03 adaq BigBox Magnet snapshot
  • 247387 11/07/08 20:03 adaq Beam Position and Current snapshot
  • 247386 11/07/08 20:03 adaq Beam Energy in A and C snapshot
  • 247385 11/07/08 20:03 a-onl Start_Run_2997,Run_type=Check out,target_type=Multi Carbon Optics/BeO,beam_e0=5.889,beam_raster=on,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=12 uA to check coin time command used T3_delay 50 140
  • 247384 11/07/08 20:02 A. Camsonne Update on HARP
  • 247383 11/07/08 20:00 a-onl End_of_Run_2996
  • 247382 11/07/08 19:45 a-onl Start_Run_2996,Run_type=Check out,target_type=Multi Carbon Optics/BeO,beam_e0=5.889,beam_raster=on,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=12 uA to chexk coin time
  • 247381 11/07/08 19:27 a-onl End_of_Run_2995
  • 247380 11/07/08 18:59 Kalyan Trigger rates at 10uA on carbon foil
  • 247379 11/07/08 18:54 a-onl Start_Run_2995,Run_type=Check out,target_type=Multi Carbon Optics/BeO,beam_e0=5.889,beam_raster=on,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=12 uA to chexk coin time
  • 247378 11/07/08 18:45 Holmstrom Harp Scan software problem
  • 247377 11/07/08 18:22 Yi Zhang Target move to BeO target to do harpscan
  • 247376 11/07/08 18:20 Yi Zhang Reference cell filled with N2(1 psi)
  • 247375 11/07/08 18:15 yez Target Moved to Refference Cell
  • 247374 11/07/08 17:39 Holmstrom Target Masking Bug?
  • 247373 11/07/08 17:23 Yi Qiang Target move to empty
  • 247372 11/07/08 17:22 a-onl End_of_Run_2994
  • 247371 11/07/08 17:22 a-onl Start_Run_2994,Run_type=Test,target_type=Multi Carbon Optics/BeO,beam_e0=5.889,beam_raster=off,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=junk
  • 247370 11/07/08 16:01 adaq Beam Time Accounting
  • 247369 11/07/08 15:55 E. Schulte Day Shift Summary, November 7, 2008
  • 247368 11/07/08 15:41 a-onl End_of_Run_2992
  • 247367 11/07/08 15:39 a-onl Start_Run_2992,Run_type=Test,target_type=Multi Carbon Optics/BeO,beam_e0=5.889,beam_raster=off,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=spot++ test
  • 247366 11/07/08 15:37 A. Camsonne Preliminary results of target survey notes by Brad
  • 247365 11/07/08 15:31 V. Sulkosky New Database Directory
  • 247364 11/07/08 15:24 A. Camsonne Fan unit added above the helicity crate in counting house
  • 247363 11/07/08 15:16 V. Sulkosky BPM Pedestal Run 2991
  • 247362 11/07/08 15:05 V. Sulkosky BPM and Raster Work II
  • 247361 11/07/08 14:34 Sulkosky/Sch BPM and Raster Work
  • 247360 11/07/08 13:55 a-onl End_of_Run_2991
  • 247359 11/07/08 13:53 a-onl Start_Run_2991,Run_type=Test,target_type=No Target,beam_e0=5.889,beam_raster=off,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=bpm pedestal run with T8
  • 247358 11/07/08 13:49 brads LHRS set to 2.35 GeV/c
  • 247357 11/07/08 13:45 brads BigBite is back at 710A
  • 247356 11/07/08 13:33 E. Cisbani / Switch network port
  • 247355 11/07/08 13:32 a-onl Start_Run_2990,Run_type=Test,target_type=No Target,beam_e0=5.889,beam_raster=off,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH test HV on
  • 247354 11/07/08 13:30 Xiaodong Jia Need to fine tune left_HRS setting, to 2.35 GeV/c.
  • 247353 11/07/08 13:30 Jack (Techs) Changed UPS
  • 247352 11/07/08 13:10 Kalyan RE: BB chamber dark current check. Beam current:4uA
  • 247351 11/07/08 12:55 A. Camsonne Transvsersity experiment status and reminder experiment meeting at 4PM in counting house 2nd floor meeting room
  • 247350 11/07/08 12:46 a-onl Start_Run_2988,Run_type=Test,target_type=No Target,beam_e0=5.889,beam_raster=off,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH test HV on
  • 247349 11/07/08 12:45 a-onl End_of_Run_2987
  • 247348 11/07/08 12:44 a-onl Start_Run_2987,Run_type=Test,target_type=No Target,beam_e0=5.889,beam_raster=off,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH test HV on
  • 247347 11/07/08 12:34 a-onl End_of_Run_2985
  • 247346 11/07/08 12:32 a-onl Start_Run_2985,Run_type=Test,target_type=No Target,beam_e0=5.889,beam_raster=off,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH test HV on
  • 247345 11/07/08 12:26 Yi Qiang Left HRS Q1 is ON (set to 2.4GeV)
  • 247344 11/07/08 12:25 a-onl Start_Run_2984,Run_type=Test,target_type=No Target,beam_e0=5.889,beam_raster=off,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH test HV on
  • 247343 11/07/08 12:23 a-onl End_of_Run_2983
  • 247342 11/07/08 12:16 a-onl Start_Run_2983,Run_type=Test,target_type=No Target,beam_e0=5.889,beam_raster=off,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH test
  • 247341 11/07/08 12:15 a-onl End_of_Run_2982
  • 247340 11/07/08 12:07 Yi Qiang Left HRS Q1 and BigBite Magnets are OFF
  • 247339 11/07/08 12:06 a-onl Start_Run_2982,Run_type=Test,target_type=No Target,beam_e0=5.889,beam_raster=off,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH test
  • 247338 11/07/08 12:05 a-onl End_of_Run_2981
  • 247337 11/07/08 12:03 adaq Radiation and Ozone snapshot
  • 247336 11/07/08 12:03 adaq Hall A Tools snapshot
  • 247335 11/07/08 12:03 adaq BigBox Magnet snapshot
  • 247334 11/07/08 12:03 adaq Beam Position and Current snapshot
  • 247333 11/07/08 12:03 adaq Beam Energy in A and C snapshot
  • 247332 11/07/08 12:00 a-onl Start_Run_2981,Run_type=Test,target_type=No Target,beam_e0=5.889,beam_raster=off,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH test
  • 247331 11/07/08 11:42 a-onl End_of_Run_2979
  • 247330 11/07/08 11:34 a-onl Start_Run_2979,Run_type=Test,target_type=Multi Carbon Optics/BeO,beam_e0=5.889,beam_raster=off,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test for Vince cont.
  • 247329 11/07/08 11:28 a-onl End_of_Run_2978
  • 247328 11/07/08 11:25 a-onl Start_Run_2978,Run_type=Test,target_type=Multi Carbon Optics/BeO,beam_e0=5.889,beam_raster=off,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test for Vince cont.
  • 247327 11/07/08 11:17 a-onl End_of_Run_2977
  • 247326 11/07/08 11:15 a-onl Start_Run_2977,Run_type=Test,target_type=Multi Carbon Optics/BeO,beam_e0=5.889,beam_raster=off,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test for Vince cont.
  • 247325 11/07/08 11:12 a-onl Start_Run_2976,Run_type=Test,target_type=Multi Carbon Optics/BeO,beam_e0=5.889,beam_raster=off,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test for Vince cont.
  • 247324 11/07/08 11:11 a-onl End_of_Run_2976
  • 247323 11/07/08 11:06 a-onl End_of_Run_2975
  • 247322 11/07/08 10:59 a-onl Start_Run_2975,Run_type=Test,target_type=Multi Carbon Optics/BeO,beam_e0=5.889,beam_raster=off,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test for Vince cont.
  • 247321 11/07/08 10:46 a-onl End_of_Run_2974
  • 247320 11/07/08 10:44 brads Raster ADC change: heads up to analysis folks
  • 247319 11/07/08 10:43 a-onl Start_Run_2974,Run_type=Test,target_type=Multi Carbon Optics/BeO,beam_e0=5.889,beam_raster=off,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test for Vince cont.
  • 247318 11/07/08 10:40 a-onl End_of_Run_2973
  • 247317 11/07/08 10:34 a-onl Start_Run_2973,Run_type=Test,target_type=Multi Carbon Optics/BeO,beam_e0=5.889,beam_raster=off,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test for Vince cont.
  • 247316 11/07/08 10:33 a-onl End_of_Run_2972
  • 247315 11/07/08 10:28 a-onl Start_Run_2972,Run_type=Test,target_type=Multi Carbon Optics/BeO,beam_e0=5.889,beam_raster=off,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test for Vince cont.
  • 247314 11/07/08 10:27 cusanno RICH dymanic library for positive polarity
  • 247313 11/07/08 10:27 a-onl End_of_Run_2971
  • 247312 11/07/08 10:24 brads Must do Bull's Eye scan when we come back
  • 247311 11/07/08 10:22 brads Hot spots found during
  • 247310 11/07/08 10:13 a-onl Start_Run_2971,Run_type=Test,target_type=Multi Carbon Optics/BeO,beam_e0=5.889,beam_raster=off,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test for Vince.
  • 247309 11/07/08 10:11 Jin Target Ladder No Respond, Reboot IOC again
  • 247308 11/07/08 10:05 Jin Target IOC Reboot
  • 247307 11/07/08 09:45 V. Sulkosky Raster Overflowing and Underflowing
  • 247306 11/07/08 09:25 Chiranjib/Ji SPIN UP @220 C
  • 247305 11/07/08 09:24 E. Schulte Requested Controlled Access to Work on Compton
  • 247304 11/07/08 09:22 adev BigBite WireChamber Measured Voltage change!
  • 247303 11/07/08 09:12 Jin/Chiranji EPR Error, Pol is low now
  • 247302 11/07/08 09:11 adev BigBite WireChamber Measured Voltage change!
  • 247301 11/07/08 08:37 brads Software problem with target ladder control
  • 247300 11/07/08 08:36 E. Schulte Tools screen for 3He
  • 247299 11/07/08 08:34 Yi Zhang 3He laser off
  • 247298 11/07/08 08:24 Kalyan Beam energy for runs 2936 to 2970 is 5.889GeV and not 2.39GeV
  • 247297 11/07/08 08:16 ycwang Turn off HV of RICH for Target reason.
  • 247296 11/07/08 08:12 E. Schulte Target Motion Control Problems
  • 247295 11/07/08 08:06 adaq Beam Time Accounting
  • 247294 11/07/08 08:06 E. Cisbani Owl Shift Summary
  • 247293 11/07/08 08:05 a-onl End_of_Run_2970
  • 247292 11/07/08 07:50 a-onl Start_Run_2970,Run_type=Test,target_type=Multi Carbon Optics/BeO,beam_e0=2.396,beam_raster=off,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=raster off,5uA,optics target check. Change to I =4uA. It works!!!
  • 247291 11/07/08 07:45 yawei run2969, ROC9 no respond.
  • 247290 11/07/08 07:42 a-onl Start_Run_2969,Run_type=Test,target_type=Multi Carbon Optics/BeO,beam_e0=2.396,beam_raster=off,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=raster off,5uA,optics target check
  • 247289 11/07/08 07:38 a-onl End_of_Run_2968
  • 247288 11/07/08 07:38 yawei run2965,2966,2967,ROC9 no respond
  • 247287 11/07/08 07:37 a-onl Start_Run_2968,Run_type=Test,target_type=Multi Carbon Optics/BeO,beam_e0=2.396,beam_raster=off,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=raster off,5uA,optics target check
  • 247286 11/07/08 07:32 a-onl Start_Run_2967,Run_type=Test,target_type=Multi Carbon Optics/BeO,beam_e0=2.396,beam_raster=off,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=raster off,5uA,optics target check
  • 247285 11/07/08 07:26 a-onl Start_Run_2966,Run_type=Test,target_type=Multi Carbon Optics/BeO,beam_e0=2.396,beam_raster=off,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=raster off,5uA,optics target check
  • 247284 11/07/08 07:21 a-onl Start_Run_2965,Run_type=Test,target_type=Multi Carbon Optics/BeO,beam_e0=2.396,beam_raster=off,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=raster off,5uA,optics target check
  • 247283 11/07/08 07:10 a-onl End_of_Run_2964
  • 247282 11/07/08 07:00 a-onl Start_Run_2964,Run_type=Test,target_type=Multi Carbon Optics/BeO,beam_e0=2.396,beam_raster=on,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=3x3 raster,5uA
  • 247281 11/07/08 06:55 a-onl Start_Run_2963,Run_type=Test,target_type=Multi Carbon Optics/BeO,beam_e0=2.396,beam_raster=on,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=3x3 raster
  • 247280 11/07/08 06:55 Huaibo no "optics with hole" selection
  • 247279 11/07/08 06:50 Huaibo no "optics with hole" seclection
  • 247278 11/07/08 06:45 a-onl End_of_Run_2962
  • 247277 11/07/08 06:37 a-onl Start_Run_2962,Run_type=Test,target_type=Multi Carbon Optics/BeO,beam_e0=2.396,beam_raster=on,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=TwoArmRICH using CramWait no limit on channel
  • 247276 11/07/08 06:34 a-onl End_of_Run_2961
  • 247275 11/07/08 06:17 a-onl Start_Run_2961,Run_type=Test,target_type=Multi Carbon Optics/BeO,beam_e0=2.396,beam_raster=on,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=TwoArmRICH using CramWait
  • 247274 11/07/08 06:16 a-onl End_of_Run_2960
  • 247273 11/07/08 06:02 a-onl Start_Run_2960,Run_type=Test,target_type=Multi Carbon Optics/BeO,beam_e0=2.396,beam_raster=on,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test only ROC14
  • 247272 11/07/08 05:58 a-onl End_of_Run_2959
  • 247271 11/07/08 05:58 a-onl Start_Run_2959,Run_type=Test,target_type=Multi Carbon Optics/BeO,beam_e0=2.396,beam_raster=on,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test only ROC14
  • 247270 11/07/08 05:54 a-onl End_of_Run_2958
  • 247269 11/07/08 05:47 a-onl Start_Run_2958,Run_type=Test,target_type=Multi Carbon Optics/BeO,beam_e0=2.396,beam_raster=on,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH run all T3 and T5 limit to 60 channels read out per ADC
  • 247268 11/07/08 05:46 ycwang A1 sum for check run#2512
  • 247267 11/07/08 05:46 a-onl End_of_Run_2957
  • 247266 11/07/08 05:30 a-onl Start_Run_2957,Run_type=Test,target_type=Multi Carbon Optics/BeO,beam_e0=2.396,beam_raster=on,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH run all T3 and T5 limit to 60 channels read out per ADC
  • 247265 11/07/08 05:29 a-onl End_of_Run_2956
  • 247264 11/07/08 05:29 A. Camsonne Limited channel readout per ADC for RICH to 60
  • 247263 11/07/08 05:14 a-onl Start_Run_2956,Run_type=Test,target_type=Multi Carbon Optics/BeO,beam_e0=2.396,beam_raster=on,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH run all T3 and T5
  • 247262 11/07/08 05:12 a-onl End_of_Run_2955
  • 247261 11/07/08 05:09 E. Cisbani Left arm HV
  • 247260 11/07/08 05:08 a-onl Start_Run_2955,Run_type=Test,target_type=Multi Carbon Optics/BeO,beam_e0=2.396,beam_raster=on,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH pedestal run T5 only while generating pedestal file
  • 247259 11/07/08 05:06 a-onl End_of_Run_2954
  • 247258 11/07/08 05:06 a-onl Start_Run_2954,Run_type=Test,target_type=Multi Carbon Optics/BeO,beam_e0=2.396,beam_raster=on,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH pedestal run T5 only
  • 247257 11/07/08 05:02 adaq Hall A tools
  • 247256 11/07/08 04:59 a-onl Start_Run_2952,Run_type=Test,target_type=Multi Carbon Optics/BeO,beam_e0=2.396,beam_raster=on,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH test
  • 247255 11/07/08 04:59 adev BigBite WireChamber Measured Voltage change!
  • 247254 11/07/08 04:51 Yi Zhang BB chamber test
  • 247253 11/07/08 04:49 adev BigBite WireChamber Measured Voltage change!
  • 247252 11/07/08 04:39 A. Camsonne recorded raster value at 6 GeV overflowing ADC ?
  • 247251 11/07/08 04:36 a-onl End_of_Run_2951
  • 247250 11/07/08 04:35 ycwang Coincidence peak in run 2946
  • 247249 11/07/08 04:34 A. Camsonne ROC9 rebooted by itself
  • 247248 11/07/08 04:32 a-onl Start_Run_2951,Run_type=Test,target_type=Multi Carbon Optics/BeO,beam_e0=2.396,beam_raster=on,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=junk
  • 247247 11/07/08 04:30 a-onl Start_Run_2950,Run_type=Test,target_type=Multi Carbon Optics/BeO,beam_e0=2.396,beam_raster=on,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=junk
  • 247246 11/07/08 04:26 a-onl Start_Run_2949,Run_type=Test,target_type=Multi Carbon Optics/BeO,beam_e0=2.396,beam_raster=on,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB chamber HV check; spot(4.5x4) MCC unit
  • 247245 11/07/08 04:17 ycwang Target position for RUN 2946
  • 247244 11/07/08 04:06 adev BigBite WireChamber Measured Voltage change!
  • 247243 11/07/08 04:03 adaq Radiation and Ozone snapshot
  • 247242 11/07/08 04:03 adaq Hall A Tools snapshot
  • 247241 11/07/08 04:03 adaq BigBox Magnet snapshot
  • 247240 11/07/08 04:03 adaq Beam Position and Current snapshot
  • 247239 11/07/08 04:02 adaq Beam Energy in A and C snapshot
  • 247238 11/07/08 04:02 a-onl Start_Run_2948,Run_type=Test,target_type=Multi Carbon Optics/BeO,beam_e0=2.396,beam_raster=on,leftarm_p=2.4,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB chamber HV check; spot(4.5x4) MCC unit,BB HV tuned
  • 247237 11/07/08 03:56 a-onl End_of_Run_2947
  • 247236 11/07/08 03:56 adev BigBite WireChamber Measured Voltage change!
  • 247235 11/07/08 03:55 ycwang lots of new warnings when replay
  • 247234 11/07/08 03:55 Yi Zhang BB chamber dark current check. Beam current:4uA
  • 247233 11/07/08 03:54 a-onl Start_Run_2947,Run_type=Test,target_type=Multi Carbon Optics/BeO,beam_e0=2.396,beam_raster=on,leftarm_p=2.4,leftarm_theta=14,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB chamber HV check; spot(4.5x4) MCC unit,BB HV tuned
  • 247232 11/07/08 03:45 adev BigBite WireChamber Measured Voltage change!
  • 247231 11/07/08 03:44 a-onl End_of_Run_2946
  • 247230 11/07/08 03:42 a-onl Start_Run_2946,Run_type=Test,target_type=Multi Carbon Optics/BeO,beam_e0=2.396,beam_raster=on,leftarm_p=2.4,leftarm_theta=14,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB chamber HV check; spot(4.5x4) MCC unit,BB HV tuned
  • 247229 11/07/08 03:40 Yi Zhang BB chamber dark current under 3 uA
  • 247228 11/07/08 03:37 a-onl End_of_Run_2945
  • 247227 11/07/08 03:36 a-onl Start_Run_2945,Run_type=Test,target_type=Multi Carbon Optics/BeO,beam_e0=2.396,beam_raster=on,leftarm_p=2.4,leftarm_theta=14,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB chamber HV check; spot(4.5x4) MCC unit
  • 247226 11/07/08 03:34 Yi Zhang BB chamber dark current under 3 uA
  • 247225 11/07/08 03:27 a-onl End_of_Run_2944
  • 247224 11/07/08 03:24 adev BigBite WireChamber Measured Voltage change!
  • 247223 11/07/08 03:24 a-onl Start_Run_2944,Run_type=Test,target_type=Multi Carbon Optics/BeO,beam_e0=2.396,beam_raster=on,leftarm_p=2.4,leftarm_theta=14,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BB chamber HV check; spot(4.5x4)
  • 247222 11/07/08 03:22 a-onl End_of_Run_2943
  • 247221 11/07/08 03:21 E. Cisbani Prescale all triggers but t3 to fix raster problem did not work
  • 247220 11/07/08 03:18 a-onl Start_Run_2943,Run_type=Test,target_type=Multi Carbon Optics/BeO,beam_e0=2.396,beam_raster=on,leftarm_p=2.4,leftarm_theta=14,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=check spot++ problem 4.5x4
  • 247219 11/07/08 03:18 Yi Zhang BB chamber dark current under 2 uA
  • 247218 11/07/08 03:13 adev BigBite WireChamber Measured Voltage change!
  • 247217 11/07/08 02:43 Huaibo move target to optics
  • 247216 11/07/08 02:36 Yi Zhang THaRaster::"rb.Raster" problem
  • 247215 11/07/08 02:18 E. Cisbani Raster Size on Holey Carbon
  • 247214 11/07/08 02:17 a-onl End_of_Run_2942
  • 247213 11/07/08 02:17 a-onl Start_Run_2942,Run_type=Test,target_type=Single Carbon w/ Hole,beam_e0=2.396,beam_raster=on,leftarm_p=2.4,leftarm_theta=14,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=check spot++, holy carbon 4.5x4
  • 247212 11/07/08 02:13 online run2838,2839,target is optics with hole,sorry
  • 247211 11/07/08 02:03 a-onl End_of_Run_2941
  • 247210 11/07/08 01:49 a-onl Start_Run_2941,Run_type=Test,target_type=Single Carbon w/ Hole,beam_e0=2.396,beam_raster=on,leftarm_p=0.6,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=check spot++, holy carbon 5x5
  • 247209 11/07/08 01:47 online target is Optics with hole,sorry
  • 247208 11/07/08 01:43 online target is optics with hole, sorry
  • 247207 11/07/08 01:37 a-onl End_of_Run_2940
  • 247206 11/07/08 01:37 a-onl Start_Run_2940,Run_type=Test,target_type=Single Carbon w/ Hole,beam_e0=2.396,beam_raster=on,leftarm_p=0.6,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=check spot++, holy carbon
  • 247205 11/07/08 01:29 a-onl End_of_Run_2939
  • 247204 11/07/08 01:28 a-onl Start_Run_2939,Run_type=Test,target_type=No Target,beam_e0=2.396,beam_raster=on,leftarm_p=0.6,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=check spot++, holy carbon
  • 247203 11/07/08 01:19 a-onl End_of_Run_2938
  • 247202 11/07/08 01:17 a-onl Start_Run_2938,Run_type=Test,target_type=No Target,beam_e0=2.396,beam_raster=on,leftarm_p=0.6,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=check spot++, holy carbon
  • 247201 11/07/08 01:13 E. Cisbani Harp Scan Ref
  • 247200 11/07/08 01:08 Huaibo Move target to optics with hole
  • 247199 11/07/08 00:58 brads Beam energy fluctuations
  • 247198 11/07/08 00:51 brads 100x to 1000x the radiation dose during early 5 pass
  • 247197 11/07/08 00:36 E. Cisbani spot++ BeO target 2uA
  • 247196 11/07/08 00:35 a-onl End_of_Run_2937
  • 247195 11/07/08 00:33 a-onl Start_Run_2937,Run_type=Test,target_type=No Target,beam_e0=2.396,beam_raster=off,leftarm_p=0.6,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=check spot++
  • 247194 11/07/08 00:30 a-onl End_of_Run_2936
  • 247193 11/07/08 00:28 a-onl Start_Run_2936,Run_type=Test,target_type=No Target,beam_e0=2.396,beam_raster=off,leftarm_p=0.6,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=check spot++
  • 247192 11/07/08 00:27 brads Update on target vs. beam positions
  • 247191 11/07/08 00:10 adaq Beam Time Accounting
  • 247190 11/07/08 00:09 Ibrahim Swing Shift Summary
  • Thursday

  • 247189 11/06/08 23:58 Huaibo Move target to optics
  • 247188 11/06/08 23:15 cusanno RICH software updated
  • 247187 11/06/08 23:07 cusanno RICH tripped, presently HV is OFF
  • 247186 11/06/08 22:17 V. Sulkosky Left HRS HV GUI
  • 247185 11/06/08 21:25 brads Wien angle setting: 90% source polarization to Hall
  • 247184 11/06/08 20:59 brads BB MWDC ethanol bubbler engaged
  • 247183 11/06/08 20:49 brads BB Shower threshold set to -80mV.
  • 247182 11/06/08 20:44 adev BigBite WireChamber Measured Voltage change!
  • 247181 11/06/08 20:36 a-onl End_of_Run_2935
  • 247180 11/06/08 20:06 doug Test of Web Entry with Photo
  • 247179 11/06/08 20:02 adaq Radiation and Ozone snapshot
  • 247178 11/06/08 20:02 adaq Hall A Tools snapshot
  • 247177 11/06/08 20:02 adaq BigBox Magnet snapshot
  • 247176 11/06/08 20:02 adaq Beam Position and Current snapshot
  • 247175 11/06/08 20:02 adaq Beam Energy in A and C snapshot
  • 247174 11/06/08 20:01 brads NOT steering through the Compton chicane
  • 247173 11/06/08 19:58 a-onl Start_Run_2935,Run_type=Cosmics,target_type=No Target,beam_e0=2.396,beam_raster=on,leftarm_p=0.6,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Cosmics Wire Chamber
  • 247172 11/06/08 19:54 a-onl End_of_Run_2934
  • 247171 11/06/08 19:53 Kalyan ROC8 pedestals updated
  • 247170 11/06/08 19:51 brads testing weblog 3
  • 247169 11/06/08 19:38 xiaohui Target moved to empty
  • 247168 11/06/08 18:13 brads Web-based halog disabled for now
  • 247167 11/06/08 17:24 V. Sulkosky Web based Halog Not Working
  • 247166 11/06/08 17:22 V. Sulkosky Timing offsets for S2m updated
  • 247165 11/06/08 17:18 V. Sulkosky Timing offsets for S1 updated
  • 247164 11/06/08 17:08 Ibrahim Left Dipole Polarity Chaged to Negative
  • 247163 11/06/08 16:11 adaq Beam Time Accounting
  • 247162 11/06/08 16:01 adev BigBite WireChamber Measured Voltage change!
  • 247161 11/06/08 15:57 a-onl Start_Run_2934,Run_type=Cosmics,target_type=No Target,beam_e0=2.396,beam_raster=on,leftarm_p=0.6,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Cosmics Wire Chamber
  • 247160 11/06/08 14:37 a-onl End_of_Run_2933
  • 247158 11/06/08 14:25 Kalyan Run 2932 is pedestals for ROC8
  • 247157 11/06/08 14:19 a-onl End_of_Run_2932
  • 247156 11/06/08 14:19 a-onl Start_Run_2932,Run_type=Test,target_type=No Target,beam_e0=2.396,beam_raster=on,leftarm_p=0.6,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Pedestals for ROC8 (pedsupress temporarily disabled for this run)
  • 247155 11/06/08 12:59 a-onl End_of_Run_2931
  • 247154 11/06/08 12:58 a-onl Start_Run_2931,Run_type=Test,target_type=No Target,beam_e0=2.396,beam_raster=on,leftarm_p=0.6,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH pedestal, T3 only
  • 247153 11/06/08 12:57 a-onl End_of_Run_2930
  • 247152 11/06/08 12:23 a-onl Start_Run_2930,Run_type=Test,target_type=No Target,beam_e0=2.396,beam_raster=on,leftarm_p=0.6,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH Clock Test (pedestals)
  • 247151 11/06/08 12:10 E. Schulte Tools Screen
  • 247150 11/06/08 12:02 adaq Radiation and Ozone snapshot
  • 247149 11/06/08 12:02 adaq Hall A Tools snapshot
  • 247148 11/06/08 12:02 adaq BigBox Magnet snapshot
  • 247147 11/06/08 12:02 adaq Beam Position and Current snapshot
  • 247146 11/06/08 12:02 adaq Beam Energy in A and C snapshot
  • 247145 11/06/08 11:32 a-onl End_of_Run_2929
  • 247144 11/06/08 11:29 a-onl Start_Run_2929,Run_type=Test,target_type=No Target,beam_e0=2.396,beam_raster=on,leftarm_p=0.6,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH Clock Test (pedestals)
  • 247143 11/06/08 11:28 a-onl End_of_Run_2928
  • 247142 11/06/08 11:27 a-onl Start_Run_2928,Run_type=Test,target_type=No Target,beam_e0=2.396,beam_raster=on,leftarm_p=0.6,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH Clock Test (pedestals)
  • 247141 11/06/08 11:23 a-onl End_of_Run_2927
  • 247140 11/06/08 11:20 a-onl Start_Run_2927,Run_type=Test,target_type=No Target,beam_e0=2.396,beam_raster=on,leftarm_p=0.6,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH Clock Test (pedestals)
  • 247139 11/06/08 11:18 a-onl End_of_Run_2926
  • 247138 11/06/08 11:16 a-onl Start_Run_2926,Run_type=Test,target_type=No Target,beam_e0=2.396,beam_raster=on,leftarm_p=0.6,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH Clock Test (pedestals)
  • 247137 11/06/08 11:11 a-onl End_of_Run_2925
  • 247136 11/06/08 11:04 brads Target -> He3 Cell
  • 247135 11/06/08 11:02 a-onl Start_Run_2925,Run_type=Test,target_type=No Target,beam_e0=2.396,beam_raster=on,leftarm_p=0.6,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH Clock Test (in pedestal mode)
  • 247134 11/06/08 11:00 brads Target -> Reference cell
  • 247133 11/06/08 10:45 E. Schulte Gas Shed Air Flow Alarm
  • 247132 11/06/08 10:41 Jin Huang Target Summary during Spin flip test @ Delta Resonance & further
  • 247131 11/06/08 10:41 Yi Qiang Target to Optics (BeO)
  • 247130 11/06/08 10:38 adaq target: Spin Flip Report: Some 1.6% holding filed drift begin from 9:55AM completely screwed up the polarization
  • 247129 11/06/08 10:34 a-onl Start_Run_2924,Run_type=Test,target_type=No Target,beam_e0=2.396,beam_raster=on,leftarm_p=0.6,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH Clock Test (in pedestal mode)
  • 247128 11/06/08 10:30 E. Schulte Target FP Loss Test Needs to be Performed Again
  • 247127 11/06/08 10:30 brads LHRS moved to 16 degrees
  • 247126 11/06/08 10:01 Yi Qiang Lasers Locked
  • 247125 11/06/08 09:26 Jin Huang Spin Signal Locked to Spin Up = 1; Spin Down =0
  • 247124 11/06/08 08:52 E. Schulte Hall to go to Controlled Access
  • 247123 11/06/08 08:52 adaq target: Spin Flip Report: End of beam time
  • 247122 11/06/08 08:49 a-onl End_of_Run_2923
  • 247121 11/06/08 08:40 gaskell HAPPEX run 30379 during run 2923
  • 247120 11/06/08 08:35 a-onl Start_Run_2923,Run_type=Calibration,target_type=Reference Cell: 3He,beam_e0=2.396,beam_raster=on,leftarm_p=0.6,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Delta Production at Transverse Pol He3 Cell. HAPPEX run # 30378
  • 247119 11/06/08 08:34 a-onl End_of_Run_2922
  • 247118 11/06/08 08:30 Jin Huang he3 vertex and Inv Mass in 2 pass beam
  • 247117 11/06/08 08:26 ycwang After RUN 2917, the ADCs of S1 and S2m have two peaks.
  • 247116 11/06/08 08:22 ycwang After RUN 2917, the ADCs of S1 and S2m have two peaks.
  • 247115 11/06/08 08:17 Jin Huang Real Spin Signal reads OK in DAQ
  • 247114 11/06/08 08:07 adaq Beam Time Accounting
  • 247113 11/06/08 08:04 Yi Zhang owl shift summay
  • 247112 11/06/08 07:50 a-onl Start_Run_2922,Run_type=Calibration,target_type=Reference Cell: 3He,beam_e0=2.396,beam_raster=on,leftarm_p=0.6,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Delta Production at Transverse Pol He3 Cell. HAPPEX run # 30378
  • 247111 11/06/08 07:49 a-onl End_of_Run_2921
  • 247110 11/06/08 07:47 Yi Zhang hall a tools screen
  • 247109 11/06/08 07:31 Jin Huang Laser Absorption Spectrum
  • 247108 11/06/08 07:05 a-onl Start_Run_2921,Run_type=Calibration,target_type=Reference Cell: 3He,beam_e0=2.396,beam_raster=on,leftarm_p=0.6,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Delta Production at Transverse Pol He3 Cell. HAPPEX run # 30377 . Start at Spin+ state
  • 247107 11/06/08 07:04 a-onl End_of_Run_2920
  • 247106 11/06/08 06:10 a-onl Start_Run_2920,Run_type=Calibration,target_type=Reference Cell: 3He,beam_e0=2.396,beam_raster=on,leftarm_p=0.6,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Delta Production at Transverse Pol He3 Cell. HAPPEX run # 30376 . Start at Spin+ state
  • 247105 11/06/08 06:08 a-onl End_of_Run_2919
  • 247104 11/06/08 05:39 jpchen Raster size for beam on pol He3
  • 247103 11/06/08 05:38 Yi Zhang BB chamber HV triped
  • 247102 11/06/08 05:38 Yi Zhang spot++ of he3 run#2919,
  • 247101 11/06/08 05:38 Jin Huang Automatic Spin Flip ON
  • 247100 11/06/08 05:35 Yi Zhang BB chamber HV triped
  • 247099 11/06/08 05:31 a-onl Start_Run_2919,Run_type=Calibration,target_type=Reference Cell: 3He,beam_e0=2.396,beam_raster=on,leftarm_p=0.6,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Delta Production at Transverse Pol He3 Cell. Auto Spin Flip enabled, Start at Spin+ state, Pol.=~50%
  • 247098 11/06/08 05:30 adev BigBite WireChamber Measured Voltage change!
  • 247097 11/06/08 05:26 ycwang RICH Pedestal RUN with beam RUN # 2913
  • 247096 11/06/08 05:23 A. Camsonne RICH runs summary and test
  • 247095 11/06/08 05:19 adev BigBite WireChamber Measured Voltage change!
  • 247094 11/06/08 05:18 adaq target: Spin Flip Report: Spin Flip System Fully Enabled with Transverse Pumping
  • 247093 11/06/08 05:13 samuel 1 atm Nytrogemn gas in reference cell.
  • 247092 11/06/08 05:09 samuel move target to He3 Position.
  • 247091 11/06/08 05:05 a-onl End_of_Run_2918
  • 247090 11/06/08 05:02 a-onl Start_Run_2918,Run_type=Optics,target_type=Reference Cell: N2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH run mixed triggers ps1=20,ps2-10,ps5=1 with pedestal subtratction
  • 247089 11/06/08 05:01 a-onl End_of_Run_2917
  • 247088 11/06/08 04:59 a-onl Start_Run_2917,Run_type=Optics,target_type=Reference Cell: N2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH run T3 only with pedestal subtraction ps3=1 T3=750 Hz
  • 247087 11/06/08 04:59 a-onl End_of_Run_2916
  • 247086 11/06/08 04:54 A. Camsonne RICH run 2916
  • 247085 11/06/08 04:51 a-onl Start_Run_2916,Run_type=Optics,target_type=Reference Cell: N2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH run T3 only with pedestal subtraction ps3=1 T3=750 Hz
  • 247084 11/06/08 04:48 a-onl End_of_Run_2915
  • 247083 11/06/08 04:48 a-onl Start_Run_2915,Run_type=Optics,target_type=Reference Cell: N2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=test RICH T1 only ps1=10 ps5=1
  • 247082 11/06/08 04:44 a-onl End_of_Run_2914
  • 247081 11/06/08 04:41 a-onl Start_Run_2914,Run_type=Optics,target_type=Reference Cell: N2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH run T3 no pedestal subtraction
  • 247080 11/06/08 04:38 a-onl End_of_Run_2913
  • 247079 11/06/08 04:38 A. Camsonne RICH good pedestal run 2913
  • 247078 11/06/08 04:37 a-onl Start_Run_2913,Run_type=Optics,target_type=Reference Cell: N2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH pedestal run
  • 247077 11/06/08 04:35 a-onl End_of_Run_2912
  • 247076 11/06/08 04:35 a-onl Start_Run_2912,Run_type=Optics,target_type=Reference Cell: N2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH test
  • 247075 11/06/08 04:32 a-onl Start_Run_2911,Run_type=Optics,target_type=Reference Cell: N2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH pedestal run
  • 247074 11/06/08 04:30 a-onl Start_Run_2910,Run_type=Optics,target_type=Reference Cell: N2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH pedestal run
  • 247073 11/06/08 04:19 a-onl Start_Run_2908,Run_type=Optics,target_type=Reference Cell: N2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH check
  • 247072 11/06/08 04:13 a-onl End_of_Run_2907
  • 247071 11/06/08 04:06 a-onl Start_Run_2907,Run_type=Optics,target_type=Reference Cell: N2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=N2 crossection checks
  • 247070 11/06/08 04:05 adaq LHRS momentum change to 600 MeV/C
  • 247069 11/06/08 04:02 jpchen Use N2 for RICH study
  • 247068 11/06/08 04:02 adaq Radiation and Ozone snapshot
  • 247067 11/06/08 04:02 adaq Hall A Tools snapshot
  • 247066 11/06/08 04:02 adaq BigBox Magnet snapshot
  • 247065 11/06/08 04:02 adaq Beam Position and Current snapshot
  • 247064 11/06/08 04:02 adaq Beam Energy in A and C snapshot
  • 247063 11/06/08 03:59 a-onl End_of_Run_2906
  • 247062 11/06/08 03:59 ycwang No signal on BB Sum TDCs etc for Run #2902
  • 247061 11/06/08 03:56 a-onl Start_Run_2906,Run_type=Optics,target_type=Reference Cell: N2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=N2 crossection checks
  • 247060 11/06/08 03:55 a-onl End_of_Run_2905
  • 247059 11/06/08 03:51 a-onl Start_Run_2905,Run_type=Optics,target_type=Reference Cell: N2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=N2 crossection checks
  • 247058 11/06/08 03:45 ycwang Run #2902 coincidence peak
  • 247057 11/06/08 03:40 samuel fill reference cell with N2 to pressure 50PSI
  • 247056 11/06/08 03:34 a-onl End_of_Run_2904
  • 247055 11/06/08 03:19 a-onl Start_Run_2904,Run_type=Optics,target_type=Reference Cell: H2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=H2 crossection checks, T1=200Hz
  • 247054 11/06/08 03:17 a-onl End_of_Run_2903
  • 247053 11/06/08 03:03 a-onl Start_Run_2903,Run_type=Optics,target_type=Reference Cell: H2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=H2 crossection checks, T1 only
  • 247052 11/06/08 03:00 a-onl End_of_Run_2902
  • 247051 11/06/08 02:59 xqian The t0 offset for wire chamber for t5.
  • 247050 11/06/08 02:43 a-onl Start_Run_2902,Run_type=Optics,target_type=Reference Cell: H2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=H2 crossection checks, T5 only
  • 247049 11/06/08 02:41 adev BigBite WireChamber Measured Voltage change!
  • 247048 11/06/08 02:34 a-onl End_of_Run_2901
  • 247047 11/06/08 02:26 a-onl Start_Run_2901,Run_type=Optics,target_type=Reference Cell: H2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=H2 crossection checks, T1 only
  • 247046 11/06/08 02:23 a-onl End_of_Run_2900
  • 247045 11/06/08 02:04 a-onl Start_Run_2900,Run_type=Optics,target_type=Reference Cell: H2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=H2 cross check, T5 only
  • 247044 11/06/08 01:58 a-onl End_of_Run_2899
  • 247043 11/06/08 01:45 xqian Delta in BigBite
  • 247042 11/06/08 01:43 xqian BigBite Wire Chamber running condition study
  • 247041 11/06/08 01:32 a-onl Start_Run_2899,Run_type=Cosmics,target_type=Reference Cell: H2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=beam trip, take a cosmics run
  • 247040 11/06/08 01:28 adev BigBite WireChamber Measured Voltage change!
  • 247039 11/06/08 01:26 kalyan late entry : T3 delay changed
  • 247038 11/06/08 01:24 a-onl End_of_Run_2898
  • 247037 11/06/08 00:58 a-onl Start_Run_2898,Run_type=Optics,target_type=Reference Cell: H2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Ref. Cell H2,Beam current 4uA raster size 3x3mm 12uA
  • 247036 11/06/08 00:51 Yi Qiang Typo on the error
  • 247035 11/06/08 00:46 adev BigBite WireChamber Measured Voltage change!
  • 247034 11/06/08 00:41 Yi Qiang Max Polarization for current Transverse and Vertical Pumping
  • 247033 11/06/08 00:41 Yi Qiang Max Polarization for current Transverse and Vertical Pumping
  • 247032 11/06/08 00:35 a-onl End_of_Run_2897
  • 247030 11/06/08 00:01 adaq Beam Time Accounting
  • 247031 11/06/08 00:01 Ibrahim Swing Shift Summary
  • Wednesday

  • 247029 11/05/08 23:47 a-onl Start_Run_2897,Run_type=Optics,target_type=Reference Cell: H2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Ref. Cell H2,Beam current 4uA raster size 3x3mm
  • 247028 11/05/08 23:45 a-onl End_of_Run_2896
  • 247027 11/05/08 23:27 Chiranjib/Xi Temperature Test
  • 247026 11/05/08 23:27 Chiranjib/Xi Temperature Test
  • 247025 11/05/08 23:04 xiaohui Temp. test done
  • 247024 11/05/08 22:49 a-onl Start_Run_2896,Run_type=Optics,target_type=Reference Cell: H2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Ref. Cell H2,Beam current 4uA raster size 3x3mm
  • 247023 11/05/08 22:47 a-onl End_of_Run_2895
  • 247022 11/05/08 22:47 a-onl Start_Run_2895,Run_type=Cosmics,target_type=Reference Cell: H2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Ref. Cell H2,Beam current 4uA raster size 3x3mm
  • 247021 11/05/08 22:45 a-onl End_of_Run_2894
  • 247020 11/05/08 22:18 a-onl Start_Run_2894,Run_type=Optics,target_type=Reference Cell: H2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Ref. Cell H2,Beam current 4uA raster size 3x3mm
  • 247019 11/05/08 22:17 a-onl End_of_Run_2893
  • 247018 11/05/08 22:12 a-onl Start_Run_2893,Run_type=Optics,target_type=Reference Cell: H2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Ref. Cell H2,Beam current 4uA raster size 3x3mm
  • 247017 11/05/08 22:10 a-onl End_of_Run_2892
  • 247016 11/05/08 22:04 a-onl Start_Run_2892,Run_type=Optics,target_type=Reference Cell: H2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Ref. Cell H2,Beam current 4uA raster size 3x3mm
  • 247015 11/05/08 22:00 a-onl End_of_Run_2891
  • 247014 11/05/08 21:58 adev BigBite WireChamber Measured Voltage change!
  • 247013 11/05/08 21:54 xiaohui 3He temp. test started
  • 247012 11/05/08 21:49 a-onl Start_Run_2891,Run_type=Optics,target_type=Reference Cell: H2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Ref. Cell H2,Beam current 2uA raster size 3x3mm
  • 247011 11/05/08 21:47 a-onl End_of_Run_2890
  • 247010 11/05/08 21:43 a-onl Start_Run_2890,Run_type=Optics,target_type=Reference Cell: H2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Ref. Cell H2,Beam current 2uA raster size 5x5mm
  • 247009 11/05/08 21:42 a-onl End_of_Run_2889
  • 247008 11/05/08 21:42 a-onl Start_Run_2889,Run_type=Optics,target_type=Reference Cell: H2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Ref. Cell H2,Beam current 2uA raster size 4x46mm
  • 247007 11/05/08 21:40 a-onl Start_Run_2888,Run_type=Optics,target_type=Reference Cell: H2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Ref. Cell H2,Beam current 2uA raster size 3x36mm
  • 247005 11/05/08 21:38 a-onl End_of_Run_2887
  • 247004 11/05/08 21:37 adev BigBite WireChamber Measured Voltage change!
  • 247003 11/05/08 21:35 a-onl Start_Run_2887,Run_type=Optics,target_type=Reference Cell: H2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Ref. Cell H2,Beam current 2uA raster size 2x2mm
  • 247002 11/05/08 21:30 xiaohui target moved to reference
  • 247001 11/05/08 21:26 a-onl End_of_Run_2886
  • 247000 11/05/08 21:16 a-onl Start_Run_2886,Run_type=Optics,target_type=Multi Carbon Optics/BeO,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=optics without hole,Beam current 2uA raster size 3x3mm
  • 246999 11/05/08 21:13 a-onl End_of_Run_2885
  • 246998 11/05/08 21:05 adev BigBite WireChamber Measured Voltage change!
  • 246997 11/05/08 21:05 Ibrahim Harp Scan Results (20:35)
  • 246996 11/05/08 21:04 a-onl Start_Run_2885,Run_type=Optics,target_type=Multi Carbon Optics/BeO,beam_e0=2.396,beam_raster=off,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=optics without hole,Beam current 2uA
  • 246995 11/05/08 20:55 adev BigBite WireChamber Measured Voltage change!
  • 246994 11/05/08 20:54 a-onl End_of_Run_2884
  • 246993 11/05/08 20:54 a-onl Start_Run_2884,Run_type=Optics,target_type=Multi Carbon Optics/BeO,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=optics without hole,Beam current 5uA
  • 246992 11/05/08 20:53 chench 3x3 raster size check
  • 246991 11/05/08 20:50 a-onl End_of_Run_2883
  • 246990 11/05/08 20:45 a-onl Start_Run_2883,Run_type=Optics,target_type=Multi Carbon Optics/BeO,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=optics without hole,Beam current 4uA
  • 246989 11/05/08 20:37 chench beam position confirm
  • 246988 11/05/08 20:35 a-onl End_of_Run_2882
  • 246987 11/05/08 20:33 adev BigBite WireChamber Measured Voltage change!
  • 246986 11/05/08 20:28 a-onl Start_Run_2882,Run_type=Optics,target_type=Multi Carbon Optics/BeO,beam_e0=2.396,beam_raster=off,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=optics without hole,Beam current 4uA
  • 246985 11/05/08 20:02 adaq Radiation and Ozone snapshot
  • 246984 11/05/08 20:02 adaq Hall A Tools snapshot
  • 246983 11/05/08 20:01 adaq BigBox Magnet snapshot
  • 246982 11/05/08 20:01 adaq Beam Position and Current snapshot
  • 246981 11/05/08 20:01 adaq Beam Energy in A and C snapshot
  • 246980 11/05/08 19:58 a-onl End_of_Run_2881
  • 246979 11/05/08 19:53 a-onl Start_Run_2881,Run_type=Optics,target_type=Multi Carbon Optics/BeO,beam_e0=2.396,beam_raster=off,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=optics without hole,5uA
  • 246978 11/05/08 19:50 a-onl End_of_Run_2880
  • 246976 11/05/08 19:47 adev SCALER = -1
  • 246975 11/05/08 19:42 a-onl End_of_Run_2879
  • 246974 11/05/08 19:38 a-onl Start_Run_2879,Run_type=Cosmics,target_type=Reference Cell: H2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=DAQ test
  • 246973 11/05/08 19:38 adaq target: Spin Flip Report: Tested for Trans Spin Flip
  • 246972 11/05/08 19:36 a-onl End_of_Run_2878
  • 246971 11/05/08 19:36 a-onl Start_Run_2878,Run_type=Cosmics,target_type=Reference Cell: H2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=DAQ test
  • 246970 11/05/08 19:34 Xiaohui Target moved to optics
  • 246969 11/05/08 19:25 a-onl End_of_Run_2877
  • 246968 11/05/08 19:24 a-onl Start_Run_2877,Run_type=Cosmics,target_type=Reference Cell: H2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=Cosmics
  • 246967 11/05/08 19:22 A. Camsonne new CODA configuration for pedestal
  • 246966 11/05/08 19:13 a-onl End_of_Run_2876
  • 246965 11/05/08 19:10 a-onl Start_Run_2876,Run_type=Cosmics,target_type=Reference Cell: H2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=Cosmics
  • 246964 11/05/08 19:09 adev BigBite WireChamber Measured Voltage change!
  • 246963 11/05/08 19:06 brads BigBite Sieve in the OUT position
  • 246962 11/05/08 19:05 xiaohui Target moved to empty
  • 246961 11/05/08 19:04 a-onl Start_Run_2875,Run_type=Cosmics,target_type=Reference Cell: H2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=RICH pedestal in RICHPedestal only T3 good run
  • 246960 11/05/08 19:02 a-onl Start_Run_2874,Run_type=Cosmics,target_type=Reference Cell: H2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=RICH pedestal in RICHPedestal only T3
  • 246959 11/05/08 19:00 a-onl End_of_Run_2873
  • 246958 11/05/08 19:00 a-onl Start_Run_2873,Run_type=Cosmics,target_type=Reference Cell: H2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=Cosmics with mutliple triggers
  • 246957 11/05/08 18:59 a-onl End_of_Run_2872
  • 246956 11/05/08 18:54 a-onl Start_Run_2872,Run_type=Cosmics,target_type=Reference Cell: H2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=Cosmics with mutliple triggers
  • 246955 11/05/08 18:49 a-onl End_of_Run_2871
  • 246954 11/05/08 18:49 a-onl Start_Run_2871,Run_type=Cosmics,target_type=Reference Cell: H2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=RICH pedestal
  • 246953 11/05/08 18:46 Ibrahim Gas Shed Airflow Alarm
  • 246952 11/05/08 18:46 a-onl End_of_Run_2870
  • 246951 11/05/08 18:44 chench ref. cell H2 cosmic run
  • 246950 11/05/08 18:29 a-onl Start_Run_2870,Run_type=Cosmics,target_type=Reference Cell: H2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=target collimators IN; raster=3x3 mm; H2 pressure=135 psi Cosmics
  • 246949 11/05/08 18:25 V. Sulkosky Harp A Scan Analysis from 16:58
  • 246948 11/05/08 18:15 a-onl End_of_Run_2869
  • 246947 11/05/08 18:12 chench To modify the halog
  • 246946 11/05/08 17:58 chench reference cell He
  • 246945 11/05/08 17:55 chench Ref. Cell He
  • 246944 11/05/08 17:47 Kalyan shift in coincidence peak by 16ns
  • 246943 11/05/08 17:45 adev BigBite WireChamber Measured Voltage change!
  • 246942 11/05/08 17:43 a-onl Start_Run_2869,Run_type=Optics,target_type=Reference Cell: H2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=target collimators IN; raster=3x3 mm; H2 pressure=135 psi
  • 246941 11/05/08 17:42 a-onl End_of_Run_2868
  • 246940 11/05/08 17:36 a-onl Start_Run_2868,Run_type=Optics,target_type=Reference Cell: H2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=target collimators IN; raster=3x3 mm; H2 pressure=135 psi
  • 246939 11/05/08 17:30 a-onl End_of_Run_2867
  • 246938 11/05/08 17:28 a-onl Start_Run_2867,Run_type=Optics,target_type=Reference Cell: H2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=target collimators IN; raster=3x3 mm; H2 pressure=135 psi
  • 246937 11/05/08 17:26 Katich BeO cross-hair position
  • 246936 11/05/08 17:14 adev BigBite WireChamber Measured Voltage change!
  • 246935 11/05/08 17:11 xiaohui target moved to reference cell
  • 246934 11/05/08 17:00 brads Harp A Scan (needs analysis)
  • 246933 11/05/08 16:59 brads Harp B scaA
  • 246932 11/05/08 16:52 brads Harp B scan
  • 246931 11/05/08 16:47 brads BPMs vs. BeO spot image
  • 246930 11/05/08 16:42 a-onl End_of_Run_2866
  • 246929 11/05/08 16:05 V. Sulkosky Preliminary Left HRS Detector Position Measurements
  • 246928 11/05/08 16:00 adaq Beam Time Accounting
  • 246927 11/05/08 15:59 a-onl Start_Run_2866,Run_type=Optics,target_type=Reference Cell: H2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=target collimators IN; raster=3x3 mm; H2 pressure=135 psi
  • 246926 11/05/08 15:59 a-onl End_of_Run_2865
  • 246925 11/05/08 15:58 E. Schulte Day Shift Summary, November 5, 2008
  • 246924 11/05/08 15:42 Chiranjib/yi EPR measurement
  • 246923 11/05/08 15:32 gaskell BB tracking OK
  • 246922 11/05/08 15:10 a-onl Start_Run_2865,Run_type=Optics,target_type=Reference Cell: H2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=target collimators IN; raster=3x3 mm; H2 pressure=135 psi
  • 246921 11/05/08 15:09 a-onl End_of_Run_2864
  • 246920 11/05/08 15:04 Kalyan Note on target-helicity gated scalers
  • 246919 11/05/08 14:54 E. Schulte Gas Shed Ariflow Alarm
  • 246918 11/05/08 14:49 gaskell Run 2863 is junk
  • 246917 11/05/08 14:48 a-onl Start_Run_2864,Run_type=Optics,target_type=Reference Cell: H2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=target collimators IN; raster=3x3 mm; H2 pressure=135 psi
  • 246916 11/05/08 14:36 a-onl Start_Run_2863,Run_type=Optics,target_type=Reference Cell: H2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=target collimators IN; raster=3x3 mm; H2 pressure=135 psi
  • 246915 11/05/08 14:28 E. Schulte Spot++ for run 2862
  • 246914 11/05/08 14:26 Yi Qiang Target to Reference (H2 135 psig)
  • 246913 11/05/08 14:23 a-onl End_of_Run_2862
  • 246912 11/05/08 14:14 Yi Qiang Filled reference cell with H2 to 135 psig
  • 246911 11/05/08 14:13 a-onl Start_Run_2862,Run_type=Optics,target_type=Multi Carbon Optics/BeO,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=target collimators IN
  • 246910 11/05/08 14:11 E. Schulte Run # 2861 No Raster Spot++
  • 246909 11/05/08 14:10 a-onl End_of_Run_2861
  • 246908 11/05/08 14:02 Kalyan T3 delay changed
  • 246907 11/05/08 13:59 a-onl Start_Run_2861,Run_type=Optics,target_type=Multi Carbon Optics/BeO,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=target collimators IN
  • 246906 11/05/08 13:57 Yi Qiang Target to Optics (BeO)
  • 246905 11/05/08 13:55 A. Camsonne RICH crl change and logic change
  • 246904 11/05/08 13:53 gaskell SPOT++ with holey target: raster=4x4
  • 246903 11/05/08 13:53 a-onl End_of_Run_2860
  • 246902 11/05/08 13:50 a-onl Start_Run_2860,Run_type=Optics,target_type=Single Carbon w/ Hole,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=Spot run on Holey carbon
  • 246901 11/05/08 13:50 a-onl End_of_Run_2859
  • 246900 11/05/08 13:44 a-onl Start_Run_2859,Run_type=Optics,target_type=Single Carbon w/ Hole,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=Spot run on Holey carbon
  • 246899 11/05/08 13:39 gaskell Harp scan 1H03B
  • 246898 11/05/08 13:39 R. Michaels xcefdmp updated
  • 246897 11/05/08 13:37 gaskell Harp scan - 1H03A
  • 246896 11/05/08 13:35 yez Target Moved to Optics w hole
  • 246895 11/05/08 13:27 E. Schulte Tools screen with new positions.
  • 246894 11/05/08 13:26 a-onl End_of_Run_2858
  • 246893 11/05/08 13:25 yez Target Moved to Empty Position
  • 246892 11/05/08 13:24 gaskell spot++ screen grab
  • 246891 11/05/08 13:20 a-onl Start_Run_2858,Run_type=Optics,target_type=No Target,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=Spot run ; reference cell; Hydrogen 110psi; collimater out; T3 delayed by 50 ns - RICH test
  • 246890 11/05/08 13:20 E. Schulte Picture after position change
  • 246889 11/05/08 13:10 E. Schulte Starting positions on BeO Target
  • 246888 11/05/08 13:05 gaskell Runs 2854-2857 DAQ tests
  • 246887 11/05/08 13:02 yez Target Moved to Optics w/o hold
  • 246886 11/05/08 13:00 E. Schulte Wien Angle Changed
  • 246885 11/05/08 12:56 Kalyan coincidence time peak
  • 246884 11/05/08 12:55 a-onl End_of_Run_2857
  • 246883 11/05/08 12:54 V. Sulkosky Added Trigger T5 to 1875 TDC
  • 246882 11/05/08 12:38 a-onl Start_Run_2857,Run_type=Optics,target_type=No Target,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=Junk run for Vince; reference cell; Hydrogen 110psi; collimater out; T3 delayed by 50 ns - RICH test
  • 246880 11/05/08 12:26 a-onl Start_Run_2856,Run_type=Optics,target_type=Reference Cell: H2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=Junk run for Vince; reference cell; Hydrogen 110psi; collimater out; T3 delayed by 50 ns
  • 246879 11/05/08 12:26 a-onl End_of_Run_2856
  • 246878 11/05/08 12:26 a-onl End_of_Run_2856
  • 246877 11/05/08 12:18 Yi Qiang Laser absorption spectrum for transverse pumping
  • 246876 11/05/08 12:14 Yi Qiang Continue Transver pumping
  • 246875 11/05/08 12:01 adaq Radiation and Ozone snapshot
  • 246874 11/05/08 12:01 adaq Hall A Tools snapshot
  • 246873 11/05/08 12:01 adaq BigBox Magnet snapshot
  • 246872 11/05/08 12:01 adaq Beam Position and Current snapshot
  • 246871 11/05/08 12:01 adaq Beam Energy in A and C snapshot
  • 246870 11/05/08 12:01 a-onl Start_Run_2855,Run_type=Optics,target_type=Reference Cell: H2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=Junk run for Vince; reference cell; Hydrogen 110psi; collimater out; T3 delayed by 50 ns
  • 246869 11/05/08 11:58 a-onl End_of_Run_2854
  • 246868 11/05/08 11:54 E. Schulte Tools Screen
  • 246867 11/05/08 11:49 E. Schulte Reference Cell Controller Back in Remote Mode
  • 246866 11/05/08 11:43 brads Confirmed target alignment with laser level
  • 246865 11/05/08 11:40 brads H2 pressure regulator swapped out
  • 246864 11/05/08 11:31 a-onl Start_Run_2854,Run_type=Optics,target_type=Reference Cell: H2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=Junk run for Vince; reference cell; Hydrogen 110psi; collimater out; T3 delayed by 50 ns
  • 246863 11/05/08 11:23 E. Schulte Target Collimators are In
  • 246862 11/05/08 11:04 V. Sulkosky Beam positions from run 2853
  • 246861 11/05/08 11:00 V. Sulkosky BPM Calibration Completed
  • 246860 11/05/08 10:47 E. Schulte Reference Cell Control Panel in Local
  • 246859 11/05/08 10:34 Yi Qiang Laser locked for openning the enclosure
  • 246858 11/05/08 10:32 E. Schulte Coincidence Time Location Tests
  • 246857 11/05/08 10:11 brads Target alignment photos from Tuesday
  • 246856 11/05/08 10:06 E. Schulte MCC to do Wein angle change
  • 246855 11/05/08 10:04 E. Schulte Hall in Controlled Access
  • 246854 11/05/08 10:03 a-onl End_of_Run_2853
  • 246853 11/05/08 10:03 a-onl Start_Run_2853,Run_type=Optics,target_type=Reference Cell: H2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=reference cell; Hydrogen 110psi; collimater out; T3 delayed by 50 ns
  • 246852 11/05/08 09:57 a-onl End_of_Run_2852
  • 246851 11/05/08 09:38 E. Schulte 2uA RTD screens
  • 246850 11/05/08 09:08 a-onl Start_Run_2852,Run_type=Optics,target_type=Reference Cell: H2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=reference cell; Hydrogen 110psi; collimater out; T3 delayed by 50 ns
  • 246849 11/05/08 09:05 a-onl End_of_Run_2851
  • 246848 11/05/08 08:59 a-onl Start_Run_2851,Run_type=Optics,target_type=Reference Cell: H2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=reference cell; Hydrogen 110psi; collimater out
  • 246847 11/05/08 08:58 a-onl End_of_Run_2850
  • 246846 11/05/08 08:55 xqian Elastic proton in Left HRS, 2pass run 2847
  • 246845 11/05/08 08:53 Yi Qiang Orientation of Spot++ plot and BeO photo
  • 246844 11/05/08 08:23 xqian elastic electron at two pass
  • 246843 11/05/08 08:05 adaq Beam Time Accounting
  • 246842 11/05/08 08:04 a-onl Start_Run_2850,Run_type=Optics,target_type=Reference Cell: H2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=reference cell; Hydrogen 110psi; collimater out
  • 246841 11/05/08 08:03 a-onl End_of_Run_2849
  • 246840 11/05/08 08:01 Yi Zhang Owl shift summary
  • 246839 11/05/08 07:55 xqian Sieve pattern with 2pass H2 data elastic
  • 246838 11/05/08 07:53 xqian Unsuccesful extraction of the coincidence peak
  • 246837 11/05/08 07:48 Yi Zhang Reference cell Raster size test
  • 246836 11/05/08 07:36 adev BigBite WireChamber Measured Voltage change!
  • 246835 11/05/08 07:33 Yi Zhang HallA main tools
  • 246834 11/05/08 07:32 Yi Zhang BCM&BPM
  • 246833 11/05/08 07:20 a-onl Start_Run_2849,Run_type=Optics,target_type=Reference Cell: H2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=reference cell; Hydrogen 110psi; collimater out
  • 246832 11/05/08 07:18 a-onl End_of_Run_2848
  • 246830 11/05/08 07:14 a-onl End_of_Run_2847
  • 246829 11/05/08 06:54 adev BigBite WireChamber Measured Voltage change!
  • 246828 11/05/08 06:48 a-onl Start_Run_2847,Run_type=Optics,target_type=Reference Cell: H2,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=reference cell; Hydrogen 110psi; collimater out
  • 246827 11/05/08 06:43 adev BigBite WireChamber Measured Voltage change!
  • 246826 11/05/08 06:21 Jin Cannot fill in H2 to Reference Cell to Required Pressure
  • 246825 11/05/08 06:17 Huaibo move target to reference cell(H2)
  • 246824 11/05/08 06:12 Jin Huang Look into the second peak on BB momentum spectrum
  • 246823 11/05/08 05:58 a-onl End_of_Run_2846
  • 246822 11/05/08 05:53 Yi Zhang raster size 1x1 mm
  • 246821 11/05/08 05:51 Yi Zhang raster 3x3 mm(4x4 in MCC unit) run#2846
  • 246820 11/05/08 05:46 a-onl Start_Run_2846,Run_type=Optics,target_type=Single Carbon w/ Hole,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=1x1 raster on; optics; 3x3mm raster size
  • 246819 11/05/08 05:44 Yi Zhang raster size
  • 246818 11/05/08 05:40 a-onl End_of_Run_2845
  • 246817 11/05/08 05:36 Jin Might be able to further center spot
  • 246816 11/05/08 05:29 adev BigBite WireChamber Measured Voltage change!
  • 246815 11/05/08 05:28 a-onl Start_Run_2845,Run_type=Optics,target_type=Single Carbon w/ Hole,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=1x1 raster on; optics; -3.0x4.2 in BPMa, -3.2x4.2 in BPMb
  • 246814 11/05/08 05:27 a-onl End_of_Run_2844
  • 246813 11/05/08 05:24 a-onl Start_Run_2844,Run_type=Optics,target_type=Single Carbon w/ Hole,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=1x1 raster on; optics; -3.0x4.2 in BPMa, -3.2x4.2 in BPMb
  • 246812 11/05/08 05:19 adev BigBite WireChamber Measured Voltage change!
  • 246811 11/05/08 05:10 Yi Zhang problem with spot++?
  • 246810 11/05/08 05:01 Huaibo Move target to optics
  • 246809 11/05/08 05:00 a-onl End_of_Run_2843
  • 246808 11/05/08 04:51 a-onl Start_Run_2843,Run_type=Check out,target_type=Single Carbon w/ Hole,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=4x4 raster on carbon with hole -3.0x4.2 in BPMa, -3.2x4.2 in BPMb
  • 246807 11/05/08 04:49 xqian Sieve at 2 pass with multi-carbon target
  • 246806 11/05/08 04:44 a-onl End_of_Run_2842
  • 246805 11/05/08 04:44 a-onl Start_Run_2842,Run_type=Check out,target_type=Single Carbon w/ Hole,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=4x4 raster on carbon with hole -3.0x2.2 in BPMa, -3.2x2.2 in BPMb
  • 246804 11/05/08 04:38 a-onl End_of_Run_2841
  • 246803 11/05/08 04:28 a-onl Start_Run_2841,Run_type=Check out,target_type=Single Carbon w/ Hole,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=4x4 raster on carbon with hole -3.0x2.2 in BPMa, -3.2x2.2 in BPMb
  • 246802 11/05/08 04:26 a-onl End_of_Run_2840
  • 246801 11/05/08 04:24 a-onl Start_Run_2840,Run_type=Check out,target_type=Single Carbon w/ Hole,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=4x4 raster on carbon with hole with RICH T3 only RICH HV ramping
  • 246800 11/05/08 04:19 xqian The charge asymmetry is amplified by factor of 10
  • 246799 11/05/08 04:17 a-onl End_of_Run_2839
  • 246798 11/05/08 04:16 xqian False asymmetry study
  • 246797 11/05/08 04:12 a-onl Start_Run_2839,Run_type=Check out,target_type=Single Carbon w/ Hole,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=4x4 raster on carbon with hole
  • 246796 11/05/08 04:09 a-onl End_of_Run_2838
  • 246795 11/05/08 04:09 Jin Hole on Carbon
  • 246794 11/05/08 04:05 brads spot++ BPMA_y is very different than 'Tools' BPMA_y
  • 246793 11/05/08 04:05 adev BigBite WireChamber Measured Voltage change!
  • 246792 11/05/08 04:01 adaq Radiation and Ozone snapshot
  • 246791 11/05/08 04:01 adaq Hall A Tools snapshot
  • 246790 11/05/08 04:01 adaq BigBox Magnet snapshot
  • 246789 11/05/08 04:01 adaq Beam Position and Current snapshot
  • 246788 11/05/08 04:01 adaq Beam Energy in A and C snapshot
  • 246787 11/05/08 03:54 adev BigBite WireChamber Measured Voltage change!
  • 246786 11/05/08 03:50 a-onl Start_Run_2838,Run_type=Check out,target_type=Single Carbon w/ Hole,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=asked for 4x4 raster on carbon with hole
  • 246785 11/05/08 03:48 Huaibo Move target to optics with hole
  • 246784 11/05/08 03:44 xqian vertex z for beam position
  • 246783 11/05/08 03:43 adev BigBite WireChamber Measured Voltage change!
  • 246782 11/05/08 03:43 a-onl End_of_Run_2837
  • 246781 11/05/08 03:36 adaq Vertex Z for spot++ run
  • 246780 11/05/08 03:36 a-onl Start_Run_2837,Run_type=Check out,target_type=Multi Carbon Optics/BeO,beam_e0=2.396,beam_raster=off,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=no raster, beam position checking
  • 246779 11/05/08 03:32 adev BigBite WireChamber Measured Voltage change!
  • 246778 11/05/08 03:28 a-onl End_of_Run_2836
  • 246777 11/05/08 03:23 a-onl Start_Run_2836,Run_type=Check out,target_type=Multi Carbon Optics/BeO,beam_e0=2.396,beam_raster=off,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=no raster
  • 246776 11/05/08 03:20 brads BeO snapshot with beam at -3.1, 3.2mm
  • 246775 11/05/08 03:16 Huaibo Move target to optics
  • 246774 11/05/08 03:10 adaq Carbon Hole Spot++ With Raster
  • 246773 11/05/08 03:08 a-onl End_of_Run_2835
  • 246772 11/05/08 03:03 a-onl Start_Run_2835,Run_type=Check out,target_type=Multi Carbon Optics/BeO,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=raster check, 4x4 in MCC side
  • 246771 11/05/08 02:45 a-onl End_of_Run_2834
  • 246770 11/05/08 02:39 a-onl Start_Run_2834,Run_type=Check out,target_type=Multi Carbon Optics/BeO,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=raster check, 4x4 in MCC side
  • 246769 11/05/08 02:37 xqian Update on longitudinal asymmetry
  • 246768 11/05/08 02:36 Huaibo Target moves to optics with hole
  • 246767 11/05/08 02:33 xqian Transverse asymmetry in the delta region
  • 246766 11/05/08 02:32 Huaibo Target moves to optics
  • 246765 11/05/08 02:25 a-onl End_of_Run_2833
  • 246764 11/05/08 02:22 a-onl Start_Run_2833,Run_type=Check out,target_type=Multi Carbon Optics/BeO,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=raster check, 4x4 in MCC side
  • 246763 11/05/08 02:19 a-onl End_of_Run_2832
  • 246762 11/05/08 02:16 a-onl Start_Run_2832,Run_type=Check out,target_type=Multi Carbon Optics/BeO,beam_e0=2.396,beam_raster=on,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=raster check, 4x4 in MCC side
  • 246761 11/05/08 02:06 Yi Zhang spot positions @ 2 pass
  • 246760 11/05/08 01:59 Huaibo Target moves to optics with hole
  • 246759 11/05/08 01:59 Yi Zhang spot++ after beam postion turning
  • 246758 11/05/08 01:49 a-onl End_of_Run_2831
  • 246757 11/05/08 01:48 a-onl Start_Run_2831,Run_type=Check out,target_type=Multi Carbon Optics/BeO,beam_e0=2.396,beam_raster=off,leftarm_p=1.233,leftarm_theta=46.366,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=spot++
  • 246756 11/05/08 01:35 Yi Zhang Harp Scan
  • 246755 11/05/08 01:01 a-onl End_of_Run_2830
  • 246754 11/05/08 01:01 a-onl Start_Run_2830,Run_type=Cosmics,target_type=No Target,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=junk
  • 246753 11/05/08 00:57 Huaibo Target move to optics
  • 246752 11/05/08 00:38 Jin Huang Couple Tests On Target
  • 246751 11/05/08 00:10 Jin Huang Transverse Pumping Status
  • 246750 11/05/08 00:02 gilman swing shift summary, Nov 4, 2008
  • 246749 11/05/08 00:02 adaq Beam Time Accounting
  • Tuesday

  • 246748 11/04/08 23:55 brads BB Shower thresholds (T1, T6) set to -80mV
  • 246747 11/04/08 23:08 adev BigBite WireChamber Measured Voltage change!
  • 246746 11/04/08 23:01 gilman beam coming back, HV trip
  • 246745 11/04/08 23:00 a-onl End_of_Run_2829
  • 246744 11/04/08 22:39 Yi Qiang Target wall collimators were taken out
  • 246743 11/04/08 22:37 Yi Qiang Target activities during controlled access
  • 246742 11/04/08 22:28 gilman controlled access over
  • 246741 11/04/08 22:12 gilman moved spectrometer
  • 246740 11/04/08 21:48 brads Closed valve on 3He bottle
  • 246739 11/04/08 21:46 Jin Huang Vertical Pumping Laser Fiber Status
  • 246738 11/04/08 21:09 gilman Ready to move spectrometer to 46.37 degr
  • 246737 11/04/08 20:19 Techs Q1 slow ramp rate
  • 246736 11/04/08 20:08 gilman Q1 problem
  • 246735 11/04/08 20:01 adaq Radiation and Ozone snapshot
  • 246734 11/04/08 20:01 adaq Hall A Tools snapshot
  • 246733 11/04/08 20:01 adaq BigBox Magnet snapshot
  • 246732 11/04/08 20:01 adaq Beam Position and Current snapshot
  • 246731 11/04/08 20:00 adaq Beam Energy in A and C snapshot
  • 246730 11/04/08 19:42 a-onl Start_Run_2829,Run_type=Cosmics,target_type=No Target,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic check of shower with T1 and T6
  • 246729 11/04/08 19:40 a-onl Start_Run_2828,Run_type=Cosmics,target_type=No Target,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=junk
  • 246728 11/04/08 19:40 a-onl End_of_Run_2828
  • 246727 11/04/08 19:31 gilman setting spectrometer to + polarity, Q1problem
  • 246726 11/04/08 18:36 V. Sulkosky Re: Restored to Previous ADC channels for S2m
  • 246725 11/04/08 18:33 brads LHRS switched to POSITIVE polarity
  • 246724 11/04/08 18:30 brads Center beam on BeO cross hairs, NOT holey carbon
  • 246723 11/04/08 18:29 xiaohui target moved to empty
  • 246722 11/04/08 18:28 brads New target ladder installed. NO upstream C foil
  • 246721 11/04/08 18:25 xiaohui Target moved to reference cell
  • 246720 11/04/08 18:16 a-onl Start_Run_2827,Run_type=Cosmics,target_type=No Target,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH Pedestal test with T3
  • 246719 11/04/08 18:09 a-onl End_of_Run_2826
  • 246718 11/04/08 18:06 a-onl Start_Run_2826,Run_type=Cosmics,target_type=No Target,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH Pedestal test with T3
  • 246717 11/04/08 18:03 a-onl End_of_Run_2825
  • 246716 11/04/08 17:59 a-onl Start_Run_2825,Run_type=Cosmics,target_type=No Target,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=ADC test with T3
  • 246715 11/04/08 17:54 a-onl End_of_Run_2824
  • 246714 11/04/08 17:49 a-onl Start_Run_2824,Run_type=Cosmics,target_type=No Target,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=ADC test with T3
  • 246713 11/04/08 17:45 a-onl End_of_Run_2823
  • 246712 11/04/08 17:41 adev BigBite WireChamber Chamber Turn on/off!
  • 246711 11/04/08 17:40 a-onl Start_Run_2823,Run_type=Cosmics,target_type=No Target,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=ADC test with T3
  • 246710 11/04/08 17:40 xiaohui target IOC rebooted
  • 246709 11/04/08 17:38 a-onl End_of_Run_2822
  • 246708 11/04/08 17:38 a-onl Start_Run_2822,Run_type=Cosmics,target_type=No Target,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=pedestal test with T8
  • 246707 11/04/08 17:38 Techs Polarity change
  • 246706 11/04/08 17:33 a-onl End_of_Run_2821
  • 246705 11/04/08 17:31 a-onl Start_Run_2821,Run_type=Cosmics,target_type=No Target,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=pedestal test with T8
  • 246704 11/04/08 17:31 adev BigBite WireChamber Chamber Turn on/off!
  • 246703 11/04/08 17:22 a-onl End_of_Run_2820
  • 246702 11/04/08 17:20 a-onl Start_Run_2820,Run_type=Cosmics,target_type=No Target,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=pedestal test with T8
  • 246701 11/04/08 17:19 a-onl End_of_Run_2819
  • 246700 11/04/08 17:18 a-onl Start_Run_2819,Run_type=Cosmics,target_type=No Target,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=pedestal test with T8
  • 246699 11/04/08 17:16 a-onl End_of_Run_2818
  • 246698 11/04/08 17:16 a-onl Start_Run_2818,Run_type=Cosmics,target_type=No Target,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=pedestal test with T8
  • 246697 11/04/08 17:12 a-onl End_of_Run_2817
  • 246696 11/04/08 17:07 a-onl Start_Run_2817,Run_type=Cosmics,target_type=No Target,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=pedestal test with T3
  • 246695 11/04/08 16:48 a-onl End_of_Run_2816
  • 246694 11/04/08 16:47 brads MWDC bubblers still bubbling
  • 246693 11/04/08 16:46 brads BB Sieve IN
  • 246692 11/04/08 16:46 a-onl Start_Run_2816,Run_type=Cosmics,target_type=No Target,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=pedestal test with T3
  • 246691 11/04/08 16:45 a-onl End_of_Run_2815
  • 246690 11/04/08 16:40 a-onl Start_Run_2815,Run_type=Cosmics,target_type=No Target,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text= pedestal test with T3
  • 246689 11/04/08 16:28 adev BigBite GUI Stopped!
  • 246688 11/04/08 16:24 a-onl End_of_Run_2814
  • 246687 11/04/08 16:23 a-onl Start_Run_2814,Run_type=Cosmics,target_type=No Target,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=pulser pedestal test with T8
  • 246685 11/04/08 16:18 adev BigBite GUI Stopped!
  • 246684 11/04/08 16:16 a-onl End_of_Run_2813
  • 246683 11/04/08 16:15 a-onl Start_Run_2813,Run_type=Cosmics,target_type=No Target,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=pulser pedestal test with T8
  • 246682 11/04/08 16:13 a-onl End_of_Run_2812
  • 246681 11/04/08 16:07 a-onl Start_Run_2812,Run_type=Cosmics,target_type=No Target,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Wire Chamber Test BigBite!
  • 246680 11/04/08 16:01 E. Schulte Day Shift Summary, November 4, 2008: Election Day!
  • 246679 11/04/08 16:00 adaq Beam Time Accounting
  • 246678 11/04/08 15:59 brads Damage probably occured between 1am--3am, Monday Nov 3
  • 246677 11/04/08 15:41 a-onl Start_Run_2811,Run_type=Cosmics,target_type=No Target,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH pedestal
  • 246676 11/04/08 15:41 a-onl End_of_Run_2811
  • 246675 11/04/08 15:40 a-onl Start_Run_2810,Run_type=Cosmics,target_type=No Target,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH pedestal
  • 246674 11/04/08 15:40 a-onl End_of_Run_2810
  • 246673 11/04/08 15:40 a-onl End_of_Run_2810
  • 246672 11/04/08 15:29 brads Damaged target ladder
  • 246671 11/04/08 15:17 Kalyan database updated
  • 246670 11/04/08 15:14 a-onl Start_Run_2809,Run_type=Cosmics,target_type=No Target,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH pedestal
  • 246669 11/04/08 15:12 a-onl End_of_Run_2808
  • 246668 11/04/08 15:11 xqian BigBite Wire Chamber Update
  • 246667 11/04/08 15:08 a-onl Start_Run_2808,Run_type=Cosmics,target_type=No Target,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH DAQ test with condition on Track and Hold
  • 246666 11/04/08 15:07 a-onl End_of_Run_2807
  • 246665 11/04/08 15:03 a-onl Start_Run_2807,Run_type=Cosmics,target_type=No Target,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH DAQ test with condition on Track and Hold
  • 246664 11/04/08 15:02 a-onl End_of_Run_2806
  • 246663 11/04/08 14:59 a-onl Start_Run_2806,Run_type=Cosmics,target_type=No Target,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH DAQ test with condition on Track and Hold
  • 246662 11/04/08 14:58 a-onl End_of_Run_2805
  • 246661 11/04/08 14:53 a-onl Start_Run_2805,Run_type=Cosmics,target_type=No Target,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH DAQ test with condition on Track and Hold
  • 246660 11/04/08 14:49 a-onl End_of_Run_2804
  • 246659 11/04/08 14:48 adev BigBite WireChamber Measured Voltage change!
  • 246658 11/04/08 14:43 a-onl Start_Run_2804,Run_type=Cosmics,target_type=No Target,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH DAQ test with condition on Track and Hold
  • 246657 11/04/08 14:42 a-onl End_of_Run_2803
  • 246656 11/04/08 14:37 a-onl Start_Run_2803,Run_type=Cosmics,target_type=No Target,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=RICH DAQ
  • 246655 11/04/08 14:31 A. Camsonne RICH DAQ test
  • 246654 11/04/08 14:30 a-onl End_of_Run_2802
  • 246653 11/04/08 13:43 a-onl Start_Run_2802,Run_type=Cosmics,target_type=No Target,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BigBite Wire Chamber calibration test!!!
  • 246652 11/04/08 13:38 a-onl End_of_Run_2801
  • 246651 11/04/08 13:37 a-onl Start_Run_2801,Run_type=Cosmics,target_type=No Target,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test!!!!
  • 246650 11/04/08 13:33 adev BigBite WireChamber Threshold change!
  • 246649 11/04/08 13:32 E. Schulte Tools Screen
  • 246648 11/04/08 13:23 adev BigBite WireChamber Threshold change!
  • 246647 11/04/08 13:12 adev BigBite WireChamber Measured Current change!
  • 246646 11/04/08 13:09 a-onl End_of_Run_2800
  • 246645 11/04/08 13:09 a-onl Start_Run_2800,Run_type=Cosmics,target_type=No Target,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test!!!!
  • 246644 11/04/08 13:01 a-onl End_of_Run_2799
  • 246643 11/04/08 13:00 a-onl Start_Run_2799,Run_type=Cosmics,target_type=No Target,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test!!!!
  • 246642 11/04/08 12:56 a-onl End_of_Run_2798
  • 246641 11/04/08 12:54 a-onl Start_Run_2798,Run_type=Cosmics,target_type=No Target,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Test!!!!
  • 246640 11/04/08 12:51 adev BigBite WireChamber Threshold change!
  • 246639 11/04/08 12:35 brads Gas shed ethane/argon mix line connected to BB MWDCs (ethanol bubbler bypassed)
  • 246638 11/04/08 12:31 brads Hall in Controlled Access since 9am (late entry)
  • 246637 11/04/08 12:28 R. Michaels R-HRS ROC1 2V busted
  • 246636 11/04/08 12:13 R. Michaels R-HRS work
  • 246635 11/04/08 12:03 V. Sulkosky Run 2794 was a Pedestal Run
  • 246634 11/04/08 12:00 adaq Radiation and Ozone snapshot
  • 246633 11/04/08 12:00 adaq Hall A Tools snapshot
  • 246632 11/04/08 12:00 adaq BigBox Magnet snapshot
  • 246631 11/04/08 12:00 adaq Beam Position and Current snapshot
  • 246630 11/04/08 12:00 adaq Beam Energy in A and C snapshot
  • 246629 11/04/08 12:00 a-onl End_of_Run_2797
  • 246628 11/04/08 11:57 a-onl Start_Run_2797,Run_type=Cosmics,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=pedestal run with T8 pulser test!!!
  • 246627 11/04/08 11:57 a-onl End_of_Run_2796
  • 246626 11/04/08 11:53 a-onl Start_Run_2796,Run_type=Cosmics,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=pedestal run with T8 pulser test!!!
  • 246625 11/04/08 11:52 a-onl End_of_Run_2795
  • 246624 11/04/08 11:52 a-onl Start_Run_2795,Run_type=Cosmics,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=pedestal run with T8 pulser
  • 246623 11/04/08 11:39 a-onl End_of_Run_2794
  • 246622 11/04/08 11:38 a-onl Start_Run_2794,Run_type=Cosmics,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=pedestal run with T8 pulser
  • 246621 11/04/08 11:22 V. Sulkosky New Analyzer Directory Created
  • 246620 11/04/08 11:18 V. Sulkosky Moved S2m Right ADC Signals
  • 246619 11/04/08 11:12 a-onl Start_Run_2793,Run_type=Cosmics,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=pedestal run with T8 pulser
  • 246618 11/04/08 11:01 R. Michaels Re: TIR helicity (L-HRS)
  • 246617 11/04/08 10:41 a-onl Start_Run_2792,Run_type=Cosmics,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmics for BB and LHRS for BOB michaels
  • 246616 11/04/08 10:40 a-onl End_of_Run_2791
  • 246615 11/04/08 10:19 Chiranjib LaserS off
  • 246614 11/04/08 10:16 a-onl Start_Run_2791,Run_type=Cosmics,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmics for BB and LHRS
  • 246613 11/04/08 10:13 a-onl Start_Run_2790,Run_type=Cosmics,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=junk
  • 246612 11/04/08 10:13 a-onl End_of_Run_2790
  • 246611 11/04/08 10:10 A. Camsonne Compton computer crashed
  • 246610 11/04/08 09:25 A. Camsonne Compton broadcast handler issue
  • 246609 11/04/08 09:19 Lingyan Zhu target: move to empty
  • 246608 11/04/08 09:15 A. Camsonne Compton centra crystal check
  • 246607 11/04/08 09:07 Jin Huang NMR -> Pol. =52% at end of transverse run
  • 246606 11/04/08 08:50 a-onl End_of_Run_2789
  • 246605 11/04/08 08:26 adaq Beam Time Accounting
  • 246604 11/04/08 08:25 a-onl Start_Run_2789,Run_type=Production,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=delta run, transverse- pumping; Happx run#:30372
  • 246603 11/04/08 08:22 a-onl End_of_Run_2788
  • 246602 11/04/08 08:10 Jin Huang Screen Shot for Transverse Minus Spin State
  • 246601 11/04/08 08:05 Yi Zhang Owl Shift summary
  • 246600 11/04/08 08:05 Jin Huang Successful Manual Spin Flip (Pol. = 50%)
  • 246599 11/04/08 08:03 zhangyi halla main tools
  • 246598 11/04/08 07:41 a-onl Start_Run_2788,Run_type=Production,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=delta run, transverse- pumping; Happx run#:30371
  • 246597 11/04/08 07:35 samuel Turn on lasers #1,3,4,5,6
  • 246596 11/04/08 07:26 samuel turn off lasers#1,3,4,5,6
  • 246595 11/04/08 07:22 a-onl End_of_Run_2787
  • 246594 11/04/08 07:02 a-onl Start_Run_2787,Run_type=Production,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=delta run, transverse+ pumping; Happx run#:30370
  • 246593 11/04/08 07:00 Huaibo G0 helicity change
  • 246592 11/04/08 06:51 a-onl End_of_Run_2786
  • 246591 11/04/08 06:39 a-onl Start_Run_2786,Run_type=Production,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=delta run, transverse+ pumping; Happx run#:30369
  • 246590 11/04/08 06:37 Jin Run#2783 is NOT junk run
  • 246589 11/04/08 06:37 a-onl End_of_Run_2785
  • 246588 11/04/08 06:36 a-onl Start_Run_2785,Run_type=Production,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=delta run, transverse+ pumping;
  • 246587 11/04/08 06:35 a-onl End_of_Run_2784
  • 246586 11/04/08 06:34 a-onl Start_Run_2784,Run_type=Production,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=delta run, transverse+ pumping;; Junk run
  • 246585 11/04/08 06:27 a-onl End_of_Run_2783
  • 246584 11/04/08 06:06 Jin NMR Measurement, 49%
  • 246583 11/04/08 06:02 a-onl Start_Run_2783,Run_type=Production,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=delta run, transverse+ pumping;; Junk run
  • 246582 11/04/08 05:57 a-onl Start_Run_2782,Run_type=Production,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=delta run, transverse+ pumping;; Junk run
  • 246581 11/04/08 05:49 a-onl End_of_Run_2781
  • 246580 11/04/08 05:43 Moffit Beam Charge Asymmetry Feedback on.
  • 246579 11/04/08 05:27 a-onl Start_Run_2781,Run_type=Production,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=delta run, transverse+ pumping
  • 246578 11/04/08 05:25 a-onl End_of_Run_2780
  • 246577 11/04/08 04:59 a-onl Start_Run_2780,Run_type=Production,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=delta run, transverse+ pumping
  • 246576 11/04/08 04:58 a-onl End_of_Run_2779
  • 246575 11/04/08 04:19 a-onl Start_Run_2779,Run_type=Production,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=delta run, transverse+ pumping
  • 246574 11/04/08 04:12 samuel move to He3 polarized target position
  • 246573 11/04/08 04:06 Yi Zhang different VDC response in delta ragion
  • 246572 11/04/08 04:00 adaq Radiation and Ozone snapshot
  • 246571 11/04/08 04:00 adaq Hall A Tools snapshot
  • 246570 11/04/08 04:00 adaq BigBox Magnet snapshot
  • 246569 11/04/08 04:00 adaq Beam Position and Current snapshot
  • 246568 11/04/08 04:00 adaq Beam Energy in A and C snapshot
  • 246567 11/04/08 03:36 Jin Huang Target Long -> Trans
  • 246566 11/04/08 03:28 samuel move target to optics position
  • 246565 11/04/08 03:20 a-onl End_of_Run_2778
  • 246564 11/04/08 03:17 a-onl Start_Run_2778,Run_type=Production,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=delta run, transverse+ pumping
  • 246563 11/04/08 03:15 a-onl End_of_Run_2777
  • 246562 11/04/08 03:04 a-onl Start_Run_2777,Run_type=Production,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=0.85,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=delta run, transverse+ pumping
  • 246561 11/04/08 02:45 samuel move target to He3 cell position.
  • 246560 11/04/08 02:43 samuel move to optics position and put lasers#1,3,4,5,6 on
  • 246559 11/04/08 02:37 a-onl End_of_Run_2776
  • 246558 11/04/08 02:35 a-onl Start_Run_2776,Run_type=Production,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=test
  • 246557 11/04/08 02:33 a-onl End_of_Run_2775
  • 246556 11/04/08 02:33 a-onl Start_Run_2775,Run_type=Production,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Junk!!!
  • 246555 11/04/08 02:26 xqian With Helicity change sign
  • 246554 11/04/08 02:25 samuel position for making NMR pol. measurement.
  • 246553 11/04/08 02:23 a-onl End_of_Run_2774
  • 246552 11/04/08 01:45 a-onl Start_Run_2774,Run_type=Production,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Longitudinal polarized 3He
  • 246551 11/04/08 01:44 a-onl End_of_Run_2773
  • 246550 11/04/08 01:11 a-onl Start_Run_2773,Run_type=Production,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Longitudinal polarized 3He
  • 246549 11/04/08 01:07 samuel move to He3 cell
  • 246548 11/04/08 01:02 samuel move to pickup coil position
  • 246547 11/04/08 01:01 xqian 3He pressure curve makes sense
  • 246546 11/04/08 00:58 a-onl End_of_Run_2772
  • 246545 11/04/08 00:47 xqian N2 pressure curve make sense.
  • 246544 11/04/08 00:47 xqian N2 pressure curve make sense.
  • 246543 11/04/08 00:45 Jin Temporary Golden Root File Run#2769 for the Night
  • 246542 11/04/08 00:18 a-onl Start_Run_2772,Run_type=Production,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Longitudinal polarized 3He
  • 246541 11/04/08 00:17 a-onl End_of_Run_2771
  • 246540 11/04/08 00:10 xqian more data from Nov. 3rd
  • 246539 11/04/08 00:01 adaq Beam Time Accounting
  • 246538 11/04/08 00:01 gilman swing shift summary Nov 3, 2008
  • Monday

  • 246537 11/03/08 23:38 yez NMR Measurement: 42.1%
  • 246536 11/03/08 23:35 a-onl Start_Run_2771,Run_type=Production,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Longitudinal polarized 3He
  • 246535 11/03/08 23:33 V. Sulkosky BCM in End-of-Run File
  • 246534 11/03/08 23:30 a-onl End_of_Run_2770
  • 246533 11/03/08 23:03 Kalyan Re: BCM gated scaler data are non-zero, but look bogus too
  • 246532 11/03/08 22:47 a-onl Start_Run_2770,Run_type=Production,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Longitudinal polarized 3He
  • 246531 11/03/08 22:47 xqian About the BCM
  • 246530 11/03/08 22:46 a-onl End_of_Run_2769
  • 246529 11/03/08 22:44 Jin Re: BCM gated scaler data are non-zero, but look bogus too
  • 246528 11/03/08 22:26 yez Target Moved to Polarized He3 before run 2769
  • 246527 11/03/08 22:16 brads BCM gated scaler data are non-zero, but look bogus too
  • 246526 11/03/08 22:01 a-onl Start_Run_2769,Run_type=Production,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Longitudinal polarized 3He
  • 246525 11/03/08 21:57 gilman wave plate inserted
  • 246524 11/03/08 21:50 yez Corrected: Target Moved to Pickup Coils for NMR Measurement
  • 246523 11/03/08 21:46 yez Target Moved to He3, Longitudinal+, NMR Measurement=45.4%
  • 246522 11/03/08 21:30 gilman pressure curves done
  • 246521 11/03/08 21:26 a-onl End_of_Run_2768
  • 246520 11/03/08 21:15 a-onl Start_Run_2768,Run_type=Check out,target_type=Reference Cell: 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Presure curve measurement 135psig
  • 246519 11/03/08 21:11 a-onl End_of_Run_2767
  • 246518 11/03/08 21:02 a-onl Start_Run_2767,Run_type=Check out,target_type=Reference Cell: 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Presure curve measurement 105psig
  • 246517 11/03/08 21:01 a-onl End_of_Run_2766
  • 246516 11/03/08 20:53 Jin Target Signal
  • 246515 11/03/08 20:50 a-onl Start_Run_2766,Run_type=Check out,target_type=Reference Cell: 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Presure curve measurement 75psig
  • 246514 11/03/08 20:48 a-onl End_of_Run_2765
  • 246513 11/03/08 20:43 xqian Polarized 3He run today will need special treatment
  • 246512 11/03/08 20:38 a-onl Start_Run_2765,Run_type=Check out,target_type=Reference Cell: 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Presure curve measurement 45psig
  • 246511 11/03/08 20:37 a-onl End_of_Run_2764
  • 246510 11/03/08 20:26 a-onl Start_Run_2764,Run_type=Check out,target_type=Reference Cell: 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Presure curve measurement 15psig
  • 246509 11/03/08 20:23 a-onl End_of_Run_2763
  • 246508 11/03/08 20:12 a-onl Start_Run_2763,Run_type=Check out,target_type=Reference Cell: 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Presure curve measurement 0torr
  • 246507 11/03/08 20:08 yez Target Moved back to Reference Cell.
  • 246506 11/03/08 20:04 yez Target Moved to He3 Cell for presure carve measurement.
  • 246505 11/03/08 20:01 adaq Radiation and Ozone snapshot
  • 246504 11/03/08 20:01 adaq Hall A Tools snapshot
  • 246503 11/03/08 20:01 adaq BigBox Magnet snapshot
  • 246501 11/03/08 20:01 a-onl End_of_Run_2762
  • 246502 11/03/08 20:01 adaq Beam Position and Current snapshot
  • 246500 11/03/08 20:01 adaq Beam Energy in A and C snapshot
  • 246499 11/03/08 20:01 yez NMR Mesearment, 43.3%
  • 246498 11/03/08 19:55 gilman Hall A tools for shift checklist
  • 246497 11/03/08 19:49 a-onl Start_Run_2762,Run_type=Check out,target_type=Reference Cell: N2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Presure curve measurement 15psig
  • 246496 11/03/08 19:44 a-onl End_of_Run_2761
  • 246495 11/03/08 19:37 a-onl Start_Run_2761,Run_type=Check out,target_type=Reference Cell: N2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Presure curve measurement 788torr
  • 246494 11/03/08 19:34 a-onl End_of_Run_2760
  • 246493 11/03/08 19:24 a-onl Start_Run_2760,Run_type=Check out,target_type=Reference Cell: N2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Presure curve measurement 621torr
  • 246492 11/03/08 19:22 a-onl End_of_Run_2759
  • 246491 11/03/08 19:06 a-onl Start_Run_2759,Run_type=Check out,target_type=Reference Cell: N2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Presure curve measurement 404torr
  • 246490 11/03/08 19:04 a-onl End_of_Run_2758
  • 246489 11/03/08 18:54 xqian BigBite Wire Chamber Noise Problem
  • 246488 11/03/08 18:50 a-onl Start_Run_2758,Run_type=Check out,target_type=Reference Cell: N2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Presure curve measurement 197torr
  • 246487 11/03/08 18:47 a-onl End_of_Run_2757
  • 246486 11/03/08 18:35 a-onl Start_Run_2757,Run_type=Check out,target_type=Reference Cell: N2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Presure curve measurement 104torr
  • 246485 11/03/08 18:31 a-onl End_of_Run_2756
  • 246484 11/03/08 18:12 V. Sulkosky Fine Tune of S2m HVs
  • 246483 11/03/08 18:07 a-onl Start_Run_2756,Run_type=Check out,target_type=Reference Cell: Empty,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Presure curve measurement vacummn
  • 246482 11/03/08 17:55 a-onl End_of_Run_2755
  • 246481 11/03/08 17:55 a-onl Start_Run_2755,Run_type=Check out,target_type=Reference Cell: Empty,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Presure curve measurement
  • 246480 11/03/08 17:51 yez Corretecd: Last NMR measurement result is 33.64%
  • 246479 11/03/08 17:50 Jin Huang replay cratemap update
  • 246478 11/03/08 17:42 yez Target at Reference Cells, Filed rotated to Transvers+, Laser on
  • 246477 11/03/08 17:20 NMR Mesearment, target location pickup, RF vertical, pickup transverse
  • 246476 11/03/08 16:59 a-onl End_of_Run_2754
  • 246475 11/03/08 16:59 a-onl Start_Run_2754,Run_type=Check out,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Wire Chamber TEst
  • 246474 11/03/08 16:57 a-onl End_of_Run_2753
  • 246473 11/03/08 16:04 a-onl Start_Run_2753,Run_type=Calibration,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=long. polarization
  • 246472 11/03/08 16:03 E. Schulte 3He Longitudinal Polarization
  • 246471 11/03/08 16:01 adaq Beam Time Accounting
  • 246470 11/03/08 16:01 E. Schulte Day Shift Summary, November 3, 2008
  • 246469 11/03/08 16:00 E. Schulte 3He Longitudinal Polarization
  • 246468 11/03/08 15:55 Katich NMR measurement
  • 246467 11/03/08 15:44 a-onl End_of_Run_2752
  • 246466 11/03/08 15:38 Jack Gas Shed air flow
  • 246465 11/03/08 15:33 A. Camsonne Event size with RICH run 2721
  • 246464 11/03/08 15:25 a-onl Start_Run_2752,Run_type=Calibration,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=long. polarization
  • 246463 11/03/08 15:13 E. Schulte N. Okay asked about the "pickup coil" target position and controlled access procedures
  • 246462 11/03/08 15:05 a-onl End_of_Run_2751
  • 246461 11/03/08 14:34 a-onl Start_Run_2751,Run_type=Calibration,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=long. polarization
  • 246460 11/03/08 14:34 a-onl End_of_Run_2750
  • 246459 11/03/08 14:25 E. Schulte Bad Fan
  • 246458 11/03/08 14:17 Katich EPR measurement
  • 246457 11/03/08 14:17 V. Sulkosky Pedestal Shifts for S2m Right ADC Channels
  • 246456 11/03/08 14:03 Katich NMR measurement
  • 246455 11/03/08 14:00 a-onl Start_Run_2750,Run_type=Calibration,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=long. polarization
  • 246454 11/03/08 13:48 a-onl End_of_Run_2749
  • 246453 11/03/08 13:45 E. Schulte Gas Shed Low Pressure Alarm
  • 246452 11/03/08 13:43 E. Schulte Gas Shed Air Flow Alarm
  • 246451 11/03/08 13:33 a-onl Start_Run_2749,Run_type=Calibration,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=long. polarization
  • 246450 11/03/08 13:32 a-onl End_of_Run_2748
  • 246449 11/03/08 13:03 a-onl Start_Run_2748,Run_type=Calibration,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=long. polarization
  • 246448 11/03/08 13:02 a-onl End_of_Run_2747
  • 246447 11/03/08 12:52 mercado Calibrated Lumi HVs
  • 246446 11/03/08 12:29 a-onl Start_Run_2747,Run_type=Calibration,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=long. polarization
  • 246445 11/03/08 12:28 a-onl End_of_Run_2746
  • 246444 11/03/08 12:20 E. Schulte Target & Reference Cell Readbacks
  • 246443 11/03/08 12:11 E. Schulte Day Shift Tools Screen
  • 246442 11/03/08 12:07 a-onl Start_Run_2746,Run_type=Calibration,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=long. polarization
  • 246441 11/03/08 12:01 adaq Radiation and Ozone snapshot
  • 246440 11/03/08 12:01 adaq Hall A Tools snapshot
  • 246439 11/03/08 12:01 adaq BigBox Magnet snapshot
  • 246438 11/03/08 12:01 adaq Beam Position and Current snapshot
  • 246437 11/03/08 12:01 adaq Beam Energy in A and C snapshot
  • 246436 11/03/08 12:00 Katich Target -> 3He Cell in beam
  • 246435 11/03/08 11:47 Lindgren laser off
  • 246434 11/03/08 11:39 jpchen Helium flow rate to beam line Be window
  • 246433 11/03/08 11:33 brads Harp scan
  • 246432 11/03/08 11:19 adaq Beam Time Accounting
  • 246431 11/03/08 11:14 mercado Setting Lumi HVs
  • 246430 11/03/08 11:14 V. Sulkosky Shortened Fast EPICS Logger Readout
  • 246429 11/03/08 11:10 a-onl End_of_Run_2745
  • 246428 11/03/08 11:10 a-onl Start_Run_2745,Run_type=Test,target_type=No Target,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=test
  • 246427 11/03/08 11:01 ludwig Beam Time Accounting
  • 246426 11/03/08 10:57 brads Added shielding around BB DAQ network switch
  • 246425 11/03/08 10:54 brads 4x4 (MCC units) raster on BeO snapshot
  • 246424 11/03/08 10:54 E. Schulte Back to Beam Permit, MCC to Deliver Beam
  • 246423 11/03/08 10:47 brads/kalyan MWDC power supplies good, but cards are oscillating...
  • 246422 11/03/08 10:40 a-onl End_of_Run_2743
  • 246421 11/03/08 10:39 brads Ethane/argon for MWDC is at 300psi
  • 246420 11/03/08 10:38 a-onl Start_Run_2743,Run_type=Test,target_type=No Target,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Pedestal run with T8 only
  • 246419 11/03/08 10:36 R. Michaels Lumi saturated
  • 246418 11/03/08 09:48 E. Schulte Controlled Access to Check Power Supplies
  • 246417 11/03/08 09:28 a-onl End_of_Run_2742
  • 246416 11/03/08 09:24 brads Set MWDC 2, 3 threshold to 8V for now
  • 246415 11/03/08 09:11 adaq Noise in BB wirechamber 2 and 3
  • 246414 11/03/08 09:01 brads High deadtime issue connected to MWDCs
  • 246413 11/03/08 08:50 a-onl Start_Run_2742,Run_type=Test,target_type=No Target,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Threshold on chamber 2 and 3 changed to 8V
  • 246412 11/03/08 08:39 E. Schulte HRS-L Q3 cryo event picture
  • 246411 11/03/08 08:32 E. Schulte HRSL Q3 Helium Liquid Level fault
  • 246410 11/03/08 08:31 a-onl End_of_Run_2741
  • 246409 11/03/08 08:14 ycwang re :A1 peak
  • 246408 11/03/08 08:11 adaq Beam Time Accounting
  • 246407 11/03/08 08:07 Ibrahim Owl Shift Summary
  • 246406 11/03/08 07:51 a-onl Start_Run_2741,Run_type=Calibration,target_type=Reference Cell: N2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=4x4 raster at 5uA, with the wrong T7 stopped, pressure = 0 torr
  • 246405 11/03/08 07:49 a-onl End_of_Run_2740
  • 246404 11/03/08 07:47 a-onl Start_Run_2740,Run_type=Calibration,target_type=Reference Cell: N2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=4x4 raster at 5uA, with the wrong T7 stopped, pressure = 0 torr.Just For test
  • 246403 11/03/08 07:45 a-onl End_of_Run_2739
  • 246402 11/03/08 07:40 a-onl Start_Run_2739,Run_type=Calibration,target_type=Reference Cell: N2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=4x4 raster at 5uA, with the wrong T7 stopped, pressure = 0 torr.
  • 246401 11/03/08 07:37 Chiranjib target STATUS
  • 246400 11/03/08 07:34 V. Sulkosky Re: A1 HV
  • 246399 11/03/08 07:32 a-onl Start_Run_2738,Run_type=Calibration,target_type=Reference Cell: N2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=4x4 raster at 5uA, with the wrong T7 stopped, pressure = 0 torr.
  • 246398 11/03/08 07:22 ycwang RICH main HV tripped during the beam tuning
  • 246397 11/03/08 07:13 ycwang A1 peak corresponding to HV
  • 246396 11/03/08 07:13 samuel on beam line is reference cell with 0 pressure N2
  • 246395 11/03/08 07:12 Chiranjib NMR measurement
  • 246394 11/03/08 07:05 samuel move to He3 cell position
  • 246393 11/03/08 06:59 Ibrahim Deadtime is High for Production
  • 246392 11/03/08 06:57 samuel move to pickup coil position
  • 246391 11/03/08 06:55 a-onl End_of_Run_2737
  • 246390 11/03/08 06:38 Ibrahim Q3 is Up again.
  • 246389 11/03/08 06:26 a-onl Start_Run_2737,Run_type=Production,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Longitudinal pumping, 4x4 raster at 5uA, with the wrong T7 stopped.
  • 246388 11/03/08 06:20 a-onl End_of_Run_2736
  • 246387 11/03/08 06:18 a-onl Start_Run_2736,Run_type=Production,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Longitudinal pumping, 4x4 raster at 5uA, with the wrong T7 stopped.
  • 246386 11/03/08 06:07 Ibrahim Q3 Tripped Again!
  • 246385 11/03/08 06:04 xqian We See 3He polarization
  • 246384 11/03/08 06:00 Chiranjib He-3 elastic peak
  • 246383 11/03/08 05:59 xqian T7 rates goes extremely high suddenlly
  • 246382 11/03/08 05:56 a-onl End_of_Run_2735
  • 246381 11/03/08 05:43 a-onl Start_Run_2735,Run_type=Production,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Longitudinal pumping, 4x4 raster at 5uA
  • 246380 11/03/08 05:42 a-onl End_of_Run_2734
  • 246379 11/03/08 05:12 a-onl Start_Run_2734,Run_type=Production,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Longitudinal polarization, 4x4 raster at 5uA
  • 246378 11/03/08 05:02 Chiranjib Raster check
  • 246377 11/03/08 04:57 a-onl End_of_Run_2733
  • 246376 11/03/08 04:57 Chiranjib TARGET ACTIVITY
  • 246375 11/03/08 04:54 a-onl Start_Run_2733,Run_type=Check out,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Beam Check 4x4 2uA, T3=4.9e3
  • 246374 11/03/08 04:53 a-onl End_of_Run_2732
  • 246373 11/03/08 04:47 a-onl Start_Run_2732,Run_type=Check out,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Beam Check 3x3 2uA, T3=4.9e3
  • 246372 11/03/08 04:42 samuel move to He3 Cell position
  • 246371 11/03/08 04:35 Ibrahim Q3 Helium Level
  • 246370 11/03/08 04:26 samuel move target to pickup coil position
  • 246369 11/03/08 04:26 a-onl End_of_Run_2731
  • 246368 11/03/08 04:20 a-onl Start_Run_2731,Run_type=Check out,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Beam Check 5*5 2uA, T3=5.4e3
  • 246367 11/03/08 04:20 Ibrahim 1H03B Scanner
  • 246366 11/03/08 04:12 a-onl End_of_Run_2729
  • 246365 11/03/08 04:07 a-onl Start_Run_2729,Run_type=Optics,target_type=Multi Carbon Optics/BeO,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Beam Check 4*4 2uA, T3=4.9e3
  • 246364 11/03/08 04:01 adaq Radiation and Ozone snapshot
  • 246363 11/03/08 04:01 adaq Hall A Tools snapshot
  • 246362 11/03/08 04:01 adaq BigBox Magnet snapshot
  • 246361 11/03/08 04:00 adaq Beam Position and Current snapshot
  • 246360 11/03/08 04:00 adaq Beam Energy in A and C snapshot
  • 246359 11/03/08 03:58 samuel move to he3 position
  • 246358 11/03/08 03:42 samuel move to reference cell H2(100PSI)
  • 246357 11/03/08 03:39 Ibrahim BeO Screen
  • 246356 11/03/08 03:38 Ibrahim Spot++ for Run 2728
  • 246355 11/03/08 03:36 Ibrahim Rebooted ROCs 8, 9 and 10.
  • 246354 11/03/08 03:33 a-onl End_of_Run_2728
  • 246353 11/03/08 03:32 a-onl Start_Run_2728,Run_type=Optics,target_type=Multi Carbon Optics/BeO,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=Beam Check
  • 246352 11/03/08 03:23 Ibrahim Tools Screen (Q3 is Up)
  • 246351 11/03/08 03:17 Ibrahim MCC Harp Scan Results (1H03B Scanner is not working)
  • 246350 11/03/08 03:06 Ibrahim Energy Monitor is Working
  • 246349 11/03/08 02:56 Ibrahim Q3 Tripped
  • 246348 11/03/08 02:50 Ibrahim MCC are still working on the Harp Scans
  • 246347 11/03/08 02:16 jinhuang Some discuss on Polarization loss during last Rotation-NMR combination
  • 246346 11/03/08 01:15 Ibrahim Q3 Lead Flows are Steady
  • 246345 11/03/08 01:06 Ibrahim Beam Tuning
  • 246344 11/03/08 00:35 adaq Radiation and Ozone snapshot
  • 246343 11/03/08 00:35 adaq Hall A Tools snapshot
  • 246342 11/03/08 00:35 adaq BigBox Magnet snapshot
  • 246341 11/03/08 00:35 adaq Beam Position and Current snapshot
  • 246340 11/03/08 00:35 adaq Beam Energy in A and C snapshot
  • 246339 11/03/08 00:31 xqian BCM scaler calibration
  • 246338 11/03/08 00:09 Jin Huang Target Expert Summary for Nov 02 Day-Swing Shift
  • 246337 11/03/08 00:01 adaq Beam Time Accounting
  • Sunday

  • 246336 11/02/08 23:56 gilman swing shift summary, 2 nov 2008
  • 246335 11/02/08 23:42 gilman Q3 liquid levels
  • 246334 11/02/08 23:35 gilman beam coming back
  • 246333 11/02/08 23:27 Jin Target Back to Trans and Start pumping
  • 246332 11/02/08 23:22 Jin Targeg Motion
  • 246331 11/02/08 23:01 E.Chudakov Moller results
  • 246330 11/02/08 22:54 Chiranjib TRANSVERSE PUMPING LASERS
  • 246329 11/02/08 22:52 Jin Huang Stablized Target Trans Pumping Spec (Correct photo)
  • 246328 11/02/08 22:49 Jin Huang Stablized Target Trans Pumping Spec
  • 246327 11/02/08 22:26 a-onl End_of_Run_2726
  • 246326 11/02/08 22:26 a-onl Start_Run_2726,Run_type=Test,target_type=No Target,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=longitudinal production, raster 4*4 - RICH -64 ns
  • 246325 11/02/08 22:15 Tareget Grou EPR Calibration @ Transverse Pumping
  • 246324 11/02/08 22:04 gilman lead flows
  • 246323 11/02/08 22:01 gilman Q3 helium
  • 246322 11/02/08 21:55 gilman Q3 trip
  • 246321 11/02/08 21:45 A. Camsonne Blaster test for RICH ROCs
  • 246320 11/02/08 21:26 a-onl End_of_Run_2725
  • 246319 11/02/08 21:26 xqian BCM calibration BCM1 and BCM2 vs OLO2
  • 246318 11/02/08 21:16 Target Group Target -> Trans Pumping
  • 246317 11/02/08 21:09 a-onl Start_Run_2725,Run_type=Test,target_type=No Target,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=longitudinal production, raster 4*4 - RICH -64 ns
  • 246316 11/02/08 20:45 a-onl End_of_Run_2724
  • 246315 11/02/08 20:26 a-onl Start_Run_2724,Run_type=Test,target_type=No Target,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=longitudinal production, raster 4*4 - RICH -64 ns
  • 246314 11/02/08 20:25 a-onl End_of_Run_2723
  • 246313 11/02/08 20:24 a-onl End_of_Run_2723
  • 246312 11/02/08 20:23 brads Beam energy readbacks are frozen
  • 246311 11/02/08 20:18 a-onl Start_Run_2723,Run_type=Test,target_type=No Target,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=longitudinal production, raster 4*4 - RICH +120 ns
  • 246310 11/02/08 20:17 A. Camsonne RICH RUN during Moeller with timing scan
  • 246309 11/02/08 20:17 a-onl End_of_Run_2722
  • 246308 11/02/08 20:11 a-onl Start_Run_2722,Run_type=Test,target_type=No Target,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=longitudinal production, raster 4*4 - RICH -120 ns
  • 246307 11/02/08 20:03 a-onl Start_Run_2721,Run_type=Test,target_type=No Target,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=longitudinal production, raster 4*4 - RICH
  • 246306 11/02/08 20:02 a-onl End_of_Run_2720
  • 246305 11/02/08 20:01 adaq Radiation and Ozone snapshot
  • 246304 11/02/08 20:01 adaq Hall A Tools snapshot
  • 246303 11/02/08 20:00 adaq BigBox Magnet snapshot
  • 246302 11/02/08 20:00 adaq Beam Position and Current snapshot
  • 246301 11/02/08 20:00 adaq Beam Energy in A and C snapshot
  • 246300 11/02/08 19:59 Jin Huang Laser Absorption
  • 246299 11/02/08 19:44 a-onl Start_Run_2720,Run_type=Test,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=longitudinal production, raster 4*4 - RICH DMA
  • 246298 11/02/08 19:39 Kalyan Beam RF time signal as seen by BigBite scintillator
  • 246297 11/02/08 19:38 a-onl Start_Run_2719,Run_type=Test,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=longitudinal production, raster 4*4 - RICH DMA transfer test
  • 246296 11/02/08 19:37 Chiranjib EPR measurement
  • 246295 11/02/08 19:35 Kalyan BB wirechamber turned off
  • 246294 11/02/08 19:34 gilman getting ready for Moller
  • 246293 11/02/08 19:23 jinhuang Target -> Vertical+
  • 246292 11/02/08 19:16 Chiranjib Vertical lasers turned ON
  • 246291 11/02/08 19:05 Jin Huang Target Movement for NMR, Moller; Pol.= 28%
  • 246290 11/02/08 19:01 gilman getting ready for Moller
  • 246289 11/02/08 18:57 a-onl End_of_Run_2718
  • 246288 11/02/08 18:56 gilman beam positions
  • 246287 11/02/08 18:50 Gilman shift plan followup 2
  • 246286 11/02/08 18:47 a-onl Start_Run_2718,Run_type=Production,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=longitudinal production, raster 4*4
  • 246285 11/02/08 18:46 a-onl End_of_Run_2717
  • 246284 11/02/08 18:31 gilman Moller to start at about 19:00
  • 246283 11/02/08 18:15 a-onl Start_Run_2717,Run_type=Production,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=longitudinal production, raster 4*4
  • 246282 11/02/08 18:11 jinhuang Target Movement for NMR; Pol.= 31%
  • 246281 11/02/08 18:01 a-onl End_of_Run_2716
  • 246280 11/02/08 17:47 a-onl Start_Run_2716,Run_type=Production,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=longitudinal production, raster 4*4
  • 246279 11/02/08 17:46 a-onl End_of_Run_2715
  • 246278 11/02/08 17:40 gilman shift plan followup
  • 246277 11/02/08 17:13 a-onl Start_Run_2715,Run_type=Production,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=longitudinal production, raster 4*4
  • 246276 11/02/08 17:11 a-onl End_of_Run_2714
  • 246275 11/02/08 17:09 gilman shift plan
  • 246274 11/02/08 17:07 gilman cryo on call going home
  • 246273 11/02/08 16:36 a-onl Start_Run_2714,Run_type=Production,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=longitudinal production, raster 4*4
  • 246272 11/02/08 16:34 a-onl End_of_Run_2713
  • 246271 11/02/08 16:29 jinhuang Before taking beam@4PM, he3 Pol. is estimated to be 38%.
  • 246270 11/02/08 16:00 adaq Beam Time Accounting
  • 246269 11/02/08 16:00 a-onl Start_Run_2713,Run_type=Test,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=longitudinal production, raster 4*4
  • 246268 11/02/08 15:59 Jin Target Move
  • 246267 11/02/08 15:58 E. Schulte Day Shift Summary, November 2, 2008
  • 246266 11/02/08 15:48 a-onl End_of_Run_2712
  • 246265 11/02/08 15:27 E. Schulte Request Cancelled
  • 246264 11/02/08 15:26 a-onl Start_Run_2712,Run_type=Test,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=longitudinal production, raster 4*4
  • 246263 11/02/08 15:12 E. Schulte Q3 Coming back online
  • 246262 11/02/08 15:10 brads Auto-snapshots of Tools, BPMs, etc started
  • 246261 11/02/08 15:08 adaq Radiation and Ozone snapshot
  • 246260 11/02/08 15:08 adaq Hall A Tools snapshot
  • 246259 11/02/08 15:08 adaq BigBox Magnet snapshot
  • 246258 11/02/08 15:08 adaq Beam Position and Current snapshot
  • 246257 11/02/08 15:08 adaq Beam Energy in A and C snapshot
  • 246256 11/02/08 14:49 E. Schulte Controlled Access Requested
  • 246255 11/02/08 14:39 E. Schulte Update on Cryo Situation
  • 246254 11/02/08 14:21 zhangyi NMR calibration mistake & polarization after 1 run
  • 246253 11/02/08 14:18 E. Schulte Just heard from Cryo On-Call
  • 246252 11/02/08 14:13 E. Schulte Polarization really about 44%
  • 246251 11/02/08 14:02 E. Schulte J. P. called, it's not so bad
  • 246250 11/02/08 14:01 E. Schulte Beam is potentially back, still have cryo problems.
  • 246249 11/02/08 14:01 brads (tech) Problem with LHe level readback on LQ3?
  • 246248 11/02/08 13:52 E. Schulte Target polarization down to 14%
  • 246247 11/02/08 13:43 eschulte Cryo-excersion tripped Q3
  • 246246 11/02/08 13:31 E. Schulte HRS-L Q3 Interlock Fault
  • 246245 11/02/08 13:14 E. Schulte Beam Down Again
  • 246244 11/02/08 13:13 a-onl End_of_Run_2711
  • 246243 11/02/08 13:06 E. Schulte BTA not recording beam use properly
  • 246242 11/02/08 13:02 E. Schulte Vince Responded
  • 246241 11/02/08 12:56 a-onl Start_Run_2711,Run_type=Test,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=longitudinal production, raster 4*4
  • 246240 11/02/08 12:55 a-onl End_of_Run_2710
  • 246239 11/02/08 12:41 E. Schulte Dead Time Estimate on DATAMON
  • 246238 11/02/08 12:29 E. Schulte Tools Screen for Day Shift
  • 246237 11/02/08 12:20 Kalyan Trigger rates vs N2 pressure in reference cell
  • 246236 11/02/08 12:17 E. Schulte Production running under way!
  • 246235 11/02/08 12:15 a-onl Start_Run_2710,Run_type=Test,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=longitudinal production, raster 4*4
  • 246234 11/02/08 12:15 brads Pass change --> 2 pass has been deferred until Tuesday
  • 246233 11/02/08 12:09 zhangyi target ->He3 cell
  • 246232 11/02/08 12:02 zhangyi target-> pick-up coil
  • 246231 11/02/08 12:00 a-onl End_of_Run_2709
  • 246230 11/02/08 11:58 a-onl Start_Run_2709,Run_type=Test,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=raster scan 6*6
  • 246229 11/02/08 11:58 E. Schulte Raster scan on 3He, 2uA Beam Current
  • 246228 11/02/08 11:56 a-onl End_of_Run_2708
  • 246227 11/02/08 11:55 a-onl Start_Run_2708,Run_type=Test,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=raster scan 5*5
  • 246226 11/02/08 11:50 a-onl End_of_Run_2707
  • 246225 11/02/08 11:49 zhangyi Temperature screen print
  • 246224 11/02/08 11:48 a-onl Start_Run_2707,Run_type=Test,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=raster scan 4*4
  • 246223 11/02/08 11:46 E. Schulte This was not the carbon optics with hole, it was the BeO one
  • 246222 11/02/08 11:45 a-onl End_of_Run_2706
  • 246221 11/02/08 11:45 zhangyi NMR Calibration for Longitudinal Polarization
  • 246220 11/02/08 11:44 a-onl Start_Run_2706,Run_type=Test,target_type=Polarized 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=raster scan 3*3
  • 246219 11/02/08 11:24 E. Schulte Spot++ Quick Check after first NMR Measurement
  • 246218 11/02/08 11:23 a-onl End_of_Run_2705
  • 246217 11/02/08 11:20 a-onl Start_Run_2705,Run_type=Test,target_type=Multi Carbon Optics/BeO,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=carbon with the hole
  • 246216 11/02/08 10:29 E. Schulte Carbon Hole Spot++ With Raster
  • 246215 11/02/08 10:29 Yi Zhang turn off laser
  • 246214 11/02/08 10:27 E. Schulte Harp Scan from MCC
  • 246213 11/02/08 10:04 a-onl End_of_Run_2704
  • 246212 11/02/08 10:01 a-onl Start_Run_2704,Run_type=Test,target_type=Single Carbon w/ Hole,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=carbon with the hole
  • 246211 11/02/08 10:01 a-onl End_of_Run_2703
  • 246210 11/02/08 09:58 a-onl Start_Run_2703,Run_type=Test,target_type=Single Carbon w/ Hole,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=carbon with the hole
  • 246209 11/02/08 09:57 a-onl End_of_Run_2702
  • 246208 11/02/08 09:54 brads Modified names in runstart.tcl target list
  • 246207 11/02/08 09:47 E. Schulte Spot++ Carbon Hole No Raster
  • 246206 11/02/08 09:45 a-onl Start_Run_2702,Run_type=Test,target_type=Single Carbon,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=carbon with the hole
  • 246205 11/02/08 09:43 a-onl End_of_Run_2701
  • 246204 11/02/08 09:26 a-onl Start_Run_2701,Run_type=Test,target_type=Single Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=carbon with the hole
  • 246203 11/02/08 09:18 a-onl End_of_Run_2700
  • 246202 11/02/08 09:16 a-onl Start_Run_2700,Run_type=Test,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=beam position check
  • 246201 11/02/08 09:15 E. Schulte MCC Called: Sending CW beam
  • 246200 11/02/08 08:24 Ibrahim Owl Shift Summary
  • 246199 11/02/08 08:23 E. Schulte MCC Called, Checking Orbits
  • 246198 11/02/08 07:56 Ibrahim No Beam
  • 246197 11/02/08 07:53 ycwang A1 high voltage adjust (Right side)
  • 246196 11/02/08 07:46 ycwang A1 high voltage adjust
  • 246195 11/02/08 07:38 Chiranjib NMR measurement and VI
  • 246194 11/02/08 07:28 adaq Beam Time Accounting
  • 246193 11/02/08 07:26 Chiranjib NMR at the end of the shift
  • 246192 11/02/08 07:19 a-onl End_of_Run_2699
  • 246191 11/02/08 07:11 adev BigBite WireChamber Measured Voltage change!
  • 246190 11/02/08 07:02 Ibrahim MCC Harp Scan Results
  • 246189 11/02/08 07:00 adev BigBite WireChamber Measured Voltage change!
  • 246188 11/02/08 06:58 Chiranjib MCC working on the beam
  • 246187 11/02/08 06:55 Ibrahim BeO Screen
  • 246186 11/02/08 06:53 brads Beam wandering over last couple hours
  • 246185 11/02/08 06:43 Chiranjib spot++ after harp scan
  • 246184 11/02/08 06:34 a-onl Start_Run_2699,Run_type=Test,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=beam position check
  • 246183 11/02/08 06:30 Chiranjib ROC 9 issue could be radiation related??
  • 246182 11/02/08 06:29 adev BigBite WireChamber Measured Voltage change!
  • 246181 11/02/08 06:25 chiranjib Harp scan
  • 246180 11/02/08 06:22 a-onl End_of_Run_2698
  • 246179 11/02/08 06:16 a-onl Start_Run_2698,Run_type=Test,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=beam position check
  • 246178 11/02/08 06:00 Chiranjib ROC 9 died again
  • 246177 11/02/08 05:58 a-onl Start_Run_2697,Run_type=Test,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=beam position check
  • 246176 11/02/08 05:55 a-onl End_of_Run_2696
  • 246175 11/02/08 05:55 Ibrahim Tools Screen
  • 246174 11/02/08 05:53 Chiranjib ROCs rebooted
  • 246173 11/02/08 05:50 a-onl Start_Run_2696,Run_type=Test,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=beam position check
  • 246172 11/02/08 05:41 Chiranjib ROC 9 problem
  • 246171 11/02/08 05:39 a-onl Start_Run_2695,Run_type=Test,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=beam position check
  • 246170 11/02/08 05:26 a-onl Start_Run_2694,Run_type=Test,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=beam position check
  • 246169 11/02/08 05:22 a-onl Start_Run_2693,Run_type=Test,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=beam position check
  • 246168 11/02/08 05:18 a-onl Start_Run_2692,Run_type=Test,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=beam position check
  • 246167 11/02/08 05:16 adev BigBite WireChamber Threshold change!
  • 246166 11/02/08 05:13 ycwang Run # 2687 2688 2689 2690 are all junks.
  • 246165 11/02/08 05:05 adev BigBite WireChamber Measured Voltage change!
  • 246164 11/02/08 04:59 Ibrahim Beam is Very Trippy (RF Problems)
  • 246163 11/02/08 04:56 chiranjib TARGET---> OPTICS
  • 246162 11/02/08 04:55 adev BigBite WireChamber Measured Voltage change!
  • 246161 11/02/08 04:44 adev BigBite WireChamber Measured Voltage change!
  • 246160 11/02/08 04:37 Ibrahim BB Wire Chambers Trips
  • 246159 11/02/08 04:33 adev BigBite WireChamber Measured Voltage change!
  • 246158 11/02/08 04:30 Chiranjib ROC 9 died again
  • 246157 11/02/08 04:23 adev BigBite WireChamber Measured Voltage change!
  • 246156 11/02/08 04:21 a-onl Start_Run_2691,Run_type=Check out,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=1700 3.0 BB ch. check
  • 246155 11/02/08 04:19 Chiranjib ROC 9 rebooted
  • 246154 11/02/08 04:13 a-onl Start_Run_2690,Run_type=Check out,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=1700 3.0 BB ch. check
  • 246153 11/02/08 04:12 adev BigBite WireChamber Measured Voltage change!
  • 246152 11/02/08 04:04 a-onl End_of_Run_2689
  • 246151 11/02/08 04:04 a-onl Start_Run_2689,Run_type=Check out,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=1700 3.0 BB ch. check
  • 246150 11/02/08 04:01 adev BigBite WireChamber Measured Voltage change!
  • 246149 11/02/08 03:59 a-onl End_of_Run_2688
  • 246148 11/02/08 03:59 a-onl Start_Run_2688,Run_type=Check out,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=1650 3.0 BB ch. check
  • 246147 11/02/08 03:54 a-onl End_of_Run_2687
  • 246146 11/02/08 03:54 a-onl Start_Run_2687,Run_type=Check out,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=junk
  • 246145 11/02/08 03:52 a-onl End_of_Run_2686
  • 246144 11/02/08 03:51 adev BigBite WireChamber Threshold change!
  • 246143 11/02/08 03:48 a-onl Start_Run_2686,Run_type=Check out,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=1650 and 3.0V ( BB Wire ch. check)
  • 246142 11/02/08 03:46 a-onl End_of_Run_2685
  • 246141 11/02/08 03:45 a-onl Start_Run_2685,Run_type=Check out,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=1650 and 3.5V ( BB Wire ch. check)
  • 246140 11/02/08 03:42 a-onl End_of_Run_2684
  • 246139 11/02/08 03:41 a-onl Start_Run_2684,Run_type=Check out,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=1650 and 4.0V ( BB Wire ch. check)
  • 246138 11/02/08 03:40 a-onl End_of_Run_2683
  • 246137 11/02/08 03:40 adev BigBite WireChamber Threshold change!
  • 246136 11/02/08 03:39 a-onl Start_Run_2683,Run_type=Check out,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=1650 and 4.5V ( BB Wire ch. check)
  • 246135 11/02/08 03:38 a-onl End_of_Run_2682
  • 246134 11/02/08 03:37 a-onl Start_Run_2682,Run_type=Check out,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=1650 and 5.0V ( BB Wire ch. check)
  • 246133 11/02/08 03:32 a-onl End_of_Run_2681
  • 246132 11/02/08 03:30 adev BigBite WireChamber Threshold change!
  • 246131 11/02/08 03:29 a-onl Start_Run_2681,Run_type=Check out,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=1550 and 5.0V ( BB Wire ch. check)
  • 246130 11/02/08 03:24 a-onl End_of_Run_2680
  • 246129 11/02/08 03:23 a-onl Start_Run_2680,Run_type=Check out,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=1550 and 4.5V ( BB Wire ch. check)
  • 246128 11/02/08 03:22 a-onl End_of_Run_2679
  • 246127 11/02/08 03:21 a-onl Start_Run_2679,Run_type=Check out,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=1550 and 4.0V ( BB Wire ch. check)
  • 246126 11/02/08 03:20 a-onl End_of_Run_2678
  • 246125 11/02/08 03:19 adev BigBite WireChamber Measured Voltage change!
  • 246124 11/02/08 03:18 a-onl Start_Run_2678,Run_type=Check out,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=1550 and 3.5V ( BB Wire ch. check)
  • 246123 11/02/08 03:17 a-onl End_of_Run_2677
  • 246122 11/02/08 03:15 a-onl Start_Run_2677,Run_type=Check out,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=1550 and 3.0V ( BB Wire ch. check)
  • 246121 11/02/08 03:12 xqian residual
  • 246120 11/02/08 03:11 a-onl End_of_Run_2676
  • 246119 11/02/08 03:10 a-onl Start_Run_2676,Run_type=Check out,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=1500 and 3.0V ( BB Wire ch. check)
  • 246118 11/02/08 03:09 a-onl End_of_Run_2675
  • 246117 11/02/08 03:08 adev BigBite WireChamber Threshold change!
  • 246116 11/02/08 03:08 a-onl Start_Run_2675,Run_type=Check out,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=1500 and 3.5V ( BB Wire ch. check)
  • 246115 11/02/08 03:07 a-onl End_of_Run_2674
  • 246114 11/02/08 03:07 xqian BigBite Wire Chamber residual, need to recalibrate
  • 246113 11/02/08 03:07 xqian BigBite Wire Chamber residual, need to recalibrate
  • 246112 11/02/08 03:06 a-onl Start_Run_2674,Run_type=Check out,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=1500 and 4.0V ( BB Wire ch. check)
  • 246111 11/02/08 03:05 a-onl End_of_Run_2673
  • 246110 11/02/08 03:04 a-onl Start_Run_2673,Run_type=Check out,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=1500 and 4.5 V ( BB Wire ch. check)
  • 246109 11/02/08 03:03 a-onl End_of_Run_2672
  • 246108 11/02/08 03:03 a-onl Start_Run_2672,Run_type=Check out,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=1500 and 5.0 V ( BB Wire ch. check)
  • 246107 11/02/08 02:59 a-onl End_of_Run_2671
  • 246106 11/02/08 02:58 adev BigBite WireChamber Threshold change!
  • 246105 11/02/08 02:57 a-onl Start_Run_2671,Run_type=Check out,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=1600 and 5.0 V ( BB Wire ch. check)
  • 246104 11/02/08 02:56 a-onl End_of_Run_2670
  • 246103 11/02/08 02:55 a-onl Start_Run_2670,Run_type=Check out,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=1600 and 4.5 V ( BB Wire ch. check)
  • 246102 11/02/08 02:54 a-onl End_of_Run_2669
  • 246101 11/02/08 02:53 a-onl Start_Run_2669,Run_type=Check out,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=1600 and 4.0 V ( BB Wire ch. check)
  • 246100 11/02/08 02:52 Ibrahim ROC9 Rebooted Itself
  • 246099 11/02/08 02:52 a-onl End_of_Run_2668
  • 246098 11/02/08 02:51 a-onl Start_Run_2668,Run_type=Check out,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=1600 and 3.0 V ( BB Wire ch. check)
  • 246097 11/02/08 02:47 adev BigBite WireChamber Threshold change!
  • 246096 11/02/08 02:47 a-onl Start_Run_2667,Run_type=Check out,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=1600 and 2.5 V ( BB Wire ch. check)
  • 246095 11/02/08 02:46 a-onl End_of_Run_2666
  • 246094 11/02/08 02:46 a-onl Start_Run_2666,Run_type=Check out,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=1600 and 3.5 V ( BB Wire ch. check)
  • 246093 11/02/08 02:43 a-onl End_of_Run_2665
  • 246092 11/02/08 02:35 Ibrahim Beam is unstable again
  • 246091 11/02/08 02:20 xqian some thinkings
  • 246090 11/02/08 02:16 adev BigBite WireChamber Measured Voltage change!
  • 246089 11/02/08 02:14 xqian Background rates and the tracking efficiency, include multi track tracking efficiency
  • 246088 11/02/08 02:11 a-onl Start_Run_2665,Run_type=Check out,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=
  • 246087 11/02/08 02:07 Chiranjib Bb HV tripped ( #2664)
  • 246086 11/02/08 02:06 a-onl End_of_Run_2664
  • 246085 11/02/08 02:05 adev BigBite WireChamber Measured Voltage change!
  • 246075 11/02/08 01:56 a-onl Start_Run_2661,Run_type=Check out,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=
  • 246074 11/02/08 01:56 a-onl End_of_Run_2660
  • 246084 11/02/08 01:54 Chiranjib BB HV tripped ( #2663)
  • 246083 11/02/08 01:53 a-onl Start_Run_2664,Run_type=Check out,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=
  • 246073 11/02/08 01:52 adev BigBite WireChamber Measured Voltage change!
  • 246072 11/02/08 01:52 a-onl Start_Run_2660,Run_type=Test,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=10 min optics
  • 246082 11/02/08 01:49 a-onl End_of_Run_2663
  • 246071 11/02/08 01:41 adev BigBite WireChamber Measured Voltage change!
  • 246070 11/02/08 01:35 samuel move to H2 target with pressure 100PSI
  • 246081 11/02/08 01:32 Ibrahim Clock Changes (End of Daylight Saving Time)
  • 246069 11/02/08 01:32 adaq target: Spin Flip Report: It's confirmed that Q Wave Plate Direction is correct. Spin up in stepper motor control code -> Spin along holding field
  • 246068 11/02/08 01:31 a-onl End_of_Run_2659
  • 246080 11/02/08 01:30 adaq target: Spin Flip Report: Spin Flip Subsystem is now calibrated and working for Vertical Pumping.
  • 246067 11/02/08 01:24 Ibrahim Beam is stable at 2 uA.
  • 246079 11/02/08 01:23 a-onl Start_Run_2663,Run_type=Check out,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=
  • 246078 11/02/08 01:22 a-onl End_of_Run_2662
  • 246077 11/02/08 01:21 a-onl End_of_Run_2662
  • 246066 11/02/08 01:21 a-onl Start_Run_2659,Run_type=Optics,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=10 min optics
  • 246065 11/02/08 01:19 a-onl End_of_Run_2658
  • 246064 11/02/08 01:16 Ibrahim Results of the MCC Harp Scans
  • 246076 11/02/08 01:10 a-onl Start_Run_2662,Run_type=Check out,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=
  • 246063 11/02/08 01:10 adev BigBite WireChamber Measured Voltage change!
  • 246062 11/02/08 01:10 a-onl Start_Run_2658,Run_type=Optics,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=10 min optics
  • 246061 11/02/08 01:09 a-onl End_of_Run_2657
  • 246060 11/02/08 01:07 a-onl Start_Run_2657,Run_type=Optics,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=10 min optics
  • 246059 11/02/08 01:04 Ibrahim BeO Screen Shot
  • 246058 11/02/08 01:04 Ibrahim BeO Screen Shot
  • 246057 11/02/08 00:59 a-onl End_of_Run_2656
  • 246056 11/02/08 00:59 Chiranjib Spot size with Be0
  • 246055 11/02/08 00:55 a-onl Start_Run_2656,Run_type=Optics,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=BeO position check
  • 246054 11/02/08 00:54 a-onl End_of_Run_2655
  • 246053 11/02/08 00:34 a-onl Start_Run_2655,Run_type=Optics,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=
  • 246052 11/02/08 00:33 a-onl End_of_Run_2654
  • 246051 11/02/08 00:22 Xiaofeng Swing shift summary for 11/01/2008
  • 246050 11/02/08 00:09 a-onl Start_Run_2654,Run_type=Optics,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=
  • 246049 11/02/08 00:09 adaq Beam Time Accounting
  • 246048 11/02/08 00:08 a-onl End_of_Run_2653
  • 246047 11/02/08 00:06 a-onl Start_Run_2653,Run_type=Optics,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=
  • Saturday

  • 246046 11/01/08 23:53 Xiaofeng snapshot of hall A tools screen
  • 246045 11/01/08 23:51 cusanno RICH work at 2200V with 35 kHz T3
  • 246044 11/01/08 23:35 Xiaofeng MCC harp scan
  • 246043 11/01/08 23:11 a-onl End_of_Run_2652
  • 246042 11/01/08 23:07 Jin Spot ++ Screen shot
  • 246041 11/01/08 23:01 a-onl Start_Run_2652,Run_type=Optics,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=
  • 246040 11/01/08 22:59 xqian BigBite Wire Chambers trip a lot
  • 246039 11/01/08 22:56 Jin Target -> Optics (BeO)
  • 246038 11/01/08 22:55 adev BigBite WireChamber Measured Voltage change!
  • 246037 11/01/08 22:48 Yi Qiang EPR calibration of vertical pumping
  • 246036 11/01/08 22:44 adev BigBite WireChamber Measured Voltage change!
  • 246035 11/01/08 22:40 xqian N2 pressure curve for RUN2513-RUN2524
  • 246034 11/01/08 22:33 Jin Huang Target -> Reference Cell
  • 246033 11/01/08 22:32 a-onl End_of_Run_2651
  • 246032 11/01/08 22:23 brads (techs) Shimmed BB sieve pole up by +4"
  • 246031 11/01/08 22:18 a-onl Start_Run_2649,Run_type=Cosmics,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=Bigbite magnet 710 uA. RICH in BB collimator in, test DMA in ROC 15
  • 246030 11/01/08 22:17 a-onl Start_Run_2651,Run_type=Optics,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=
  • 246029 11/01/08 22:16 a-onl End_of_Run_2650
  • 246028 11/01/08 22:14 a-onl Start_Run_2650,Run_type=Optics,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=
  • 246027 11/01/08 22:03 a-onl End_of_Run_2649
  • 246026 11/01/08 22:00 brads MWDC gas supply at 400 psi
  • 246025 11/01/08 21:59 brads ROC8, 9, 10 resets plugged back in
  • 246024 11/01/08 21:58 brads BB Sieve OUT
  • 246023 11/01/08 21:50 Techs/spiege Sieve movement
  • 246022 11/01/08 21:34 Xiaofeng Hall switched to controlled access
  • 246021 11/01/08 21:32 a-onl End_of_Run_2648
  • 246020 11/01/08 21:32 a-onl Start_Run_2648,Run_type=Cosmics,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=Bigbite magnet 710 uA. RICH in BB collimator in, test DMA in ROC 15
  • 246019 11/01/08 21:13 Jin Huang MWDC Position survey with Bigbite magnets off run # 2635
  • 246018 11/01/08 21:08 A. Camsonne RICH HV to 2200V
  • 246017 11/01/08 21:05 A. Camsonne RICH screens
  • 246016 11/01/08 20:59 a-onl End_of_Run_2647
  • 246015 11/01/08 20:44 Xiaofeng beam current dropped from 2 uA to 0.5 uA from 20:30 to 20:40.
  • 246014 11/01/08 20:39 Kalyan preshower vs shower energy
  • 246013 11/01/08 20:22 a-onl Start_Run_2647,Run_type=Optics,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=Bigbite magnet 710 uA. RICH in BB collimator in
  • 246012 11/01/08 20:19 Xiaofeng bigbite collimator was installed at ~11:00 am this morning.
  • 246011 11/01/08 20:17 a-onl End_of_Run_2646
  • 246010 11/01/08 20:06 a-onl Start_Run_2646,Run_type=Optics,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=Bigbite magnet 710 uA. RICH out. BB collimator in
  • 246009 11/01/08 20:04 Xiaofeng bigbite magnet current at 710 A
  • 246008 11/01/08 19:55 a-onl End_of_Run_2645
  • 246007 11/01/08 19:46 a-onl Start_Run_2645,Run_type=Optics,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=600,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=Bigbite magnet 600 uA. RICH out
  • 246006 11/01/08 19:45 Xiaofeng Bigbite Magnet current: 600 A
  • 246005 11/01/08 19:40 a-onl End_of_Run_2644
  • 246004 11/01/08 19:28 cusanno RICH stands at 2300V
  • 246003 11/01/08 19:28 a-onl Start_Run_2644,Run_type=Optics,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=400,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=Bigbite magnet 400 uA. RICH out
  • 246002 11/01/08 19:27 Xiaofeng bigbite magnet current ramp up to 400 A
  • 246001 11/01/08 19:23 a-onl End_of_Run_2643
  • 246000 11/01/08 19:14 a-onl Start_Run_2643,Run_type=Optics,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=200,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=Bigbite magnet 200 uA. RICH out
  • 245999 11/01/08 19:13 Xiaofeng Bigbite magnet current: 200A
  • 245998 11/01/08 19:12 a-onl End_of_Run_2642
  • 245997 11/01/08 19:11 a-onl Start_Run_2642,Run_type=Optics,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=200,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=Bigbite magnet 200 uA. RICH out
  • 245996 11/01/08 19:03 a-onl End_of_Run_2641
  • 245995 11/01/08 19:01 V. Sulkosky Analysis of Bulls eye Scan
  • 245994 11/01/08 19:00 Kalyan DAQ is back again
  • 245993 11/01/08 18:51 a-onl Start_Run_2641,Run_type=Optics,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=0,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=Bigbite magnet off. RICH out
  • 245992 11/01/08 18:45 Kalyan ROC 8, 9 , 10 reboots
  • 245991 11/01/08 18:35 a-onl Start_Run_2640,Run_type=Optics,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=0,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=Bigbite magnet off. RICH out
  • 245990 11/01/08 18:30 a-onl Start_Run_2639,Run_type=Optics,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=0,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=Bigbite magnet off. RICH out
  • 245989 11/01/08 18:28 brads More ROC9 investigation next CA
  • 245988 11/01/08 18:25 a-onl Start_Run_2638,Run_type=Optics,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=0,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=Bigbite magnet off. RICH in
  • 245987 11/01/08 18:23 a-onl End_of_Run_2637
  • 245986 11/01/08 18:19 a-onl Start_Run_2637,Run_type=Optics,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=0,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=Bigbite magnet off. RICH in
  • 245985 11/01/08 18:15 Jin Target Temperature
  • 245984 11/01/08 18:15 Yi Qiang Laser absorption for vertical line
  • 245983 11/01/08 17:43 a-onl End_of_Run_2636
  • 245982 11/01/08 17:41 a-onl Start_Run_2636,Run_type=Optics,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=Bigbite magnet off
  • 245981 11/01/08 17:32 a-onl End_of_Run_2635
  • 245980 11/01/08 17:31 Xiaofeng bigbite magnet off
  • 245979 11/01/08 17:31 a-onl Start_Run_2635,Run_type=Optics,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=Bigbite magnet off
  • 245978 11/01/08 17:23 Jin Target -> Optics
  • 245977 11/01/08 17:23 a-onl End_of_Run_2634
  • 245976 11/01/08 17:20 adev BigBite WireChamber Measured Voltage change!
  • 245975 11/01/08 17:11 a-onl Start_Run_2634,Run_type=Optics,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=increase beam current to 2 uA
  • 245974 11/01/08 17:09 adev BigBite WireChamber Measured Voltage change!
  • 245973 11/01/08 17:07 a-onl End_of_Run_2633
  • 245972 11/01/08 16:55 Xiaofeng beam current increased to 2 uA
  • 245971 11/01/08 16:53 a-onl Start_Run_2633,Run_type=Optics,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=increase beam current to 2 uA
  • 245970 11/01/08 16:51 a-onl End_of_Run_2632
  • 245969 11/01/08 16:50 ycwang rich pedestal with beam in run2623
  • 245968 11/01/08 16:44 a-onl Start_Run_2632,Run_type=Optics,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=
  • 245967 11/01/08 16:41 a-onl End_of_Run_2631
  • 245966 11/01/08 16:34 ycwang rich pedestal with beam in run 2623
  • 245965 11/01/08 16:16 a-onl Start_Run_2631,Run_type=Optics,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=
  • 245964 11/01/08 16:14 a-onl End_of_Run_2630
  • 245963 11/01/08 16:04 adaq Beam Time Accounting
  • 245962 11/01/08 16:01 E. Schulte Day Shift Summary for November 1, 2008
  • 245961 11/01/08 15:56 adev BigBite WireChamber Measured Voltage change!
  • 245960 11/01/08 15:50 a-onl Start_Run_2630,Run_type=Optics,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=
  • 245959 11/01/08 15:46 adev BigBite WireChamber Measured Voltage change!
  • 245958 11/01/08 15:42 a-onl End_of_Run_2629
  • 245957 11/01/08 15:35 adev BigBite WireChamber Measured Voltage change!
  • 245956 11/01/08 15:28 a-onl Start_Run_2629,Run_type=Optics,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=
  • 245955 11/01/08 15:25 adev BigBite WireChamber Measured Voltage change!
  • 245954 11/01/08 15:19 E. Schulte BigBite Wire chamber channels tripped again.
  • 245953 11/01/08 15:15 E. Schulte MCC to send 1uA.
  • 245952 11/01/08 15:07 E. Schulte MCC has taken beam to tune to Hall B
  • 245951 11/01/08 15:07 a-onl End_of_Run_2628
  • 245950 11/01/08 15:04 adev BigBite WireChamber Measured Voltage change!
  • 245949 11/01/08 15:02 a-onl Start_Run_2628,Run_type=Optics,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=
  • 245948 11/01/08 15:00 a-onl End_of_Run_2627
  • 245947 11/01/08 15:00 a-onl Start_Run_2627,Run_type=Optics,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=
  • 245946 11/01/08 14:59 E. Schulte BigBite tripped again, and other status related issues
  • 245945 11/01/08 14:56 E. Schulte CODA end of run failed
  • 245944 11/01/08 14:54 E. Schulte Another BigBite wire chamber channel trip.
  • 245943 11/01/08 14:54 a-onl Start_Run_2626,Run_type=Optics,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=
  • 245942 11/01/08 14:53 adev BigBite WireChamber Measured Voltage change!
  • 245941 11/01/08 14:51 E. Schulte BigBite Wire Chamber channel L14.7 tripped again
  • 245940 11/01/08 14:47 a-onl End_of_Run_2625
  • 245939 11/01/08 14:37 a-onl Start_Run_2625,Run_type=Optics,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=
  • 245938 11/01/08 14:35 a-onl End_of_Run_2624
  • 245937 11/01/08 14:34 E. Schulte Beam's Back!
  • 245936 11/01/08 14:22 a-onl Start_Run_2624,Run_type=Optics,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=rich pedastal run
  • 245935 11/01/08 14:21 a-onl End_of_Run_2623
  • 245934 11/01/08 14:05 E. Schulte RICH pedestal run for Alexandre
  • 245933 11/01/08 14:04 a-onl Start_Run_2623,Run_type=Optics,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=rich pedastal run
  • 245932 11/01/08 13:53 E. Schulte MCC: Problem with an RF cavity
  • 245931 11/01/08 13:44 adaq target: Spin Flip Report: No Shift Worker Comment
  • 245930 11/01/08 13:41 a-onl End_of_Run_2622
  • 245929 11/01/08 13:11 Kalyan BB wire chamber HV
  • 245928 11/01/08 13:11 E. Schulte Beam Back, BigBit On!
  • 245927 11/01/08 13:05 a-onl Start_Run_2622,Run_type=Optics,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=
  • 245926 11/01/08 12:58 adev BigBite WireChamber Measured Voltage change!
  • 245925 11/01/08 12:53 Kalyan BB Chamber HV trips
  • 245924 11/01/08 12:50 a-onl End_of_Run_2621
  • 245923 11/01/08 12:47 adev BigBite WireChamber Measured Voltage change!
  • 245922 11/01/08 12:46 E. Schulte Hall A tools for Day Shift
  • 245921 11/01/08 12:46 Kalyan wirechamber HV tripped
  • 245920 11/01/08 12:37 a-onl Start_Run_2621,Run_type=Optics,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=
  • 245919 11/01/08 12:32 E. Schulte BigBite channels tripped again, and the Beam is Back!
  • 245918 11/01/08 12:26 adev BigBite WireChamber Measured Voltage change!
  • 245917 11/01/08 12:21 E. Schulte Another BigBite wire chamber trip
  • 245916 11/01/08 12:19 E. Schutle A few BigBite wire chamber channels tripped
  • 245915 11/01/08 12:16 adev BigBite WireChamber Measured Voltage change!
  • 245914 11/01/08 12:16 a-onl End_of_Run_2620
  • 245913 11/01/08 12:10 Katich H2 valve is currently open
  • 245912 11/01/08 12:07 Katich reference cell pressure = 99psig
  • 245911 11/01/08 12:06 a-onl Start_Run_2620,Run_type=Optics,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=pedestals
  • 245910 11/01/08 12:02 Kalyan additions to db_DL.dat
  • 245909 11/01/08 11:56 a-onl End_of_Run_2619
  • 245908 11/01/08 11:54 a-onl Start_Run_2619,Run_type=Optics,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=pedestals
  • 245907 11/01/08 11:50 E. Schulte BigBite Wire Chamber Tripped
  • 245906 11/01/08 11:47 E. Schulte Changed back to H2 Reference Cell
  • 245905 11/01/08 11:27 Katich Target moved to Optics position
  • 245904 11/01/08 11:25 E. Schulte Finished with Controlled Access & Starting with BeO target
  • 245903 11/01/08 11:16 a-onl End_of_Run_2618
  • 245902 11/01/08 11:15 a-onl Start_Run_2618,Run_type=Test,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=pedestals
  • 245901 11/01/08 11:13 a-onl End_of_Run_2617
  • 245900 11/01/08 11:05 Kalyan Fixed the prescaled T1 channel in TDC
  • 245899 11/01/08 10:55 a-onl Start_Run_2617,Run_type=Test,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=H2 optics calibration data TEST
  • 245898 11/01/08 10:54 a-onl End_of_Run_2616
  • 245897 11/01/08 10:47 V. Sulkosky BCM U10 Working
  • 245896 11/01/08 10:30 V. Sulkosky Left HRS Gas Cerenkov HV
  • 245895 11/01/08 10:21 a-onl Start_Run_2616,Run_type=Test,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=H2 optics calibration data TEST
  • 245894 11/01/08 10:20 cusanno RICH trips
  • 245893 11/01/08 10:19 E. Schulte Compton Slot 15 Channel 10 was tripped off
  • 245892 11/01/08 10:19 cusanno RICH trips
  • 245891 11/01/08 10:19 a-onl End_of_Run_2615
  • 245890 11/01/08 10:11 a-onl Start_Run_2615,Run_type=Test,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=H2 optics calibration data TEST
  • 245889 11/01/08 09:52 E. Schulte H2 Reference Cell Valve Issue
  • 245888 11/01/08 09:48 E. Schulte Controlled Access: Insert Target Collimator
  • 245887 11/01/08 09:43 a-onl End_of_Run_2614
  • 245886 11/01/08 09:39 E. Schulte Awaiting the arrival of the RC for advice on run plan
  • 245883 11/01/08 09:31 xqian ROC9 seems to be healed
  • 245882 11/01/08 09:21 a-onl Start_Run_2614,Run_type=Calibration,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=H2 optics calibration data
  • 245881 11/01/08 09:21 a-onl End_of_Run_2613
  • 245880 11/01/08 09:02 a-onl Start_Run_2613,Run_type=Calibration,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=H2 optics calibration data
  • 245879 11/01/08 09:01 a-onl End_of_Run_2612
  • 245878 11/01/08 08:42 a-onl Start_Run_2612,Run_type=Calibration,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=H2 optics calibration data
  • 245877 11/01/08 08:41 a-onl End_of_Run_2611
  • 245876 11/01/08 08:30 E. Schulte Beam Trippy
  • 245875 11/01/08 08:22 a-onl Start_Run_2611,Run_type=Calibration,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=H2 optics calibration data
  • 245874 11/01/08 08:20 a-onl End_of_Run_2610
  • 245873 11/01/08 08:14 eschulte BigBite Magnet Setting at Start of Shift
  • 245872 11/01/08 08:04 Ibrahim Owl Shift Summary
  • 245871 11/01/08 08:03 adaq Beam Time Accounting
  • 245870 11/01/08 08:01 Jin Huang Transverse pumping laser pol. measurement
  • 245869 11/01/08 07:58 a-onl Start_Run_2610,Run_type=Calibration,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=H2 optics calibration data
  • 245868 11/01/08 07:57 a-onl End_of_Run_2609
  • 245867 11/01/08 07:52 Chiranjib Couple of issues
  • 245866 11/01/08 07:47 riordan BigBite Momentum Reconstruction
  • 245865 11/01/08 07:36 Jin & Chiran EPR Pol. = 31% (still pumping)
  • 245864 11/01/08 07:35 adev BigBite WireChamber Measured Voltage change!
  • 245863 11/01/08 07:34 a-onl Start_Run_2609,Run_type=Calibration,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=H2 optics calibration data
  • 245862 11/01/08 07:33 a-onl End_of_Run_2608
  • 245861 11/01/08 07:31 xqian Wire Chamber Trip frequently when the beam is back
  • 245860 11/01/08 07:24 adev BigBite WireChamber Measured Voltage change!
  • 245859 11/01/08 07:17 a-onl Start_Run_2608,Run_type=Calibration,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=H2 optics calibration data
  • 245858 11/01/08 07:16 xqian Vertex z and the sieve pattern for the LH2 run 2605
  • 245857 11/01/08 07:14 a-onl End_of_Run_2607
  • 245856 11/01/08 07:13 adev BigBite WireChamber Measured Voltage change!
  • 245855 11/01/08 07:08 riordan Frequent BigBite MWDC Trips
  • 245854 11/01/08 07:03 adev BigBite WireChamber Measured Voltage change!
  • 245853 11/01/08 06:53 xqian Scaler T1 trigger seems to not working for the prescaled one
  • 245852 11/01/08 06:52 adev BigBite WireChamber Measured Voltage change!
  • 245851 11/01/08 06:52 a-onl Start_Run_2607,Run_type=Calibration,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=H2 optics calibration data
  • 245850 11/01/08 06:51 a-onl End_of_Run_2606
  • 245849 11/01/08 06:50 ycwang Run # 2581-2602 are all Junk
  • 245848 11/01/08 06:42 xqian F1 TDC OUTPUT FIFO overflow IN run 2604 and 2605
  • 245847 11/01/08 06:41 a-onl End_of_Run_2606
  • 245846 11/01/08 06:41 a-onl Start_Run_2606,Run_type=Calibration,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=H2 optics calibration data
  • 245845 11/01/08 06:36 riordan H2 Replays F1 FIFO overflow
  • 245844 11/01/08 06:16 xqian see the vertex too!
  • 245843 11/01/08 06:16 a-onl Start_Run_2605,Run_type=Calibration,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=H2 optics calibration data
  • 245842 11/01/08 06:15 xqian A plot with more stat.
  • 245841 11/01/08 06:15 a-onl End_of_Run_2604
  • 245840 11/01/08 06:12 a-onl Start_Run_2604,Run_type=Calibration,target_type=Reference Cell: H2,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=Finding H2 prescales
  • 245839 11/01/08 06:11 Ibrahim Beam is back
  • 245838 11/01/08 06:10 a-onl End_of_Run_2603
  • 245837 11/01/08 06:10 a-onl Start_Run_2603,Run_type=Calibration,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=pedestal run for rich with beam
  • 245836 11/01/08 06:02 adaq target: Spin Flip Report: Test
  • 245835 11/01/08 06:00 a-onl End_of_Run_2602
  • 245834 11/01/08 06:00 a-onl Start_Run_2602,Run_type=Calibration,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=cosmic data
  • 245833 11/01/08 06:00 riordan BigBite Sieve Plate
  • 245832 11/01/08 05:59 samuel move to H2(100psi) reference cell
  • 245831 11/01/08 05:56 a-onl End_of_Run_2601
  • 245830 11/01/08 05:52 Jin & Chiran EPR AFP working
  • 245829 11/01/08 05:49 a-onl Start_Run_2601,Run_type=Calibration,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=cosmic data
  • 245828 11/01/08 05:49 a-onl End_of_Run_2600
  • 245827 11/01/08 05:47 a-onl Start_Run_2600,Run_type=Calibration,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=cosmic data
  • 245826 11/01/08 05:46 a-onl End_of_Run_2599
  • 245825 11/01/08 05:45 a-onl Start_Run_2599,Run_type=Calibration,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=cosmic data
  • 245824 11/01/08 05:44 a-onl End_of_Run_2598
  • 245823 11/01/08 05:43 a-onl Start_Run_2598,Run_type=Calibration,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=cosmic data
  • 245822 11/01/08 05:41 a-onl End_of_Run_2597
  • 245821 11/01/08 05:40 a-onl Start_Run_2597,Run_type=Calibration,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=cosmic data
  • 245820 11/01/08 05:39 a-onl End_of_Run_2596
  • 245819 11/01/08 05:39 a-onl Start_Run_2596,Run_type=Calibration,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=cosmic data
  • 245818 11/01/08 05:38 a-onl End_of_Run_2595
  • 245817 11/01/08 05:32 a-onl Start_Run_2595,Run_type=Calibration,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=cosmic data
  • 245816 11/01/08 05:30 a-onl End_of_Run_2594
  • 245815 11/01/08 05:26 a-onl Start_Run_2594,Run_type=Calibration,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=test with rich in cosmic status
  • 245814 11/01/08 05:24 a-onl End_of_Run_2593
  • 245813 11/01/08 05:24 a-onl Start_Run_2593,Run_type=Calibration,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=test
  • 245812 11/01/08 05:22 a-onl End_of_Run_2592
  • 245811 11/01/08 05:21 Ibrahim Controlled Access
  • 245810 11/01/08 05:21 Ibrahim ROCs 8,9,10 cannot be rebooted remotely
  • 245809 11/01/08 05:18 a-onl Start_Run_2592,Run_type=Calibration,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=test
  • 245808 11/01/08 05:17 a-onl End_of_Run_2591
  • 245807 11/01/08 05:16 a-onl Start_Run_2591,Run_type=Calibration,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=test
  • 245806 11/01/08 05:09 a-onl Start_Run_2590,Run_type=Calibration,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=go back to
  • 245805 11/01/08 05:05 Jin & Chiran Got EPR signal during FM Sweep
  • 245804 11/01/08 04:59 a-onl Start_Run_2588,Run_type=Calibration,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=pedestal run for RICH with Beam (last run ROC 9 stuck)
  • 245803 11/01/08 04:54 a-onl Start_Run_2587,Run_type=Calibration,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=pedestal run for RICH with Beam (last run ROC 9 stuck)
  • 245802 11/01/08 04:49 a-onl Start_Run_2586,Run_type=Calibration,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=pedestal run for RICH with Beam
  • 245801 11/01/08 04:37 xqian BBNORM class is half working
  • 245800 11/01/08 04:37 a-onl End_of_Run_2585
  • 245799 11/01/08 04:25 riordan ROC 9 problems
  • 245798 11/01/08 04:23 a-onl Start_Run_2585,Run_type=Calibration,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=Optics Foils - Raster Off
  • 245797 11/01/08 04:20 a-onl Start_Run_2584,Run_type=Calibration,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=Optics Foils - Raster Off
  • 245796 11/01/08 04:17 a-onl Start_Run_2583,Run_type=Calibration,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=Optics Foils - Raster Off
  • 245795 11/01/08 04:15 a-onl End_of_Run_2582
  • 245794 11/01/08 04:05 adev BigBite WireChamber Measured Voltage change!
  • 245793 11/01/08 04:05 Jin & Chiran NMR signal after vertical pumping
  • 245792 11/01/08 04:05 a-onl Start_Run_2582,Run_type=Calibration,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=Optics Foils - Raster 4x4
  • 245791 11/01/08 04:02 Ibrahim BeO Spot Shot
  • 245790 11/01/08 04:01 riordan ROC 9 Stopped Responding
  • 245789 11/01/08 04:00 Ibrahim Tools Screen
  • 245788 11/01/08 03:57 a-onl Start_Run_2581,Run_type=Check out,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=Optics Foils - Raster 4x4
  • 245787 11/01/08 03:56 a-onl End_of_Run_2580
  • 245786 11/01/08 03:54 adev BigBite WireChamber Measured Voltage change!
  • 245785 11/01/08 03:52 a-onl Start_Run_2580,Run_type=Check out,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=Check prescales
  • 245784 11/01/08 03:45 riordan Carbon Hole spot++
  • 245783 11/01/08 03:44 a-onl End_of_Run_2579
  • 245782 11/01/08 03:39 a-onl Start_Run_2579,Run_type=Check out,target_type=Single Carbon,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=cosmic data BB SC Hv change based on run 2574
  • 245781 11/01/08 03:33 a-onl End_of_Run_2578
  • 245780 11/01/08 03:30 jpchen He3 gas bottle valve is open, ready for reference cell run
  • 245779 11/01/08 03:29 a-onl Start_Run_2578,Run_type=Check out,target_type=Single Carbon,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=cosmic data BB SC Hv change based on run 2574
  • 245778 11/01/08 03:29 Chiranjib Laser absorption for vertical line
  • 245777 11/01/08 03:26 jpchen target work
  • 245776 11/01/08 03:25 samuel move to optics hole position.
  • 245775 11/01/08 03:25 adaq spot++ on BeO
  • 245774 11/01/08 03:23 a-onl End_of_Run_2577
  • 245773 11/01/08 03:21 Chiranjib Vertical Laser Pumping started
  • 245772 11/01/08 03:21 Chiranjib Vertical Laser Pumping started
  • 245771 11/01/08 03:09 a-onl Start_Run_2577,Run_type=Check out,target_type=Multi Carbon,beam_e0=1.23060,beam_raster=off,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=sieve,bbarm_zoff=0,comment_text=cosmic data BB SC Hv change based on run 2574
  • 245770 11/01/08 03:06 a-onl End_of_Run_2576
  • 245769 11/01/08 03:06 a-onl Start_Run_2576,Run_type=Cosmics,target_type=Reference Cell: 3He,beam_e0=1.23060,beam_raster=on,leftarm_p=1.21269,leftarm_theta=16,leftarm_sieve=out,bbarm_i=710,bbarm_theta=30,bbarm_sieve=out,bbarm_zoff=0,comment_text=cosmic data BB SC Hv change based on run 2574
  • 245768 11/01/08 02:53 samuel move to optics position
  • 245767 11/01/08 02:51 a-onl End_of_Run_2575
  • 245766 11/01/08 00:36 adev BigBite WireChamber Measured Voltage change!
  • 245765 11/01/08 00:08 adaq Beam Time Accounting
  • 245764 11/01/08 00:05 Gao swing shift summary

  • Earlier months at hallaweb.jlab.org/adaq (archive).  The HALOG was generated by version HA1.42
    If you have any comments or problems, please contact:
    Robert Michaels
    rom@jlab.org
    (757) 269-7410